Autoconf

Creating Automatic Configuration Scripts for version 2.66, 2 July 2010

David MacKenzie Ben Elliston Akim Demaille

This manual (2 July 2010) is for GNU Autoconf (version 2.66), a package for creating scripts to configure source code packages using templates and an M4 macro package. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1992, 1993, 1994, 1995, 1996, 1998, 1999, 2000, 2001, 2002, 2003, 2004, 2005, 2006, 2007, 2008, 2009, 2010 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, with the Front-Cover texts being “A GNU Manual,” and with the Back-Cover Texts as in (a) below. A copy of the license is included in the section entitled “GNU Free Documentation License.”

(a) The FSF’s Back-Cover Text is: “You have the freedom to copy and modify this GNU manual. Buying copies from the FSF supports it in developing GNU and promoting software freedom.”

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 The GNU Build System. . . . . . . . . . . . . . . . . . . . . . . . 3

2.1 Automake. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Gnulib . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.3 Libtool. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.4 Pointers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3 Making configure Scripts . . . . . . . . . . . . . . . . . . . . . . 7

3.1 Writing ‘configure.ac’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.1.1 A Shell Script Compiler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.1.2 The Autoconf Language . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.1.3 Standard ‘configure.ac’ Layout . . . . . . . . . . . . . . . . . . . . . . . . . 11

3.2 Using autoscan to Create ‘configure.ac’. . . . . . . . . . . . . . . . . . . . . 12 3.3 Using ifnames to List Conditionals. . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.4 Using autoconf to Create configure . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.5 Using autoreconf to Update configure Scripts . . . . . . . . . . . . . . . 15

4 Initialization and Output Files . . . . . . . . . . . . . . . 19

4.1 Initializing configure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 4.2 Dealing with Autoconf versions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 4.3 Notices in configure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 4.4 Finding configure Input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.5 Outputting Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.6 Performing Configuration Actions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.7 Creating Configuration Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 4.8 Substitutions in Makefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

4.8.1 Preset Output Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 4.8.2 Installation Directory Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . 29 4.8.3 Changed Directory Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 4.8.4 Build Directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.8.5 Automatic Remaking. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

4.9 Configuration Header Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 4.9.1 Configuration Header Templates. . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.9.2 Using autoheader to Create ‘config.h.in’. . . . . . . . . . . . . . . 37 4.9.3 Autoheader Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

4.10 Running Arbitrary Configuration Commands . . . . . . . . . . . . . . . . . 39 4.11 Creating Configuration Links. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 4.12 Configuring Other Packages in Subdirectories. . . . . . . . . . . . . . . . . 40 4.13 Default Prefix. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

5 Existing Tests. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

5.1 Common Behavior . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 5.1.1 Standard Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 5.1.2 Default Includes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

5.2 Alternative Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.2.1 Particular Program Checks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.2.2 Generic Program and File Checks . . . . . . . . . . . . . . . . . . . . . . . . 48

5.3 Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 5.4 Library Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 5.5 Library Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

5.5.1 Portability of C Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 5.5.2 Particular Function Checks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 5.5.3 Generic Function Checks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

5.6 Header Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 5.6.1 Portability of Headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 5.6.2 Particular Header Checks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 5.6.3 Generic Header Checks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

5.7 Declarations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 5.7.1 Particular Declaration Checks . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 5.7.2 Generic Declaration Checks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

5.8 Structures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 5.8.1 Particular Structure Checks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 5.8.2 Generic Structure Checks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

5.9 Types. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 5.9.1 Particular Type Checks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 5.9.2 Generic Type Checks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

5.10 Compilers and Preprocessors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 5.10.1 Specific Compiler Characteristics . . . . . . . . . . . . . . . . . . . . . . . . 79 5.10.2 Generic Compiler Characteristics . . . . . . . . . . . . . . . . . . . . . . . . 79 5.10.3 C Compiler Characteristics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 5.10.4 C++ Compiler Characteristics . . . . . . . . . . . . . . . . . . . . . . . . . . 87 5.10.5 Objective C Compiler Characteristics. . . . . . . . . . . . . . . . . . . . 87 5.10.6 Objective C++ Compiler Characteristics . . . . . . . . . . . . . . . . 88 5.10.7 Erlang Compiler and Interpreter Characteristics . . . . . . . . . 88 5.10.8 Fortran Compiler Characteristics . . . . . . . . . . . . . . . . . . . . . . . . 89

5.11 System Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 5.12 Posix Variants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 5.13 Erlang Libraries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

6 Writing Tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

6.1 Language Choice . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 6.2 Writing Test Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

6.2.1 Guidelines for Test Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 6.2.2 Test Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 6.2.3 Generating Sources. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

6.3 Running the Preprocessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 6.4 Running the Compiler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 6.5 Running the Linker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 6.6 Checking Runtime Behavior . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 6.7 Systemology . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 6.8 Multiple Cases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109

7 Results of Tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

7.1 Defining C Preprocessor Symbols. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 7.2 Setting Output Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 7.3 Special Characters in Output Variables. . . . . . . . . . . . . . . . . . . . . . . 114 7.4 Caching Results. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115

7.4.1 Cache Variable Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 7.4.2 Cache Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 7.4.3 Cache Checkpointing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117

7.5 Printing Messages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118

8 Programming in M4 . . . . . . . . . . . . . . . . . . . . . . . . . 121

8.1 M4 Quotation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 8.1.1 Active Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 8.1.2 One Macro Call . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 8.1.3 Quoting and Parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 8.1.4 Quotation and Nested Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 8.1.5 changequote is Evil. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 8.1.6 Quadrigraphs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 8.1.7 Dealing with unbalanced parentheses . . . . . . . . . . . . . . . . . . . . 127 8.1.8 Quotation Rule Of Thumb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

8.2 Using autom4te. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 8.2.1 Invoking autom4te . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 8.2.2 Customizing autom4te . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134

8.3 Programming in M4sugar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 8.3.1 Redefined M4 Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 8.3.2 Diagnostic messages from M4sugar . . . . . . . . . . . . . . . . . . . . . . 138 8.3.3 Diversion support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 8.3.4 Conditional constructs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 8.3.5 Looping constructs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 8.3.6 Evaluation Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 8.3.7 String manipulation in M4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 8.3.8 Arithmetic computation in M4 . . . . . . . . . . . . . . . . . . . . . . . . . . 156 8.3.9 Set manipulation in M4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 8.3.10 Forbidden Patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162

8.4 Debugging via autom4te . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162

9 Programming in M4sh . . . . . . . . . . . . . . . . . . . . . . . 165

9.1 Common Shell Constructs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 9.2 Support for indirect variable names . . . . . . . . . . . . . . . . . . . . . . . . . . 168 9.3 Initialization Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 9.4 File Descriptor Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172

10 Writing Autoconf Macros . . . . . . . . . . . . . . . . . . 175

10.1 Macro Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 10.2 Macro Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 10.3 Reporting Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 10.4 Dependencies Between Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178

10.4.1 Prerequisite Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 10.4.2 Suggested Ordering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 10.4.3 One-Shot Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181

10.5 Obsoleting Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 10.6 Coding Style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182

11 Portable Shell Programming . . . . . . . . . . . . . . . 187

11.1 Shellology . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 11.2 Here-Documents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 11.3 File Descriptors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 11.4 File System Conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 11.5 Shell Pattern Matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 11.6 Shell Substitutions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 11.7 Assignments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 11.8 Parentheses in Shell Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 11.9 Slashes in Shell Scripts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 11.10 Special Shell Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 11.11 Shell Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 11.12 Limitations of Shell Builtins . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 11.13 Limitations of Usual Tools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224

12 Portable Make Programming . . . . . . . . . . . . . . 241

12.1 $< in Ordinary Make Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 12.2 Failure in Make Rules. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 12.3 Special Characters in Make Macro Names . . . . . . . . . . . . . . . . . . . 241 12.4 Backslash-Newline-Newline in Make Macro Values . . . . . . . . . . . 242 12.5 Backslash-Newline in Make Comments . . . . . . . . . . . . . . . . . . . . . . 242 12.6 Long Lines in Makefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 12.7 make macro=value and Submakes . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 12.8 The Make Macro MAKEFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 12.9 The Make Macro SHELL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 12.10 Parallel Make. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 12.11 Comments in Make Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 12.12 Newlines in Make Rules. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 12.13 The ‘obj/’ Subdirectory and Make . . . . . . . . . . . . . . . . . . . . . . . . . 247 12.14 Exit Status of make -k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 12.15 VPATH and Make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247

12.15.1 Variables listed in VPATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 12.15.2 VPATH and Double-colon Rules . . . . . . . . . . . . . . . . . . . . . . . . 248 12.15.3 $< Not Supported in Explicit Rules . . . . . . . . . . . . . . . . . . . 248 12.15.4 Automatic Rule Rewriting . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 12.15.5 Tru64 make Creates Prerequisite Directories Magically

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 12.15.6 Make Target Lookup. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251

12.16 Single Suffix Rules and Separated Dependencies . . . . . . . . . . . . 254 12.17 Timestamp Resolution and Make . . . . . . . . . . . . . . . . . . . . . . . . . . 254

13 Portable C and C++ Programming . . . . . . . 257

13.1 Varieties of Unportability . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 13.2 Integer Overflow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258

13.2.1 Basics of Integer Overflow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 13.2.2 Examples of Code Assuming Wraparound Overflow . . . . . 258 13.2.3 Optimizations That Break Wraparound Arithmetic . . . . . 260 13.2.4 Practical Advice for Signed Overflow Issues. . . . . . . . . . . . . 261 13.2.5 Signed Integer Division and Integer Overflow . . . . . . . . . . . 262

13.3 Preprocessor Arithmetic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 13.4 Properties of Null Pointers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 13.5 Buffer Overruns and Subscript Errors . . . . . . . . . . . . . . . . . . . . . . . 262 13.6 Volatile Objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 13.7 Floating Point Portability. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 13.8 Exiting Portably . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265

14 Manual Configuration . . . . . . . . . . . . . . . . . . . . . . 267

14.1 Specifying target triplets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 14.2 Getting the Canonical System Type . . . . . . . . . . . . . . . . . . . . . . . . . 268 14.3 Using the System Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269

15 Site Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . 271

15.1 Controlling Help Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 15.2 Working With External Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 15.3 Choosing Package Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 15.4 Making Your Help Strings Look Pretty . . . . . . . . . . . . . . . . . . . . . . 274 15.5 Controlling Checking of configure Options . . . . . . . . . . . . . . . . . 275 15.6 Configuring Site Details. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 15.7 Transforming Program Names When Installing . . . . . . . . . . . . . . 276

15.7.1 Transformation Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 15.7.2 Transformation Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 15.7.3 Transformation Rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277

15.8 Setting Site Defaults . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278

16 Running configure Scripts . . . . . . . . . . . . . . . . . 281

16.1 Basic Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 16.2 Compilers and Options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 16.3 Compiling For Multiple Architectures . . . . . . . . . . . . . . . . . . . . . . . 282 16.4 Installation Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 16.5 Optional Features. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 16.6 Particular systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 16.7 Specifying the System Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 16.8 Sharing Defaults . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 16.9 Defining Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 16.10 configure Invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285

17 config.status Invocation . . . . . . . . . . . . . . . . . . . . 287

18 Obsolete Constructs . . . . . . . . . . . . . . . . . . . . . . . . 291

18.1 Obsolete ‘config.status’ Invocation . . . . . . . . . . . . . . . . . . . . . . . 291 18.2 ‘acconfig.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 18.3 Using autoupdate to Modernize ‘configure.ac’ . . . . . . . . . . . . 292 18.4 Obsolete Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 18.5 Upgrading From Version 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306

18.5.1 Changed File Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 18.5.2 Changed Makefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 18.5.3 Changed Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 18.5.4 Changed Results . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 18.5.5 Changed Macro Writing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308

18.6 Upgrading From Version 2.13 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 18.6.1 Changed Quotation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 18.6.2 New Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 18.6.3 Hosts and Cross-Compilation. . . . . . . . . . . . . . . . . . . . . . . . . . . 310 18.6.4 AC_LIBOBJ vs. LIBOBJS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 18.6.5 AC_ACT_IFELSE vs. AC_TRY_ACT . . . . . . . . . . . . . . . . . . . . . . . . 312

19 Generating Test Suites with Autotest. . . . . 315

19.1 Using an Autotest Test Suite. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 19.1.1 testsuite Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 19.1.2 Autotest Logs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317

19.2 Writing ‘testsuite.at’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 19.3 Running testsuite Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 19.4 Making testsuite Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325

20 Frequent Autoconf Questions, with answers

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329

20.1 Distributing configure Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 20.2 Why Require GNU M4? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 20.3 How Can I Bootstrap? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 20.4 Why Not Imake? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 330 20.5 How Do I #define Installation Directories? . . . . . . . . . . . . . . . . . 331 20.6 What is ‘autom4te.cache’? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 332 20.7 Header Present But Cannot Be Compiled . . . . . . . . . . . . . . . . . . . 332 20.8 Expanded Before Required . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 20.9 Debugging configure scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 336

21 History of Autoconf . . . . . . . . . . . . . . . . . . . . . . . . 339

21.1 Genesis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 21.2 Exodus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 21.3 Leviticus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 340 21.4 Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 340 21.5 Deuteronomy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 341

Appendix A GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 343

Appendix B Indices . . . . . . . . . . . . . . . . . . . . . . . . . . . 351

B.1 Environment Variable Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 351 B.2 Output Variable Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 352 B.3 Preprocessor Symbol Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 354 B.4 Cache Variable Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 357 B.5 Autoconf Macro Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 B.6 M4 Macro Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 B.7 Autotest Macro Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 365 B.8 Program and Function Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 365 B.9 Concept Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 368

1 Introduction

A physicist, an engineer, and a computer scientist were discussing the nature of God. “Surely a Physicist,” said the physicist, “because early in the Creation, God made Light; and you know, Maxwell’s equations, the dual nature of electromagnetic waves, the relativistic consequences…” “An Engineer!,” said the engineer, “because

before making Light, God split the Chaos into Land and Water; it takes a

hell of an engineer to handle that big amount of mud, and orderly

separation of solids from liquids…” The computer scientist

shouted: “And the Chaos, where do you think it was coming from, hmm?”

—Anonymous

Autoconf is a tool for producing shell scripts that automatically configure software source code packages to adapt to many kinds of Posix-like systems. The configuration scripts produced by Autoconf are independent of Autoconf when they are run, so their users do not need to have Autoconf.

The configuration scripts produced by Autoconf require no manual user intervention when run; they do not normally even need an argument specifying the system type. Instead, they individually test for the presence of each feature that the software package they are for might need. (Before each check, they print a one-line message stating what they are checking for, so the user doesn’t get too bored while waiting for the script to finish.) As a result, they deal well with systems that are hybrids or customized from the more common Posix variants. There is no need to maintain files that list the features supported by each release of each variant of Posix.

For each software package that Autoconf is used with, it creates a configuration script from a template file that lists the system features that the package needs or can use. After the shell code to recognize and respond to a system feature has been written, Autoconf allows it to be shared by many software packages that can use (or need) that feature. If it later turns out that the shell code needs adjustment for some reason, it needs to be changed in only one place; all of the configuration scripts can be regenerated automatically to take advantage of the updated code.

Those who do not understand Autoconf are condemned to reinvent it, poorly. The primary goal of Autoconf is making the user’s life easier; making the maintainer’s life easier is only a secondary goal. Put another way, the primary goal is not to make the generation of ‘configure’ automatic for package maintainers (although patches along that front are welcome, since package maintainers form the user base of Autoconf); rather, the goal is to make ‘configure’ painless, portable, and predictable for the end user of each autoconfiscated package. And to this degree, Autoconf is highly successful at its goal — most complaints to the Autoconf list are about difficulties in writing Autoconf input, and not in the behavior of the resulting ‘configure’. Even packages that don’t use Autoconf will generally provide a ‘configure’ script, and the most common complaint about these alternative home-grown scripts is that they fail to meet one or more of the GNU Coding Standards that users have come to expect from Autoconf-generated ‘configure’ scripts.

The Metaconfig package is similar in purpose to Autoconf, but the scripts it produces require manual user intervention, which is quite inconvenient when configuring large source trees. Unlike Metaconfig scripts, Autoconf scripts can support cross-compiling, if some care is taken in writing them.

Autoconf does not solve all problems related to making portable software packages—for a more complete solution, it should be used in concert with other GNU build tools like Automake and Libtool. These other tools take on jobs like the creation of a portable, recursive makefile with all of the standard targets, linking of shared libraries, and so on. See Chapter 2 [The GNU Build System], page 3, for more information.

Autoconf imposes some restrictions on the names of macros used with #if in C programs (see Section B.3 [Preprocessor Symbol Index], page 354).

Autoconf requires GNU M4 version 1.4.6 or later in order to generate the scripts. It uses features that some versions of M4, including GNU M4 1.3, do not have. Autoconf works better with GNU M4 version 1.4.14 or later, though this is not required.

See Section 18.5 [Autoconf 1], page 306, for information about upgrading from version 1. See Chapter 21 [History], page 339, for the story of Autoconf’s development. See Chapter 20 [FAQ], page 329, for answers to some common questions about Autoconf.

   See the Autoconf web page (http://www.gnu.org/software/autoconf/) for up-to-date

information, details on the mailing lists, pointers to a list of known bugs, etc.

Mail suggestions to the Autoconf mailing list (autoconf@gnu.org). Past suggestions are

archived (http://lists.gnu.org/archive/html/autoconf/).

Mail bug reports to the Autoconf Bugs mailing list (bug-autoconf@gnu.org). Past bug

reports are archived (http://lists.gnu.org/archive/html/bug-autoconf/).

If possible, first check that your bug is not already solved in current development versions, and that it has not been reported yet. Be sure to include all the needed information and a short ‘configure.ac’ that demonstrates the problem.

Autoconf’s development tree is accessible via git; see the Autoconf Summary (http:// savannah.gnu.org/projects/autoconf/) for details, or view the actual repository

(http://git.sv.gnu.org/gitweb/?p=autoconf.git). Anonymous CVS access is also

available, see ‘README’ for more details. Patches relative to the current git version can be sent for review to the Autoconf Patches mailing list (autoconf-patches@gnu.org), with

discussion on prior patches archived (http://lists.gnu.org/archive/html/autoconf-

patches/); and all commits are posted in the read-only Autoconf Commit mailing list

(autoconf-commit@gnu.org), which is also archived (http://lists.gnu.org/archive/
html/autoconf-commit/).

Because of its mission, the Autoconf package itself includes only a set of often-used macros that have already demonstrated their usefulness. Nevertheless, if you wish to share your macros, or find existing ones, see the Autoconf Macro Archive (http://www.gnu.org/

software/autoconf-archive/), which is kindly run by Peter Simons (simons@cryp.to).

2 The GNU Build System

Autoconf solves an important problem—reliable discovery of system-specific build and runtime information—but this is only one piece of the puzzle for the development of portable software. To this end, the GNU project has developed a suite of integrated utilities to finish the job Autoconf started: the GNU build system, whose most important components are Autoconf, Automake, and Libtool. In this chapter, we introduce you to those tools, point you to sources of more information, and try to convince you to use the entire GNU build system for your software.

2.1 Automake

The ubiquity of make means that a makefile is almost the only viable way to distribute automatic build rules for software, but one quickly runs into its numerous limitations. Its lack of support for automatic dependency tracking, recursive builds in subdirectories, reliable timestamps (e.g., for network file systems), and so on, mean that developers must painfully (and often incorrectly) reinvent the wheel for each project. Portability is nontrivial, thanks to the quirks of make on many systems. On top of all this is the manual labor required to implement the many standard targets that users have come to expect

(make install, make distclean, make uninstall, etc.). Since you are, of course, using

Autoconf, you also have to insert repetitive code in your ‘Makefile.in’ to recognize @CC@, @CFLAGS@, and other substitutions provided by configure. Into this mess steps Automake.

Automake allows you to specify your build needs in a ‘Makefile.am’ file with a vastly simpler and more powerful syntax than that of a plain makefile, and then generates a portable ‘Makefile.in’ for use with Autoconf. For example, the ‘Makefile.am’ to build and install a simple “Hello world” program might look like:

     bin_PROGRAMS = hello
     hello_SOURCES = hello.c

The resulting ‘Makefile.in’ (~400 lines) automatically supports all the standard targets, the substitutions provided by Autoconf, automatic dependency tracking, VPATH building, and so on. make builds the hello program, and make install installs it in ‘/usr/local/bin’ (or whatever prefix was given to configure, if not ‘/usr/local’).

The benefits of Automake increase for larger packages (especially ones with subdirectories), but even for small programs the added convenience and portability can be substantial. And that’s not all…

2.2 Gnulib

GNU software has a well-deserved reputation for running on many different types of systems. While our primary goal is to write software for the GNU system, many users and developers have been introduced to us through the systems that they were already using.

Gnulib is a central location for common GNU code, intended to be shared among free software packages. Its components are typically shared at the source level, rather than being a library that gets built, installed, and linked against. The idea is to copy files from Gnulib into your own source tree. There is no distribution tarball; developers should just grab source modules from the repository. The source files are available online, under various licenses, mostly GNU GPL or GNU LGPL.

Gnulib modules typically contain C source code along with Autoconf macros used to configure the source code. For example, the Gnulib stdbool module implements a ‘stdbool.h’ header that nearly conforms to C99, even on old-fashioned hosts that lack ‘stdbool.h’. This module contains a source file for the replacement header, along with an Autoconf macro that arranges to use the replacement header on old-fashioned systems.

2.3 Libtool

Often, one wants to build not only programs, but libraries, so that other programs can benefit from the fruits of your labor. Ideally, one would like to produce shared (dynamically linked) libraries, which can be used by multiple programs without duplication on disk or in memory and can be updated independently of the linked programs. Producing shared libraries portably, however, is the stuff of nightmares—each system has its own incompatible tools, compiler flags, and magic incantations. Fortunately, GNU provides a solution: Libtool.

Libtool handles all the requirements of building shared libraries for you, and at this time seems to be the only way to do so with any portability. It also handles many other headaches, such as: the interaction of Make rules with the variable suffixes of shared libraries, linking reliably with shared libraries before they are installed by the superuser, and supplying a consistent versioning system (so that different versions of a library can be installed or upgraded without breaking binary compatibility). Although Libtool, like Autoconf, can be used without Automake, it is most simply utilized in conjunction with Automake—there, Libtool is used automatically whenever shared libraries are needed, and you need not know its syntax.

2.4 Pointers

Developers who are used to the simplicity of make for small projects on a single system might be daunted at the prospect of learning to use Automake and Autoconf. As your software is distributed to more and more users, however, you otherwise quickly find yourself putting lots of effort into reinventing the services that the GNU build tools provide, and making the same mistakes that they once made and overcame. (Besides, since you’re already learning Autoconf, Automake is a piece of cake.)

There are a number of places that you can go to for more information on the GNU build tools.

− Web

    The project home pages for Autoconf (http://www.gnu.org/software/autoconf/),
    Automake (http://www.gnu.org/software/automake/), Gnulib (http://www.gnu
    .org/software/gnulib/), and Libtool (http://www.gnu.org/software/libtool/).

− Automake Manual

See Section “Automake” in GNU Automake, for more information on Automake.

− Books The book GNU Autoconf, Automake and Libtool^{1} describes the complete GNU build environment. You can also find the entire book on-line (http://sources.redhat.com/

autobook/).

1 GNU Autoconf, Automake and Libtool, by G. V. Vaughan, B. Elliston, T. Tromey, and I. L. Taylor.

SAMS (originally New Riders), 2000, ISBN 1578701902.

3 Making configure Scripts

The configuration scripts that Autoconf produces are by convention called configure. When run, configure creates several files, replacing configuration parameters in them with appropriate values. The files that configure creates are:

− one or more ‘Makefile’ files, usually one in each subdirectory of the package (see

Section 4.8 [Makefile Substitutions], page 25); − optionally, a C header file, the name of which is configurable, containing #define

directives (see Section 4.9 [Configuration Headers], page 35); − a shell script called ‘config.status’ that, when run, recreates the files listed above

(see Chapter 17 [config.status Invocation], page 287); − an optional shell script normally called ‘config.cache’ (created when using ‘configure

--config-cache’) that saves the results of running many of the tests (see Section 7.4.2 [Cache Files], page 117);

− a file called ‘config.log’ containing any messages produced by compilers, to help debugging if configure makes a mistake.

To create a configure script with Autoconf, you need to write an Autoconf input file ‘configure.ac’ (or ‘configure.in’) and run autoconf on it. If you write your own feature tests to supplement those that come with Autoconf, you might also write files called ‘aclocal.m4’ and ‘acsite.m4’. If you use a C header file to contain #define directives, you might also run autoheader, and you can distribute the generated file ‘config.h.in’ with the package.

Here is a diagram showing how the files that can be used in configuration are produced. Programs that are executed are suffixed by ‘*’. Optional files are enclosed in square brackets (‘[]’). autoconf and autoheader also read the installed Autoconf macro files (by reading

‘autoconf.m4’).

Files used in preparing a software package for distribution, when using just Autoconf:

     your source files --> [autoscan*] --> [configure.scan] --> configure.ac
     configure.ac --.
                    |   .------> autoconf* -----> configure
     [aclocal.m4] --+---+
                    |   `-----> [autoheader*] --> [config.h.in]
     [acsite.m4] ---'
     Makefile.in

Additionally, if you use Automake, the following additional productions come into play:

     [acinclude.m4] --.
                      |
     [local macros] --+--> aclocal* --> aclocal.m4
                      |
     configure.ac ----'

     configure.ac --.
                    +--> automake* --> Makefile.in
     Makefile.am ---'

Files used in configuring a software package:

                            .-------------> [config.cache]
     configure* ------------+-------------> config.log
                            |
     [config.h.in] -.       v            .-> [config.h] -.
                    +--> config.status* -+               +--> make*
     Makefile.in ---'                    `-> Makefile ---'
3.1 Writing ‘configure.ac’

To produce a configure script for a software package, create a file called ‘configure.ac’ that contains invocations of the Autoconf macros that test the system features your package needs or can use. Autoconf macros already exist to check for many features; see Chapter 5 [Existing Tests], page 43, for their descriptions. For most other features, you can use Autoconf template macros to produce custom checks; see Chapter 6 [Writing Tests], page 99, for information about them. For especially tricky or specialized features, ‘configure.ac’ might need to contain some hand-crafted shell commands; see Chapter 11 [Portable Shell Programming], page 187. The autoscan program can give you a good start in writing ‘configure.ac’ (see Section 3.2 [autoscan Invocation], page 12, for more information).

Previous versions of Autoconf promoted the name ‘configure.in’, which is somewhat ambiguous (the tool needed to process this file is not described by its extension), and introduces a slight confusion with ‘config.h.in’ and so on (for which ‘.in’ means “to be processed by configure”). Using ‘configure.ac’ is now preferred.

3.1.1 A Shell Script Compiler

Just as for any other computer language, in order to properly program ‘configure.ac’ in Autoconf you must understand what problem the language tries to address and how it does so.

The problem Autoconf addresses is that the world is a mess. After all, you are using Autoconf in order to have your package compile easily on all sorts of different systems, some of them being extremely hostile. Autoconf itself bears the price for these differences: configure must run on all those systems, and thus configure must limit itself to their lowest common denominator of features.

Naturally, you might then think of shell scripts; who needs autoconf? A set of properly written shell functions is enough to make it easy to write configure scripts by hand. Sigh! Unfortunately, even in 2008, where shells without any function support are far and few between, there are pitfalls to avoid when making use of them. Also, finding a Bourne shell that accepts shell functions is not trivial, even though there is almost always one on interesting porting targets.

So, what is really needed is some kind of compiler, autoconf, that takes an Autoconf program, ‘configure.ac’, and transforms it into a portable shell script, configure.

How does autoconf perform this task?

There are two obvious possibilities: creating a brand new language or extending an existing one. The former option is attractive: all sorts of optimizations could easily be implemented in the compiler and many rigorous checks could be performed on the Autoconf program (e.g., rejecting any non-portable construct). Alternatively, you can extend an existing language, such as the sh (Bourne shell) language.

Autoconf does the latter: it is a layer on top of sh. It was therefore most convenient to implement autoconf as a macro expander: a program that repeatedly performs macro expansions on text input, replacing macro calls with macro bodies and producing a pure sh script in the end. Instead of implementing a dedicated Autoconf macro expander, it is natural to use an existing general-purpose macro language, such as M4, and implement the extensions as a set of M4 macros.

3.1.2 The Autoconf Language

The Autoconf language differs from many other computer languages because it treats actual code the same as plain text. Whereas in C, for instance, data and instructions have different syntactic status, in Autoconf their status is rigorously the same. Therefore, we need a means to distinguish literal strings from text to be expanded: quotation.

When calling macros that take arguments, there must not be any white space between the macro name and the open parenthesis.

     AC_INIT ([oops], [1.0]) # incorrect
     AC_INIT([hello], [1.0]) # good

Arguments should be enclosed within the quote characters ‘[’ and ‘]’, and be separated by commas. Any leading blanks or newlines in arguments are ignored, unless they are quoted. You should always quote an argument that might contain a macro name, comma, parenthesis, or a leading blank or newline. This rule applies recursively for every macro call, including macros called from other macros. For more details on quoting rules, see Chapter 8 [Programming in M4], page 121.

For instance:

     AC_CHECK_HEADER([stdio.h],
                     [AC_DEFINE([HAVE_STDIO_H], [1],
                        [Define to 1 if you have <stdio.h>.])],
                     [AC_MSG_ERROR([sorry, can't do anything for you])])

is quoted properly. You may safely simplify its quotation to:

     AC_CHECK_HEADER([stdio.h],
                     [AC_DEFINE([HAVE_STDIO_H], 1,
                        [Define to 1 if you have <stdio.h>.])],
                     [AC_MSG_ERROR([sorry, can't do anything for you])])

because ‘1’ cannot contain a macro call. Here, the argument of AC_MSG_ERROR must be quoted; otherwise, its comma would be interpreted as an argument separator. Also, the second and third arguments of ‘AC_CHECK_HEADER’ must be quoted, since they contain macro calls. The three arguments ‘HAVE_STDIO_H’, ‘stdio.h’, and ‘Define to 1 if you have <stdio.h>.’ do not need quoting, but if you unwisely defined a macro with a name like ‘Define’ or ‘stdio’ then they would need quoting. Cautious Autoconf users would keep the quotes, but many Autoconf users find such precautions annoying, and would rewrite the example as follows:

     AC_CHECK_HEADER(stdio.h,
                     [AC_DEFINE(HAVE_STDIO_H, 1,
                        [Define to 1 if you have <stdio.h>.])],
                     [AC_MSG_ERROR([sorry, can't do anything for you])])

This is safe, so long as you adopt good naming conventions and do not define macros with names like ‘HAVE_STDIO_H’, ‘stdio’, or ‘h’. Though it is also safe here to omit the quotes around ‘Define to 1 if you have <stdio.h>.’ this is not recommended, as message strings are more likely to inadvertently contain commas.

The following example is wrong and dangerous, as it is underquoted:

     AC_CHECK_HEADER(stdio.h,
                     AC_DEFINE(HAVE_STDIO_H, 1,
                        Define to 1 if you have <stdio.h>.),
                     AC_MSG_ERROR([sorry, can't do anything for you]))

In other cases, you may have to use text that also resembles a macro call. You must quote that text even when it is not passed as a macro argument. For example, these two approaches in ‘configure.ac’ (quoting just the potential problems, or quoting the entire line) will protect your script in case autoconf ever adds a macro AC_DC:

     echo "Hard rock was here!  --[AC_DC]"
     [echo "Hard rock was here!  --AC_DC"]

which results in this text in ‘configure’:

     echo "Hard rock was here!  --AC_DC"
     echo "Hard rock was here!  --AC_DC"

When you use the same text in a macro argument, you must therefore have an extra quotation level (since one is stripped away by the macro substitution). In general, then, it is a good idea to use double quoting for all literal string arguments, either around just the problematic portions, or over the entire argument:

     AC_MSG_WARN([[AC_DC] stinks  --Iron Maiden])
     AC_MSG_WARN([[AC_DC stinks  --Iron Maiden]])

However, the above example triggers a warning about a possibly unexpanded macro when running autoconf, because it collides with the namespace of macros reserved for the Autoconf language. To be really safe, you can use additional escaping (either a quadrigraph, or creative shell constructs) to silence that particular warning:

     echo "Hard rock was here!  --AC""_DC"
     AC_MSG_WARN([[AC@&t@_DC stinks  --Iron Maiden]])

You are now able to understand one of the constructs of Autoconf that has been continually misunderstood… The rule of thumb is that whenever you expect macro expansion, expect quote expansion; i.e., expect one level of quotes to be lost. For instance:

     AC_COMPILE_IFELSE([char b[10];], [], [AC_MSG_ERROR([you lose])])

is incorrect: here, the first argument of AC_COMPILE_IFELSE is ‘char b[10];’ and is expanded once, which results in ‘char b10;’. (There was an idiom common in Autoconf’s past to address this issue via the M4 changequote primitive, but do not use it!) Let’s take a closer look: the author meant the first argument to be understood as a literal, and therefore it must be quoted twice:

     AC_COMPILE_IFELSE([[char b[10];]], [], [AC_MSG_ERROR([you lose])])

Voilà, you actually produce ‘char b[10];’ this time!

On the other hand, descriptions (e.g., the last parameter of AC_DEFINE or AS_HELP_ STRING) are not literals—they are subject to line breaking, for example—and should not be double quoted. Even if these descriptions are short and are not actually broken, double quoting them yields weird results.

Some macros take optional arguments, which this documentation represents as [arg] (not to be confused with the quote characters). You may just leave them empty, or use ‘[]’ to make the emptiness of the argument explicit, or you may simply omit the trailing commas. The three lines below are equivalent:

     AC_CHECK_HEADERS([stdio.h], [], [], [])
     AC_CHECK_HEADERS([stdio.h],,,)
     AC_CHECK_HEADERS([stdio.h])

It is best to put each macro call on its own line in ‘configure.ac’. Most of the macros don’t add extra newlines; they rely on the newline after the macro call to terminate the commands. This approach makes the generated configure script a little easier to read by not inserting lots of blank lines. It is generally safe to set shell variables on the same line as a macro call, because the shell allows assignments without intervening newlines.

You can include comments in ‘configure.ac’ files by starting them with the ‘#’. For example, it is helpful to begin ‘configure.ac’ files with a line like this:

     # Process this file with autoconf to produce a configure script.

3.1.3 Standard ‘configure.ac’ Layout

The order in which ‘configure.ac’ calls the Autoconf macros is not important, with a few exceptions. Every ‘configure.ac’ must contain a call to AC_INIT before the checks, and a call to AC_OUTPUT at the end (see Section 4.5 [Output], page 22). Additionally, some macros rely on other macros having been called first, because they check previously set values of some variables to decide what to do. These macros are noted in the individual descriptions (see Chapter 5 [Existing Tests], page 43), and they also warn you when configure is created if they are called out of order.

To encourage consistency, here is a suggested order for calling the Autoconf macros. Generally speaking, the things near the end of this list are those that could depend on things earlier in it. For example, library functions could be affected by types and libraries.

Autoconf requirements

     AC_INIT(package, version, bug-report-address)

information on the package

checks for programs checks for libraries checks for header files checks for types checks for structures checks for compiler characteristics checks for library functions checks for system services

     AC_CONFIG_FILES([file...])
     AC_OUTPUT
3.2 Using autoscan to Create ‘configure.ac’

The autoscan program can help you create and/or maintain a ‘configure.ac’ file for a software package. autoscan examines source files in the directory tree rooted at a directory given as a command line argument, or the current directory if none is given. It searches the source files for common portability problems and creates a file ‘configure.scan’ which is a preliminary ‘configure.ac’ for that package, and checks a possibly existing ‘configure.ac’ for completeness.

When using autoscan to create a ‘configure.ac’, you should manually examine ‘configure.scan’ before renaming it to ‘configure.ac’; it probably needs some adjustments. Occasionally, autoscan outputs a macro in the wrong order relative to another macro, so that autoconf produces a warning; you need to move such macros manually. Also, if you want the package to use a configuration header file, you must add a call to AC_CONFIG_HEADERS (see Section 4.9 [Configuration Headers], page 35). You might also have to change or add some #if directives to your program in order to make it work with Autoconf (see Section 3.3 [ifnames Invocation], page 13, for information about a program that can help with that job).

When using autoscan to maintain a ‘configure.ac’, simply consider adding its suggestions. The file ‘autoscan.log’ contains detailed information on why a macro is requested.

autoscan uses several data files (installed along with Autoconf) to determine which macros to output when it finds particular symbols in a package’s source files. These data files all have the same format: each line consists of a symbol, one or more blanks, and the Autoconf macro to output if that symbol is encountered. Lines starting with ‘#’ are comments.

autoscan accepts the following options:

‘--help’

‘-h’ Print a summary of the command line options and exit.

‘--version’

‘-V’ Print the version number of Autoconf and exit.

‘--verbose’

‘-v’ Print the names of the files it examines and the potentially interesting symbols it finds in them. This output can be voluminous.

‘--debug’

‘-d’ Don’t remove temporary files.

‘--include=dir’

‘-I dir’ Append dir to the include path. Multiple invocations accumulate.

‘--prepend-include=dir’

‘-B dir’ Prepend dir to the include path. Multiple invocations accumulate.

3.3 Using ifnames to List Conditionals

ifnames can help you write ‘configure.ac’ for a software package. It prints the identifiers that the package already uses in C preprocessor conditionals. If a package has already been set up to have some portability, ifnames can thus help you figure out what its configure needs to check for. It may help fill in some gaps in a ‘configure.ac’ generated by autoscan (see Section 3.2 [autoscan Invocation], page 12).

ifnames scans all of the C source files named on the command line (or the standard input, if none are given) and writes to the standard output a sorted list of all the identifiers that appear in those files in #if, #elif, #ifdef, or #ifndef directives. It prints each identifier on a line, followed by a space-separated list of the files in which that identifier occurs. ifnames accepts the following options:

‘--help’

‘-h’ Print a summary of the command line options and exit.

‘--version’

‘-V’ Print the version number of Autoconf and exit.

3.4 Using autoconf to Create configure

To create configure from ‘configure.ac’, run the autoconf program with no arguments. autoconf processes ‘configure.ac’ with the M4 macro processor, using the Autoconf macros. If you give autoconf an argument, it reads that file instead of ‘configure.ac’ and writes the configuration script to the standard output instead of to configure. If you give autoconf the argument ‘-’, it reads from the standard input instead of ‘configure.ac’ and writes the configuration script to the standard output.

The Autoconf macros are defined in several files. Some of the files are distributed with Autoconf; autoconf reads them first. Then it looks for the optional file ‘acsite.m4’ in the directory that contains the distributed Autoconf macro files, and for the optional file ‘aclocal.m4’ in the current directory. Those files can contain your site’s or the package’s own Autoconf macro definitions (see Chapter 10 [Writing Autoconf Macros], page 175, for more information). If a macro is defined in more than one of the files that autoconf reads, the last definition it reads overrides the earlier ones.

autoconf accepts the following options:

‘--help’

‘-h’ Print a summary of the command line options and exit.

‘--version’

‘-V’ Print the version number of Autoconf and exit.

‘--verbose’

‘-v’ Report processing steps.

‘--debug’

‘-d’ Don’t remove the temporary files.

‘--force’

‘-f’ Remake ‘configure’ even if newer than its input files.

‘--include=dir’

‘-I dir’ Append dir to the include path. Multiple invocations accumulate.

‘--prepend-include=dir’

‘-B dir’ Prepend dir to the include path. Multiple invocations accumulate.

‘--output=file’

‘-o file’ Save output (script or trace) to file. The file ‘-’ stands for the standard output.

‘--warnings=category’
‘-W category’

Report the warnings related to category (which can actually be a comma separated list). See Section 10.3 [Reporting Messages], page 177, macro AC_ DIAGNOSE, for a comprehensive list of categories. Special values include:

‘all’ report all the warnings

‘none’ report none

‘error’ treats warnings as errors

‘no-category’

disable warnings falling into category

Warnings about ‘syntax’ are enabled by default, and the environment variable WARNINGS, a comma separated list of categories, is honored as well. Passing ‘-W category’ actually behaves as if you had passed ‘--warnings

syntax,$WARNINGS,category’. To disable the defaults and WARNINGS, and

then enable warnings about obsolete constructs, use ‘-W none,obsolete’.

Because autoconf uses autom4te behind the scenes, it displays a back trace for errors, but not for warnings; if you want them, just pass ‘-W error’. See Section 8.2.1 [autom4te Invocation], page 130, for some examples.

‘--trace=macro[:format]’
‘-t macro[:format]’

Do not create the configure script, but list the calls to macro according to the format. Multiple ‘--trace’ arguments can be used to list several macros. Multiple ‘--trace’ arguments for a single macro are not cumulative; instead, you should just make format as long as needed.

The format is a regular string, with newlines if desired, and several special escape codes. It defaults to ‘$f:$l:$n:$%’; see Section 8.2.1 [autom4te Invocation], page 130, for details on the format.

‘--initialization’

‘-i’ By default, ‘--trace’ does not trace the initialization of the Autoconf macros (typically the AC_DEFUN definitions). This results in a noticeable speedup, but can be disabled by this option.

It is often necessary to check the content of a ‘configure.ac’ file, but parsing it yourself is extremely fragile and error-prone. It is suggested that you rely upon ‘--trace’ to scan ‘configure.ac’. For instance, to find the list of variables that are substituted, use:

$ autoconf -t AC_SUBST
configure.ac:2:AC_SUBST:ECHO_C
configure.ac:2:AC_SUBST:ECHO_N
configure.ac:2:AC_SUBST:ECHO_T
More traces deleted

The example below highlights the difference between ‘$@’, ‘$*’, and ‘$%’.

$ cat configure.ac
AC_DEFINE(This, is, [an
[example]])
$ autoconf -t 'AC_DEFINE:@: $@
*: $*
%: $%'
@: [This],[is],[an
[example]]
*: This,is,an
[example]
%: This:is:an [example]

The format gives you a lot of freedom:

$ autoconf -t 'AC_SUBST:$$ac_subst{"$1"} = "$f:$l";'
$ac_subst{"ECHO_C"} = "configure.ac:2";
$ac_subst{"ECHO_N"} = "configure.ac:2";
$ac_subst{"ECHO_T"} = "configure.ac:2";
More traces deleted

A long separator can be used to improve the readability of complex structures, and to ease their parsing (for instance when no single character is suitable as a separator):

$ autoconf -t 'AM_MISSING_PROG:${|:::::|}*'
ACLOCAL|:::::|aclocal|:::::|$missing_dir
AUTOCONF|:::::|autoconf|:::::|$missing_dir
AUTOMAKE|:::::|automake|:::::|$missing_dir
More traces deleted

3.5 Using autoreconf to Update configure Scripts

Installing the various components of the GNU Build System can be tedious: running autopoint for Gettext, automake for ‘Makefile.in’ etc. in each directory. It may be needed either because some tools such as automake have been updated on your system, or because some of the sources such as ‘configure.ac’ have been updated, or finally, simply in order to install the GNU Build System in a fresh tree.

   autoreconf runs autoconf, autoheader, aclocal, automake, libtoolize, and

autopoint (when appropriate) repeatedly to update the GNU Build System in the specified directories and their subdirectories (see Section 4.12 [Subdirectories], page 40). By default, it only remakes those files that are older than their sources. The environment

variables AUTOCONF, AUTOHEADER, AUTOMAKE, ACLOCAL, AUTOPOINT, LIBTOOLIZE, M4, and

MAKE may be used to override the invocation of the respective tools.

If you install a new version of some tool, you can make autoreconf remake all of the files by giving it the ‘--force’ option.

See Section 4.8.5 [Automatic Remaking], page 34, for Make rules to automatically rebuild configure scripts when their source files change. That method handles the timestamps of configuration header templates properly, but does not pass ‘--autoconf-dir=dir’ or

‘--localdir=dir’.

Gettext supplies the autopoint command to add translation infrastructure to a source package. If you use autopoint, your ‘configure.ac’ should invoke both AM_GNU_GETTEXT

and AM_GNU_GETTEXT_VERSION(gettext-version). See Section “Invoking the autopoint

Program” in GNU gettext utilities, for further details.

autoreconf accepts the following options:

‘--help’

‘-h’ Print a summary of the command line options and exit.

‘--version’

‘-V’ Print the version number of Autoconf and exit.

‘--verbose’

‘-V’ Print the name of each directory autoreconf examines and the commands it runs. If given two or more times, pass ‘--verbose’ to subordinate tools that support it.

‘--debug’

‘-d’ Don’t remove the temporary files.

‘--force’

‘-f’ Remake even ‘configure’ scripts and configuration headers that are newer than their input files (‘configure.ac’ and, if present, ‘aclocal.m4’).

‘--install’

‘-i’ Install the missing auxiliary files in the package. By default, files are copied; this can be changed with ‘--symlink’. If deemed appropriate, this option triggers calls to ‘automake --add-missing’,

          ‘libtoolize’, ‘autopoint’, etc.
‘--no-recursive’

Do not rebuild files in subdirectories to configure (see Section 4.12 [Subdirectories], page 40, macro AC_CONFIG_SUBDIRS).

‘--symlink’

‘-s’ When used with ‘--install’, install symbolic links to the missing auxiliary files instead of copying them.

‘--make’

‘-m’ When the directories were configured, update the configuration by running

‘./config.status --recheck && ./config.status’, and then run ‘make’.
‘--include=dir’

‘-I dir’ Append dir to the include path. Multiple invocations accumulate. Passed on

to aclocal, autoconf and autoheader internally.
‘--prepend-include=dir’

‘-B dir’ Prepend dir to the include path. Multiple invocations accumulate. Passed on

to autoconf and autoheader internally.
‘--warnings=category’
‘-W category’

Report the warnings related to category (which can actually be a comma separated list).

‘cross’ related to cross compilation issues.

‘obsolete’

report the uses of obsolete constructs.

‘portability’

portability issues

‘syntax’ dubious syntactic constructs.

‘all’ report all the warnings

‘none’ report none

‘error’ treats warnings as errors

‘no-category’

disable warnings falling into category

Warnings about ‘syntax’ are enabled by default, and the environment variable WARNINGS, a comma separated list of categories, is honored as well. Passing ‘-W category’ actually behaves as if you had passed ‘--warnings

syntax,$WARNINGS,category’. To disable the defaults and WARNINGS, and

then enable warnings about obsolete constructs, use ‘-W none,obsolete’.

If you want autoreconf to pass flags that are not listed here on to aclocal, set ACLOCAL_ AMFLAGS in your ‘Makefile.am’. Due to a limitation in the Autoconf implementation these flags currently must be set on a single line in ‘Makefile.am’, without any backslash-newlines.

4 Initialization and Output Files

Autoconf-generated configure scripts need some information about how to initialize, such as how to find the package’s source files and about the output files to produce. The following sections describe the initialization and the creation of output files.

4.1 Initializing configure

Every configure script must call AC_INIT before doing anything else that produces output. Calls to silent macros, such as AC_DEFUN, may also occur prior to AC_INIT, although these are generally used via ‘aclocal.m4’, since that is implicitly included before the start of ‘configure.ac’. The only other required macro is AC_OUTPUT (see Section 4.5 [Output], page 22).

[Macro] AC_INIT (package, version, [bug-report], [tarname], [url])

Process any command-line arguments and perform various initializations and verifications. Set the name of the package and its version. These are typically used in ‘--version’ support, including that of configure. The optional argument bug-report should be the email to which users should send bug reports. The package tarname differs from package: the latter designates the full package name (e.g., ‘GNU Autoconf’), while the former is meant for distribution tar ball names (e.g., ‘autoconf’). It defaults to package with ‘GNU ’ stripped, lower-cased, and all characters other than alphanumerics and underscores are changed to ‘-’. If provided, url should be the home page for the package. It is preferable that the arguments of AC_INIT be static, i.e., there should not be any shell computation, but they can be computed by M4. The following M4 macros (e.g., AC_PACKAGE_NAME), output variables (e.g., PACKAGE_ NAME), and preprocessor symbols (e.g., PACKAGE_NAME), are defined by AC_INIT:

AC_PACKAGE_NAME, PACKAGE_NAME

Exactly package.

AC_PACKAGE_TARNAME, PACKAGE_TARNAME

Exactly tarname, possibly generated from package.

AC_PACKAGE_VERSION, PACKAGE_VERSION

Exactly version.

AC_PACKAGE_STRING, PACKAGE_STRING
          Exactly ‘package version’.
AC_PACKAGE_BUGREPORT, PACKAGE_BUGREPORT

Exactly bug-report, if one was provided.

AC_PACKAGE_URL, PACKAGE_URL

Exactly url, if one was provided. If url was empty, but package be-

          gins with ‘GNU ’, then this defaults to ‘http://www.gnu.org/software/

tarname/’, otherwise, no URL is assumed.

If your configure script does its own option processing, it should inspect ‘$@’ or ‘$*’ immediately after calling AC_INIT, because other Autoconf macros liberally use the set command to process strings, and this has the side effect of updating ‘$@’ and ‘$*’. However, we suggest that you use standard macros like AC_ARG_ENABLE instead of attempting to implement your own option processing. See Chapter 15 [Site Configuration], page 271.

4.2 Dealing with Autoconf versions

The following optional macros can be used to help choose the minimum version of Autoconf that can successfully compile a given ‘configure.ac’.

[Macro] AC_PREREQ (version)

Ensure that a recent enough version of Autoconf is being used. If the version of Autoconf being used to create configure is earlier than version, print an error message to the standard error output and exit with failure (exit status is 63). For example:

     AC_PREREQ([2.66])

This macro may be used before AC_INIT.

[Macro] AC_AUTOCONF_VERSION

This macro was introduced in Autoconf 2.62. It identifies the version of Autoconf that is currently parsing the input file, in a format suitable for m4_version_compare (see [m4 version compare], page 157); in other words, for this release of Autoconf, its value is ‘2.66’. One potential use of this macro is for writing conditional fallbacks based on when a feature was added to Autoconf, rather than using AC_PREREQ to require the newer version of Autoconf. However, remember that the Autoconf philosophy favors feature checks over version checks. You should not expand this macro directly; use ‘m4_defn([AC_AUTOCONF_VERSION])’ instead. This is because some users might have a beta version of Autoconf installed, with arbitrary letters included in its version string. This means it is possible for the version string to contain the name of a defined macro, such that expanding AC_ AUTOCONF_VERSION would trigger the expansion of that macro during rescanning, and change the version string to be different than what you intended to check.

4.3 Notices in configure

The following macros manage version numbers for configure scripts. Using them is optional.

[Macro] AC_COPYRIGHT (copyright-notice)

State that, in addition to the Free Software Foundation’s copyright on the Autoconf macros, parts of your configure are covered by the copyright-notice. The copyright-notice shows up in both the head of configure and in ‘configure

--version’.
[Macro] AC_REVISION (revision-info)

Copy revision stamp revision-info into the configure script, with any dollar signs or double-quotes removed. This macro lets you put a revision stamp from ‘configure.ac’ into configure without RCS or CVS changing it when you check in configure. That way, you can determine easily which revision of ‘configure.ac’ a particular configure corresponds to. For example, this line in ‘configure.ac’:

     AC_REVISION([$Revision: 1.30 $])

produces this in configure:

     #!/bin/sh
     # From configure.ac Revision: 1.30

4.4 Finding configure Input

[Macro] AC_CONFIG_SRCDIR (unique-file-in-source-dir)

unique-file-in-source-dir is some file that is in the package’s source directory; configure checks for this file’s existence to make sure that the directory that it is told contains the source code in fact does. Occasionally people accidentally specify the wrong directory with ‘--srcdir’; this is a safety check. See Section 16.10 [configure Invocation], page 285, for more information.

Packages that do manual configuration or use the install program might need to tell configure where to find some other shell scripts by calling AC_CONFIG_AUX_DIR, though the default places it looks are correct for most cases.

[Macro] AC_CONFIG_AUX_DIR (dir)

Use the auxiliary build tools (e.g., ‘install-sh’, ‘config.sub’, ‘config.guess’, Cygnus configure, Automake and Libtool scripts, etc.) that are in directory dir. These are auxiliary files used in configuration. dir can be either absolute or relative to ‘srcdir’. The default is ‘srcdir’ or ‘srcdir/..’ or ‘srcdir/../..’, whichever is the first that contains ‘install-sh’. The other files are not checked for, so that using AC_PROG_INSTALL does not automatically require distributing the other auxiliary files. It checks for ‘install.sh’ also, but that name is obsolete because some make have a rule that creates ‘install’ from it if there is no makefile. The auxiliary directory is commonly named ‘build-aux’. If you need portability to DOS variants, do not name the auxiliary directory ‘aux’. See Section 11.4 [File System Conventions], page 193.

[Macro] AC_REQUIRE_AUX_FILE (file)

Declares that file is expected in the directory defined above. In Autoconf proper, this macro does nothing: its sole purpose is to be traced by third-party tools to produce a list of expected auxiliary files. For instance it is called by macros like AC_PROG_ INSTALL (see Section 5.2.1 [Particular Programs], page 45) or AC_CANONICAL_BUILD (see Section 14.2 [Canonicalizing], page 268) to register the auxiliary files they need.

Similarly, packages that use aclocal should declare where local macros can be found

using AC_CONFIG_MACRO_DIR.
[Macro] AC_CONFIG_MACRO_DIR (dir)

Specify dir as the location of additional local Autoconf macros. This macro is intended for use by future versions of commands like autoreconf that trace macro calls. It should be called directly from ‘configure.ac’ so that tools that install macros for aclocal can find the macros’ declarations.

Note that if you use aclocal from Automake to generate ‘aclocal.m4’, you must also set ACLOCAL_AMFLAGS = -I dir in your top-level ‘Makefile.am’. Due to a limitation in the Autoconf implementation of autoreconf, these include directives currently must be set on a single line in ‘Makefile.am’, without any backslash-newlines.

4.5 Outputting Files

Every Autoconf script, e.g., ‘configure.ac’, should finish by calling AC_OUTPUT. That is the macro that generates and runs ‘config.status’, which in turn creates the makefiles and any other files resulting from configuration. This is the only required macro besides AC_INIT (see Section 4.4 [Input], page 21).

[Macro] AC_OUTPUT

Generate ‘config.status’ and launch it. Call this macro once, at the end of

‘configure.ac’.

‘config.status’ performs all the configuration actions: all the output files (see Section 4.7 [Configuration Files], page 25, macro AC_CONFIG_FILES), header files (see Section 4.9 [Configuration Headers], page 35, macro AC_CONFIG_HEADERS), commands (see Section 4.10 [Configuration Commands], page 39, macro AC_CONFIG_COMMANDS), links (see Section 4.11 [Configuration Links], page 40, macro AC_CONFIG_LINKS), subdirectories to configure (see Section 4.12 [Subdirectories], page 40, macro AC_CONFIG_ SUBDIRS) are honored.

The location of your AC_OUTPUT invocation is the exact point where configuration actions are taken: any code afterwards is executed by configure once config.status was run. If you want to bind actions to config.status itself (independently of whether configure is being run), see Section 4.10 [Running Arbitrary Configuration Commands], page 39.

Historically, the usage of AC_OUTPUT was somewhat different. See Section 18.4 [Obsolete Macros], page 293, for a description of the arguments that AC_OUTPUT used to support.

If you run make in subdirectories, you should run it using the make variable MAKE. Most versions of make set MAKE to the name of the make program plus any options it was given. (But many do not include in it the values of any variables set on the command line, so those are not passed on automatically.) Some old versions of make do not set this variable. The following macro allows you to use it even with those versions.

[Macro] AC_PROG_MAKE_SET

If the Make command, $MAKE if set or else ‘make’, predefines $(MAKE), define output variable SET_MAKE to be empty. Otherwise, define SET_MAKE to a macro definition

that sets $(MAKE), such as ‘MAKE=make’. Calls AC_SUBST for SET_MAKE.

If you use this macro, place a line like this in each ‘Makefile.in’ that runs MAKE on other directories:

@SET_MAKE@

4.6 Performing Configuration Actions

‘configure’ is designed so that it appears to do everything itself, but there is actually a hidden slave: ‘config.status’. ‘configure’ is in charge of examining your system, but it is ‘config.status’ that actually takes the proper actions based on the results of ‘configure’. The most typical task of ‘config.status’ is to instantiate files.

This section describes the common behavior of the four standard instantiating macros:

AC_CONFIG_FILES, AC_CONFIG_HEADERS, AC_CONFIG_COMMANDS and AC_CONFIG_LINKS.

They all have this prototype:

AC_CONFIG_ITEMS(tag..., [commands], [init-cmds])

where the arguments are:

tag… A blank-or-newline-separated list of tags, which are typically the names of the files to instantiate.

You are encouraged to use literals as tags. In particular, you should avoid

     ... && my_foos="$my_foos fooo"
     ... && my_foos="$my_foos foooo"
     AC_CONFIG_ITEMS([$my_foos])

and use this instead:

     ... && AC_CONFIG_ITEMS([fooo])
     ... && AC_CONFIG_ITEMS([foooo])
The macros AC_CONFIG_FILES and AC_CONFIG_HEADERS use special tag values:

they may have the form ‘output’ or ‘output:inputs’. The file output is instantiated from its templates, inputs (defaulting to ‘output.in’).

‘AC_CONFIG_FILES([Makefile:boiler/top.mk:boiler/bot.mk)]’, for exam-

ple, asks for the creation of the file ‘Makefile’ that contains the expansion of the output variables in the concatenation of ‘boiler/top.mk’ and ‘boiler/bot.mk’.

The special value ‘-’ might be used to denote the standard output when used in output, or the standard input when used in the inputs. You most probably don’t need to use this in ‘configure.ac’, but it is convenient when using the command line interface of ‘./config.status’, see Chapter 17 [config.status Invocation], page 287, for more details.

The inputs may be absolute or relative file names. In the latter case they are first looked for in the build tree, and then in the source tree. Input files should be text files, and a line length below 2000 bytes should be safe.

commands Shell commands output literally into ‘config.status’, and associated with a tag that the user can use to tell ‘config.status’ which commands to run. The commands are run each time a tag request is given to ‘config.status’, typically each time the file ‘tag’ is created.

The variables set during the execution of configure are not available here: you first need to set them via the init-cmds. Nonetheless the following variables are precomputed: srcdir The name of the top source directory, assuming that the working

directory is the top build directory. This is what the configure option ‘--srcdir’ sets.

ac_top_srcdir

The name of the top source directory, assuming that the working directory is the current build directory.

ac_top_build_prefix

The name of the top build directory, assuming that the working directory is the current build directory. It can be empty, or else ends with a slash, so that you may concatenate it.

ac_srcdir

The name of the corresponding source directory, assuming that the working directory is the current build directory.

tmp The name of a temporary directory within the build tree, which you can use if you need to create additional temporary files. The directory is cleaned up when config.status is done or interrupted. Please use package-specific file name prefixes to avoid clashing with files that config.status may use internally.

The current directory refers to the directory (or pseudo-directory) containing the input part of tags. For instance, running

     AC_CONFIG_COMMANDS([deep/dir/out:in/in.in], [...], [...])

with ‘--srcdir=../package’ produces the following values:

     # Argument of --srcdir
     srcdir='../package'
     # Reversing deep/dir
     ac_top_build_prefix='../../'
     # Concatenation of $ac_top_build_prefix and srcdir
     ac_top_srcdir='../../../package'
     # Concatenation of $ac_top_srcdir and deep/dir
     ac_srcdir='../../../package/deep/dir'

independently of ‘in/in.in’.

init-cmds Shell commands output unquoted near the beginning of ‘config.status’, and executed each time ‘config.status’ runs (regardless of the tag). Because they are unquoted, for example, ‘$var’ is output as the value of var. init-cmds is typically used by ‘configure’ to give ‘config.status’ some variables it needs to run the commands.

You should be extremely cautious in your variable names: all the init-cmds share the same name space and may overwrite each other in unpredictable ways. Sorry…

All these macros can be called multiple times, with different tag values, of course!

4.7 Creating Configuration Files

Be sure to read the previous section, Section 4.6 [Configuration Actions], page 23.

[Macro] AC_CONFIG_FILES (file . . ., [cmds], [init-cmds])

Make AC_OUTPUT create each ‘file’ by copying an input file (by default ‘file.in’), substituting the output variable values. This macro is one of the instantiating macros; see Section 4.6 [Configuration Actions], page 23. See Section 4.8 [Makefile Substitutions], page 25, for more information on using output variables. See Section 7.2 [Setting Output Variables], page 112, for more information on creating them. This macro creates the directory that the file is in if it doesn’t exist. Usually, makefiles are created this way, but other files, such as ‘.gdbinit’, can be specified as well. Typical calls to AC_CONFIG_FILES look like this:

          AC_CONFIG_FILES([Makefile src/Makefile man/Makefile X/Imakefile])
          AC_CONFIG_FILES([autoconf], [chmod +x autoconf])

You can override an input file name by appending to file a colon-separated list of input files. Examples:

          AC_CONFIG_FILES([Makefile:boiler/top.mk:boiler/bot.mk]
                          [lib/Makefile:boiler/lib.mk])

Doing this allows you to keep your file names acceptable to DOS variants, or to prepend and/or append boilerplate to the file.

4.8 Substitutions in Makefiles

Each subdirectory in a distribution that contains something to be compiled or installed should come with a file ‘Makefile.in’, from which configure creates a file ‘Makefile’ in that directory. To create ‘Makefile’, configure performs a simple variable substitution, replacing occurrences of ‘@variable@’ in ‘Makefile.in’ with the value that configure has determined for that variable. Variables that are substituted into output files in this way are called output variables. They are ordinary shell variables that are set in configure. To make configure substitute a particular variable into the output files, the macro AC_SUBST must be called with that variable name as an argument. Any occurrences of ‘@variable@’ for other variables are left unchanged. See Section 7.2 [Setting Output Variables], page 112, for more information on creating output variables with AC_SUBST.

A software package that uses a configure script should be distributed with a file ‘Makefile.in’, but no makefile; that way, the user has to properly configure the package for the local system before compiling it.

See Section “Makefile Conventions” in The GNU Coding Standards, for more information on what to put in makefiles.

4.8.1 Preset Output Variables

Some output variables are preset by the Autoconf macros. Some of the Autoconf macros set additional output variables, which are mentioned in the descriptions for those macros. See Section B.2 [Output Variable Index], page 352, for a complete list of output variables. See Section 4.8.2 [Installation Directory Variables], page 29, for the list of the preset ones related to installation directories. Below are listed the other preset ones, many of which are precious variables (see Section 7.2 [Setting Output Variables], page 112, AC_ARG_VAR).

The preset variables which are available during ‘config.status’ (see Section 4.6 [Configuration Actions], page 23) may also be used during configure tests. For example, it is permissible to reference ‘$srcdir’ when constructing a list of directories to pass via option ‘-I’ during a compiler feature check. When used in this manner, coupled with the fact that configure is always run from the top build directory, it is sufficient to use just ‘$srcdir’ instead of ‘$top_srcdir’.

[Variable] CFLAGS

Debugging and optimization options for the C compiler. If it is not set in the environment when configure runs, the default value is set when you call AC_PROG_CC (or empty if you don’t). configure uses this variable when compiling or linking programs to test for C features. If a compiler option affects only the behavior of the preprocessor (e.g., ‘-D name’), it should be put into CPPFLAGS instead. If it affects only the linker (e.g., ‘-L directory’), it should be put into LDFLAGS instead. If it affects only the compiler proper, CFLAGS is the natural home for it. If an option affects multiple phases of the compiler, though, matters get tricky. One approach to put such options directly into CC, e.g., CC='gcc -m64'. Another is to put them into both CPPFLAGS and LDFLAGS, but not

into CFLAGS.

However, remember that some ‘Makefile’ variables are reserved by the GNU Coding Standards for the use of the “user”—the person building the package. For instance, CFLAGS is one such variable. Sometimes package developers are tempted to set user variables such as CFLAGS because it appears to make their job easier. However, the package itself should never set a user variable, particularly not to include switches that are required for proper compilation of the package. Since these variables are documented as being for the package builder, that person rightfully expects to be able to override any of these variables at build time. If the package developer needs to add switches without interfering with the user, the proper way to do that is to introduce an additional variable. Automake makes this easy by introducing AM_CFLAGS (see Section “Flag Variables Ordering” in GNU Automake), but the concept is the same even if Automake is not used.

[Variable] configure_input

A comment saying that the file was generated automatically by configure and giving the name of the input file. AC_OUTPUT adds a comment line containing this variable to the top of every makefile it creates. For other files, you should reference this variable in a comment at the top of each input file. For example, an input shell script should begin like this:

     #!/bin/sh
     # @configure_input@

The presence of that line also reminds people editing the file that it needs to be processed by configure in order to be used.

[Variable] CPPFLAGS

Preprocessor options for the C, C++, Objective C, and Objective C++ preprocessors and compilers. If it is not set in the environment when configure runs, the default value is empty. configure uses this variable when preprocessing or compiling programs to test for C, C++, Objective C, and Objective C++ features. This variable’s contents should contain options like ‘-I’, ‘-D’, and ‘-U’ that affect only the behavior of the preprocessor. Please see the explanation of CFLAGS for what you can do if an option affects other phases of the compiler as well. Currently, configure always links as part of a single invocation of the compiler that also preprocesses and compiles, so it uses this variable also when linking programs. However, it is unwise to depend on this behavior because the GNU Coding Standards do not require it and many packages do not use CPPFLAGS when linking programs. See Section 7.3 [Special Chars in Variables], page 114, for limitations that CPPFLAGS might run into.

[Variable] CXXFLAGS

Debugging and optimization options for the C++ compiler. It acts like CFLAGS, but for C++ instead of C.

[Variable] DEFS

‘-D’ options to pass to the C compiler. If AC_CONFIG_HEADERS is called, configure replaces ‘@DEFS@’ with ‘-DHAVE_CONFIG_H’ instead (see Section 4.9 [Configuration Headers], page 35). This variable is not defined while configure is performing its tests, only when creating the output files. See Section 7.2 [Setting Output Variables], page 112, for how to check the results of previous tests.

[Variable] ECHO_C

[Variable] ECHO_N

[Variable] ECHO_T

How does one suppress the trailing newline from echo for question-answer message pairs? These variables provide a way:

     echo $ECHO_N "And the winner is... $ECHO_C"
     sleep 100000000000
     echo "${ECHO_T}dead."

Some old and uncommon echo implementations offer no means to achieve this, in which case ECHO_T is set to tab. You might not want to use it.

[Variable] ERLCFLAGS

Debugging and optimization options for the Erlang compiler. If it is not set in the environment when configure runs, the default value is empty. configure uses this variable when compiling programs to test for Erlang features.

[Variable] FCFLAGS

Debugging and optimization options for the Fortran compiler. If it is not set in the environment when configure runs, the default value is set when you call AC_PROG_ FC (or empty if you don’t). configure uses this variable when compiling or linking programs to test for Fortran features.

[Variable] FFLAGS

Debugging and optimization options for the Fortran 77 compiler. If it is not set in the environment when configure runs, the default value is set when you call AC_PROG_ F77 (or empty if you don’t). configure uses this variable when compiling or linking programs to test for Fortran 77 features.

[Variable] LDFLAGS

Options for the linker. If it is not set in the environment when configure runs, the default value is empty. configure uses this variable when linking programs to test for C, C++, Objective C, Objective C++, and Fortran features. This variable’s contents should contain options like ‘-s’ and ‘-L’ that affect only the behavior of the linker. Please see the explanation of CFLAGS for what you can do if an option also affects other phases of the compiler. Don’t use this variable to pass library names (‘-l’) to the linker; use LIBS instead.

[Variable] LIBS

‘-l’ options to pass to the linker. The default value is empty, but some Autoconf macros may prepend extra libraries to this variable if those libraries are found and provide necessary functions, see Section 5.4 [Libraries], page 51. configure uses this variable when linking programs to test for C, C++, Objective C, Objective C++, and Fortran features.

[Variable] OBJCFLAGS

Debugging and optimization options for the Objective C compiler. It acts like CFLAGS, but for Objective C instead of C.

[Variable] OBJCXXFLAGS

Debugging and optimization options for the Objective C++ compiler. It acts like CXXFLAGS, but for Objective C++ instead of C++.

[Variable] builddir

Rigorously equal to ‘.’. Added for symmetry only.

[Variable] abs_builddir

Absolute name of builddir.

[Variable] top_builddir

The relative name of the top level of the current build tree. In the top-level directory, this is the same as builddir.

[Variable] top_build_prefix

The relative name of the top level of the current build tree with final slash if nonemtpy. This is the same as top_builddir, except that it contains zero or more runs of ../, so it should not be appended with a slash for concatenation. This helps for make implementations that otherwise do not treat ‘./file’ and ‘file’ as equal in the toplevel build directory.

[Variable] abs_top_builddir

Absolute name of top_builddir.

[Variable] srcdir

The name of the directory that contains the source code for that makefile.

[Variable] abs_srcdir

Absolute name of srcdir.

[Variable] top_srcdir

The name of the top-level source code directory for the package. In the top-level directory, this is the same as srcdir.

[Variable] abs_top_srcdir

Absolute name of top_srcdir.

4.8.2 Installation Directory Variables

The following variables specify the directories for package installation, see Section “Variables for Installation Directories” in The GNU Coding Standards, for more information. Each variable corresponds to an argument of configure; trailing slashes are stripped so that expressions such as ‘${prefix}/lib’ expand with only one slash between directory names. See the end of this section for details on when and how to use these variables.

[Variable] bindir

The directory for installing executables that users run.

[Variable] datadir

The directory for installing idiosyncratic read-only architecture-independent data.

[Variable] datarootdir

The root of the directory tree for read-only architecture-independent data files.

[Variable] docdir

The directory for installing documentation files (other than Info and man).

[Variable] dvidir

The directory for installing documentation files in DVI format.

[Variable] exec_prefix

The installation prefix for architecture-dependent files. By default it’s the same as prefix. You should avoid installing anything directly to exec_prefix. However, the default value for directories containing architecture-dependent files should be relative

to exec_prefix.

[Variable] htmldir

The directory for installing HTML documentation.

[Variable] includedir

The directory for installing C header files.

[Variable] infodir

The directory for installing documentation in Info format.

[Variable] libdir

The directory for installing object code libraries.

[Variable] libexecdir

The directory for installing executables that other programs run.

[Variable] localedir

The directory for installing locale-dependent but architecture-independent data, such as message catalogs. This directory usually has a subdirectory per locale.

[Variable] localstatedir

The directory for installing modifiable single-machine data.

[Variable] mandir

The top-level directory for installing documentation in man format.

[Variable] oldincludedir

The directory for installing C header files for non-GCC compilers.

[Variable] pdfdir

The directory for installing PDF documentation.

[Variable] prefix

The common installation prefix for all files. If exec_prefix is defined to a different value, prefix is used only for architecture-independent files.

[Variable] psdir

The directory for installing PostScript documentation.

[Variable] sbindir

The directory for installing executables that system administrators run.

[Variable] sharedstatedir

The directory for installing modifiable architecture-independent data.

[Variable] sysconfdir

The directory for installing read-only single-machine data.

Most of these variables have values that rely on prefix or exec_prefix. It is deliberate that the directory output variables keep them unexpanded: typically ‘@datarootdir@’ is

replaced by ‘${prefix}/share’, not ‘/usr/local/share’, and ‘@datadir@’ is replaced by
‘${datarootdir}’.

This behavior is mandated by the GNU Coding Standards, so that when the user runs:

‘make’ she can still specify a different prefix from the one specified to configure, in which case, if needed, the package should hard code dependencies corresponding to the make-specified prefix.

‘make install’

she can specify a different installation location, in which case the package must still depend on the location which was compiled in (i.e., never recompile when ‘make install’ is run). This is an extremely important feature, as many people may decide to install all the files of a package grouped together, and then install links from the final locations to there.

In order to support these features, it is essential that datarootdir remains defined as ‘${prefix}/share’, so that its value can be expanded based on the current value of prefix.

A corollary is that you should not use these variables except in makefiles. For instance, instead of trying to evaluate datadir in ‘configure’ and hard-coding it in makefiles using e.g.,

‘AC_DEFINE_UNQUOTED([DATADIR], ["$datadir"], [Data directory.])’, you should add
‘-DDATADIR='$(datadir)'’ to your makefile’s definition of CPPFLAGS (AM_CPPFLAGS if you

are also using Automake).

Similarly, you should not rely on AC_CONFIG_FILES to replace bindir and friends in your shell scripts and other files; instead, let make manage their replacement. For instance Autoconf ships templates of its shell scripts ending with ‘.in’, and uses a makefile snippet similar to the following to build scripts like autoheader and autom4te:

     edit = sed \
             -e 's|@bindir[@]|$(bindir)|g' \
             -e 's|@pkgdatadir[@]|$(pkgdatadir)|g' \
             -e 's|@prefix[@]|$(prefix)|g'
     autoheader autom4te: Makefile
             rm -f $@ $@.tmp
             srcdir=''; \
               test -f ./$@.in || srcdir=$(srcdir)/; \
               $(edit) $${srcdir}$@.in >$@.tmp
             chmod +x $@.tmp
             chmod a-w $@.tmp
             mv $@.tmp $@
     autoheader: $(srcdir)/autoheader.in
     autom4te: $(srcdir)/autom4te.in

Some details are noteworthy:

‘@bindir[@]’

The brackets prevent configure from replacing ‘@bindir@’ in the Sed expression itself. Brackets are preferable to a backslash here, since Posix says ‘\@’ is not portable.

‘$(bindir)’

Don’t use ‘@bindir@’! Use the matching makefile variable instead.

‘$(pkgdatadir)’

The example takes advantage of the variable ‘$(pkgdatadir)’ provided by Automake; it is equivalent to ‘$(datadir)/$(PACKAGE)’.

‘/’ Don’t use ‘/’ in the Sed expressions that replace file names since most likely the variables you use, such as ‘$(bindir)’, contain ‘/’. Use a shell metacharacter instead, such as ‘|’.

special characters File names, file name components, and the value of VPATH should not contain shell metacharacters or white space. See Section 7.3 [Special Chars in Variables], page 114. dependency on ‘Makefile’

Since edit uses values that depend on the configuration specific values (prefix, etc.) and not only on VERSION and so forth, the output depends on ‘Makefile’,

          not ‘configure.ac’.

‘$@’ The main rule is generic, and uses ‘$@’ extensively to avoid the need for multiple copies of the rule.

Separated dependencies and single suffix rules

You can’t use them! The above snippet cannot be (portably) rewritten as:

               autoconf autoheader: Makefile
               .in:
                       rm -f $@ $@.tmp
                       $(edit) $< >$@.tmp
                       chmod +x $@.tmp
                       mv $@.tmp $@

See Section 12.16 [Single Suffix Rules], page 254, for details.

‘$(srcdir)’

Be sure to specify the name of the source directory, otherwise the package won’t support separated builds.

For the more specific installation of Erlang libraries, the following variables are defined:

[Variable] ERLANG_INSTALL_LIB_DIR

The common parent directory of Erlang library installation directories. This variable

     is set by calling the AC_ERLANG_SUBST_INSTALL_LIB_DIR macro in ‘configure.ac’.
[Variable] ERLANG_INSTALL_LIB_DIR_library

The installation directory for Erlang library library. This variable is set by using the

     ‘AC_ERLANG_SUBST_INSTALL_LIB_SUBDIR’ macro in ‘configure.ac’.

See Section 5.13 [Erlang Libraries], page 96, for details.

4.8.3 Changed Directory Variables

In Autoconf 2.60, the set of directory variables has changed, and the defaults of some variables have been adjusted (see Section 4.8.2 [Installation Directory Variables], page 29) to changes in the GNU Coding Standards. Notably, ‘datadir’, ‘infodir’, and ‘mandir’ are now expressed in terms of ‘datarootdir’. If you are upgrading from an earlier Autoconf version, you may need to adjust your files to ensure that the directory variables are substituted correctly (see Section 20.5 [Defining Directories], page 331), and that a definition of ‘datarootdir’ is in place. For example, in a ‘Makefile.in’, adding

     datarootdir = @datarootdir@

is usually sufficient. If you use Automake to create ‘Makefile.in’, it will add this for you.

To help with the transition, Autoconf warns about files that seem to use datarootdir without defining it. In some cases, it then expands the value of $datarootdir in substitutions of the directory variables. The following example shows such a warning:

$ cat configure.ac
AC_INIT
AC_CONFIG_FILES([Makefile])
AC_OUTPUT
$ cat Makefile.in
prefix = @prefix@
datadir = @datadir@
$ autoconf
$ configure
configure: creating ./config.status
config.status: creating Makefile
config.status: WARNING:
               Makefile.in seems to ignore the --datarootdir setting
$ cat Makefile
prefix = /usr/local
datadir = ${prefix}/share

Usually one can easily change the file to accommodate both older and newer Autoconf releases:

$ cat Makefile.in
prefix = @prefix@
datarootdir = @datarootdir@
datadir = @datadir@
$ configure
configure: creating ./config.status
config.status: creating Makefile
$ cat Makefile
prefix = /usr/local
datarootdir = ${prefix}/share
datadir = ${datarootdir}

In some cases, however, the checks may not be able to detect that a suitable initialization of datarootdir is in place, or they may fail to detect that such an initialization is necessary in the output file. If, after auditing your package, there are still spurious ‘configure’ warnings about datarootdir, you may add the line

AC_DEFUN([AC_DATAROOTDIR_CHECKED])

to your ‘configure.ac’ to disable the warnings. This is an exception to the usual rule that you should not define a macro whose name begins with AC_(see Section 10.2 [Macro Names], page 176).

4.8.4 Build Directories

You can support compiling a software package for several architectures simultaneously from the same copy of the source code. The object files for each architecture are kept in their own directory.

To support doing this, make uses the VPATH variable to find the files that are in the source directory. GNU Make can do this. Most other recent make programs can do this as well, though they may have difficulties and it is often simpler to recommend GNU make (see Section 12.15 [VPATH and Make], page 247). Older make programs do not support VPATH; when using them, the source code must be in the same directory as the object files.

If you are using GNU Automake, the remaining details in this section are already covered for you, based on the contents of your ‘Makefile.am’. But if you are using Autoconf in isolation, then supporting VPATH requires the following in your ‘Makefile.in’:

     srcdir = @srcdir@
     VPATH = @srcdir@

Do not set VPATH to the value of another variable (see Section 12.15.1 [Variables listed in VPATH], page 248.

configure substitutes the correct value for srcdir when it produces ‘Makefile’.

Do not use the make variable $<, which expands to the file name of the file in the source directory (found with VPATH), except in implicit rules. (An implicit rule is one such as ‘.c.o’, which tells how to create a ‘.o’ file from a ‘.c’ file.) Some versions of make do not set $< in explicit rules; they expand it to an empty value.

Instead, Make command lines should always refer to source files by prefixing them with ‘$(srcdir)/’. For example:

     time.info: time.texinfo
             $(MAKEINFO) '$(srcdir)/time.texinfo'

4.8.5 Automatic Remaking

You can put rules like the following in the top-level ‘Makefile.in’ for a package to automatically update the configuration information when you change the configuration files. This example includes all of the optional files, such as ‘aclocal.m4’ and those related to configuration header files. Omit from the ‘Makefile.in’ rules for any of these files that your package does not use.

The ‘$(srcdir)/’ prefix is included because of limitations in the VPATH mechanism.

The ‘stamp-’ files are necessary because the timestamps of ‘config.h.in’ and ‘config.h’ are not changed if remaking them does not change their contents. This feature avoids unnecessary recompilation. You should include the file ‘stamp-h.in’ in your package’s distribution, so that make considers ‘config.h.in’ up to date. Don’t use touch (see [Limitations of Usual Tools], page 238); instead, use echo (using date would cause needless differences, hence CVS conflicts, etc.).

     $(srcdir)/configure: configure.ac aclocal.m4
             cd '$(srcdir)' && autoconf
     # autoheader might not change config.h.in, so touch a stamp file.
     $(srcdir)/config.h.in: stamp-h.in
     $(srcdir)/stamp-h.in: configure.ac aclocal.m4
             cd '$(srcdir)' && autoheader
             echo timestamp > '$(srcdir)/stamp-h.in'
     config.h: stamp-h
     stamp-h: config.h.in config.status
             ./config.status
     Makefile: Makefile.in config.status
             ./config.status
     config.status: configure
             ./config.status --recheck

(Be careful if you copy these lines directly into your makefile, as you need to convert the indented lines to start with the tab character.)

In addition, you should use

     AC_CONFIG_FILES([stamp-h], [echo timestamp > stamp-h])

so ‘config.status’ ensures that ‘config.h’ is considered up to date. See Section 4.5 [Output], page 22, for more information about AC_OUTPUT.

See Chapter 17 [config.status Invocation], page 287, for more examples of handling configuration-related dependencies.

4.9 Configuration Header Files

When a package contains more than a few tests that define C preprocessor symbols, the command lines to pass ‘-D’ options to the compiler can get quite long. This causes two problems. One is that the make output is hard to visually scan for errors. More seriously, the command lines can exceed the length limits of some operating systems. As an alternative to passing ‘-D’ options to the compiler, configure scripts can create a C header file containing ‘#define’ directives. The AC_CONFIG_HEADERS macro selects this kind of output. Though it can be called anywhere between AC_INIT and AC_OUTPUT, it is customary to call it right

after AC_INIT.

The package should ‘#include’ the configuration header file before any other header files, to prevent inconsistencies in declarations (for example, if it redefines const).

To provide for VPATH builds, remember to pass the C compiler a ‘-I.’ option (or ‘-I..’; whichever directory contains ‘config.h’). Even if you use ‘#include "config.h"’, the preprocessor searches only the directory of the currently read file, i.e., the source directory, not the build directory.

With the appropriate ‘-I’ option, you can use ‘#include <config.h>’. Actually, it’s a good habit to use it, because in the rare case when the source directory contains another ‘config.h’, the build directory should be searched first.

[Macro] AC_CONFIG_HEADERS (header . . ., [cmds], [init-cmds])

This macro is one of the instantiating macros; see Section 4.6 [Configuration Actions], page 23. Make AC_OUTPUT create the file(s) in the blank-or-newline-separated list header containing C preprocessor #define statements, and replace ‘@DEFS@’ in generated files with ‘-DHAVE_CONFIG_H’ instead of the value of DEFS. The usual name for header is ‘config.h’. If header already exists and its contents are identical to what AC_OUTPUT would put in it, it is left alone. Doing this allows making some changes in the configuration without needlessly causing object files that depend on the header file to be recompiled. Usually the input file is named ‘header.in’; however, you can override the input file name by appending to header a colon-separated list of input files. For example, you might need to make the input file name acceptable to DOS variants:

     AC_CONFIG_HEADERS([config.h:config.hin])
[Macro] AH_HEADER

This macro is defined as the name of the first declared config header and undefined if no config headers have been declared up to this point. A third-party macro may, for example, require use of a config header without invoking AC CONFIG HEADERS twice, like this:

     AC_CONFIG_COMMANDS_PRE(
             [m4_ifndef([AH_HEADER], [AC_CONFIG_HEADERS([config.h])])])

See Section 4.6 [Configuration Actions], page 23, for more details on header.

4.9.1 Configuration Header Templates

Your distribution should contain a template file that looks as you want the final header file to look, including comments, with #undef statements which are used as hooks. For example, suppose your ‘configure.ac’ makes these calls:

AC_CONFIG_HEADERS([conf.h])
AC_CHECK_HEADERS([unistd.h])

Then you could have code like the following in ‘conf.h.in’. The ‘conf.h’ created by configure defines ‘HAVE_UNISTD_H’ to 1, if and only if the system has ‘unistd.h’.

/* Define as 1 if you have unistd.h.  */
#undef HAVE_UNISTD_H

The format of the template file is stricter than what the C preprocessor is required to accept. A directive line should contain only whitespace, ‘#undef’, and ‘HAVE_UNISTD_H’. The use of ‘#define’ instead of ‘#undef’, or of comments on the same line as ‘#undef’, is strongly discouraged. Each hook should only be listed once. Other preprocessor lines, such as ‘#ifdef’ or ‘#include’, are copied verbatim from the template into the generated header.

Since it is a tedious task to keep a template header up to date, you may use autoheader to generate it, see Section 4.9.2 [autoheader Invocation], page 37.

During the instantiation of the header, each ‘#undef’ line in the template file for each symbol defined by ‘AC_DEFINE’ is changed to an appropriate ‘#define’. If the corresponding ‘AC_DEFINE’ has not been executed during the configure run, the ‘#undef’ line is commented out. (This is important, e.g., for ‘_POSIX_SOURCE’: on many systems, it can be implicitly defined by the compiler, and undefining it in the header would then break compilation of subsequent headers.)

Currently, all remaining ‘#undef’ lines in the header template are commented out, whether or not there was a corresponding ‘AC_DEFINE’ for the macro name; but this behavior is not guaranteed for future releases of Autoconf.

Generally speaking, since you should not use ‘#define’, and you cannot guarantee whether a ‘#undef’ directive in the header template will be converted to a ‘#define’ or commented out in the generated header file, the template file cannot be used for conditional definition effects. Consequently, if you need to use the construct

     #ifdef THIS
     # define THAT
     #endif

you must place it outside of the template. If you absolutely need to hook it to the config header itself, please put the directives to a separate file, and ‘#include’ that file from the config header template. If you are using autoheader, you would probably use ‘AH_BOTTOM’ to append the ‘#include’ directive.

4.9.2 Using autoheader to Create ‘config.h.in’

The autoheader program can create a template file of C ‘#define’ statements for configure to use. It searches for the first invocation of AC_CONFIG_HEADERS in ‘configure’ sources to determine the name of the template. (If the first call of AC_CONFIG_HEADERS specifies more than one input file name, autoheader uses the first one.)

It is recommended that only one input file is used. If you want to append a boilerplate

code, it is preferable to use ‘AH_BOTTOM([#include <conf_post.h>])’. File ‘conf_post.h’

is not processed during the configuration then, which make things clearer. Analogically, AH_TOP can be used to prepend a boilerplate code.

In order to do its job, autoheader needs you to document all of the symbols that you might use. Typically this is done via an AC_DEFINE or AC_DEFINE_UNQUOTED call whose first argument is a literal symbol and whose third argument describes the symbol (see Section 7.1 [Defining Symbols], page 111). Alternatively, you can use AH_TEMPLATE (see Section 4.9.3 [Autoheader Macros], page 38), or you can supply a suitable input file for a subsequent configuration header file. Symbols defined by Autoconf’s builtin tests are already documented properly; you need to document only those that you define yourself.

You might wonder why autoheader is needed: after all, why would configure need to “patch” a ‘config.h.in’ to produce a ‘config.h’ instead of just creating ‘config.h’ from scratch? Well, when everything rocks, the answer is just that we are wasting our time maintaining autoheader: generating ‘config.h’ directly is all that is needed. When things go wrong, however, you’ll be thankful for the existence of autoheader.

The fact that the symbols are documented is important in order to check that ‘config.h’ makes sense. The fact that there is a well-defined list of symbols that should be defined (or not) is also important for people who are porting packages to environments where configure cannot be run: they just have to fill in the blanks.

But let’s come back to the point: the invocation of autoheader…

If you give autoheader an argument, it uses that file instead of ‘configure.ac’ and writes the header file to the standard output instead of to ‘config.h.in’. If you give autoheader an argument of ‘-’, it reads the standard input instead of ‘configure.ac’ and writes the header file to the standard output.

autoheader accepts the following options:

‘--help’

‘-h’ Print a summary of the command line options and exit.

‘--version’

‘-V’ Print the version number of Autoconf and exit.

‘--verbose’

‘-v’ Report processing steps.

‘--debug’

‘-d’ Don’t remove the temporary files.

‘--force’

‘-f’ Remake the template file even if newer than its input files.

‘--include=dir’

‘-I dir’ Append dir to the include path. Multiple invocations accumulate.

‘--prepend-include=dir’

‘-B dir’ Prepend dir to the include path. Multiple invocations accumulate.

‘--warnings=category’
‘-W category’

Report the warnings related to category (which can actually be a comma separated list). Current categories include:

          ‘obsolete’

report the uses of obsolete constructs

‘all’ report all the warnings

‘none’ report none

‘error’ treats warnings as errors

          ‘no-category’

disable warnings falling into category

4.9.3 Autoheader Macros

autoheader scans ‘configure.ac’ and figures out which C preprocessor symbols it might define. It knows how to generate templates for symbols defined by AC_CHECK_HEADERS, AC_ CHECK_FUNCS etc., but if you AC_DEFINE any additional symbol, you must define a template for it. If there are missing templates, autoheader fails with an error message.

The template for a symbol is created by autoheader from the description argument to an AC_DEFINE; see Section 7.1 [Defining Symbols], page 111.

For special needs, you can use the following macros.

[Macro] AH_TEMPLATE (key, description)

Tell autoheader to generate a template for key. This macro generates standard templates just like AC_DEFINE when a description is given.

For example:

     AH_TEMPLATE([CRAY_STACKSEG_END],
                 [Define to one of _getb67, GETB67, getb67
                  for Cray-2 and Cray-YMP systems.  This
                  function is required for alloca.c support
                  on those systems.])

generates the following template, with the description properly justified.

     /* Define to one of _getb67, GETB67, getb67 for Cray-2 and
        Cray-YMP systems.  This function is required for alloca.c
        support on those systems.  */
     #undef CRAY_STACKSEG_END
[Macro] AH_VERBATIM (key, template)

Tell autoheader to include the template as-is in the header template file. This template is associated with the key, which is used to sort all the different templates and guarantee their uniqueness. It should be a symbol that can be defined via AC_

DEFINE.
[Macro] AH_TOP (text)

Include text at the top of the header template file.

[Macro] AH_BOTTOM (text)

Include text at the bottom of the header template file.

Please note that text gets included “verbatim” to the template file, not to the resulting config header, so it can easily get mangled when the template is processed. There is rarely a need for something other than

AH_BOTTOM([#include <custom.h>])

4.10 Running Arbitrary Configuration Commands

You can execute arbitrary commands before, during, and after ‘config.status’ is run. The three following macros accumulate the commands to run when they are called multiple times. AC_CONFIG_COMMANDS replaces the obsolete macro AC_OUTPUT_COMMANDS; see Section 18.4 [Obsolete Macros], page 293, for details.

[Macro] AC_CONFIG_COMMANDS (tag . . ., [cmds], [init-cmds])

Specify additional shell commands to run at the end of ‘config.status’, and shell commands to initialize any variables from configure. Associate the commands with tag. Since typically the cmds create a file, tag should naturally be the name of that file. If needed, the directory hosting tag is created. This macro is one of the instantiating macros; see Section 4.6 [Configuration Actions], page 23.

Here is an unrealistic example:

     fubar=42
     AC_CONFIG_COMMANDS([fubar],
                        [echo this is extra $fubar, and so on.],
                        [fubar=$fubar])

Here is a better one:

     AC_CONFIG_COMMANDS([timestamp], [date >timestamp])

The following two macros look similar, but in fact they are not of the same breed: they are executed directly by ‘configure’, so you cannot use ‘config.status’ to rerun them.

[Macro] AC_CONFIG_COMMANDS_PRE (cmds)

Execute the cmds right before creating ‘config.status’. This macro presents the last opportunity to call AC_SUBST, AC_DEFINE, or AC_CONFIG_ ITEMS macros.

[Macro] AC_CONFIG_COMMANDS_POST (cmds)

Execute the cmds right after creating ‘config.status’.

4.11 Creating Configuration Links

You may find it convenient to create links whose destinations depend upon results of tests. One can use AC_CONFIG_COMMANDS but the creation of relative symbolic links can be delicate when the package is built in a directory different from the source directory.

[Macro] AC_CONFIG_LINKS (dest:source . . ., [cmds], [init-cmds])

Make AC_OUTPUT link each of the existing files source to the corresponding link name dest. Makes a symbolic link if possible, otherwise a hard link if possible, otherwise a copy. The dest and source names should be relative to the top level source or build directory. This macro is one of the instantiating macros; see Section 4.6 [Configuration Actions], page 23. For example, this call:

     AC_CONFIG_LINKS([host.h:config/$machine.h
                     object.h:config/$obj_format.h])

creates in the current directory ‘host.h’ as a link to ‘srcdir/config/$machine.h’,

and ‘object.h’ as a link to ‘srcdir/config/$obj_format.h’.

The tempting value ‘.’ for dest is invalid: it makes it impossible for ‘config.status’ to guess the links to establish. One can then run:

     ./config.status host.h object.h

to create the links.

4.12 Configuring Other Packages in Subdirectories

In most situations, calling AC_OUTPUT is sufficient to produce makefiles in subdirectories. However, configure scripts that control more than one independent package can use AC_ CONFIG_SUBDIRS to run configure scripts for other packages in subdirectories.

[Macro] AC_CONFIG_SUBDIRS (dir . . .)

Make AC_OUTPUT run configure in each subdirectory dir in the given blank-or-newline-separated list. Each dir should be a literal, i.e., please do not use:

     if test "x$package_foo_enabled" = xyes; then
       my_subdirs="$my_subdirs foo"
     fi
     AC_CONFIG_SUBDIRS([$my_subdirs])

because this prevents ‘./configure --help=recursive’ from displaying the options of the package foo. Instead, you should write:

     if test "x$package_foo_enabled" = xyes; then
       AC_CONFIG_SUBDIRS([foo])
     fi

If a given dir is not found, an error is reported: if the subdirectory is optional, write:

     if test -d "$srcdir/foo"; then
       AC_CONFIG_SUBDIRS([foo])
     fi

If a given dir contains configure.gnu, it is run instead of configure. This is for packages that might use a non-Autoconf script Configure, which can’t be called through a wrapper configure since it would be the same file on case-insensitive file systems. Likewise, if a dir contains ‘configure.in’ but no configure, the Cygnus

configure script found by AC_CONFIG_AUX_DIR is used.

The subdirectory configure scripts are given the same command line options that were given to this configure script, with minor changes if needed, which include: − adjusting a relative name for the cache file; − adjusting a relative name for the source directory; − propagating the current value of $prefix, including if it was defaulted, and if

the default values of the top level and of the subdirectory ‘configure’ differ.

This macro also sets the output variable subdirs to the list of directories ‘dir ...’. Make rules can use this variable to determine which subdirectories to recurse into. This macro may be called multiple times.

4.13 Default Prefix

By default, configure sets the prefix for files it installs to ‘/usr/local’. The user of configure can select a different prefix using the ‘--prefix’ and ‘--exec-prefix’ options. There are two ways to change the default: when creating configure, and when running it.

Some software packages might want to install in a directory other than ‘/usr/local’ by default. To accomplish that, use the AC_PREFIX_DEFAULT macro.

[Macro] AC_PREFIX_DEFAULT (prefix)

Set the default installation prefix to prefix instead of ‘/usr/local’.

It may be convenient for users to have configure guess the installation prefix from the location of a related program that they have already installed. If you wish to do that, you

can call AC_PREFIX_PROGRAM.

[Macro] AC_PREFIX_PROGRAM (program)

If the user did not specify an installation prefix (using the ‘--prefix’ option), guess a value for it by looking for program in PATH, the way the shell does. If program is found, set the prefix to the parent of the directory containing program, else default the prefix as described above (‘/usr/local’ or AC_PREFIX_DEFAULT). For example, if program is gcc and the PATH contains ‘/usr/local/gnu/bin/gcc’, set the prefix

to ‘/usr/local/gnu’.

5 Existing Tests

These macros test for particular system features that packages might need or want to use. If you need to test for a kind of feature that none of these macros check for, you can probably do it by calling primitive test macros with appropriate arguments (see Chapter 6 [Writing Tests], page 99).

These tests print messages telling the user which feature they’re checking for, and what they find. They cache their results for future configure runs (see Section 7.4 [Caching Results], page 115).

Some of these macros set output variables. See Section 4.8 [Makefile Substitutions], page 25, for how to get their values. The phrase “define name” is used below as a shorthand to mean “define the C preprocessor symbol name to the value 1”. See Section 7.1 [Defining Symbols], page 111, for how to get those symbol definitions into your program.

5.1 Common Behavior

Much effort has been expended to make Autoconf easy to learn. The most obvious way to reach this goal is simply to enforce standard interfaces and behaviors, avoiding exceptions as much as possible. Because of history and inertia, unfortunately, there are still too many exceptions in Autoconf; nevertheless, this section describes some of the common rules.

5.1.1 Standard Symbols

All the generic macros that AC_DEFINE a symbol as a result of their test transform their argument values to a standard alphabet. First, argument is converted to upper case and any asterisks (‘*’) are each converted to ‘P’. Any remaining characters that are not alphanumeric are converted to underscores.

For instance,

     AC_CHECK_TYPES([struct $Expensive*])

defines the symbol ‘HAVE_STRUCT__EXPENSIVEP’ if the check succeeds.

5.1.2 Default Includes

Several tests depend upon a set of header files. Since these headers are not universally available, tests actually have to provide a set of protected includes, such as:

     #ifdef TIME_WITH_SYS_TIME
     # include <sys/time.h>
     # include <time.h>
     #else
     # ifdef HAVE_SYS_TIME_H
     #  include <sys/time.h>
     # else
     #  include <time.h>
     # endif
     #endif

Unless you know exactly what you are doing, you should avoid using unconditional includes, and check the existence of the headers you include beforehand (see Section 5.6 [Header Files], page 64).

Most generic macros use the following macro to provide the default set of includes:

[Macro] AC_INCLUDES_DEFAULT ([include-directives])

Expand to include-directives if defined, otherwise to:

#include <stdio.h>
#ifdef HAVE_SYS_TYPES_H
# include <sys/types.h>
#endif
#ifdef HAVE_SYS_STAT_H
# include <sys/stat.h>
#endif
#ifdef STDC_HEADERS
# include <stdlib.h>
# include <stddef.h>
#else
# ifdef HAVE_STDLIB_H
#  include <stdlib.h>
# endif
#endif
#ifdef HAVE_STRING_H
# if !defined STDC_HEADERS && defined HAVE_MEMORY_H
#  include <memory.h>
# endif
# include <string.h>
#endif
#ifdef HAVE_STRINGS_H
# include <strings.h>
#endif
#ifdef HAVE_INTTYPES_H
# include <inttypes.h>
#endif
#ifdef HAVE_STDINT_H
# include <stdint.h>
#endif
#ifdef HAVE_UNISTD_H
# include <unistd.h>
#endif

If the default includes are used, then check for the presence of these headers and their compatibility, i.e., you don’t need to run AC_HEADER_STDC, nor check for ‘stdlib.h’ etc.

These headers are checked for in the same order as they are included. For instance, on some systems ‘string.h’ and ‘strings.h’ both exist, but conflict. Then HAVE_

STRING_H is defined, not HAVE_STRINGS_H.

5.2 Alternative Programs

These macros check for the presence or behavior of particular programs. They are used to choose between several alternative programs and to decide what to do once one has been chosen. If there is no macro specifically defined to check for a program you need, and you don’t need to check for any special properties of it, then you can use one of the general program-check macros.

5.2.1 Particular Program Checks

These macros check for particular programs—whether they exist, and in some cases whether they support certain features.

[Macro] AC_PROG_AWK

Check for gawk, mawk, nawk, and awk, in that order, and set output variable AWK to the first one that is found. It tries gawk first because that is reported to be the best implementation. The result can be overridden by setting the variable AWK or the cache

variable ac_cv_prog_AWK.
[Macro] AC_PROG_GREP

Look for the best available grep or ggrep that accepts the longest input lines possible, and that supports multiple ‘-e’ options. Set the output variable GREP to whatever is chosen. See [Limitations of Usual Tools], page 231, for more information about portability problems with the grep command family. The result can be overridden by setting the GREP variable and is cached in the ac_cv_path_GREP variable.

[Macro] AC_PROG_EGREP

Check whether $GREP -E works, or else look for the best available egrep or gegrep that accepts the longest input lines possible. Set the output variable EGREP to whatever is chosen. The result can be overridden by setting the EGREP variable and is cached in the ac_cv_path_EGREP variable.

[Macro] AC_PROG_FGREP

Check whether $GREP -F works, or else look for the best available fgrep or gfgrep that accepts the longest input lines possible. Set the output variable FGREP to whatever is chosen. The result can be overridden by setting the FGREP variable and is cached in the ac_cv_path_FGREP variable.

[Macro] AC_PROG_INSTALL

Set output variable INSTALL to the name of a BSD-compatible install program, if one is found in the current PATH. Otherwise, set INSTALL to ‘dir/install-sh -c’, checking the directories specified to AC_CONFIG_AUX_DIR (or its default directories) to determine dir (see Section 4.5 [Output], page 22). Also set the variables INSTALL_

PROGRAM and INSTALL_SCRIPT to ‘${INSTALL}’ and INSTALL_DATA to ‘${INSTALL}
-m 644’.

‘@INSTALL@’ is special, as its value may vary for different configuration files. This macro screens out various instances of install known not to work. It prefers to find a C program rather than a shell script, for speed. Instead of ‘install-sh’, it can also use ‘install.sh’, but that name is obsolete because some make programs have a rule that creates ‘install’ from it if there is no makefile. Further, this macro requires install to be able to install multiple files into a target directory in a single invocation. Autoconf comes with a copy of ‘install-sh’ that you can use. If you use AC_PROG_ INSTALL, you must include either ‘install-sh’ or ‘install.sh’ in your distribution; otherwise configure produces an error message saying it can’t find them—even if the system you’re on has a good install program. This check is a safety measure to prevent you from accidentally leaving that file out, which would prevent your package from installing on systems that don’t have a BSD-compatible install program. If you need to use your own installation program because it has features not found in standard install programs, there is no reason to use AC_PROG_INSTALL; just put the file name of your program into your ‘Makefile.in’ files. The result of the test can be overridden by setting the variable INSTALL or the cache

variable ac_cv_path_install.
[Macro] AC_PROG_MKDIR_P

Set output variable MKDIR_P to a program that ensures that for each argument, a directory named by this argument exists, creating it and its parent directories if needed, and without race conditions when two instances of the program attempt to make the same directory at nearly the same time. This macro uses the ‘mkdir -p’ command if possible. Otherwise, it falls back on invoking install-sh with the ‘-d’ option, so your package should contain ‘install-sh’ as described under AC_PROG_INSTALL. An ‘install-sh’ file that predates Autoconf 2.60 or Automake 1.10 is vulnerable to race conditions, so if you want to support parallel installs from different packages into the same directory you need to make sure you have an up-to-date ‘install-sh’. In particular, be careful about using ‘autoreconf -if’ if your Automake predates Automake 1.10. This macro is related to the AS_MKDIR_P macro (see Chapter 9 [Programming in M4sh], page 165), but it sets an output variable intended for use in other files, whereas AS_MKDIR_P is intended for use in scripts like configure. Also, AS_MKDIR_P does not accept options, but MKDIR_P supports the ‘-m’ option, e.g., a makefile might invoke $(MKDIR_P) -m 0 dir to create an inaccessible directory, and conversely a makefile should use $(MKDIR_P) -- $(FOO) if FOO might yield a value that begins with ‘-’. Finally, AS_MKDIR_P does not check for race condition vulnerability, whereas AC_

PROG_MKDIR_P does.

‘@MKDIR_P@’ is special, as its value may vary for different configuration files. The result of the test can be overridden by setting the variable MKDIR_P or the cache

variable ac_cv_path_mkdir.
[Macro] AC_PROG_LEX

If flex is found, set output variable LEX to ‘flex’ and LEXLIB to ‘-lfl’, if that library is in a standard place. Otherwise set LEX to ‘lex’ and LEXLIB to ‘-ll’, if found. If neither variant is available, set LEX to ‘:’; for packages that ship the generated ‘file.yy.c’ alongside the source ‘file.l’, this default allows users without a lexer generator to still build the package even if the timestamp for ‘file.l’ is inadvertantly changed.

Define YYTEXT_POINTER if yytext defaults to ‘char *’ instead of to ‘char []’. Also set output variable LEX_OUTPUT_ROOT to the base of the file name that the lexer generates; usually ‘lex.yy’, but sometimes something else. These results vary according to whether lex or flex is being used. You are encouraged to use Flex in your sources, since it is both more pleasant to use than plain Lex and the C source it produces is portable. In order to ensure portability, however, you must either provide a function yywrap or, if you don’t use it (e.g., your scanner has no ‘#include’-like feature), simply include a ‘%noyywrap’ statement in the scanner’s source. Once this done, the scanner is portable (unless you felt free to use nonportable constructs) and does not depend on any library. In this case, and in this case only, it is suggested that you use this Autoconf snippet:

     AC_PROG_LEX
     if test "x$LEX" != xflex; then
       LEX="$SHELL $missing_dir/missing flex"
       AC_SUBST([LEX_OUTPUT_ROOT], [lex.yy])
       AC_SUBST([LEXLIB], [''])
     fi

The shell script missing can be found in the Automake distribution. Remember that the user may have supplied an alternate location in LEX, so if Flex is required, it is better to check that the user provided something sufficient by parsing the output of ‘$LEX --version’ than by simply relying on test "x$LEX" = xflex. To ensure backward compatibility, Automake’s AM_PROG_LEX invokes (indirectly) this macro twice, which causes an annoying but benign “AC_PROG_LEX invoked multiple times” warning. Future versions of Automake will fix this issue; meanwhile, just ignore this message. As part of running the test, this macro may delete any file in the configuration directory named ‘lex.yy.c’ or ‘lexyy.c’. The result of this test can be influenced by setting the variable LEX or the cache

variable ac_cv_prog_LEX.
[Macro] AC_PROG_LN_S

If ‘ln -s’ works on the current file system (the operating system and file system support symbolic links), set the output variable LN_S to ‘ln -s’; otherwise, if ‘ln’ works, set LN_S to ‘ln’, and otherwise set it to ‘cp -p’. If you make a link in a directory other than the current directory, its meaning depends on whether ‘ln’ or ‘ln -s’ is used. To safely create links using ‘$(LN_S)’, either find out which form is used and adjust the arguments, or always invoke ln in the directory where the link is to be created. In other words, it does not work to do:

     $(LN_S) foo /x/bar

Instead, do:

     (cd /x && $(LN_S) foo bar)
[Macro] AC_PROG_RANLIB

Set output variable RANLIB to ‘ranlib’ if ranlib is found, and otherwise to ‘:’ (do nothing).

[Macro] AC_PROG_SED

Set output variable SED to a Sed implementation that conforms to Posix and does not have arbitrary length limits. Report an error if no acceptable Sed is found. See [Limitations of Usual Tools], page 234, for more information about portability problems with Sed.

The result of this test can be overridden by setting the SED variable and is cached in

the ac_cv_path_SED variable.
[Macro] AC_PROG_YACC

If bison is found, set output variable YACC to ‘bison -y’. Otherwise, if byacc is found, set YACC to ‘byacc’. Otherwise set YACC to ‘yacc’. The result of this test can be influenced by setting the variable YACC or the cache variable ac_cv_prog_YACC.

5.2.2 Generic Program and File Checks

These macros are used to find programs not covered by the “particular” test macros. If you need to check the behavior of a program as well as find out whether it is present, you have to write your own test for it (see Chapter 6 [Writing Tests], page 99). By default, these macros use the environment variable PATH. If you need to check for a program that might not be in the user’s PATH, you can pass a modified path to use instead, like this:

AC_PATH_PROG([INETD], [inetd], [/usr/libexec/inetd],
             [$PATH$PATH_SEPARATOR/usr/libexec$PATH_SEPARATOR]dnl
[/usr/sbin$PATH_SEPARATOR/usr/etc$PATH_SEPARATOR/etc])

You are strongly encouraged to declare the variable passed to AC_CHECK_PROG etc. as precious, See Section 7.2 [Setting Output Variables], page 112, AC_ARG_VAR, for more details.

[Macro] AC_CHECK_PROG (variable, prog-to-check-for, value-if-found,
   [value-if-not-found], [path = ‘$PATH’], [reject])

Check whether program prog-to-check-for exists in path. If it is found, set variable to value-if-found, otherwise to value-if-not-found, if given. Always pass over reject (an absolute file name) even if it is the first found in the search path; in that case, set variable using the absolute file name of the prog-to-check-for found that is not reject. If variable was already set, do nothing. Calls AC_SUBST for variable. The result of this test can be overridden by setting the variable variable or the cache variable

ac_cv_prog_variable.
[Macro] AC_CHECK_PROGS (variable, progs-to-check-for,
   [value-if-not-found], [path = ‘$PATH’])

Check for each program in the blank-separated list progs-to-check-for existing in the path. If one is found, set variable to the name of that program. Otherwise, continue checking the next program in the list. If none of the programs in the list are found, set variable to value-if-not-found; if value-if-not-found is not specified, the value of variable is not changed. Calls AC_SUBST for variable. The result of this test can be overridden by setting the variable variable or the cache variable ac_cv_prog_

variable.

[Macro] AC_CHECK_TARGET_TOOL (variable, prog-to-check-for,
   [value-if-not-found], [path = ‘$PATH’])

Like AC_CHECK_PROG, but first looks for prog-to-check-for with a prefix of the target type as determined by AC_CANONICAL_TARGET, followed by a dash (see Section 14.2 [Canonicalizing], page 268). If the tool cannot be found with a prefix, and if the build and target types are equal, then it is also searched for without a prefix. As noted in Section 14.1 [Specifying Target Triplets], page 267, the target is rarely specified, because most of the time it is the same as the host: it is the type of system for which any compiler tool in the package produces code. What this macro looks for is, for example, a tool (assembler, linker, etc.) that the compiler driver (gcc for the GNU C Compiler) uses to produce objects, archives or executables.

[Macro] AC_CHECK_TOOL (variable, prog-to-check-for,
   [value-if-not-found], [path = ‘$PATH’])

Like AC_CHECK_PROG, but first looks for prog-to-check-for with a prefix of the host type as specified by ‘--host’, followed by a dash. For example, if the user runs

‘configure --build=x86_64-gnu --host=i386-gnu’, then this call:
     AC_CHECK_TOOL([RANLIB], [ranlib], [:])

sets RANLIB to ‘i386-gnu-ranlib’ if that program exists in path, or otherwise to ‘ranlib’ if that program exists in path, or to ‘:’ if neither program exists. When cross-compiling, this macro will issue a warning if no program prefixed with the host type could be found. For more information, see Section 14.1 [Specifying Target Triplets], page 267.

[Macro] AC_CHECK_TARGET_TOOLS (variable, progs-to-check-for,
   [value-if-not-found], [path = ‘$PATH’])

Like AC_CHECK_TARGET_TOOL, each of the tools in the list progs-to-check-for are checked with a prefix of the target type as determined by AC_CANONICAL_TARGET, followed by a dash (see Section 14.2 [Canonicalizing], page 268). If none of the tools can be found with a prefix, and if the build and target types are equal, then the first one without a prefix is used. If a tool is found, set variable to the name of that program. If none of the tools in the list are found, set variable to value-if-not-found; if value-if-not-found is not specified, the value of variable is not changed. Calls AC_SUBST for variable.

[Macro] AC_CHECK_TOOLS (variable, progs-to-check-for,
   [value-if-not-found], [path = ‘$PATH’])

Like AC_CHECK_TOOL, each of the tools in the list progs-to-check-for are checked with a prefix of the host type as determined by AC_CANONICAL_HOST, followed by a dash (see Section 14.2 [Canonicalizing], page 268). If none of the tools can be found with a prefix, then the first one without a prefix is used. If a tool is found, set variable to the name of that program. If none of the tools in the list are found, set variable to value-if-not-found; if value-if-not-found is not specified, the value of variable is not changed. Calls AC_SUBST for variable. When cross-compiling, this macro will issue a warning if no program prefixed with the host type could be found. For more information, see Section 14.1 [Specifying Target Triplets], page 267.

[Macro] AC_PATH_PROG (variable, prog-to-check-for,
   [value-if-not-found], [path = ‘$PATH’])

Like AC_CHECK_PROG, but set variable to the absolute name of prog-to-check-for if found. The result of this test can be overridden by setting the variable variable. A positive result of this test is cached in the ac_cv_path_variable variable.

[Macro] AC_PATH_PROGS (variable, progs-to-check-for,
   [value-if-not-found], [path = ‘$PATH’])

Like AC_CHECK_PROGS, but if any of progs-to-check-for are found, set variable to the absolute name of the program found. The result of this test can be overridden by setting the variable variable. A positive result of this test is cached in the ac_cv_

path_variable variable.
[Macro] AC_PATH_PROGS_FEATURE_CHECK (variable, progs-to-check-for,
   feature-test, [action-if-not-found], [path = ‘$PATH’])

This macro was introduced in Autoconf 2.62. If variable is not empty, then set the cache variable ac_cv_path_variable to its value. Otherwise, check for each program in the blank-separated list progs-to-check-for existing in path. For each program found, execute feature-test with ac_path_variable set to the absolute name of the candidate program. If no invocation of feature-test sets the shell variable ac_cv_ path_variable, then action-if-not-found is executed. feature-test will be run even when ac_cv_path_variable is set, to provide the ability to choose a better candidate found later in path; to accept the current setting and bypass all futher checks, feature-

test can execute ac_path_variable_found=:.

Note that this macro has some subtle differences from AC_CHECK_PROGS. It is designed to be run inside AC_CACHE_VAL, therefore, it should have no side effects. In particular, variable is not set to the final value of ac_cv_path_variable, nor is AC_ SUBST automatically run. Also, on failure, any action can be performed, whereas

AC_CHECK_PROGS only performs variable=value-if-not-found.

Here is an example, similar to what Autoconf uses in its own configure script. It will search for an implementation of m4 that supports the indir builtin, even if it goes by the name gm4 or is not the first implementation on PATH.

     AC_CACHE_CHECK([for m4 that supports indir], [ac_cv_path_M4],
       [AC_PATH_PROGS_FEATURE_CHECK([M4], [m4 gm4],
         [[m4out=`echo 'changequote([,])indir([divnum])' | $ac_path_M4`
           test "x$m4out" = x0 \
           && ac_cv_path_M4=$ac_path_M4 ac_path_M4_found=:]],
         [AC_MSG_ERROR([could not find m4 that supports indir])])])
     AC_SUBST([M4], [$ac_cv_path_M4])
[Macro] AC_PATH_TARGET_TOOL (variable, prog-to-check-for,
   [value-if-not-found], [path = ‘$PATH’])

Like AC_CHECK_TARGET_TOOL, but set variable to the absolute name of the program if it is found.

[Macro] AC_PATH_TOOL (variable, prog-to-check-for,
   [value-if-not-found], [path = ‘$PATH’])

Like AC_CHECK_TOOL, but set variable to the absolute name of the program if it is found. When cross-compiling, this macro will issue a warning if no program prefixed with the host type could be found. For more information, see Section 14.1 [Specifying Target Triplets], page 267.

5.3 Files

You might also need to check for the existence of files. Before using these macros, ask yourself whether a runtime test might not be a better solution. Be aware that, like most Autoconf macros, they test a feature of the host machine, and therefore, they die when cross-compiling.

[Macro] AC_CHECK_FILE (file, [action-if-found], [action-if-not-found])

Check whether file file exists on the native system. If it is found, execute actionif-found, otherwise do action-if-not-found, if given. The result of this test is cached in the ac_cv_file_file variable, with characters not suitable for a variable name mapped to underscores.

[Macro] AC_CHECK_FILES (files, [action-if-found], [action-if-not-found])

Executes AC_CHECK_FILE once for each file listed in files. Additionally, defines ‘HAVE_file’ (see Section 5.1.1 [Standard Symbols], page 43) for each file found. The results of each test are cached in the ac_cv_file_file variable, with characters not suitable for a variable name mapped to underscores.

5.4 Library Files

The following macros check for the presence of certain C, C++, or Fortran library archive files.

[Macro] AC_CHECK_LIB (library, function, [action-if-found],
   [action-if-not-found], [other-libraries])

Test whether the library library is available by trying to link a test program that calls function function with the library. function should be a function provided by the library. Use the base name of the library; e.g., to check for ‘-lmp’, use ‘mp’ as the library argument. action-if-found is a list of shell commands to run if the link with the library succeeds; action-if-not-found is a list of shell commands to run if the link fails. If action-if-found is not specified, the default action prepends ‘-llibrary’ to LIBS and defines ‘HAVE_LIBlibrary’ (in all capitals). This macro is intended to support building LIBS in a right-to-left (least-dependent to most-dependent) fashion such that library dependencies are satisfied as a natural side effect of consecutive tests. Linkers are sensitive to library ordering so the order in which LIBS is generated is important to reliable detection of libraries. If linking with library results in unresolved symbols that would be resolved by linking with additional libraries, give those libraries as the other-libraries argument, separated by spaces: e.g., ‘-lXt -lX11’. Otherwise, this macro may fail to detect that library is present, because linking the test program can fail with unresolved symbols. The other-libraries argument should be limited to cases where it is desirable to test for one library in the presence of another that is not already in LIBS.

AC_CHECK_LIB requires some care in usage, and should be avoided in some common cases. Many standard functions like gethostbyname appear in the standard C library on some hosts, and in special libraries like nsl on other hosts. On some hosts the special libraries contain variant implementations that you may not want to use. These

days it is normally better to use AC_SEARCH_LIBS([gethostbyname], [nsl]) instead
of AC_CHECK_LIB([nsl], [gethostbyname]).

The result of this test is cached in the ac_cv_lib_library_function variable.

[Macro] AC_SEARCH_LIBS (function, search-libs, [action-if-found],
   [action-if-not-found], [other-libraries])

Search for a library defining function if it’s not already available. This equates to call-

ing ‘AC_LINK_IFELSE([AC_LANG_CALL([], [function])])’ first with no libraries,

then for each library listed in search-libs.

Prepend ‘-llibrary’ to LIBS for the first library found to contain function, and run action-if-found. If the function is not found, run action-if-not-found.

If linking with library results in unresolved symbols that would be resolved by linking with additional libraries, give those libraries as the other-libraries argument, separated by spaces: e.g., ‘-lXt -lX11’. Otherwise, this macro fails to detect that function is present, because linking the test program always fails with unresolved symbols.

The result of this test is cached in the ac_cv_search_function variable as ‘none required’ if function is already available, as ‘no’ if no library containing function was found, otherwise as the ‘-llibrary’ option that needs to be prepended to LIBS.

5.5 Library Functions

The following macros check for particular C library functions. If there is no macro specifically defined to check for a function you need, and you don’t need to check for any special properties of it, then you can use one of the general function-check macros.

5.5.1 Portability of C Functions

Most usual functions can either be missing, or be buggy, or be limited on some architectures. This section tries to make an inventory of these portability issues. By definition, this list always requires additions. Please help us keeping it as complete as possible.

exit On ancient hosts, exit returned int. This is because exit predates void, and there was a long tradition of it returning int. On current hosts, the problem more likely is that exit is not declared, due to C++ problems of some sort or another. For this reason we suggest that test programs not invoke exit, but return from main instead.

free The C standard says a call free (NULL) does nothing, but some old systems don’t support this (e.g., NextStep).

isinf

isnan The C99 standard says that isinf and isnan are macros. On some systems just macros are available (e.g., HP-UX and Solaris 10), on some systems both macros and functions (e.g., glibc 2.3.2), and on some systems only functions (e.g., IRIX 6 and Solaris 9). In some cases these functions are declared in nonstandard headers like <sunmath.h> and defined in non-default libraries like

‘-lm’ or ‘-lsunmath’.

The C99 isinf and isnan macros work correctly with long double arguments, but pre-C99 systems that use functions typically assume double arguments. On such a system, isinf incorrectly returns true for a finite long double argument that is outside the range of double.

The best workaround for these issues is to use gnulib modules isinf and isnan (see Section 2.2 [Gnulib], page 3). But a lighter weight solution involves code like the following.

      #include <math.h>
      #ifndef isnan
      # define isnan(x) \
          (sizeof (x) == sizeof (long double) ? isnan_ld (x) \
           : sizeof (x) == sizeof (double) ? isnan_d (x) \
           : isnan_f (x))
      static inline int isnan_f  (float       x) { return x != x; }
      static inline int isnan_d  (double      x) { return x != x; }
      static inline int isnan_ld (long double x) { return x != x; }
      #endif
      #ifndef isinf
      # define isinf(x) \
          (sizeof (x) == sizeof (long double) ? isinf_ld (x) \
           : sizeof (x) == sizeof (double) ? isinf_d (x) \
           : isinf_f (x))
      static inline int isinf_f  (float       x)
      { return !isnan (x) && isnan (x - x); }
      static inline int isinf_d  (double      x)
      { return !isnan (x) && isnan (x - x); }
      static inline int isinf_ld (long double x)
      { return !isnan (x) && isnan (x - x); }
      #endif

Use AC_C_INLINE (see Section 5.10.3 [C Compiler], page 81) so that this code works on compilers that lack the inline keyword. Some optimizing compilers mishandle these definitions, but systems with that bug typically have many other floating point corner-case compliance problems anyway, so it’s probably not worth worrying about.

malloc The C standard says a call malloc (0) is implementation dependent. It can return either NULL or a new non-null pointer. The latter is more common (e.g., the GNU C Library) but is by no means universal. AC_FUNC_MALLOC can be used to insist on non-NULL (see Section 5.5.2 [Particular Functions], page 55).

putenv Posix prefers setenv to putenv; among other things, putenv is not required of all Posix implementations, but setenv is.

Posix specifies that putenv puts the given string directly in environ, but some systems make a copy of it instead (e.g., glibc 2.0, or BSD). And when a copy is made, unsetenv might not free it, causing a memory leak (e.g., FreeBSD 4). On some systems putenv ("FOO") removes ‘FOO’ from the environment, but this is not standard usage and it dumps core on some systems (e.g., AIX). On MinGW, a call putenv ("FOO=") removes ‘FOO’ from the environment, rather than inserting it with an empty value.

realloc The C standard says a call realloc (NULL, size) is equivalent to malloc (size), but some old systems don’t support this (e.g., NextStep).

signal handler Normally signal takes a handler function with a return type of void, but some old systems required int instead. Any actual int value returned is not used; this is only a difference in the function prototype demanded. All systems we know of in current use return void. The int was to support K&R C, where of course void is not available. The obsolete macro AC_TYPE_ SIGNAL (see [AC TYPE SIGNAL], page 304) can be used to establish the correct type in all cases. In most cases, it is more robust to use sigaction when it is available, rather

than signal.

snprintf The C99 standard says that if the output array isn’t big enough and if no other errors occur, snprintf and vsnprintf truncate the output and return the number of bytes that ought to have been produced. Some older systems return the truncated length (e.g., GNU C Library 2.0.x or IRIX 6.5), some a negative value (e.g., earlier GNU C Library versions), and some the buffer length without truncation (e.g., 32-bit Solaris 7). Also, some buggy older systems ignore the length and overrun the buffer (e.g., 64-bit Solaris 7).

sprintf The C standard says sprintf and vsprintf return the number of bytes written. On some ancient systems (SunOS 4 for instance) they return the buffer pointer instead, but these no longer need to be worried about.

sscanf On various old systems, e.g., HP-UX 9, sscanf requires that its input string be writable (though it doesn’t actually change it). This can be a problem when using gcc since it normally puts constant strings in read-only memory (see Section “Incompatibilities” in Using and Porting the GNU Compiler Collection). Apparently in some cases even having format strings read-only can be a problem.

strerror_r

Posix specifies that strerror_r returns an int, but many systems (e.g., GNU C Library version 2.2.4) provide a different version returning a char *. AC_FUNC_ STRERROR_R can detect which is in use (see Section 5.5.2 [Particular Functions], page 55).

strnlen AIX 4.3 provides a broken version which produces the following results:

     strnlen ("foobar", 0) = 0
     strnlen ("foobar", 1) = 3

     strnlen ("foobar", 2) = 2
     strnlen ("foobar", 3) = 1
     strnlen ("foobar", 4) = 0
     strnlen ("foobar", 5) = 6
     strnlen ("foobar", 6) = 6
     strnlen ("foobar", 7) = 6
     strnlen ("foobar", 8) = 6
     strnlen ("foobar", 9) = 6

sysconf _SC_PAGESIZE is standard, but some older systems (e.g., HP-UX 9) have _SC_ PAGE_SIZE instead. This can be tested with #ifdef.

unlink The Posix spec says that unlink causes the given file to be removed only after there are no more open file handles for it. Some non-Posix hosts have trouble with this requirement, though, and some DOS variants even corrupt the file system.

unsetenv On MinGW, unsetenv is not available, but a variable ‘FOO’ can be removed with a call putenv ("FOO="), as described under putenv above.

va_copy The C99 standard provides va_copy for copying va_list variables. It may be available in older environments too, though possibly as __va_copy (e.g., gcc in strict pre-C99 mode). These can be tested with #ifdef. A fallback to memcpy

(&dst, &src, sizeof (va_list)) gives maximum portability.

va_list va_list is not necessarily just a pointer. It can be a struct (e.g., gcc on Alpha), which means NULL is not portable. Or it can be an array (e.g., gcc in some PowerPC configurations), which means as a function parameter it can be effectively call-by-reference and library routines might modify the value back in the caller (e.g., vsnprintf in the GNU C Library 2.1).

Signed >> Normally the C >> right shift of a signed type replicates the high bit, giving a so-called “arithmetic” shift. But care should be taken since Standard C doesn’t require that behavior. On those few processors without a native arithmetic shift (for instance Cray vector systems) zero bits may be shifted in, the same as a shift of an unsigned type.

Integer / C divides signed integers by truncating their quotient toward zero, yielding the same result as Fortran. However, before C99 the standard allowed C implementations to take the floor or ceiling of the quotient in some cases. Hardly any implementations took advantage of this freedom, though, and it’s probably not worth worrying about this issue nowadays.

5.5.2 Particular Function Checks

These macros check for particular C functions—whether they exist, and in some cases how they respond when given certain arguments.

[Macro] AC_FUNC_ALLOCA

Check how to get alloca. Tries to get a builtin version by checking for ‘alloca.h’ or the predefined C preprocessor macros __GNUC__and _AIX. If this macro finds

‘alloca.h’, it defines HAVE_ALLOCA_H.

If those attempts fail, it looks for the function in the standard C library. If any of those methods succeed, it defines HAVE_ALLOCA. Otherwise, it sets the output variable ALLOCA to ‘${LIBOBJDIR}alloca.o’ and defines C_ALLOCA (so programs can periodically call ‘alloca (0)’ to garbage collect). This variable is separate from LIBOBJS so multiple programs can share the value of ALLOCA without needing to create an actual library, in case only some of them use the code in LIBOBJS. The ‘${LIBOBJDIR}’ prefix serves the same purpose as in LIBOBJS (see Section 18.6.4 [AC LIBOBJ vs LIBOBJS], page 312).

This macro does not try to get alloca from the System V R3 ‘libPW’ or the System V R4 ‘libucb’ because those libraries contain some incompatible functions that cause trouble. Some versions do not even contain alloca or contain a buggy version. If you still want to use their alloca, use ar to extract ‘alloca.o’ from them instead of compiling ‘alloca.c’.

Source files that use alloca should start with a piece of code like the following, to declare it properly.

     #ifdef HAVE_ALLOCA_H
     # include <alloca.h>
     #elif defined __GNUC__
     # define alloca __builtin_alloca
     #elif defined _AIX
     # define alloca __alloca
     #elif defined _MSC_VER
     # include <malloc.h>
     # define alloca _alloca
     #else
     # include <stddef.h>
     # ifdef  __cplusplus
     extern "C"
     # endif
     void *alloca (size_t);
     #endif
[Macro] AC_FUNC_CHOWN

If the chown function is available and works (in particular, it should accept ‘-1’ for uid and gid), define HAVE_CHOWN. The result of this macro is cached in the ac_cv_

func_chown_works variable.
[Macro] AC_FUNC_CLOSEDIR_VOID

If the closedir function does not return a meaningful value, define CLOSEDIR_VOID. Otherwise, callers ought to check its return value for an error indicator.

Currently this test is implemented by running a test program. When cross compiling the pessimistic assumption that closedir does not return a meaningful value is made.

The result of this macro is cached in the ac_cv_func_closedir_void variable.

This macro is obsolescent, as closedir returns a meaningful value on current systems. New programs need not use this macro.

[Macro] AC_FUNC_ERROR_AT_LINE

If the error_at_line function is not found, require an AC_LIBOBJ replacement of

‘error’.

The result of this macro is cached in the ac_cv_lib_error_at_line variable.

[Macro] AC_FUNC_FNMATCH

If the fnmatch function conforms to Posix, define HAVE_FNMATCH. Detect common implementation bugs, for example, the bugs in Solaris 2.4. Unlike the other specific AC_FUNC macros, AC_FUNC_FNMATCH does not replace a broken/missing fnmatch. This is for historical reasons. See AC_REPLACE_FNMATCH below. The result of this macro is cached in the ac_cv_func_fnmatch_works variable. This macro is obsolescent. New programs should use Gnulib’s fnmatch-posix module. See Section 2.2 [Gnulib], page 3.

[Macro] AC_FUNC_FNMATCH_GNU

Behave like AC_REPLACE_FNMATCH (replace) but also test whether fnmatch supports GNU extensions. Detect common implementation bugs, for example, the bugs in the GNU C Library 2.1. The result of this macro is cached in the ac_cv_func_fnmatch_gnu variable. This macro is obsolescent. New programs should use Gnulib’s fnmatch-gnu module. See Section 2.2 [Gnulib], page 3.

[Macro] AC_FUNC_FORK

This macro checks for the fork and vfork functions. If a working fork is found, define HAVE_WORKING_FORK. This macro checks whether fork is just a stub by trying to run it. If ‘vfork.h’ is found, define HAVE_VFORK_H. If a working vfork is found, define HAVE_ WORKING_VFORK. Otherwise, define vfork to be fork for backward compatibility with previous versions of autoconf. This macro checks for several known errors in implementations of vfork and considers the system to not have a working vfork if it detects any of them. It is not considered to be an implementation error if a child’s invocation of signal modifies the parent’s signal handler, since child processes rarely change their signal handlers. Since this macro defines vfork only for backward compatibility with previous versions of autoconf you’re encouraged to define it yourself in new code:

     #ifndef HAVE_WORKING_VFORK
     # define vfork fork
     #endif
[Macro] AC_FUNC_FSEEKO

If the fseeko function is available, define HAVE_FSEEKO. Define _LARGEFILE_SOURCE if necessary to make the prototype visible on some systems (e.g., glibc 2.2). Otherwise linkage problems may occur when compiling with AC_SYS_LARGEFILE on largefilesensitive systems where off_t does not default to a 64bit entity. All systems with

fseeko also supply ftello.

[Macro] AC_FUNC_GETGROUPS

If the getgroups function is available and works (unlike on Ultrix 4.3, where

‘getgroups (0, 0)’ always fails), define HAVE_GETGROUPS. Set GETGROUPS_LIBS to

any libraries needed to get that function. This macro runs AC_TYPE_GETGROUPS.

[Macro] AC_FUNC_GETLOADAVG

Check how to get the system load averages. To perform its tests properly, this macro needs the file ‘getloadavg.c’; therefore, be sure to set the AC_LIBOBJ replacement directory properly (see Section 5.5.3 [Generic Functions], page 62, AC_CONFIG_LIBOBJ_

DIR).

If the system has the getloadavg function, define HAVE_GETLOADAVG, and set GETLOADAVG_LIBS to any libraries necessary to get that function. Also add GETLOADAVG_LIBS to LIBS. Otherwise, require an AC_LIBOBJ replacement for ‘getloadavg’ with source code in ‘dir/getloadavg.c’, and possibly define several other C preprocessor macros and output variables:

 1. Define C_GETLOADAVG.

2. Define SVR4, DGUX, UMAX, or UMAX4_3 if on those systems.

3. If ‘nlist.h’ is found, define HAVE_NLIST_H.

 4. If ‘struct nlist’ has an ‘n_un.n_name’ member, define HAVE_STRUCT_NLIST_N_

UN_N_NAME. The obsolete symbol NLIST_NAME_UNION is still defined, but do not depend upon it.

5. Programs may need to be installed set-group-ID (or set-user-ID) for getloadavg to work. In this case, define GETLOADAVG_PRIVILEGED, set the output variable NEED_SETGID to ‘true’ (and otherwise to ‘false’), and set KMEM_GROUP to the name of the group that should own the installed program.

The AC_FUNC_GETLOADAVG macro is obsolescent. New programs should use Gnulib’s getloadavg module. See Section 2.2 [Gnulib], page 3.

[Macro] AC_FUNC_GETMNTENT

Check for getmntent in the standard C library, and then in the ‘sun’, ‘seq’, and ‘gen’ libraries, for UNICOS, IRIX 4, PTX, and UnixWare, respectively. Then, if

getmntent is available, define HAVE_GETMNTENT and set ac_cv_func_getmntent to
yes. Otherwise set ac_cv_func_getmntent to no.

The result of this macro can be overridden by setting the cache variable ac_cv_

search_getmntent.
[Macro] AC_FUNC_GETPGRP

Define GETPGRP_VOID if it is an error to pass 0 to getpgrp; this is the Posix behavior. On older BSD systems, you must pass 0 to getpgrp, as it takes an argument and behaves like Posix’s getpgid.

     #ifdef GETPGRP_VOID
       pid = getpgrp ();
     #else
       pid = getpgrp (0);
     #endif

This macro does not check whether getpgrp exists at all; if you need to work in that situation, first call AC_CHECK_FUNC for getpgrp. The result of this macro is cached in the ac_cv_func_getpgrp_void variable. This macro is obsolescent, as current systems have a getpgrp whose signature conforms to Posix. New programs need not use this macro.

[Macro] AC_FUNC_LSTAT_FOLLOWS_SLASHED_SYMLINK

If ‘link’ is a symbolic link, then lstat should treat ‘link/’ the same as ‘link/.’. However, many older lstat implementations incorrectly ignore trailing slashes. It is safe to assume that if lstat incorrectly ignores trailing slashes, then other symbolic-link-aware functions like unlink also incorrectly ignore trailing slashes. If lstat behaves properly, define LSTAT_FOLLOWS_SLASHED_SYMLINK, otherwise require an AC_LIBOBJ replacement of lstat. The result of this macro is cached in the ac_cv_func_lstat_dereferences_

slashed_symlink variable.
[Macro] AC_FUNC_MALLOC

If the malloc function is compatible with the GNU C library malloc (i.e., ‘malloc (0)’ returns a valid pointer), define HAVE_MALLOC to 1. Otherwise define HAVE_MALLOC to 0, ask for an AC_LIBOBJ replacement for ‘malloc’, and define malloc to rpl_malloc so that the native malloc is not used in the main project. Typically, the replacement file ‘malloc.c’ should look like (note the ‘#undef malloc’):

#include <config.h>
#undef malloc
#include <sys/types.h>
void *malloc ();
/* Allocate an N-byte block of memory from the heap.
   If N is zero, allocate a 1-byte block.  */
void *
rpl_malloc (size_t n)
{
  if (n == 0)
    n = 1;
  return malloc (n);
}

The result of this macro is cached in the ac_cv_func_malloc_0_nonnull variable.

[Macro] AC_FUNC_MBRTOWC

Define HAVE_MBRTOWC to 1 if the function mbrtowc and the type mbstate_t are properly declared. The result of this macro is cached in the ac_cv_func_mbrtowc variable.

[Macro] AC_FUNC_MEMCMP

If the memcmp function is not available, or does not work on 8-bit data (like the one on SunOS 4.1.3), or fails when comparing 16 bytes or more and with at least one buffer not starting on a 4-byte boundary (such as the one on NeXT x86 OpenStep), require an AC_LIBOBJ replacement for ‘memcmp’. The result of this macro is cached in the ac_cv_func_memcmp_working variable. This macro is obsolescent, as current systems have a working memcmp. New programs need not use this macro.

[Macro] AC_FUNC_MKTIME

If the mktime function is not available, or does not work correctly, require an AC_ LIBOBJ replacement for ‘mktime’. For the purposes of this test, mktime should conform to the Posix standard and should be the inverse of localtime. The result of this macro is cached in the ac_cv_func_working_mktime variable.

[Macro] AC_FUNC_MMAP

If the mmap function exists and works correctly, define HAVE_MMAP. This checks only private fixed mapping of already-mapped memory. The result of this macro is cached in the ac_cv_func_mmap_fixed_mapped variable.

[Macro] AC_FUNC_OBSTACK

If the obstacks are found, define HAVE_OBSTACK, else require an AC_LIBOBJ replacement for ‘obstack’. The result of this macro is cached in the ac_cv_func_obstack variable.

[Macro] AC_FUNC_REALLOC

If the realloc function is compatible with the GNU C library realloc (i.e., ‘realloc (NULL, 0)’ returns a valid pointer), define HAVE_REALLOC to 1. Otherwise define HAVE_REALLOC to 0, ask for an AC_LIBOBJ replacement for ‘realloc’, and define realloc to rpl_realloc so that the native realloc is not used in the main project.

See AC_FUNC_MALLOC for details.

The result of this macro is cached in the ac_cv_func_realloc_0_nonnull variable.

[Macro] AC_FUNC_SELECT_ARGTYPES

Determines the correct type to be passed for each of the select function’s arguments,

and defines those types in SELECT_TYPE_ARG1, SELECT_TYPE_ARG234, and SELECT_
TYPE_ARG5 respectively. SELECT_TYPE_ARG1 defaults to ‘int’, SELECT_TYPE_ARG234
defaults to ‘int *’, and SELECT_TYPE_ARG5 defaults to ‘struct timeval *’.

This macro is obsolescent, as current systems have a select whose signature conforms to Posix. New programs need not use this macro.

[Macro] AC_FUNC_SETPGRP

If setpgrp takes no argument (the Posix version), define SETPGRP_VOID. Otherwise, it is the BSD version, which takes two process IDs as arguments. This macro does not check whether setpgrp exists at all; if you need to work in that situation, first

call AC_CHECK_FUNC for setpgrp.

The result of this macro is cached in the ac_cv_func_setpgrp_void variable.

This macro is obsolescent, as current systems have a setpgrp whose signature conforms to Posix. New programs need not use this macro.

[Macro] AC_FUNC_STAT
[Macro] AC_FUNC_LSTAT

Determine whether stat or lstat have the bug that it succeeds when given the zerolength file name as argument. The stat and lstat from SunOS 4.1.4 and the Hurd (as of 1998-11-01) do this.

If it does, then define HAVE_STAT_EMPTY_STRING_BUG (or HAVE_LSTAT_EMPTY_

STRING_BUG) and ask for an AC_LIBOBJ replacement of it. The results of these macros are cached in the ac_cv_func_stat_empty_string_bug

and the ac_cv_func_lstat_empty_string_bug variables, respectively.

These macros are obsolescent, as no current systems have the bug. New programs need not use these macros.

[Macro] AC_FUNC_STRCOLL

If the strcoll function exists and works correctly, define HAVE_STRCOLL. This does a bit more than ‘AC_CHECK_FUNCS(strcoll)’, because some systems have incorrect definitions of strcoll that should not be used. The result of this macro is cached in the ac_cv_func_strcoll_works variable.

[Macro] AC_FUNC_STRERROR_R

If strerror_r is available, define HAVE_STRERROR_R, and if it is declared, define

HAVE_DECL_STRERROR_R. If it returns a char * message, define STRERROR_R_CHAR_

P; otherwise it returns an int error number. The Thread-Safe Functions option of Posix requires strerror_r to return int, but many systems (including, for example, version 2.2.4 of the GNU C Library) return a char * value that is not necessarily equal to the buffer argument. The result of this macro is cached in the ac_cv_func_strerror_r_char_p variable.

[Macro] AC_FUNC_STRFTIME

Check for strftime in the ‘intl’ library, for SCO Unix. Then, if strftime is avail-

able, define HAVE_STRFTIME.

This macro is obsolescent, as no current systems require the ‘intl’ library for strftime. New programs need not use this macro.

[Macro] AC_FUNC_STRTOD

If the strtod function does not exist or doesn’t work correctly, ask for an AC_LIBOBJ replacement of ‘strtod’. In this case, because ‘strtod.c’ is likely to need ‘pow’, set the output variable POW_LIB to the extra library needed. This macro caches its result in the ac_cv_func_strtod variable and depends upon the result in the ac_cv_func_pow variable.

[Macro] AC_FUNC_STRTOLD

If the strtold function exists and conforms to C99, define HAVE_STRTOLD. This macro caches its result in the ac_cv_func_strtold variable.

[Macro] AC_FUNC_STRNLEN

If the strnlen function is not available, or is buggy (like the one from AIX 4.3), require an AC_LIBOBJ replacement for it. This macro caches its result in the ac_cv_func_strnlen_working variable.

[Macro] AC_FUNC_UTIME_NULL

If ‘utime (file, NULL)’ sets file’s timestamp to the present, define HAVE_UTIME_

NULL.

This macro caches its result in the ac_cv_func_utime_null variable. This macro is obsolescent, as all current systems have a utime that behaves this way. New programs need not use this macro.

[Macro] AC_FUNC_VPRINTF

If vprintf is found, define HAVE_VPRINTF. Otherwise, if _doprnt is found, define HAVE_DOPRNT. (If vprintf is available, you may assume that vfprintf and vsprintf are also available.) This macro is obsolescent, as all current systems have vprintf. New programs need not use this macro.

[Macro] AC_REPLACE_FNMATCH

If the fnmatch function does not conform to Posix (see AC_FUNC_FNMATCH), ask for its AC_LIBOBJ replacement.

The files ‘fnmatch.c’, ‘fnmatch_loop.c’, and ‘fnmatch_.h’ in the AC_LIBOBJ replace-

ment directory are assumed to contain a copy of the source code of GNU fnmatch. If necessary, this source code is compiled as an AC_LIBOBJ replacement, and the ‘fnmatch_.h’ file is linked to ‘fnmatch.h’ so that it can be included in place of the

system <fnmatch.h>.

This macro caches its result in the ac_cv_func_fnmatch_works variable. This macro is obsolescent, as it assumes the use of particular source files. New programs should use Gnulib’s fnmatch-posix module, which provides this macro along with the source files. See Section 2.2 [Gnulib], page 3.

5.5.3 Generic Function Checks

These macros are used to find functions not covered by the “particular” test macros. If the functions might be in libraries other than the default C library, first call AC_CHECK_LIB for those libraries. If you need to check the behavior of a function as well as find out whether it is present, you have to write your own test for it (see Chapter 6 [Writing Tests], page 99).

[Macro] AC_CHECK_FUNC (function, [action-if-found],
   [action-if-not-found])

If C function function is available, run shell commands action-if-found, otherwise action-if-not-found. If you just want to define a symbol if the function is available, consider using AC_CHECK_FUNCS instead. This macro checks for functions with C linkage even when AC_LANG(C++) has been called, since C is more standardized than C++. (see Section 6.1 [Language Choice], page 99, for more information about selecting the language for checks.) This macro caches its result in the ac_cv_func_function variable.

[Macro] AC_CHECK_FUNCS (function . . ., [action-if-found],
   [action-if-not-found])

For each function enumerated in the blank-or-newline-separated argument list, define HAVE_function (in all capitals) if it is available. If action-if-found is given, it is additional shell code to execute when one of the functions is found. You can give it a value of ‘break’ to break out of the loop on the first match. If action-if-not-found is given, it is executed when one of the functions is not found. Results are cached for each function as in AC_CHECK_FUNC.

[Macro] AC_CHECK_FUNCS_ONCE (function . . .)

For each function enumerated in the blank-or-newline-separated argument list, define HAVE_function (in all capitals) if it is available. This is a once-only variant of AC_ CHECK_FUNCS. It generates the checking code at most once, so that configure is smaller and faster; but the checks cannot be conditionalized and are always done once, early during the configure run.

Autoconf follows a philosophy that was formed over the years by those who have struggled for portability: isolate the portability issues in specific files, and then program as if you were in a Posix environment. Some functions may be missing or unfixable, and your package must be ready to replace them.

Suitable replacements for many such problem functions are available from Gnulib (see Section 2.2 [Gnulib], page 3).

[Macro] AC_LIBOBJ (function)

Specify that ‘function.c’ must be included in the executables to replace a missing or broken implementation of function. Technically, it adds ‘function.$ac_objext’ to the output variable LIBOBJS if it is not already in, and calls AC_LIBSOURCE for ‘function.c’. You should not directly change LIBOBJS, since this is not traceable.

[Macro] AC_LIBSOURCE (file)

Specify that file might be needed to compile the project. If you need to know what files might be needed by a ‘configure.ac’, you should trace AC_LIBSOURCE. file must be a literal. This macro is called automatically from AC_LIBOBJ, but you must call it explicitly if you pass a shell variable to AC_LIBOBJ. In that case, since shell variables cannot be traced statically, you must pass to AC_LIBSOURCE any possible files that the shell variable might cause AC_LIBOBJ to need. For example, if you want to pass a variable $foo_or_bar to AC_LIBOBJ that holds either "foo" or "bar", you should do:

     AC_LIBSOURCE([foo.c])
     AC_LIBSOURCE([bar.c])
     AC_LIBOBJ([$foo_or_bar])

There is usually a way to avoid this, however, and you are encouraged to simply call AC_LIBOBJ with literal arguments. Note that this macro replaces the obsolete AC_LIBOBJ_DECL, with slightly different semantics: the old macro took the function name, e.g., foo, as its argument rather than the file name.

[Macro] AC_LIBSOURCES (files)

Like AC_LIBSOURCE, but accepts one or more files in a comma-separated M4 list. Thus, the above example might be rewritten:

     AC_LIBSOURCES([foo.c, bar.c])
     AC_LIBOBJ([$foo_or_bar])
[Macro] AC_CONFIG_LIBOBJ_DIR (directory)

Specify that AC_LIBOBJ replacement files are to be found in directory, a name relative to the top level of the source tree. The replacement directory defaults to ‘.’, the top level directory, and the most typical value is ‘lib’, corresponding to

‘AC_CONFIG_LIBOBJ_DIR([lib])’.

configure might need to know the replacement directory for the following reasons:

(i) some checks use the replacement files, (ii) some macros bypass broken system headers by installing links to the replacement headers (iii) when used in conjunction with Automake, within each makefile, directory is used as a relative path from $(top_ srcdir) to each object named in LIBOBJS and LTLIBOBJS, etc.

It is common to merely check for the existence of a function, and ask for its AC_LIBOBJ replacement if missing. The following macro is a convenient shorthand.

[Macro] AC_REPLACE_FUNCS (function . . .)

Like AC_CHECK_FUNCS, but uses ‘AC_LIBOBJ(function)’ as action-if-not-found. You can declare your replacement function by enclosing the prototype in ‘#ifndef HAVE_function’. If the system has the function, it probably declares it in a header file you should be including, so you shouldn’t redeclare it lest your declaration conflict.

5.6 Header Files

The following macros check for the presence of certain C header files. If there is no macro specifically defined to check for a header file you need, and you don’t need to check for any special properties of it, then you can use one of the general header-file check macros.

5.6.1 Portability of Headers

This section tries to collect knowledge about common headers, and the problems they cause. By definition, this list always requires additions. Please help us keeping it as complete as possible.

‘limits.h’
     C99 says that ‘limits.h’ defines LLONG_MIN, LLONG_MAX, and ULLONG_MAX, but

many almost-C99 environments (e.g., default GCC 4.0.2 + glibc 2.4) do not define them.

‘inttypes.h’ vs. ‘stdint.h’

The C99 standard says that ‘inttypes.h’ includes ‘stdint.h’, so there’s no need to include ‘stdint.h’ separately in a standard environment. Some implementations have ‘inttypes.h’ but not ‘stdint.h’ (e.g., Solaris 7), but we don’t know of any implementation that has ‘stdint.h’ but not ‘inttypes.h’.

‘linux/irda.h’
It requires ‘linux/types.h’ and ‘sys/socket.h’.
‘linux/random.h’
It requires ‘linux/types.h’.
‘net/if.h’

On Darwin, this file requires that ‘sys/socket.h’ be included beforehand. One should run:

AC_CHECK_HEADERS([sys/socket.h])
AC_CHECK_HEADERS([net/if.h], [], [],
[#include <stdio.h>
#ifdef STDC_HEADERS
# include <stdlib.h>
# include <stddef.h>
#else
# ifdef HAVE_STDLIB_H
#  include <stdlib.h>
# endif
#endif
#ifdef HAVE_SYS_SOCKET_H
# include <sys/socket.h>
#endif
])
‘netinet/if_ether.h’

On Darwin, this file requires that ‘stdio.h’ and ‘sys/socket.h’ be included beforehand. One should run:

AC_CHECK_HEADERS([sys/socket.h])
AC_CHECK_HEADERS([netinet/if_ether.h], [], [],
[#include <stdio.h>
#ifdef STDC_HEADERS
# include <stdlib.h>
# include <stddef.h>
#else
# ifdef HAVE_STDLIB_H
#  include <stdlib.h>
# endif
#endif
#ifdef HAVE_SYS_SOCKET_H
# include <sys/socket.h>
#endif
])
‘stdint.h’

See above, item ‘inttypes.h’ vs. ‘stdint.h’.

‘stdlib.h’

On many systems (e.g., Darwin), ‘stdio.h’ is a prerequisite.

‘sys/mount.h’

On FreeBSD 4.8 on ia32 and using gcc version 2.95.4, ‘sys/params.h’ is a prerequisite.

‘sys/ptem.h’

On Solaris 8, ‘sys/stream.h’ is a prerequisite.

‘sys/socket.h’

On Darwin, ‘stdlib.h’ is a prerequisite.

‘sys/ucred.h’

On Tru64 5.1, ‘sys/types.h’ is a prerequisite.

‘X11/extensions/scrnsaver.h’

Using XFree86, this header requires ‘X11/Xlib.h’, which is probably so required that you might not even consider looking for it.

          AC_CHECK_HEADERS([X11/extensions/scrnsaver.h], [], [],
          [[#include <X11/Xlib.h>
          ]])

5.6.2 Particular Header Checks

These macros check for particular system header files—whether they exist, and in some cases whether they declare certain symbols.

[Macro] AC_HEADER_ASSERT

Check whether to enable assertions in the style of ‘assert.h’. Assertions are enabled by default, but the user can override this by invoking configure with the

‘--disable-assert’ option.

[Macro] AC_HEADER_DIRENT

Check for the following header files. For the first one that is found and defines ‘DIR’, define the listed C preprocessor macro:

‘dirent.h’  HAVE_DIRENT_H
‘sys/ndir.h’ HAVE_SYS_NDIR_H
‘sys/dir.h’ HAVE_SYS_DIR_H
‘ndir.h’  HAVE_NDIR_H

The directory-library declarations in your source code should look something like the following:

     #include <sys/types.h>
     #ifdef HAVE_DIRENT_H
     # include <dirent.h>
     # define NAMLEN(dirent) strlen ((dirent)->d_name)
     #else
     # define dirent direct
     # define NAMLEN(dirent) ((dirent)->d_namlen)
     # ifdef HAVE_SYS_NDIR_H
     #  include <sys/ndir.h>
     # endif
     # ifdef HAVE_SYS_DIR_H
     #  include <sys/dir.h>
     # endif
     # ifdef HAVE_NDIR_H
     #  include <ndir.h>
     # endif
     #endif

Using the above declarations, the program would declare variables to be of type struct dirent, not struct direct, and would access the length of a directory entry name by passing a pointer to a struct dirent to the NAMLEN macro. This macro also checks for the SCO Xenix ‘dir’ and ‘x’ libraries. This macro is obsolescent, as all current systems with directory libraries have <dirent.h>. New programs need not use this macro.

Also see AC_STRUCT_DIRENT_D_INO and AC_STRUCT_DIRENT_D_TYPE (see Section 5.8.1

[Particular Structures], page 74).

[Macro] AC_HEADER_MAJOR
If ‘sys/types.h’ does not define major, minor, and makedev, but ‘sys/mkdev.h’
does, define MAJOR_IN_MKDEV; otherwise, if ‘sys/sysmacros.h’ does, define MAJOR_
IN_SYSMACROS.
[Macro] AC_HEADER_RESOLV

Checks for header ‘resolv.h’, checking for prerequisites first. To properly use ‘resolv.h’, your code should contain something like the following:

#ifdef HAVE_SYS_TYPES_H
#  include <sys/types.h>
#endif
#ifdef HAVE_NETINET_IN_H
#  include <netinet/in.h>   /* inet_functions / structs */
#endif
#ifdef HAVE_ARPA_NAMESER_H
#  include <arpa/nameser.h> /* DNS HEADER struct */
#endif
#ifdef HAVE_NETDB_H
#  include <netdb.h>
#endif
#include <resolv.h>

[Macro] AC_HEADER_STAT

If the macros S_ISDIR, S_ISREG, etc. defined in ‘sys/stat.h’ do not work properly (returning false positives), define STAT_MACROS_BROKEN. This is the case on Tektronix UTekV, Amdahl UTS and Motorola System V/88. This macro is obsolescent, as no current systems have the bug. New programs need not use this macro.

[Macro] AC_HEADER_STDBOOL

If ‘stdbool.h’ exists and conforms to C99, define HAVE_STDBOOL_H to 1; if the type _Bool is defined, define HAVE__BOOL to 1. To fulfill the C99 requirements, your ‘system.h’ could contain the following code:

#ifdef HAVE_STDBOOL_H
# include <stdbool.h>
#else
# ifndef HAVE__BOOL
#  ifdef __cplusplus
typedef bool _Bool;
#  else
#   define _Bool signed char
#  endif
# endif
# define bool _Bool
# define false 0
# define true 1
# define __bool_true_false_are_defined 1
#endif

Alternatively you can use the ‘stdbool’ package of Gnulib (see Section 2.2 [Gnulib], page 3); it packages the above code into a replacement header and contains a few other bells and whistles. This macro caches its result in the ac_cv_header_stdbool_h variable.

[Macro] AC_HEADER_STDC

Define STDC_HEADERS if the system has C header files conforming to ANSI C89 (ISO C90). Specifically, this macro checks for ‘stdlib.h’, ‘stdarg.h’, ‘string.h’, and ‘float.h’; if the system has those, it probably has the rest of the C89 header files. This macro also checks whether ‘string.h’ declares memchr (and thus presumably the other mem functions), whether ‘stdlib.h’ declare free (and thus presumably malloc and other related functions), and whether the ‘ctype.h’ macros work on characters with the high bit set, as the C standard requires. If you use this macro, your code can refer to STDC_HEADERS to determine whether the system has conforming header files (and probably C library functions). This macro caches its result in the ac_cv_header_stdc variable. This macro is obsolescent, as current systems have conforming header files. New programs need not use this macro. Nowadays ‘string.h’ is part of the C standard and declares functions like strcpy, and ‘strings.h’ is standardized by Posix and declares BSD functions like bcopy; but historically, string functions were a major sticking point in this area. If you still want to worry about portability to ancient systems without standard headers, there is so much variation that it is probably easier to declare the functions you use than to figure out exactly what the system header files declare. Some ancient systems contained a mix of functions from the C standard and from BSD; some were mostly standard but lacked ‘memmove’; some defined the BSD functions as macros in ‘string.h’ or ‘strings.h’; some had only the BSD functions but ‘string.h’; some declared the memory functions in ‘memory.h’, some in ‘string.h’; etc. It is probably sufficient to check for one string function and one memory function; if the library had the standard versions of those then it probably had most of the others. If you put the following in

‘configure.ac’:
     # This example is obsolescent.
     # Nowadays you can omit these macro calls.
     AC_HEADER_STDC
     AC_CHECK_FUNCS([strchr memcpy])

then, in your code, you can use declarations like this:

     /* This example is obsolescent.
        Nowadays you can just #include <string.h>.  */
     #ifdef STDC_HEADERS
     # include <string.h>
     #else
     # ifndef HAVE_STRCHR
     #  define strchr index
     #  define strrchr rindex
     # endif
     char *strchr (), *strrchr ();
     # ifndef HAVE_MEMCPY
     #  define memcpy(d, s, n) bcopy ((s), (d), (n))
     #  define memmove(d, s, n) bcopy ((s), (d), (n))
     # endif
     #endif

If you use a function like memchr, memset, strtok, or strspn, which have no BSD equivalent, then macros don’t suffice to port to ancient hosts; you must provide an implementation of each function. An easy way to incorporate your implementations only when needed (since the ones in system C libraries may be hand optimized) is to, taking

memchr for example, put it in ‘memchr.c’ and use ‘AC_REPLACE_FUNCS([memchr])’.

[Macro] AC_HEADER_SYS_WAIT

If ‘sys/wait.h’ exists and is compatible with Posix, define HAVE_SYS_WAIT_H. Incompatibility can occur if ‘sys/wait.h’ does not exist, or if it uses the old BSD union wait instead of int to store a status value. If ‘sys/wait.h’ is not Posix compatible, then instead of including it, define the Posix macros with their usual interpretations. Here is an example:

#include <sys/types.h>
#ifdef HAVE_SYS_WAIT_H
# include <sys/wait.h>
#endif
#ifndef WEXITSTATUS
# define WEXITSTATUS(stat_val) ((unsigned int) (stat_val) >> 8)
#endif
#ifndef WIFEXITED
# define WIFEXITED(stat_val) (((stat_val) & 255) == 0)
#endif

This macro caches its result in the ac_cv_header_sys_wait_h variable. This macro is obsolescent, as current systems are compatible with Posix. New programs need not use this macro.

_POSIX_VERSION is defined when ‘unistd.h’ is included on Posix systems. If there is no ‘unistd.h’, it is definitely not a Posix system. However, some non-Posix systems do have

‘unistd.h’.

The way to check whether the system supports Posix is:

#ifdef HAVE_UNISTD_H
# include <sys/types.h>
# include <unistd.h>
#endif
#ifdef _POSIX_VERSION
/* Code for Posix systems.  */
#endif
[Macro] AC_HEADER_TIME

If a program may include both ‘time.h’ and ‘sys/time.h’, define TIME_WITH_SYS_ TIME. On some ancient systems, ‘sys/time.h’ included ‘time.h’, but ‘time.h’ was not protected against multiple inclusion, so programs could not explicitly include both files. This macro is useful in programs that use, for example, struct timeval as well as struct tm. It is best used in conjunction with HAVE_SYS_TIME_H, which can be

checked for using AC_CHECK_HEADERS([sys/time.h]).
#ifdef TIME_WITH_SYS_TIME
# include <sys/time.h>
# include <time.h>
#else
# ifdef HAVE_SYS_TIME_H
#  include <sys/time.h>
# else
#  include <time.h>
# endif
#endif

This macro caches its result in the ac_cv_header_time variable. This macro is obsolescent, as current systems can include both files when they exist. New programs need not use this macro.

[Macro] AC_HEADER_TIOCGWINSZ
If the use of TIOCGWINSZ requires ‘<sys/ioctl.h>’, then define GWINSZ_IN_SYS_
IOCTL. Otherwise TIOCGWINSZ can be found in ‘<termios.h>’.

Use:

     #ifdef HAVE_TERMIOS_H
     # include <termios.h>
     #endif
     #ifdef GWINSZ_IN_SYS_IOCTL
     # include <sys/ioctl.h>
     #endif

5.6.3 Generic Header Checks

These macros are used to find system header files not covered by the “particular” test macros. If you need to check the contents of a header as well as find out whether it is present, you have to write your own test for it (see Chapter 6 [Writing Tests], page 99).

[Macro] AC_CHECK_HEADER (header-file, [action-if-found],
   [action-if-not-found], [includes])

If the system header file header-file is compilable, execute shell commands action-if-found, otherwise execute action-if-not-found. If you just want to define a symbol if the header file is available, consider using AC_CHECK_HEADERS instead. includes is decoded to determine the appropriate include directives. If omitted or empty, ‘configure’ will check for both header existence (with the preprocessor) and usability (with the compiler), using AC_INCLUDES_DEFAULT for the compile test. If there is a discrepancy between the results, a warning is issued to the user, and the compiler results are favored (see Section 20.7 [Present But Cannot Be Compiled], page 332). In general, favoring the compiler results means that a header will be treated as not found even though the file exists, because you did not provide enough prerequisites. Providing a non-empty includes argument allows the code to provide any prerequisites prior to including the header under test; it is common to use the argument AC_ INCLUDES_DEFAULT (see Section 5.1.2 [Default Includes], page 43). With an explicit fourth argument, no preprocessor test is needed. As a special case, an includes of exactly ‘-’ triggers the older preprocessor check, which merely determines existence of the file in the preprocessor search path; this should only be used as a last resort (it is safer to determine the actual prerequisites and perform a compiler check, or else use AC_PREPROC_IFELSE to make it obvious that only a preprocessor check is desired). This macro caches its result in the ac_cv_header_header-file variable, with characters not suitable for a variable name mapped to underscores.

[Macro] AC_CHECK_HEADERS (header-file . . ., [action-if-found],
   [action-if-not-found], [includes])

For each given system header file header-file in the blank-separated argument list that exists, define HAVE_header-file (in all capitals). If action-if-found is given, it is additional shell code to execute when one of the header files is found. You can give it a value of ‘break’ to break out of the loop on the first match. If action-if-not-found is given, it is executed when one of the header files is not found. includes is interpreted as in AC_CHECK_HEADER, in order to choose the set of preprocessor directives supplied before the header under test. This macro caches its result in the ac_cv_header_header-file variable, with characters not suitable for a variable name mapped to underscores.

Previous versions of Autoconf merely checked whether the header was accepted by the preprocessor. This was changed because the old test was inappropriate for typical uses. Headers are typically used to compile, not merely to preprocess, and the old behavior sometimes accepted headers that clashed at compile-time (see Section 20.7 [Present But Cannot Be Compiled], page 332). If you need to check whether a header is preprocessable, you can use AC_PREPROC_IFELSE (see Section 6.3 [Running the Preprocessor], page 104).

Actually requiring a header to compile improves the robustness of the test, but it also requires that you make sure that headers that must be included before the header-file be part of the includes, (see Section 5.1.2 [Default Includes], page 43). If looking for ‘bar.h’, which requires that ‘foo.h’ be included before if it exists, we suggest the following scheme:

AC_CHECK_HEADERS([foo.h])
AC_CHECK_HEADERS([bar.h], [], [],
[#ifdef HAVE_FOO_H
# include <foo.h>
#endif
])

The following variant generates smaller, faster configure files if you do not need the

full power of AC_CHECK_HEADERS.
[Macro] AC_CHECK_HEADERS_ONCE (header-file . . .)

For each given system header file header-file in the blank-separated argument list that exists, define HAVE_header-file (in all capitals). This is a once-only variant of AC_CHECK_HEADERS. It generates the checking code at most once, so that configure is smaller and faster; but the checks cannot be conditionalized and are always done once, early during the configure run. Thus, this macro is only safe for checking headers that do not have prerequisites beyond what AC_INCLUDES_DEFAULT provides.

5.7 Declarations

The following macros check for the declaration of variables and functions. If there is no macro specifically defined to check for a symbol you need, then you can use the general macros (see Section 5.7.2 [Generic Declarations], page 72) or, for more complex tests, you may use AC_COMPILE_IFELSE (see Section 6.4 [Running the Compiler], page 106).

5.7.1 Particular Declaration Checks

There are no specific macros for declarations.

5.7.2 Generic Declaration Checks

These macros are used to find declarations not covered by the “particular” test macros.

[Macro] AC_CHECK_DECL (symbol, [action-if-found], [action-if-not-found],
   [includes = ‘AC_INCLUDES_DEFAULT’])

If symbol (a function, variable, or constant) is not declared in includes and a declaration is needed, run the shell commands action-if-not-found, otherwise action-if-found. includes is a series of include directives, defaulting to AC_INCLUDES_DEFAULT (see Section 5.1.2 [Default Includes], page 43), which are used prior to the declaration under test.

This macro actually tests whether symbol is defined as a macro or can be used as an r-value, not whether it is really declared, because it is much safer to avoid introducing extra declarations when they are not needed. In order to facilitate use of C++ and overloaded function declarations, it is possible to specify function argument types in parentheses for types which can be zero-initialized:

     AC_CHECK_DECL([basename(char *)])

This macro caches its result in the ac_cv_have_decl_symbol variable, with characters not suitable for a variable name mapped to underscores.

[Macro] AC_CHECK_DECLS (symbols, [action-if-found],
   [action-if-not-found], [includes = ‘AC_INCLUDES_DEFAULT’])

For each of the symbols (comma-separated list with optional function argument types for C++ overloads), define HAVE_DECL_symbol (in all capitals) to ‘1’ if symbol is declared, otherwise to ‘0’. If action-if-not-found is given, it is additional shell code to execute when one of the function declarations is needed, otherwise action-if-found is executed.

includes is a series of include directives, defaulting to AC_INCLUDES_DEFAULT (see Section 5.1.2 [Default Includes], page 43), which are used prior to the declarations under test.

This macro uses an M4 list as first argument:

     AC_CHECK_DECLS([strdup])
     AC_CHECK_DECLS([strlen])
     AC_CHECK_DECLS([malloc, realloc, calloc, free])
     AC_CHECK_DECLS([j0], [], [], [[#include <math.h>]])
     AC_CHECK_DECLS([[basename(char *)], [dirname(char *)]])

Unlike the other ‘AC_CHECK_*S’ macros, when a symbol is not declared, HAVE_DECL_ symbol is defined to ‘0’ instead of leaving HAVE_DECL_symbol undeclared. When you are sure that the check was performed, use HAVE_DECL_symbol in #if:

     #if !HAVE_DECL_SYMBOL
     extern char *symbol;
     #endif

If the test may have not been performed, however, because it is safer not to declare a symbol than to use a declaration that conflicts with the system’s one, you should use:

     #if defined HAVE_DECL_MALLOC && !HAVE_DECL_MALLOC
     void *malloc (size_t *s);
     #endif

You fall into the second category only in extreme situations: either your files may be used without being configured, or they are used during the configuration. In most cases the traditional approach is enough. This macro caches its results in ac_cv_have_decl_symbol variables, with characters not suitable for a variable name mapped to underscores.

[Macro] AC_CHECK_DECLS_ONCE (symbols)

For each of the symbols (comma-separated list), define HAVE_DECL_symbol (in all capitals) to ‘1’ if symbol is declared in the default include files, otherwise to ‘0’. This is a once-only variant of AC_CHECK_DECLS. It generates the checking code at most once, so that configure is smaller and faster; but the checks cannot be conditionalized and are always done once, early during the configure run.

5.8 Structures

The following macros check for the presence of certain members in C structures. If there is no macro specifically defined to check for a member you need, then you can use the general structure-member macros (see Section 5.8.2 [Generic Structures], page 75) or, for more complex tests, you may use AC_COMPILE_IFELSE (see Section 6.4 [Running the Compiler], page 106).

5.8.1 Particular Structure Checks

The following macros check for certain structures or structure members.

[Macro] AC_STRUCT_DIRENT_D_INO

Perform all the actions of AC_HEADER_DIRENT (see Section 5.6.2 [Particular Headers], page 66). Then, if struct dirent contains a d_ino member, define HAVE_STRUCT_

DIRENT_D_INO.

HAVE_STRUCT_DIRENT_D_INO indicates only the presence of d_ino, not whether its contents are always reliable. Traditionally, a zero d_ino indicated a deleted directory entry, though current systems hide this detail from the user and never return zero d_ino values. Many current systems report an incorrect d_ino for a directory entry that is a mount point.

[Macro] AC_STRUCT_DIRENT_D_TYPE

Perform all the actions of AC_HEADER_DIRENT (see Section 5.6.2 [Particular Headers], page 66). Then, if struct dirent contains a d_type member, define HAVE_STRUCT_

DIRENT_D_TYPE.
[Macro] AC_STRUCT_ST_BLOCKS
If struct stat contains an st_blocks member, define HAVE_STRUCT_STAT_ST_

BLOCKS. Otherwise, require an AC_LIBOBJ replacement of ‘fileblocks’. The former name, HAVE_ST_BLOCKS is to be avoided, as its support will cease in the future. This macro caches its result in the ac_cv_member_struct_stat_st_blocks variable.

[Macro] AC_STRUCT_TM

If ‘time.h’ does not define struct tm, define TM_IN_SYS_TIME, which means that including ‘sys/time.h’ had better define struct tm.

This macro is obsolescent, as ‘time.h’ defines struct tm in current systems. New programs need not use this macro.

[Macro] AC_STRUCT_TIMEZONE

Figure out how to get the current timezone. If struct tm has a tm_zone member, define HAVE_STRUCT_TM_TM_ZONE (and the obsoleted HAVE_TM_ZONE). Otherwise, if the external array tzname is found, define HAVE_TZNAME; if it is declared, define HAVE_

DECL_TZNAME.

5.8.2 Generic Structure Checks

These macros are used to find structure members not covered by the “particular” test macros.

[Macro] AC_CHECK_MEMBER (aggregate.member, [action-if-found],
   [action-if-not-found], [includes = ‘AC_INCLUDES_DEFAULT’])

Check whether member is a member of the aggregate aggregate. If no includes are specified, the default includes are used (see Section 5.1.2 [Default Includes], page 43).

     AC_CHECK_MEMBER([struct passwd.pw_gecos], [],
                     [AC_MSG_ERROR([we need `passwd.pw_gecos'])],
                     [[#include <pwd.h>]])

You can use this macro for submembers:

     AC_CHECK_MEMBER(struct top.middle.bot)

This macro caches its result in the av_cv_member_aggregate_member variable, with characters not suitable for a variable name mapped to underscores.

[Macro] AC_CHECK_MEMBERS (members, [action-if-found],
   [action-if-not-found], [includes = ‘AC_INCLUDES_DEFAULT’])

Check for the existence of each ‘aggregate.member’ of members using the previous macro. When member belongs to aggregate, define HAVE_aggregate_member (in all capitals, with spaces and dots replaced by underscores). If action-if-found is given, it is executed for each of the found members. If action-if-not-found is given, it is executed for each of the members that could not be found.

includes is a series of include directives, defaulting to AC_INCLUDES_DEFAULT (see Section 5.1.2 [Default Includes], page 43), which are used prior to the members under test.

This macro uses M4 lists:

     AC_CHECK_MEMBERS([struct stat.st_rdev, struct stat.st_blksize])

5.9 Types

The following macros check for C types, either builtin or typedefs. If there is no macro specifically defined to check for a type you need, and you don’t need to check for any special properties of it, then you can use a general type-check macro.

5.9.1 Particular Type Checks

These macros check for particular C types in ‘sys/types.h’, ‘stdlib.h’, ‘stdint.h’, ‘inttypes.h’ and others, if they exist.

The Gnulib stdint module is an alternate way to define many of these symbols; it is useful if you prefer your code to assume a C99-or-better environment. See Section 2.2 [Gnulib], page 3.

[Macro] AC_TYPE_GETGROUPS

Define GETGROUPS_T to be whichever of gid_t or int is the base type of the array argument to getgroups. This macro caches the base type in the ac_cv_type_getgroups variable.

[Macro] AC_TYPE_INT8_T

If ‘stdint.h’ or ‘inttypes.h’ does not define the type int8_t, define int8_t to a signed integer type that is exactly 8 bits wide and that uses two’s complement representation, if such a type exists. If you are worried about porting to hosts that lack such a type, you can use the results of this macro in C89-or-later code as follows:

     #if HAVE_STDINT_H
     # include <stdint.h>
     #endif
     #if defined INT8_MAX || defined int8_t
      code using int8_t
     #else
      complicated alternative using >8-bit 'signed char'
     #endif

This macro caches the type in the ac_cv_c_int8_t variable.

[Macro] AC_TYPE_INT16_T

This is like AC_TYPE_INT8_T, except for 16-bit integers.

[Macro] AC_TYPE_INT32_T

This is like AC_TYPE_INT8_T, except for 32-bit integers.

[Macro] AC_TYPE_INT64_T

This is like AC_TYPE_INT8_T, except for 64-bit integers.

[Macro] AC_TYPE_INTMAX_T
If ‘stdint.h’ or ‘inttypes.h’ defines the type intmax_t, define HAVE_INTMAX_T.

Otherwise, define intmax_t to the widest signed integer type.

[Macro] AC_TYPE_INTPTR_T
If ‘stdint.h’ or ‘inttypes.h’ defines the type intptr_t, define HAVE_INTPTR_T.

Otherwise, define intptr_t to a signed integer type wide enough to hold a pointer, if such a type exists.

[Macro] AC_TYPE_LONG_DOUBLE

If the C compiler supports a working long double type, define HAVE_LONG_DOUBLE. The long double type might have the same range and precision as double.

This macro caches its result in the ac_cv_type_long_double variable. This macro is obsolescent, as current C compilers support long double. New programs need not use this macro.

[Macro] AC_TYPE_LONG_DOUBLE_WIDER

If the C compiler supports a working long double type with more range or precision

than the double type, define HAVE_LONG_DOUBLE_WIDER.

This macro caches its result in the ac_cv_type_long_double_wider variable.

[Macro] AC_TYPE_LONG_LONG_INT

If the C compiler supports a working long long int type, define HAVE_LONG_LONG_ INT. However, this test does not test long long int values in preprocessor #if expressions, because too many compilers mishandle such expressions. See Section 13.3 [Preprocessor Arithmetic], page 262. This macro caches its result in the ac_cv_type_long_long_int variable.

[Macro] AC_TYPE_MBSTATE_T

Define HAVE_MBSTATE_T if <wchar.h> declares the mbstate_t type. Also, define mbstate_t to be a type if <wchar.h> does not declare it. This macro caches its result in the ac_cv_type_mbstate_t variable.

[Macro] AC_TYPE_MODE_T

Define mode_t to a suitable type, if standard headers do not define it. This macro caches its result in the ac_cv_type_mode_t variable.

[Macro] AC_TYPE_OFF_T

Define off_t to a suitable type, if standard headers do not define it. This macro caches its result in the ac_cv_type_off_t variable.

[Macro] AC_TYPE_PID_T

Define pid_t to a suitable type, if standard headers do not define it. This macro caches its result in the ac_cv_type_pid_t variable.

[Macro] AC_TYPE_SIZE_T

Define size_t to a suitable type, if standard headers do not define it. This macro caches its result in the ac_cv_type_size_t variable.

[Macro] AC_TYPE_SSIZE_T

Define ssize_t to a suitable type, if standard headers do not define it. This macro caches its result in the ac_cv_type_ssize_t variable.

[Macro] AC_TYPE_UID_T

Define uid_t and gid_t to suitable types, if standard headers do not define them. This macro caches its result in the ac_cv_type_uid_t variable.

[Macro] AC_TYPE_UINT8_T

If ‘stdint.h’ or ‘inttypes.h’ does not define the type uint8_t, define uint8_t to an unsigned integer type that is exactly 8 bits wide, if such a type exists. This is like AC_TYPE_INT8_T, except for unsigned integers.

[Macro] AC_TYPE_UINT16_T

This is like AC_TYPE_UINT8_T, except for 16-bit integers.

[Macro] AC_TYPE_UINT32_T

This is like AC_TYPE_UINT8_T, except for 32-bit integers.

[Macro] AC_TYPE_UINT64_T

This is like AC_TYPE_UINT8_T, except for 64-bit integers.

[Macro] AC_TYPE_UINTMAX_T
If ‘stdint.h’ or ‘inttypes.h’ defines the type uintmax_t, define HAVE_UINTMAX_T.

Otherwise, define uintmax_t to the widest unsigned integer type.

[Macro] AC_TYPE_UINTPTR_T
If ‘stdint.h’ or ‘inttypes.h’ defines the type uintptr_t, define HAVE_UINTPTR_

T. Otherwise, define uintptr_t to an unsigned integer type wide enough to hold a pointer, if such a type exists.

[Macro] AC_TYPE_UNSIGNED_LONG_LONG_INT

If the C compiler supports a working unsigned long long int type, define HAVE_

UNSIGNED_LONG_LONG_INT. However, this test does not test unsigned long long

int values in preprocessor #if expressions, because too many compilers mishandle such expressions. See Section 13.3 [Preprocessor Arithmetic], page 262. This macro caches its result in the ac_cv_type_unsigned_long_long_int variable.

5.9.2 Generic Type Checks

These macros are used to check for types not covered by the “particular” test macros.

[Macro] AC_CHECK_TYPE (type, [action-if-found], [action-if-not-found],
   [includes = ‘AC_INCLUDES_DEFAULT’])

Check whether type is defined. It may be a compiler builtin type or defined by the includes. includes is a series of include directives, defaulting to AC_INCLUDES_DEFAULT (see Section 5.1.2 [Default Includes], page 43), which are used prior to the type under test. In C, type must be a type-name, so that the expression ‘sizeof (type)’ is valid (but ‘sizeof ((type))’ is not). The same test is applied when compiling for C++, which means that in C++ type should be a type-id and should not be an anonymous

‘struct’ or ‘union’.

This macro caches its result in the ac_cv_type_type variable, with ‘*’ mapped to ‘p’ and other characters not suitable for a variable name mapped to underscores.

[Macro] AC_CHECK_TYPES (types, [action-if-found], [action-if-not-found],
   [includes = ‘AC_INCLUDES_DEFAULT’])

For each type of the types that is defined, define HAVE_type (in all capitals). Each type must follow the rules of AC_CHECK_TYPE. If no includes are specified, the default includes are used (see Section 5.1.2 [Default Includes], page 43). If action-if-found is given, it is additional shell code to execute when one of the types is found. If action-if-not-found is given, it is executed when one of the types is not found. This macro uses M4 lists:

          AC_CHECK_TYPES([ptrdiff_t])
          AC_CHECK_TYPES([unsigned long long int, uintmax_t])
          AC_CHECK_TYPES([float_t], [], [], [[#include <math.h>]])

Autoconf, up to 2.13, used to provide to another version of AC_CHECK_TYPE, broken by design. In order to keep backward compatibility, a simple heuristic, quite safe but not totally, is implemented. In case of doubt, read the documentation of the former AC_CHECK_ TYPE, see Section 18.4 [Obsolete Macros], page 293.

5.10 Compilers and Preprocessors

All the tests for compilers (AC_PROG_CC, AC_PROG_CXX, AC_PROG_F77) define the output variable EXEEXT based on the output of the compiler, typically to the empty string if Posix and ‘.exe’ if a DOS variant.

They also define the output variable OBJEXT based on the output of the compiler, after ‘.c’ files have been excluded, typically to ‘o’ if Posix, ‘obj’ if a DOS variant.

If the compiler being used does not produce executables, the tests fail. If the executables can’t be run, and cross-compilation is not enabled, they fail too. See Chapter 14 [Manual Configuration], page 267, for more on support for cross compiling.

5.10.1 Specific Compiler Characteristics

Some compilers exhibit different behaviors.

Static/Dynamic Expressions Autoconf relies on a trick to extract one bit of information from the C compiler: using negative array sizes. For instance the following excerpt of a C source demonstrates how to test whether ‘int’ objects are 4 bytes wide:

               static int test_array[sizeof (int) == 4 ? 1 : -1];

To our knowledge, there is a single compiler that does not support this trick: the HP C compilers (the real ones, not only the “bundled”) on HP-UX 11.00. They incorrectly reject the above program with the diagnostic “Variable-length arrays cannot have static storage.” This bug comes from HP compilers’ mishandling of sizeof (int), not from the ? 1 : -1, and Autoconf works around this problem by casting sizeof (int) to long int before comparing it.

5.10.2 Generic Compiler Characteristics

[Macro] AC_CHECK_SIZEOF (type-or-expr, [unused], [includes =
        ‘AC_INCLUDES_DEFAULT’])

Define SIZEOF_type-or-expr (see Section 5.1.1 [Standard Symbols], page 43) to be the size in bytes of type-or-expr, which may be either a type or an expression returning a value that has a size. If the expression ‘sizeof (type-or-expr)’ is invalid, the result is 0. includes is a series of include directives, defaulting to AC_INCLUDES_ DEFAULT (see Section 5.1.2 [Default Includes], page 43), which are used prior to the expression under test. This macro now works even when cross-compiling. The unused argument was used when cross-compiling. For example, the call

     AC_CHECK_SIZEOF([int *])

defines SIZEOF_INT_P to be 8 on DEC Alpha AXP systems. This macro caches its result in the ac_cv_sizeof_type-or-expr variable, with ‘*’ mapped to ‘p’ and other characters not suitable for a variable name mapped to underscores.

[Macro] AC_CHECK_ALIGNOF (type, [includes = ‘AC_INCLUDES_DEFAULT’])

Define ALIGNOF_type (see Section 5.1.1 [Standard Symbols], page 43) to be the alignment in bytes of type. ‘type y;’ must be valid as a structure member declaration. If ‘type’ is unknown, the result is 0. If no includes are specified, the default includes are used (see Section 5.1.2 [Default Includes], page 43). This macro caches its result in the ac_cv_alignof_type-or-expr variable, with ‘*’ mapped to ‘p’ and other characters not suitable for a variable name mapped to underscores.

[Macro] AC_COMPUTE_INT (var, expression, [includes =
   ‘AC_INCLUDES_DEFAULT’], [action-if-fails])

Store into the shell variable var the value of the integer expression. The value should fit in an initializer in a C variable of type signed long. To support cross compilation (in which case, the macro only works on hosts that use twos-complement arithmetic), it should be possible to evaluate the expression at compile-time. If no includes are specified, the default includes are used (see Section 5.1.2 [Default Includes], page 43). Execute action-if-fails if the value cannot be determined correctly.

[Macro] AC_LANG_WERROR

Normally Autoconf ignores warnings generated by the compiler, linker, and preprocessor. If this macro is used, warnings count as fatal errors for the current language. This macro is useful when the results of configuration are used where warnings are unacceptable; for instance, if parts of a program are built with the GCC ‘-Werror’ option. If the whole program is built using ‘-Werror’ it is often simpler to put ‘-Werror’ in the compiler flags (CFLAGS, etc.).

[Macro] AC_OPENMP

OpenMP (http://www.openmp.org/) specifies extensions of C, C++, and Fortran that simplify optimization of shared memory parallelism, which is a common problem on multicore CPUs. If the current language is C, the macro AC_OPENMP sets the variable OPENMP_CFLAGS to the C compiler flags needed for supporting OpenMP. OPENMP_CFLAGS is set to empty if the compiler already supports OpenMP, if it has no way to activate OpenMP support, or if the user rejects OpenMP support by invoking ‘configure’ with the

‘--disable-openmp’ option.

OPENMP_CFLAGS needs to be used when compiling programs, when preprocessing program source, and when linking programs. Therefore you need to add $(OPENMP_ CFLAGS) to the CFLAGS of C programs that use OpenMP. If you preprocess OpenMPspecific C code, you also need to add $(OPENMP_CFLAGS) to CPPFLAGS. The presence of OpenMP support is revealed at compile time by the preprocessor macro _OPENMP.

Linking a program with OPENMP_CFLAGS typically adds one more shared library to the program’s dependencies, so its use is recommended only on programs that actually require OpenMP. If the current language is C++, AC_OPENMP sets the variable OPENMP_CXXFLAGS, suitably for the C++ compiler. The same remarks hold as for C. If the current language is Fortran 77 or Fortran, AC_OPENMP sets the variable OPENMP_ FFLAGS or OPENMP_FCFLAGS, respectively. Similar remarks as for C hold, except that CPPFLAGS is not used for Fortran, and no preprocessor macro signals OpenMP support. For portability, it is best to avoid spaces between ‘#’ and ‘pragma omp’. That is, write ‘#pragma omp’, not ‘# pragma omp’. The Sun WorkShop 6.2 C compiler chokes on the latter.

5.10.3 C Compiler Characteristics

The following macros provide ways to find and exercise a C Compiler. There are a few constructs that ought to be avoided, but do not deserve being checked for, since they can easily be worked around.

Don’t use lines containing solitary backslashes They tickle a bug in the HP-UX C compiler (checked on HP-UX 10.20, 11.00, and 11i). When given the following source:

#ifdef __STDC__
/\
* A comment with backslash-newlines in it.  %{ %} *\
\
/
char str[] = "\\
" A string with backslash-newlines in it %{ %} \\
"";
char apostrophe = '\\
\
'\
';
#endif

the compiler incorrectly fails with the diagnostics “Non-terminating comment at end of file” and “Missing ‘#endif’ at end of file.” Removing the lines with solitary backslashes solves the problem.

Don’t compile several files at once if output matters to you Some compilers, such as HP’s, report names of files being compiled when given more than one file operand. For instance:

$ cc a.c b.c
a.c:
b.c:

This can cause problems if you observe the output of the compiler to detect failures. Invoking ‘cc -c a.c && cc -c b.c && cc -o c a.o b.o’ solves the issue.

Don’t rely on #error failing The IRIX C compiler does not fail when #error is preprocessed; it simply emits a diagnostic and continues, exiting successfully. So, instead of an error directive like #error "Unsupported word size" it is more portable to use an invalid directive like #Unsupported word size in Autoconf tests. In ordinary source code, #error is OK, since installers with inadequate compilers like IRIX can simply examine these compilers’ diagnostic output.

Don’t rely on correct #line support On Solaris, c89 (at least Sun C 5.3 through 5.8) diagnoses #line directives whose line numbers are greater than 32767. Nothing in Posix makes this invalid. That is why Autoconf stopped issuing #line directives.

[Macro] AC_PROG_CC ([compiler-search-list])

Determine a C compiler to use. If CC is not already set in the environment, check for gcc and cc, then for other C compilers. Set output variable CC to the name of the compiler found. This macro may, however, be invoked with an optional first argument which, if specified, must be a blank-separated list of C compilers to search for. This just gives the user an opportunity to specify an alternative search list for the C compiler. For example, if you didn’t like the default order, then you could invoke AC_PROG_CC like this:

     AC_PROG_CC([gcc cl cc])

If the C compiler does not handle function prototypes correctly by default, try to add an option to output variable CC to make it so. This macro tries various options that select standard-conformance modes on various systems. After calling this macro you can check whether the C compiler has been set to accept ANSI C89 (ISO C90); if not, the shell variable ac_cv_prog_cc_c89 is set to ‘no’. See

also AC_C_PROTOTYPES below.

If using the GNU C compiler, set shell variable GCC to ‘yes’. If output variable CFLAGS was not already set, set it to ‘-g -O2’ for the GNU C compiler (‘-O2’ on systems where GCC does not accept ‘-g’), or ‘-g’ for other compilers. If your package does not like this default, then it is acceptable to insert the line ‘: ${CFLAGS=""}’ after AC_INIT and before AC_PROG_CC to select an empty default instead. Many Autoconf macros use a compiler, and thus call ‘AC_REQUIRE([AC_PROG_CC])’ to ensure that the compiler has been determined before the body of the outermost AC_DEFUN macro. Although AC_PROG_CC is safe to directly expand multiple times, it performs certain checks (such as the proper value of EXEEXT) only on the first invocation. Therefore, care must be used when invoking this macro from within another macro rather than at the top level (see Section 20.8 [Expanded Before Required], page 334).

[Macro] AC_PROG_CC_C_O

If the C compiler does not accept the ‘-c’ and ‘-o’ options simultaneously, define NO_MINUS_C_MINUS_O. This macro actually tests both the compiler found by AC_ PROG_CC, and, if different, the first cc in the path. The test fails if one fails. This macro was created for GNU Make to choose the default C compilation rule.

For the compiler compiler, this macro caches its result in the ac_cv_prog_cc_

compiler_c_o variable.
[Macro] AC_PROG_CPP

Set output variable CPP to a command that runs the C preprocessor. If ‘$CC -E’ doesn’t work, ‘/lib/cpp’ is used. It is only portable to run CPP on files with a ‘.c’ extension. Some preprocessors don’t indicate missing include files by the error status. For such preprocessors an internal variable is set that causes other macros to check the standard error from the preprocessor and consider the test failed if any warnings have been reported. For most preprocessors, though, warnings do not cause include-file tests to fail unless AC_PROG_CPP_WERROR is also specified.

[Macro] AC_PROG_CPP_WERROR

This acts like AC_PROG_CPP, except it treats warnings from the preprocessor as errors even if the preprocessor exit status indicates success. This is useful for avoiding headers that generate mandatory warnings, such as deprecation notices.

The following macros check for C compiler or machine architecture features. To check for characteristics not listed here, use AC_COMPILE_IFELSE (see Section 6.4 [Running the Compiler], page 106) or AC_RUN_IFELSE (see Section 6.6 [Runtime], page 107).

[Macro] AC_PROG_CC_STDC

If the C compiler cannot compile ISO Standard C (currently C99), try to add an option to output variable CC to make it work. If the compiler does not support C99, fall back to supporting ANSI C89 (ISO C90). After calling this macro you can check whether the C compiler has been set to accept Standard C; if not, the shell variable ac_cv_prog_cc_stdc is set to ‘no’.

[Macro] AC_PROG_CC_C89

If the C compiler is not in ANSI C89 (ISO C90) mode by default, try to add an option to output variable CC to make it so. This macro tries various options that select ANSI C89 on some system or another, preferring extended functionality modes over strict conformance modes. It considers the compiler to be in ANSI C89 mode if it handles function prototypes correctly. After calling this macro you can check whether the C compiler has been set to accept ANSI C89; if not, the shell variable ac_cv_prog_cc_c89 is set to ‘no’. This macro is called automatically by AC_PROG_CC.

[Macro] AC_PROG_CC_C99

If the C compiler is not in C99 mode by default, try to add an option to output variable CC to make it so. This macro tries various options that select C99 on some system or another, preferring extended functionality modes over strict conformance modes. It considers the compiler to be in C99 mode if it handles _Bool, // comments, flexible array members, inline, signed and unsigned long long int, mixed code and declarations, named initialization of structs, restrict, va_copy, varargs macros, variable declarations in for loops, and variable length arrays. After calling this macro you can check whether the C compiler has been set to accept C99; if not, the shell variable ac_cv_prog_cc_c99 is set to ‘no’.

[Macro] AC_C_BACKSLASH_A

Define ‘HAVE_C_BACKSLASH_A’ to 1 if the C compiler understands ‘\a’. This macro is obsolescent, as current C compilers understand ‘\a’. New programs need not use this macro.

[Macro] AC_C_BIGENDIAN ([action-if-true], [action-if-false],
   [action-if-unknown], [action-if-universal])

If words are stored with the most significant byte first (like Motorola and SPARC CPUs), execute action-if-true. If words are stored with the least significant byte first (like Intel and VAX CPUs), execute action-if-false. This macro runs a test-case if endianness cannot be determined from the system header files. When cross-compiling, the test-case is not run but grep’ed for some magic values. action-if-unknown is executed if the latter case fails to determine the byte sex of the host system. In some cases a single run of a compiler can generate code for multiple architectures. This can happen, for example, when generating Mac OS X universal binary files, which work on both PowerPC and Intel architectures. In this case, the different variants might be for different architectures whose endiannesses differ. If configure detects this, it executes action-if-universal instead of action-if-unknown. The default for action-if-true is to define ‘WORDS_BIGENDIAN’. The default for actionif-false is to do nothing. The default for action-if-unknown is to abort configure and tell the installer how to bypass this test. And finally, the default for action-if-universal is to ensure that ‘WORDS_BIGENDIAN’ is defined if and only if a universal build is detected and the current code is big-endian; this default works only if autoheader is used (see Section 4.9.2 [autoheader Invocation], page 37). If you use this macro without specifying action-if-universal, you should also use AC_ CONFIG_HEADERS; otherwise ‘WORDS_BIGENDIAN’ may be set incorrectly for Mac OS X universal binary files.

[Macro] AC_C_CONST

If the C compiler does not fully support the const keyword, define const to be empty. Some C compilers that do not define __STDC__do support const; some compilers that define __STDC__do not completely support const. Programs can simply use const as if every C compiler supported it; for those that don’t, the makefile or configuration header file defines it as empty. Occasionally installers use a C++ compiler to compile C code, typically because they lack a C compiler. This causes problems with const, because C and C++ treat const differently. For example:

     const int foo;

is valid in C but not in C++. These differences unfortunately cannot be papered over by defining const to be empty. If autoconf detects this situation, it leaves const alone, as this generally yields better results in practice. However, using a C++ compiler to compile C code is not recommended or supported, and installers who run into trouble in this area should get a C compiler like GCC to compile their C code.

This macro caches its result in the ac_cv_c_const variable. This macro is obsolescent, as current C compilers support const. New programs need not use this macro.

[Macro] AC_C_RESTRICT

If the C compiler recognizes a variant spelling for the restrict keyword (__restrict, __restrict__, or _Restrict), then define restrict to that; this is more likely to do the right thing with compilers that support language variants where plain restrict is not a keyword. Otherwise, if the C compiler recognizes the restrict keyword, don’t do anything. Otherwise, define restrict to be empty. Thus, programs may simply use restrict as if every C compiler supported it; for those that do not, the makefile or configuration header defines it away. Although support in C++ for the restrict keyword is not required, several C++ compilers do accept the keyword. This macro works for them, too. This macro caches ‘no’ in the ac_cv_c_restrict variable if restrict is not supported, and a supported spelling otherwise.

[Macro] AC_C_VOLATILE

If the C compiler does not understand the keyword volatile, define volatile to be empty. Programs can simply use volatile as if every C compiler supported it; for those that do not, the makefile or configuration header defines it as empty. If the correctness of your program depends on the semantics of volatile, simply defining it to be empty does, in a sense, break your code. However, given that the compiler does not support volatile, you are at its mercy anyway. At least your program compiles, when it wouldn’t before. See Section 13.6 [Volatile Objects], page 263, for more about volatile. In general, the volatile keyword is a standard C feature, so you might expect that volatile is available only when __STDC__is defined. However, Ultrix 4.3’s native compiler does support volatile, but does not define __STDC__. This macro is obsolescent, as current C compilers support volatile. New programs need not use this macro.

[Macro] AC_C_INLINE

If the C compiler supports the keyword inline, do nothing. Otherwise define inline to __inline__or __inline if it accepts one of those, otherwise define inline to be empty.

[Macro] AC_C_CHAR_UNSIGNED

If the C type char is unsigned, define __CHAR_UNSIGNED__, unless the C compiler predefines it. These days, using this macro is not necessary. The same information can be determined by this portable alternative, thus avoiding the use of preprocessor macros in the namespace reserved for the implementation.

     #include <limits.h>
     #if CHAR_MIN == 0
     # define CHAR_UNSIGNED 1
     #endif

[Macro] AC_C_STRINGIZE

If the C preprocessor supports the stringizing operator, define HAVE_STRINGIZE. The stringizing operator is ‘#’ and is found in macros such as this:

     #define x(y) #y

This macro is obsolescent, as current C compilers support the stringizing operator. New programs need not use this macro.

[Macro] AC_C_FLEXIBLE_ARRAY_MEMBER

If the C compiler supports flexible array members, define FLEXIBLE_ARRAY_MEMBER to nothing; otherwise define it to 1. That way, a declaration like this:

     struct s
       {
         size_t n_vals;
         double val[FLEXIBLE_ARRAY_MEMBER];
       };

will let applications use the “struct hack” even with compilers that do not support flexible array members. To allocate and use such an object, you can use code like this:

     size_t i;
     size_t n = compute_value_count ();
     struct s *p =
        malloc (offsetof (struct s, val)
                + n * sizeof (double));
     p->n_vals = n;
     for (i = 0; i < n; i++)
       p->val[i] = compute_value (i);
[Macro] AC_C_VARARRAYS

If the C compiler supports variable-length arrays, define HAVE_C_VARARRAYS. A variable-length array is an array of automatic storage duration whose length is determined at run time, when the array is declared.

[Macro] AC_C_TYPEOF

If the C compiler supports GCC’s typeof syntax either directly or through a different spelling of the keyword (e.g., __typeof__), define HAVE_TYPEOF. If the support is available only through a different spelling, define typeof to that spelling.

[Macro] AC_C_PROTOTYPES

If function prototypes are understood by the compiler (as determined by AC_PROG_

CC), define PROTOTYPES and __PROTOTYPES. Defining __PROTOTYPES is for the benefit

of header files that cannot use macros that infringe on user name space. This macro is obsolescent, as current C compilers support prototypes. New programs need not use this macro.

[Macro] AC_PROG_GCC_TRADITIONAL

Add ‘-traditional’ to output variable CC if using the GNU C compiler and ioctl does not work properly without ‘-traditional’. That usually happens when the fixed header files have not been installed on an old system.

This macro is obsolescent, since current versions of the GNU C compiler fix the header files automatically when installed.

5.10.4 C++ Compiler Characteristics

[Macro] AC_PROG_CXX ([compiler-search-list])

Determine a C++ compiler to use. Check whether the environment variable CXX or CCC (in that order) is set; if so, then set output variable CXX to its value. Otherwise, if the macro is invoked without an argument, then search for a C++ compiler under the likely names (first g++ and c++ then other names). If none of those checks succeed, then as a last resort set CXX to g++. This macro may, however, be invoked with an optional first argument which, if specified, must be a blank-separated list of C++ compilers to search for. This just gives the user an opportunity to specify an alternative search list for the C++ compiler. For example, if you didn’t like the default order, then you could invoke AC_PROG_CXX like this:

     AC_PROG_CXX([gcc cl KCC CC cxx cc++ xlC aCC c++ g++])

If using the GNU C++ compiler, set shell variable GXX to ‘yes’. If output variable CXXFLAGS was not already set, set it to ‘-g -O2’ for the GNU C++ compiler (‘-O2’ on systems where G++ does not accept ‘-g’), or ‘-g’ for other compilers. If your package does not like this default, then it is acceptable to insert the line ‘: ${CXXFLAGS=""}’ after AC_INIT and before AC_PROG_CXX to select an empty default instead.

[Macro] AC_PROG_CXXCPP

Set output variable CXXCPP to a command that runs the C++ preprocessor. If ‘$CXX -E’ doesn’t work, ‘/lib/cpp’ is used. It is portable to run CXXCPP only on files with a ‘.c’, ‘.C’, ‘.cc’, or ‘.cpp’ extension. Some preprocessors don’t indicate missing include files by the error status. For such preprocessors an internal variable is set that causes other macros to check the standard error from the preprocessor and consider the test failed if any warnings have been reported. However, it is not known whether such broken preprocessors exist for C++.

[Macro] AC_PROG_CXX_C_O

Test whether the C++ compiler accepts the options ‘-c’ and ‘-o’ simultaneously, and

define CXX_NO_MINUS_C_MINUS_O, if it does not.

5.10.5 Objective C Compiler Characteristics

[Macro] AC_PROG_OBJC ([compiler-search-list])

Determine an Objective C compiler to use. If OBJC is not already set in the environment, check for Objective C compilers. Set output variable OBJC to the name of the compiler found. This macro may, however, be invoked with an optional first argument which, if specified, must be a blank-separated list of Objective C compilers to search for. This just gives the user an opportunity to specify an alternative search list for the Objective C compiler. For example, if you didn’t like the default order, then you could invoke

AC_PROG_OBJC like this:

     AC_PROG_OBJC([gcc objcc objc])

If using the GNU Objective C compiler, set shell variable GOBJC to ‘yes’. If output variable OBJCFLAGS was not already set, set it to ‘-g -O2’ for the GNU Objective C compiler (‘-O2’ on systems where gcc does not accept ‘-g’), or ‘-g’ for other compilers.

[Macro] AC_PROG_OBJCPP

Set output variable OBJCPP to a command that runs the Objective C preprocessor. If ‘$OBJC -E’ doesn’t work, ‘/lib/cpp’ is used.

5.10.6 Objective C++ Compiler Characteristics

[Macro] AC_PROG_OBJCXX ([compiler-search-list])

Determine an Objective C++ compiler to use. If OBJCXX is not already set in the environment, check for Objective C++ compilers. Set output variable OBJCXX to the name of the compiler found. This macro may, however, be invoked with an optional first argument which, if specified, must be a blank-separated list of Objective C++ compilers to search for. This just gives the user an opportunity to specify an alternative search list for the Objective C++ compiler. For example, if you didn’t like the default order, then you could invoke AC_PROG_OBJCXX like this:

     AC_PROG_OBJCXX([gcc g++ objcc++ objcxx])

If using the GNU Objective C++ compiler, set shell variable GOBJCXX to ‘yes’. If output variable OBJCXXFLAGS was not already set, set it to ‘-g -O2’ for the GNU Objective C++ compiler (‘-O2’ on systems where gcc does not accept ‘-g’), or ‘-g’ for other compilers.

[Macro] AC_PROG_OBJCXXCPP

Set output variable OBJCXXCPP to a command that runs the Objective C++ preprocessor. If ‘$OBJCXX -E’ doesn’t work, ‘/lib/cpp’ is used.

5.10.7 Erlang Compiler and Interpreter Characteristics

Autoconf defines the following macros for determining paths to the essential Erlang/OTP programs:

[Macro] AC_ERLANG_PATH_ERLC ([value-if-not-found], [path = ‘$PATH’])

Determine an Erlang compiler to use. If ERLC is not already set in the environment, check for erlc. Set output variable ERLC to the complete path of the compiler command found. In addition, if ERLCFLAGS is not set in the environment, set it to an empty value. The two optional arguments have the same meaning as the two last arguments of macro AC_PROG_PATH for looking for the erlc program. For example, to look for

erlc only in the ‘/usr/lib/erlang/bin’ directory:
     AC_ERLANG_PATH_ERLC([not found], [/usr/lib/erlang/bin])
[Macro] AC_ERLANG_NEED_ERLC ([path = ‘$PATH’])

A simplified variant of the AC_ERLANG_PATH_ERLC macro, that prints an error message and exits the configure script if the erlc program is not found.

[Macro] AC_ERLANG_PATH_ERL ([value-if-not-found], [path = ‘$PATH’])

Determine an Erlang interpreter to use. If ERL is not already set in the environment, check for erl. Set output variable ERL to the complete path of the interpreter command found. The two optional arguments have the same meaning as the two last arguments of macro AC_PROG_PATH for looking for the erl program. For example, to look for erl only in the ‘/usr/lib/erlang/bin’ directory:

     AC_ERLANG_PATH_ERL([not found], [/usr/lib/erlang/bin])
[Macro] AC_ERLANG_NEED_ERL ([path = ‘$PATH’])

A simplified variant of the AC_ERLANG_PATH_ERL macro, that prints an error message and exits the configure script if the erl program is not found.

5.10.8 Fortran Compiler Characteristics

The Autoconf Fortran support is divided into two categories: legacy Fortran 77 macros (F77), and modern Fortran macros (FC). The former are intended for traditional Fortran 77 code, and have output variables like F77, FFLAGS, and FLIBS. The latter are for newer programs that can (or must) compile under the newer Fortran standards, and have output variables like FC, FCFLAGS, and FCLIBS.

Except for the macros AC_FC_SRCEXT, AC_FC_FREEFORM, and AC_FC_FIXEDFORM (see be-

low), the FC and F77 macros behave almost identically, and so they are documented together in this section.

[Macro] AC_PROG_F77 ([compiler-search-list])

Determine a Fortran 77 compiler to use. If F77 is not already set in the environment, then check for g77 and f77, and then some other names. Set the output variable F77 to the name of the compiler found. This macro may, however, be invoked with an optional first argument which, if specified, must be a blank-separated list of Fortran 77 compilers to search for. This just gives the user an opportunity to specify an alternative search list for the Fortran 77 compiler. For example, if you didn’t like the default order, then you could invoke

AC_PROG_F77 like this:
     AC_PROG_F77([fl32 f77 fort77 xlf g77 f90 xlf90])

If using g77 (the GNU Fortran 77 compiler), then set the shell variable G77 to ‘yes’. If the output variable FFLAGS was not already set in the environment, then set it to ‘-g -02’ for g77 (or ‘-O2’ where g77 does not accept ‘-g’). Otherwise, set FFLAGS to ‘-g’ for all other Fortran 77 compilers.

[Macro] AC_PROG_FC ([compiler-search-list], [dialect])

Determine a Fortran compiler to use. If FC is not already set in the environment, then dialect is a hint to indicate what Fortran dialect to search for; the default is to search for the newest available dialect. Set the output variable FC to the name of the compiler found. By default, newer dialects are preferred over older dialects, but if dialect is specified then older dialects are preferred starting with the specified dialect. dialect can currently be one of Fortran 77, Fortran 90, or Fortran 95. However, this is only a hint of which compiler name to prefer (e.g., f90 or f95), and no attempt is made to guarantee that a particular language standard is actually supported. Thus, it is preferable that you avoid the dialect option, and use AC PROG FC only for code compatible with the latest Fortran standard.

This macro may, alternatively, be invoked with an optional first argument which, if specified, must be a blank-separated list of Fortran compilers to search for, just as in

AC_PROG_F77.

If the output variable FCFLAGS was not already set in the environment, then set it to ‘-g -02’ for GNU g77 (or ‘-O2’ where g77 does not accept ‘-g’). Otherwise, set FCFLAGS to ‘-g’ for all other Fortran compilers.

[Macro] AC_PROG_F77_C_O
[Macro] AC_PROG_FC_C_O

Test whether the Fortran compiler accepts the options ‘-c’ and ‘-o’ simultaneously,

and define F77_NO_MINUS_C_MINUS_O or FC_NO_MINUS_C_MINUS_O, respectively, if it

does not.

The following macros check for Fortran compiler characteristics. To check for characteristics not listed here, use AC_COMPILE_IFELSE (see Section 6.4 [Running the Compiler], page 106) or AC_RUN_IFELSE (see Section 6.6 [Runtime], page 107), making sure to first set the current language to Fortran 77 or Fortran via AC_LANG([Fortran 77]) or AC_LANG(Fortran) (see Section 6.1 [Language Choice], page 99).

[Macro] AC_F77_LIBRARY_LDFLAGS
[Macro] AC_FC_LIBRARY_LDFLAGS

Determine the linker flags (e.g., ‘-L’ and ‘-l’) for the Fortran intrinsic and runtime libraries that are required to successfully link a Fortran program or shared library. The output variable FLIBS or FCLIBS is set to these flags (which should be included after LIBS when linking).

This macro is intended to be used in those situations when it is necessary to mix, e.g., C++ and Fortran source code in a single program or shared library (see Section “Mixing Fortran 77 With C and C++” in GNU Automake).

For example, if object files from a C++ and Fortran compiler must be linked together, then the C++ compiler/linker must be used for linking (since special C++-ish things need to happen at link time like calling global constructors, instantiating templates, enabling exception support, etc.).

However, the Fortran intrinsic and runtime libraries must be linked in as well, but the C++ compiler/linker doesn’t know by default how to add these Fortran 77 libraries. Hence, this macro was created to determine these Fortran libraries.

The macros AC_F77_DUMMY_MAIN and AC_FC_DUMMY_MAIN or AC_F77_MAIN and AC_FC_

MAIN are probably also necessary to link C/C++ with Fortran; see below. Further, it is highly recommended that you use AC_CONFIG_HEADERS (see Section 4.9 [Configuration Headers], page 35) because the complex defines that the function wrapper macros create may not work with C/C++ compiler drivers.

[Macro] AC_F77_DUMMY_MAIN ([action-if-found], [action-if-not-found])
[Macro] AC_FC_DUMMY_MAIN ([action-if-found], [action-if-not-found])

With many compilers, the Fortran libraries detected by AC_F77_LIBRARY_LDFLAGS or AC_FC_LIBRARY_LDFLAGS provide their own main entry function that initializes things like Fortran I/O, and which then calls a user-provided entry function named (say)

MAIN__to run the user’s program. The AC_F77_DUMMY_MAIN and AC_FC_DUMMY_MAIN

or AC_F77_MAIN and AC_FC_MAIN macros figure out how to deal with this interaction. When using Fortran for purely numerical functions (no I/O, etc.) often one prefers to provide one’s own main and skip the Fortran library initializations. In this case, however, one may still need to provide a dummy MAIN__routine in order to prevent linking errors on some systems. AC_F77_DUMMY_MAIN or AC_FC_DUMMY_MAIN detects whether any such routine is required for linking, and what its name is; the shell vari-

able F77_DUMMY_MAIN or FC_DUMMY_MAIN holds this name, unknown when no solution

was found, and none when no such dummy main is needed. By default, action-if-found defines F77_DUMMY_MAIN or FC_DUMMY_MAIN to the name of this routine (e.g., MAIN__) if it is required. action-if-not-found defaults to exiting with an error. In order to link with Fortran routines, the user’s C/C++ program should then include the following code to define the dummy main if it is needed:

     #ifdef F77_DUMMY_MAIN
     #  ifdef __cplusplus
          extern "C"
     #  endif
        int F77_DUMMY_MAIN () { return 1; }
     #endif

(Replace F77 with FC for Fortran instead of Fortran 77.) Note that this macro is called automatically from AC_F77_WRAPPERS or AC_FC_ WRAPPERS; there is generally no need to call it explicitly unless one wants to change the default actions.

[Macro] AC_F77_MAIN
[Macro] AC_FC_MAIN

As discussed above, many Fortran libraries allow you to provide an entry point called (say) MAIN__instead of the usual main, which is then called by a main function in the Fortran libraries that initializes things like Fortran I/O. The AC_F77_MAIN and AC_FC_MAIN macros detect whether it is possible to utilize such an alternate main function, and defines F77_MAIN and FC_MAIN to the name of the function. (If no alternate main function name is found, F77_MAIN and FC_MAIN are simply defined to

main.)

Thus, when calling Fortran routines from C that perform things like I/O, one should use this macro and declare the "main" function like so:

     #ifdef __cplusplus
       extern "C"
     #endif
     int F77_MAIN (int argc, char *argv[]);

(Again, replace F77 with FC for Fortran instead of Fortran 77.)

[Macro] AC_F77_WRAPPERS
[Macro] AC_FC_WRAPPERS
Defines C macros F77_FUNC (name, NAME), FC_FUNC (name, NAME), F77_FUNC_

(name, NAME), and FC_FUNC_(name, NAME) to properly mangle the names of C/C++ identifiers, and identifiers with underscores, respectively, so that they match the name-mangling scheme used by the Fortran compiler.

Fortran is case-insensitive, and in order to achieve this the Fortran compiler converts all identifiers into a canonical case and format. To call a Fortran subroutine from C or to write a C function that is callable from Fortran, the C program must explicitly use identifiers in the format expected by the Fortran compiler. In order to do this, one simply wraps all C identifiers in one of the macros provided by AC_F77_WRAPPERS or AC_FC_WRAPPERS. For example, suppose you have the following Fortran 77 subroutine:

           subroutine foobar (x, y)
           double precision x, y
           y = 3.14159 * x
           return
           end

You would then declare its prototype in C or C++ as:

     #define FOOBAR_F77 F77_FUNC (foobar, FOOBAR)
     #ifdef __cplusplus
     extern "C"  /* prevent C++ name mangling */
     #endif
     void FOOBAR_F77 (double *x, double *y);

Note that we pass both the lowercase and uppercase versions of the function name to F77_FUNC so that it can select the right one. Note also that all parameters to Fortran 77 routines are passed as pointers (see Section “Mixing Fortran 77 With C and C++” in GNU Automake).

(Replace F77 with FC for Fortran instead of Fortran 77.)

Although Autoconf tries to be intelligent about detecting the name-mangling scheme of the Fortran compiler, there may be Fortran compilers that it doesn’t support yet. In this case, the above code generates a compile-time error, but some other behavior (e.g., disabling Fortran-related features) can be induced by checking whether F77_

FUNC or FC_FUNC is defined.

Now, to call that routine from a C program, we would do something like:

     {
         double x = 2.7183, y;
         FOOBAR_F77 (&x, &y);
     }

If the Fortran identifier contains an underscore (e.g., foo_bar), you should use F77_ FUNC_or FC_FUNC_instead of F77_FUNC or FC_FUNC (with the same arguments). This is because some Fortran compilers mangle names differently if they contain an underscore.

[Macro] AC_F77_FUNC (name, [shellvar])
[Macro] AC_FC_FUNC (name, [shellvar])

Given an identifier name, set the shell variable shellvar to hold the mangled version name according to the rules of the Fortran linker (see also AC_F77_WRAPPERS or AC_ FC_WRAPPERS). shellvar is optional; if it is not supplied, the shell variable is simply name. The purpose of this macro is to give the caller a way to access the namemangling information other than through the C preprocessor as above, for example, to call Fortran routines from some language other than C/C++.

[Macro] AC_FC_SRCEXT (ext, [action-if-success], [action-if-failure])

By default, the FC macros perform their tests using a ‘.f’ extension for source-code files. Some compilers, however, only enable newer language features for appropriately named files, e.g., Fortran 90 features only for ‘.f90’ files. On the other hand, some other compilers expect all source files to end in ‘.f’ and require special flags to support other file name extensions. The AC_FC_SRCEXT macro deals with both of these issues.

The AC_FC_SRCEXT tries to get the FC compiler to accept files ending with the extension .ext (i.e., ext does not contain the dot). If any special compiler flags are needed for this, it stores them in the output variable FCFLAGS_ext. This extension and these flags are then used for all subsequent FC tests (until AC_FC_SRCEXT is called again).

For example, you would use AC_FC_SRCEXT(f90) to employ the ‘.f90’ extension in future tests, and it would set the FCFLAGS_f90 output variable with any extra flags that are needed to compile such files.

The FCFLAGS_ext can not be simply absorbed into FCFLAGS, for two reasons based on the limitations of some compilers. First, only one FCFLAGS_ext can be used at a time, so files with different extensions must be compiled separately. Second, FCFLAGS_ ext must appear immediately before the source-code file name when compiling. So, continuing the example above, you might compile a ‘foo.f90’ file in your makefile with the command:

     foo.o: foo.f90
          $(FC) -c $(FCFLAGS) $(FCFLAGS_f90) '$(srcdir)/foo.f90'

If AC_FC_SRCEXT succeeds in compiling files with the ext extension, it calls actionif-success (defaults to nothing). If it fails, and cannot find a way to make the FC compiler accept such files, it calls action-if-failure (defaults to exiting with an error message).

[Macro] AC_FC_FREEFORM ([action-if-success], [action-if-failure])

The AC_FC_FREEFORM tries to ensure that the Fortran compiler ($FC) allows freeformat source code (as opposed to the older fixed-format style from Fortran 77). If necessary, it may add some additional flags to FCFLAGS.

This macro is most important if you are using the default ‘.f’ extension, since many compilers interpret this extension as indicating fixed-format source unless an additional flag is supplied. If you specify a different extension with AC_FC_SRCEXT, such as ‘.f90’, then AC_FC_FREEFORM ordinarily succeeds without modifying FCFLAGS. For extensions which the compiler does not know about, the flag set by the AC_FC_SRCEXT macro might let the compiler assume Fortran 77 by default, however.

If AC_FC_FREEFORM succeeds in compiling free-form source, it calls action-if-success (defaults to nothing). If it fails, it calls action-if-failure (defaults to exiting with an error message).

[Macro] AC_FC_FIXEDFORM ([action-if-success], [action-if-failure])

The AC_FC_FIXEDFORM tries to ensure that the Fortran compiler ($FC) allows the old fixed-format source code (as opposed to free-format style). If necessary, it may add some additional flags to FCFLAGS. This macro is needed for some compilers alias names like xlf95 which assume freeform source code by default, and in case you want to use fixed-form source with an extension like ‘.f90’ which many compilers interpret as free-form by default. If you specify a different extension with AC_FC_SRCEXT, such as ‘.f’, then AC_FC_FIXEDFORM ordinarily succeeds without modifying FCFLAGS. If AC_FC_FIXEDFORM succeeds in compiling fixed-form source, it calls action-if-success (defaults to nothing). If it fails, it calls action-if-failure (defaults to exiting with an error message).

5.11 System Services

The following macros check for operating system services or capabilities.

[Macro] AC_PATH_X

Try to locate the X Window System include files and libraries. If the user gave the command line options ‘--x-includes=dir’ and ‘--x-libraries=dir’, use those directories. If either or both were not given, get the missing values by running xmkmf (or an executable pointed to by the XMKMF environment variable) on a trivial ‘Imakefile’ and examining the makefile that it produces. Setting XMKMF to ‘false’ disables this method. If this method fails to find the X Window System, configure looks for the files in several directories where they often reside. If either method is successful, set the shell variables x_includes and x_libraries to their locations, unless they are in directories the compiler searches by default. If both methods fail, or the user gave the command line option ‘--without-x’, set the shell variable no_x to ‘yes’; otherwise set it to the empty string.

[Macro] AC_PATH_XTRA

An enhanced version of AC_PATH_X. It adds the C compiler flags that X needs to output variable X_CFLAGS, and the X linker flags to X_LIBS. Define X_DISPLAY_ MISSING if X is not available. This macro also checks for special libraries that some systems need in order to compile X programs. It adds any that the system needs to output variable X_EXTRA_LIBS. And it checks for special X11R6 libraries that need to be linked with before ‘-lX11’, and adds any found to the output variable X_PRE_LIBS.

[Macro] AC_SYS_INTERPRETER

Check whether the system supports starting scripts with a line of the form ‘#!/bin/sh’ to select the interpreter to use for the script. After running this macro, shell code in ‘configure.ac’ can check the shell variable interpval; it is set to ‘yes’ if the system supports ‘#!’, ‘no’ if not.

[Macro] AC_SYS_LARGEFILE

Arrange for 64-bit file offsets, known as large-file support (http://www.unix-systems

.org/version2/whatsnew/lfs20mar.html). On some hosts, one must use special

compiler options to build programs that can access large files. Append any such options to the output variable CC. Define _FILE_OFFSET_BITS and _LARGE_FILES if necessary. Large-file support can be disabled by configuring with the ‘--disable-largefile’ option. If you use this macro, check that your program works even when off_t is wider than long int, since this is common when large-file support is enabled. For example, it is not correct to print an arbitrary off_t value X with printf ("%ld", (long int) X). The LFS introduced the fseeko and ftello functions to replace their C counterparts fseek and ftell that do not use off_t. Take care to use AC_FUNC_FSEEKO to make their prototypes available when using them and large-file support is enabled.

[Macro] AC_SYS_LONG_FILE_NAMES

If the system supports file names longer than 14 characters, define HAVE_LONG_FILE_

NAMES.
[Macro] AC_SYS_POSIX_TERMIOS

Check to see if the Posix termios headers and functions are available on the system. If so, set the shell variable ac_cv_sys_posix_termios to ‘yes’. If not, set the variable to ‘no’.

5.12 Posix Variants

The following macro makes it possible to use features of Posix that are extensions to C, as well as platform extensions not defined by Posix.

[Macro] AC_USE_SYSTEM_EXTENSIONS

This macro was introduced in Autoconf 2.60. If possible, enable extensions to C or Posix on hosts that normally disable the extensions, typically due to standardsconformance namespace issues. This should be called before any macros that run the C compiler. The following preprocessor macros are defined where appropriate:

_GNU_SOURCE

Enable extensions on GNU/Linux.

__EXTENSIONS__

Enable general extensions on Solaris.

_POSIX_PTHREAD_SEMANTICS

Enable threading extensions on Solaris.

_TANDEM_SOURCE

Enable extensions for the HP NonStop platform.

_ALL_SOURCE

Enable extensions for AIX 3, and for Interix.

_POSIX_SOURCE

Enable Posix functions for Minix.

_POSIX_1_SOURCE

Enable additional Posix functions for Minix.

_MINIX Identify Minix platform. This particular preprocessor macro is obsolescent, and may be removed in a future release of Autoconf.

5.13 Erlang Libraries

The following macros check for an installation of Erlang/OTP, and for the presence of certain Erlang libraries. All those macros require the configuration of an Erlang interpreter and an Erlang compiler (see Section 5.10.7 [Erlang Compiler and Interpreter], page 88).

[Macro] AC_ERLANG_SUBST_ERTS_VER

Set the output variable ERLANG_ERTS_VER to the version of the Erlang runtime system (as returned by Erlang’s erlang:system_info(version) function). The result of this test is cached if caching is enabled when running configure. The ERLANG_ERTS_VER variable is not intended to be used for testing for features of specific ERTS versions, but to be used for substituting the ERTS version in Erlang/OTP release resource files (.rel files), as shown below.

[Macro] AC_ERLANG_SUBST_ROOT_DIR

Set the output variable ERLANG_ROOT_DIR to the path to the base directory in which Erlang/OTP is installed (as returned by Erlang’s code:root_dir/0 function). The result of this test is cached if caching is enabled when running configure.

[Macro] AC_ERLANG_SUBST_LIB_DIR

Set the output variable ERLANG_LIB_DIR to the path of the library directory of Erlang/OTP (as returned by Erlang’s code:lib_dir/0 function), which subdirectories each contain an installed Erlang/OTP library. The result of this test is cached if caching is enabled when running configure.

[Macro] AC_ERLANG_CHECK_LIB (library, [action-if-found],
   [action-if-not-found])

Test whether the Erlang/OTP library library is installed by calling Erlang’s code:lib_dir/1 function. The result of this test is cached if caching is enabled when running configure. action-if-found is a list of shell commands to run if the library is installed; action-if-not-found is a list of shell commands to run if it is not. Additionally, if the library is installed, the output variable ‘ERLANG_LIB_DIR_library’ is set to the path to the library installation directory, and the output variable ‘ERLANG_LIB_VER_library’ is set to the version number that is part of the subdirectory name, if it is in the standard form (library-version). If the directory name does not have a version part, ‘ERLANG_LIB_VER_library’ is set to the empty string. If the library is not installed, ‘ERLANG_LIB_DIR_library’ and ‘ERLANG_LIB_VER_library’ are set to "not found". For example, to check if library stdlib is installed:

     AC_ERLANG_CHECK_LIB([stdlib],
       [echo "stdlib version \"$ERLANG_LIB_VER_stdlib\""
        echo "is installed in \"$ERLANG_LIB_DIR_stdlib\""],
       [AC_MSG_ERROR([stdlib was not found!])])
The ‘ERLANG_LIB_VER_library’ variables (set by AC_ERLANG_CHECK_LIB) and the
ERLANG_ERTS_VER variable (set by AC_ERLANG_SUBST_ERTS_VER) are not intended to

be used for testing for features of specific versions of libraries or of the Erlang runtime system. Those variables are intended to be substituted in Erlang release resource files (.rel files). For instance, to generate a ‘example.rel’ file for an application depending on the stdlib library, ‘configure.ac’ could contain:

     AC_ERLANG_SUBST_ERTS_VER
     AC_ERLANG_CHECK_LIB([stdlib],
       [],
       [AC_MSG_ERROR([stdlib was not found!])])
     AC_CONFIG_FILES([example.rel])

The ‘example.rel.in’ file used to generate ‘example.rel’ should contain:

     {release,
         {"@PACKAGE@", "@VERSION@"},
         {erts, "@ERLANG_ERTS_VER@"},
         [{stdlib, "@ERLANG_LIB_VER_stdlib@"},
          {@PACKAGE@, "@VERSION@"}]}.

In addition to the above macros, which test installed Erlang libraries, the following macros determine the paths to the directories into which newly built Erlang libraries are to be installed:

[Macro] AC_ERLANG_SUBST_INSTALL_LIB_DIR

Set the ERLANG_INSTALL_LIB_DIR output variable to the directory into which every built Erlang library should be installed in a separate subdirectory. If this variable is not set in the environment when configure runs, its default value is

${libdir}/erlang/lib.
[Macro] AC_ERLANG_SUBST_INSTALL_LIB_SUBDIR (library, version)

Set the ‘ERLANG_INSTALL_LIB_DIR_library’ output variable to the directory into which the built Erlang library library version version should be installed. If this variable is not set in the environment when configure runs, its default value is

‘$ERLANG_INSTALL_LIB_DIR/library-version’, the value of the ERLANG_INSTALL_
LIB_DIR variable being set by the AC_ERLANG_SUBST_INSTALL_LIB_DIR macro.

6 Writing Tests

If the existing feature tests don’t do something you need, you have to write new ones. These macros are the building blocks. They provide ways for other macros to check whether various kinds of features are available and report the results.

This chapter contains some suggestions and some of the reasons why the existing tests are written the way they are. You can also learn a lot about how to write Autoconf tests by looking at the existing ones. If something goes wrong in one or more of the Autoconf tests, this information can help you understand the assumptions behind them, which might help you figure out how to best solve the problem.

These macros check the output of the compiler system of the current language (see Section 6.1 [Language Choice], page 99). They do not cache the results of their tests for future use (see Section 7.4 [Caching Results], page 115), because they don’t know enough about the information they are checking for to generate a cache variable name. They also do not print any messages, for the same reason. The checks for particular kinds of features call these macros and do cache their results and print messages about what they’re checking for.

When you write a feature test that could be applicable to more than one software package, the best thing to do is encapsulate it in a new macro. See Chapter 10 [Writing Autoconf Macros], page 175, for how to do that.

6.1 Language Choice

Autoconf-generated configure scripts check for the C compiler and its features by default. Packages that use other programming languages (maybe more than one, e.g., C and C++) need to test features of the compilers for the respective languages. The following macros determine which programming language is used in the subsequent tests in ‘configure.ac’.

[Macro] AC_LANG (language)

Do compilation tests using the compiler, preprocessor, and file extensions for the specified language. Supported languages are:

‘C’ Do compilation tests using CC and CPP and use extension ‘.c’ for test programs. Use compilation flags: CPPFLAGS with CPP, and both CPPFLAGS

               and CFLAGS with CC.

‘C++’ Do compilation tests using CXX and CXXCPP and use extension ‘.C’ for test programs. Use compilation flags: CPPFLAGS with CXXCPP, and both

               CPPFLAGS and CXXFLAGS with CXX.
     ‘Fortran 77’

Do compilation tests using F77 and use extension ‘.f’ for test programs. Use compilation flags: FFLAGS.

‘Fortran’ Do compilation tests using FC and use extension ‘.f’ (or whatever has been set by AC_FC_SRCEXT) for test programs. Use compilation flags:

               FCFLAGS.

‘Erlang’ Compile and execute tests using ERLC and ERL and use extension ‘.erl’ for test Erlang modules. Use compilation flags: ERLCFLAGS.

‘Objective C’

Do compilation tests using OBJC and OBJCPP and use extension ‘.m’ for test programs. Use compilation flags: CPPFLAGS with OBJCPP, and both

          CPPFLAGS and OBJCFLAGS with OBJC.
‘Objective C++’

Do compilation tests using OBJCXX and OBJCXXCPP and use extension ‘.mm’ for test programs. Use compilation flags: CPPFLAGS with OBJCXXCPP, and

          both CPPFLAGS and OBJCXXFLAGS with OBJCXX.
[Macro] AC_LANG_PUSH (language)

Remember the current language (as set by AC_LANG) on a stack, and then select the language. Use this macro and AC_LANG_POP in macros that need to temporarily switch to a particular language.

[Macro] AC_LANG_POP ([language])

Select the language that is saved on the top of the stack, as set by AC_LANG_PUSH, and remove it from the stack.

If given, language specifies the language we just quit. It is a good idea to specify it when it’s known (which should be the case…), since Autoconf detects inconsistencies.

     AC_LANG_PUSH([Fortran 77])
     # Perform some tests on Fortran 77.
     # ...
     AC_LANG_POP([Fortran 77])
[Macro] AC_LANG_ASSERT (language)

Check statically that the current language is language. You should use this in your language specific macros to avoid that they be called with an inappropriate language.

This macro runs only at autoconf time, and incurs no cost at configure time. Sadly enough and because Autoconf is a two layer language^{1}, the macros AC_LANG_PUSH and AC_LANG_POP cannot be “optimizing”, therefore as much as possible you ought to avoid using them to wrap your code, rather, require from the user to run the macro with a correct current language, and check it with AC_LANG_ASSERT. And anyway, that may help the user understand she is running a Fortran macro while expecting a result about her Fortran 77 compiler…

[Macro] AC_REQUIRE_CPP

Ensure that whichever preprocessor would currently be used for tests has been found. Calls AC_REQUIRE (see Section 10.4.1 [Prerequisite Macros], page 178) with an argument of either AC_PROG_CPP or AC_PROG_CXXCPP, depending on which language is current.

1 Because M4 is not aware of Sh code, especially conditionals, some optimizations that look nice statically

may produce incorrect results at runtime.

6.2 Writing Test Programs

Autoconf tests follow a common scheme: feed some program with some input, and most of the time, feed a compiler with some source file. This section is dedicated to these source samples.

6.2.1 Guidelines for Test Programs

The most important rule to follow when writing testing samples is:

Look for realism.

This motto means that testing samples must be written with the same strictness as real programs are written. In particular, you should avoid “shortcuts” and simplifications.

Don’t just play with the preprocessor if you want to prepare a compilation. For instance, using cpp to check whether a header is functional might let your configure accept a header which causes some compiler error. Do not hesitate to check a header with other headers included before, especially required headers.

Make sure the symbols you use are properly defined, i.e., refrain from simply declaring a function yourself instead of including the proper header.

Test programs should not write to standard output. They should exit with status 0 if the test succeeds, and with status 1 otherwise, so that success can be distinguished easily from a core dump or other failure; segmentation violations and other failures produce a nonzero exit status. Unless you arrange for exit to be declared, test programs should return, not exit, from main, because on many systems exit is not declared by default.

Test programs can use #if or #ifdef to check the values of preprocessor macros defined by tests that have already run. For example, if you call AC_HEADER_STDBOOL, then later on in ‘configure.ac’ you can have a test program that includes ‘stdbool.h’ conditionally:

     #ifdef HAVE_STDBOOL_H
     # include <stdbool.h>
     #endif
   Both #if HAVE_STDBOOL_H and #ifdef HAVE_STDBOOL_H will work with any standard C

compiler. Some developers prefer #if because it is easier to read, while others prefer #ifdef because it avoids diagnostics with picky compilers like GCC with the ‘-Wundef’ option.

If a test program needs to use or create a data file, give it a name that starts with ‘conftest’, such as ‘conftest.data’. The configure script cleans up by running ‘rm -f -r conftest*’ after running test programs and if the script is interrupted.

6.2.2 Test Functions

These days it’s safe to assume support for function prototypes (introduced in C89).

Functions that test programs declare should also be conditionalized for C++, which requires ‘extern "C"’ prototypes. Make sure to not include any header files containing clashing prototypes.

     #ifdef __cplusplus
     extern "C"
     #endif
     void *valloc (size_t);

If a test program calls a function with invalid parameters (just to see whether it exists), organize the program to ensure that it never invokes that function. You can do this by calling it in another function that is never invoked. You can’t do it by putting it after a call to exit, because GCC version 2 knows that exit never returns and optimizes out any code that follows it in the same block.

If you include any header files, be sure to call the functions relevant to them with the correct number of arguments, even if they are just 0, to avoid compilation errors due to prototypes. GCC version 2 has internal prototypes for several functions that it automatically inlines; for example, memcpy. To avoid errors when checking for them, either pass them the correct number of arguments or redeclare them with a different return type (such as char).

6.2.3 Generating Sources

Autoconf provides a set of macros that can be used to generate test source files. They are written to be language generic, i.e., they actually depend on the current language (see Section 6.1 [Language Choice], page 99) to “format” the output properly.

[Macro] AC_LANG_CONFTEST (source)

Save the source text in the current test source file: ‘conftest.extension’ where the extension depends on the current language. As of Autoconf 2.63b, the source file also contains the results of all of the AC_DEFINE performed so far. Note that the source is evaluated exactly once, like regular Autoconf macro arguments, and therefore (i) you may pass a macro invocation, (ii) if not, be sure to double quote if needed.

[Macro] AC_LANG_SOURCE (source)

Expands into the source, with the definition of all the AC_DEFINE performed so far.

For instance executing (observe the double quotation!):

AC_INIT([Hello], [1.0], [bug-hello@example.org], [],
        [http://www.example.org/])
AC_DEFINE([HELLO_WORLD], ["Hello, World\n"],
  [Greetings string.])
AC_LANG([C])
AC_LANG_CONFTEST(
   [AC_LANG_SOURCE([[const char hw[] = "Hello, World\n";]])])
gcc -E -dD -o - conftest.c

on a system with gcc installed, results in:

...
# 1 "conftest.c"
#define PACKAGE_NAME "Hello"
#define PACKAGE_TARNAME "hello"
#define PACKAGE_VERSION "1.0"
#define PACKAGE_STRING "Hello 1.0"
#define PACKAGE_BUGREPORT "bug-hello@example.org"
#define PACKAGE_URL "http://www.example.org/"
#define HELLO_WORLD "Hello, World\n"

const char hw[] = "Hello, World\n";

When the test language is Fortran or Erlang, the AC_DEFINE definitions are not automatically translated into constants in the source code by this macro.

[Macro] AC_LANG_PROGRAM (prologue, body)

Expands into a source file which consists of the prologue, and then body as body of the main function (e.g., main in C). Since it uses AC_LANG_SOURCE, the features of the latter are available.

For instance:

AC_INIT([Hello], [1.0], [bug-hello@example.org], [],
        [http://www.example.org/])
AC_DEFINE([HELLO_WORLD], ["Hello, World\n"],
  [Greetings string.])
AC_LANG_CONFTEST(
[AC_LANG_PROGRAM([[const char hw[] = "Hello, World\n";]],
                 [[fputs (hw, stdout);]])])
gcc -E -dD -o - conftest.c

on a system with gcc installed, results in:

...
# 1 "conftest.c"
#define PACKAGE_NAME "Hello"
#define PACKAGE_TARNAME "hello"
#define PACKAGE_VERSION "1.0"
#define PACKAGE_STRING "Hello 1.0"
#define PACKAGE_BUGREPORT "bug-hello@example.org"
#define PACKAGE_URL "http://www.example.org/"
#define HELLO_WORLD "Hello, World\n"
const char hw[] = "Hello, World\n";
int
main ()
{
fputs (hw, stdout);
  ;
  return 0;
}

In Erlang tests, the created source file is that of an Erlang module called conftest (‘conftest.erl’). This module defines and exports at least one start/0 function, which is called to perform the test. The prologue is optional code that is inserted between the module header and the start/0 function definition. body is the body of the start/0 function without the final period (see Section 6.6 [Runtime], page 107, about constraints on this function’s behavior).

For instance:

AC_INIT([Hello], [1.0], [bug-hello@example.org])
AC_LANG(Erlang)
AC_LANG_CONFTEST(
[AC_LANG_PROGRAM([[-define(HELLO_WORLD, "Hello, world!").]],
                 [[io:format("~s~n", [?HELLO_WORLD])]])])
cat conftest.erl

results in:

-module(conftest).
-export([start/0]).
-define(HELLO_WORLD, "Hello, world!").
start() ->
io:format("~s~n", [?HELLO_WORLD])
.
[Macro] AC_LANG_CALL (prologue, function)

Expands into a source file which consists of the prologue, and then a call to the function as body of the main function (e.g., main in C). Since it uses AC_LANG_ PROGRAM, the feature of the latter are available. This function will probably be replaced in the future by a version which would enable specifying the arguments. The use of this macro is not encouraged, as it violates strongly the typing system. This macro cannot be used for Erlang tests.

[Macro] AC_LANG_FUNC_LINK_TRY (function)

Expands into a source file which uses the function in the body of the main function (e.g., main in C). Since it uses AC_LANG_PROGRAM, the features of the latter are available. As AC_LANG_CALL, this macro is documented only for completeness. It is considered to be severely broken, and in the future will be removed in favor of actual function calls (with properly typed arguments). This macro cannot be used for Erlang tests.

6.3 Running the Preprocessor

Sometimes one might need to run the preprocessor on some source file. Usually it is a bad idea, as you typically need to compile your project, not merely run the preprocessor on it; therefore you certainly want to run the compiler, not the preprocessor. Resist the temptation of following the easiest path.

Nevertheless, if you need to run the preprocessor, then use AC_PREPROC_IFELSE. The macros described in this section cannot be used for tests in Erlang or Fortran, since

those languages require no preprocessor.

[Macro] AC_PREPROC_IFELSE (input, [action-if-true], [action-if-false])

Run the preprocessor of the current language (see Section 6.1 [Language Choice], page 99) on the input, run the shell commands action-if-true on success, action-if-false otherwise. The input can be made by AC_LANG_PROGRAM and friends.

This macro uses CPPFLAGS, but not CFLAGS, because ‘-g’, ‘-O’, etc. are not valid options to many C preprocessors.

It is customary to report unexpected failures with AC_MSG_FAILURE.

For instance:

AC_INIT([Hello], [1.0], [bug-hello@example.org])
AC_DEFINE([HELLO_WORLD], ["Hello, World\n"],
  [Greetings string.])
AC_PREPROC_IFELSE(
   [AC_LANG_PROGRAM([[const char hw[] = "Hello, World\n";]],
                    [[fputs (hw, stdout);]])],
   [AC_MSG_RESULT([OK])],
   [AC_MSG_FAILURE([unexpected preprocessor failure])])

results in:

checking for gcc... gcc
checking for C compiler default output file name... a.out
checking whether the C compiler works... yes
checking whether we are cross compiling... no
checking for suffix of executables...
checking for suffix of object files... o
checking whether we are using the GNU C compiler... yes
checking whether gcc accepts -g... yes
checking for gcc option to accept ISO C89... none needed
checking how to run the C preprocessor... gcc -E
OK

The macro AC_TRY_CPP (see Section 18.4 [Obsolete Macros], page 293) used to play the role of AC_PREPROC_IFELSE, but double quotes its argument, making it impossible to use it to elaborate sources. You are encouraged to get rid of your old use of the macro AC_TRY_CPP in favor of AC_PREPROC_IFELSE, but, in the first place, are you sure you need to run the preprocessor and not the compiler?

[Macro] AC_EGREP_HEADER (pattern, header-file, action-if-found,
   [action-if-not-found])

If the output of running the preprocessor on the system header file header-file matches the extended regular expression pattern, execute shell commands action-if-found, otherwise execute action-if-not-found.

[Macro] AC_EGREP_CPP (pattern, program, [action-if-found],
   [action-if-not-found])

program is the text of a C or C++ program, on which shell variable, back quote, and backslash substitutions are performed. If the output of running the preprocessor on program matches the extended regular expression pattern, execute shell commands action-if-found, otherwise execute action-if-not-found.

6.4 Running the Compiler

To check for a syntax feature of the current language’s (see Section 6.1 [Language Choice], page 99) compiler, such as whether it recognizes a certain keyword, or simply to try some library feature, use AC_COMPILE_IFELSE to try to compile a small program that uses that feature.

[Macro] AC_COMPILE_IFELSE (input, [action-if-true], [action-if-false])

Run the compiler and compilation flags of the current language (see Section 6.1 [Language Choice], page 99) on the input, run the shell commands action-if-true on success, action-if-false otherwise. The input can be made by AC_LANG_PROGRAM and friends. It is customary to report unexpected failures with AC_MSG_FAILURE. This macro does not try to link; use AC_LINK_IFELSE if you need to do that (see Section 6.5 [Running the Linker], page 106). If needed, action-if-true can further access the just-compiled

     object file ‘conftest.$OBJEXT’.

This macro uses AC_REQUIRE for the compiler associated with the current language, which means that if the compiler has not yet been determined, the compiler determination will be made prior to the body of the outermust AC_DEFUN macro that triggered this macro to expand (see Section 20.8 [Expanded Before Required], page 334).

For tests in Erlang, the input must be the source code of a module named conftest. AC_ COMPILE_IFELSE generates a ‘conftest.beam’ file that can be interpreted by the Erlang virtual machine (ERL). It is recommended to use AC_LANG_PROGRAM to specify the test program, to ensure that the Erlang module has the right name.

6.5 Running the Linker

To check for a library, a function, or a global variable, Autoconf configure scripts try to compile and link a small program that uses it. This is unlike Metaconfig, which by default uses nm or ar on the C library to try to figure out which functions are available. Trying to link with the function is usually a more reliable approach because it avoids dealing with the variations in the options and output formats of nm and ar and in the location of the standard libraries. It also allows configuring for cross-compilation or checking a function’s runtime behavior if needed. On the other hand, it can be slower than scanning the libraries once, but accuracy is more important than speed.

AC_LINK_IFELSE is used to compile test programs to test for functions and global variables. It is also used by AC_CHECK_LIB to check for libraries (see Section 5.4 [Libraries], page 51), by adding the library being checked for to LIBS temporarily and trying to link a small program.

[Macro] AC_LINK_IFELSE (input, [action-if-true], [action-if-false])

Run the compiler (and compilation flags) and the linker of the current language (see Section 6.1 [Language Choice], page 99) on the input, run the shell commands action-if-true on success, action-if-false otherwise. The input can be made by AC_ LANG_PROGRAM and friends. If needed, action-if-true can further access the just-linked

     program file ‘conftest$EXEEXT’.

LDFLAGS and LIBS are used for linking, in addition to the current compilation flags.

It is customary to report unexpected failures with AC_MSG_FAILURE. This macro does not try to execute the program; use AC_RUN_IFELSE if you need to do that (see Section 6.6 [Runtime], page 107).

The AC_LINK_IFELSE macro cannot be used for Erlang tests, since Erlang programs are interpreted and do not require linking.

6.6 Checking Runtime Behavior

Sometimes you need to find out how a system performs at runtime, such as whether a given function has a certain capability or bug. If you can, make such checks when your program runs instead of when it is configured. You can check for things like the machine’s endianness when your program initializes itself.

If you really need to test for a runtime behavior while configuring, you can write a test program to determine the result, and compile and run it using AC_RUN_IFELSE. Avoid running test programs if possible, because this prevents people from configuring your package for cross-compiling.

[Macro] AC_RUN_IFELSE (input, [action-if-true], [action-if-false],
   [action-if-cross-compiling])

If program compiles and links successfully and returns an exit status of 0 when executed, run shell commands action-if-true. Otherwise, run shell commands action-if-false. The input can be made by AC_LANG_PROGRAM and friends. LDFLAGS and LIBS are used for linking, in addition to the compilation flags of the current language (see Section 6.1 [Language Choice], page 99). Additionally, action-if-true can run ./conftest$EXEEXT for further testing. If the compiler being used does not produce executables that run on the system where configure is being run, then the test program is not run. If the optional shell commands action-if-cross-compiling are given, they are run instead. Otherwise, configure prints an error message and exits. In the action-if-false section, the failing exit status is available in the shell variable ‘$?’. This exit status might be that of a failed compilation, or it might be that of a failed program execution. It is customary to report unexpected failures with AC_MSG_FAILURE.

Try to provide a pessimistic default value to use when cross-compiling makes runtime tests impossible. You do this by passing the optional last argument to AC_RUN_IFELSE. autoconf prints a warning message when creating configure each time it encounters a call to AC_RUN_IFELSE with no action-if-cross-compiling argument given. You may ignore the warning, though users cannot configure your package for cross-compiling. A few of the macros distributed with Autoconf produce this warning message.

To configure for cross-compiling you can also choose a value for those parameters based on the canonical system name (see Chapter 14 [Manual Configuration], page 267). Alternatively, set up a test results cache file with the correct values for the host system (see Section 7.4 [Caching Results], page 115).

To provide a default for calls of AC_RUN_IFELSE that are embedded in other macros, including a few of the ones that come with Autoconf, you can test whether the shell variable cross_compiling is set to ‘yes’, and then use an alternate method to get the results instead of calling the macros.

It is also permissible to temporarily assign to cross_compiling in order to force tests to behave as though they are in a cross-compilation environment, particularly since this provides a way to test your action-if-cross-compiling even when you are not using a crosscompiler.

     # We temporarily set cross-compile mode to force AC_COMPUTE_INT
     # to use the slow link-only method
     save_cross_compiling=$cross_compiling
     cross_compiling=yes
     AC_COMPUTE_INT([...])
     cross_compiling=$save_cross_compiling

A C or C++ runtime test should be portable. See Chapter 13 [Portable C and C++], page 257.

Erlang tests must exit themselves the Erlang VM by calling the halt/1 function: the given status code is used to determine the success of the test (status is 0) or its failure (status is different than 0), as explained above. It must be noted that data output through the standard output (e.g., using io:format/2) may be truncated when halting the VM. Therefore, if a test must output configuration information, it is recommended to create and to output data into the temporary file named ‘conftest.out’, using the functions of module file. The conftest.out file is automatically deleted by the AC_RUN_IFELSE macro. For instance, a simplified implementation of Autoconf’s AC_ERLANG_SUBST_LIB_DIR macro is:

     AC_INIT([LibdirTest], [1.0], [bug-libdirtest@example.org])
     AC_ERLANG_NEED_ERL
     AC_LANG(Erlang)
     AC_RUN_IFELSE(
       [AC_LANG_PROGRAM([], [dnl
         file:write_file("conftest.out", code:lib_dir()),
         halt(0)])],
       [echo "code:lib_dir() returned: `cat conftest.out`"],
       [AC_MSG_FAILURE([test Erlang program execution failed])])

6.7 Systemology

This section aims at presenting some systems and pointers to documentation. It may help you addressing particular problems reported by users.

Posix-conforming systems (http://www.opengroup.org/susv3) are derived from the

Unix operating system (http://www.bell-labs.com/history/unix/).

The Rosetta Stone for Unix (http://bhami.com/rosetta.html) contains a table correlating the features of various Posix-conforming systems. Unix History (http:// www.levenez.com/unix/) is a simplified diagram of how many Unix systems were derived from each other.

The Heirloom Project (http://heirloom.sourceforge.net/) provides some variants of traditional implementations of Unix utilities.

Darwin Darwin is also known as Mac OS X. Beware that the file system can be casepreserving, but case insensitive. This can cause nasty problems, since for instance the installation attempt for a package having an ‘INSTALL’ file can result in ‘make install’ report that nothing was to be done!

That’s all dependent on whether the file system is a UFS (case sensitive) or HFS+ (case preserving). By default Apple wants you to install the OS on HFS+. Unfortunately, there are some pieces of software which really need to be built on UFS. We may want to rebuild Darwin to have both UFS and HFS+ available (and put the /local/build tree on the UFS).

QNX 4.25 QNX is a realtime operating system running on Intel architecture meant to be scalable from the small embedded systems to the hundred processor supercomputer. It claims to be Posix certified. More information is available on the

QNX home page (http://www.qnx.com/).

Tru64 Documentation of several versions of Tru64 (http://h30097.www3.hp.com/ docs/) is available in different formats.

Unix version 7 Officially this was called the “Seventh Edition” of “the UNIX time-sharing system” but we use the more-common name “Unix version 7”. Documentation is available in the Unix Seventh Edition Manual (http://plan9.bell-labs.com/ 7thEdMan/). Previous versions of Unix are called “Unix version 6”, etc., but they were not as widely used.

6.8 Multiple Cases

Some operations are accomplished in several possible ways, depending on the OS variant. Checking for them essentially requires a “case statement”. Autoconf does not directly provide one; however, it is easy to simulate by using a shell variable to keep track of whether a way to perform the operation has been found yet.

Here is an example that uses the shell variable fstype to keep track of whether the remaining cases need to be checked. Note that since the value of fstype is under our control, we don’t have to use the longer ‘test "x$fstype" = xno’.

AC_MSG_CHECKING([how to get file system type])
fstype=no
# The order of these tests is important.
AC_COMPILE_IFELSE([AC_LANG_PROGRAM([[#include <sys/statvfs.h>
#include <sys/fstyp.h>]])],
                  [AC_DEFINE([FSTYPE_STATVFS], [1],
                     [Define if statvfs exists.])
                   fstype=SVR4])
if test $fstype = no; then
  AC_COMPILE_IFELSE([AC_LANG_PROGRAM([[#include <sys/statfs.h>
#include <sys/fstyp.h>]])],
                  [AC_DEFINE([FSTYPE_USG_STATFS], [1],
                     [Define if USG statfs.])
                   fstype=SVR3])
fi
if test $fstype = no; then
  AC_COMPILE_IFELSE([AC_LANG_PROGRAM([[#include <sys/statfs.h>
#include <sys/vmount.h>]])]),
                  [AC_DEFINE([FSTYPE_AIX_STATFS], [1],
                     [Define if AIX statfs.])
                   fstype=AIX])
fi
# (more cases omitted here)
AC_MSG_RESULT([$fstype])

7 Results of Tests

Once configure has determined whether a feature exists, what can it do to record that information? There are four sorts of things it can do: define a C preprocessor symbol, set a variable in the output files, save the result in a cache file for future configure runs, and print a message letting the user know the result of the test.

7.1 Defining C Preprocessor Symbols

A common action to take in response to a feature test is to define a C preprocessor symbol indicating the results of the test. That is done by calling AC_DEFINE or AC_DEFINE_UNQUOTED.

By default, AC_OUTPUT places the symbols defined by these macros into the output variable DEFS, which contains an option ‘-Dsymbol=value’ for each symbol defined. Unlike in Autoconf version 1, there is no variable DEFS defined while configure is running. To check whether Autoconf macros have already defined a certain C preprocessor symbol, test the value of the appropriate cache variable, as in this example:

AC_CHECK_FUNC([vprintf], [AC_DEFINE([HAVE_VPRINTF], [1],
                          [Define if vprintf exists.])])
if test "x$ac_cv_func_vprintf" != xyes; then
  AC_CHECK_FUNC([_doprnt], [AC_DEFINE([HAVE_DOPRNT], [1],
                            [Define if _doprnt exists.])])
fi

If AC_CONFIG_HEADERS has been called, then instead of creating DEFS, AC_OUTPUT creates a header file by substituting the correct values into #define statements in a template file. See Section 4.9 [Configuration Headers], page 35, for more information about this kind of output.

[Macro] AC_DEFINE (variable, value, [description])
[Macro] AC_DEFINE (variable)

Define variable to value (verbatim), by defining a C preprocessor macro for variable. variable should be a C identifier, optionally suffixed by a parenthesized argument list to define a C preprocessor macro with arguments. The macro argument list, if present, should be a comma-separated list of C identifiers, possibly terminated by an ellipsis ‘...’ if C99 syntax is employed. variable should not contain comments, white space, trigraphs, backslash-newlines, universal character names, or non-ASCII characters. value may contain backslash-escaped newlines, which will be preserved if you use AC_ CONFIG_HEADERS but flattened if passed via @DEFS@ (with no effect on the compilation, since the preprocessor sees only one line in the first place). value should not contain raw newlines. If you are not using AC_CONFIG_HEADERS, value should not contain any ‘#’ characters, as make tends to eat them. To use a shell variable, use AC_DEFINE_

UNQUOTED instead.

description is only useful if you are using AC_CONFIG_HEADERS. In this case, description is put into the generated ‘config.h.in’ as the comment before the macro define. The following example defines the C preprocessor variable EQUATION to be the string constant ‘"$a > $b"’:

     AC_DEFINE([EQUATION], ["$a > $b"],
       [Equation string.])

If neither value nor description are given, then value defaults to 1 instead of to the empty string. This is for backwards compatibility with older versions of Autoconf, but this usage is obsolescent and may be withdrawn in future versions of Autoconf.

If the variable is a literal string, it is passed to m4_pattern_allow (see Section 8.3.10 [Forbidden Patterns], page 162).

If multiple AC_DEFINE statements are executed for the same variable name (not counting any parenthesized argument list), the last one wins.

[Macro] AC_DEFINE_UNQUOTED (variable, value, [description])
[Macro] AC_DEFINE_UNQUOTED (variable)

Like AC_DEFINE, but three shell expansions are performed—once—on variable and value: variable expansion (‘$’), command substitution (‘`’), and backslash escaping (‘\’), as if in an unquoted here-document. Single and double quote characters in the value have no special meaning. Use this macro instead of AC_DEFINE when variable or value is a shell variable. Examples:

     AC_DEFINE_UNQUOTED([config_machfile], ["$machfile"],
       [Configuration machine file.])
     AC_DEFINE_UNQUOTED([GETGROUPS_T], [$ac_cv_type_getgroups],
       [getgroups return type.])
     AC_DEFINE_UNQUOTED([$ac_tr_hdr], [1],
       [Translated header name.])

Due to a syntactical bizarreness of the Bourne shell, do not use semicolons to separate AC_DEFINE or AC_DEFINE_UNQUOTED calls from other macro calls or shell code; that can cause syntax errors in the resulting configure script. Use either blanks or newlines. That is, do this:

AC_CHECK_HEADER([elf.h],
  [AC_DEFINE([SVR4], [1], [System V Release 4]) LIBS="-lelf $LIBS"])

or this:

AC_CHECK_HEADER([elf.h],
  [AC_DEFINE([SVR4], [1], [System V Release 4])
   LIBS="-lelf $LIBS"])

instead of this:

AC_CHECK_HEADER([elf.h],
  [AC_DEFINE([SVR4], [1], [System V Release 4]); LIBS="-lelf $LIBS"])

7.2 Setting Output Variables

Another way to record the results of tests is to set output variables, which are shell variables whose values are substituted into files that configure outputs. The two macros below create new output variables. See Section 4.8.1 [Preset Output Variables], page 25, for a list of output variables that are always available.

[Macro] AC_SUBST (variable, [value])

Create an output variable from a shell variable. Make AC_OUTPUT substitute the variable variable into output files (typically one or more makefiles). This means that AC_OUTPUT replaces instances of ‘@variable@’ in input files with the value that the shell variable variable has when AC_OUTPUT is called. The value can contain any non-NUL character, including newline. If you are using Automake 1.11 or newer, for newlines in values you might want to consider using AM_SUBST_NOTMAKE to prevent

automake from adding a line variable = @variable@ to the ‘Makefile.in’ files (see

Section “Automake” in Other things Automake recognizes). Variable occurrences should not overlap: e.g., an input file should not contain ‘@var1@var2@’ if var1 and var2 are variable names. The substituted value is not rescanned for more output variables; occurrences of ‘@variable@’ in the value are inserted literally into the output file. (The algorithm uses the special marker |#_!!_#| internally, so neither the substituted value nor the output file may contain

|#_!!_#|.)

If value is given, in addition assign it to variable. The string variable is passed to m4_pattern_allow (see Section 8.3.10 [Forbidden Patterns], page 162).

[Macro] AC_SUBST_FILE (variable)

Another way to create an output variable from a shell variable. Make AC_OUTPUT insert (without substitutions) the contents of the file named by shell variable variable into output files. This means that AC_OUTPUT replaces instances of ‘@variable@’ in output files (such as ‘Makefile.in’) with the contents of the file that the shell variable variable names when AC_OUTPUT is called. Set the variable to ‘/dev/null’ for cases that do not have a file to insert. This substitution occurs only when the ‘@variable@’ is on a line by itself, optionally surrounded by spaces and tabs. The substitution replaces the whole line, including the spaces, tabs, and the terminating newline. This macro is useful for inserting makefile fragments containing special dependencies or other make directives for particular host or target types into makefiles. For example, ‘configure.ac’ could contain:

     AC_SUBST_FILE([host_frag])
     host_frag=$srcdir/conf/sun4.mh

and then a ‘Makefile.in’ could contain:

     @host_frag@

The string variable is passed to m4_pattern_allow (see Section 8.3.10 [Forbidden Patterns], page 162).

Running configure in varying environments can be extremely dangerous. If for instance the user runs ‘CC=bizarre-cc ./configure’, then the cache, ‘config.h’, and many other output files depend upon bizarre-cc being the C compiler. If for some reason the user runs ./configure again, or if it is run via ‘./config.status --recheck’, (See Section 4.8.5 [Automatic Remaking], page 34, and see Chapter 17 [config.status Invocation], page 287), then the configuration can be inconsistent, composed of results depending upon two different compilers.

Environment variables that affect this situation, such as ‘CC’ above, are called precious variables, and can be declared as such by AC_ARG_VAR.

[Macro] AC_ARG_VAR (variable, description)

Declare variable is a precious variable, and include its description in the variable

section of ‘./configure --help’.

Being precious means that − variable is substituted via AC_SUBST. − The value of variable when configure was launched is saved in the cache,

including if it was not specified on the command line but via the environment. Indeed, while configure can notice the definition of CC in ‘./configure

CC=bizarre-cc’, it is impossible to notice it in ‘CC=bizarre-cc ./configure’,

which, unfortunately, is what most users do. We emphasize that it is the initial value of variable which is saved, not that found during the execution of configure. Indeed, specifying ‘./configure FOO=foo’ and letting ‘./configure’ guess that FOO is foo can be two different things.

− variable is checked for consistency between two configure runs. For instance:

     $ ./configure --silent --config-cache
     $ CC=cc ./configure --silent --config-cache
     configure: error: `CC' was not set in the previous run
     configure: error: changes in the environment can compromise \
     the build
     configure: error: run `make distclean' and/or \
     `rm config.cache' and start over

and similarly if the variable is unset, or if its content is changed. If the content has white space changes only, then the error is degraded to a warning only, but the old value is reused.

− variable is kept during automatic reconfiguration (see Chapter 17 [config.status Invocation], page 287) as if it had been passed as a command line argument, including when no cache is used:

     $ CC=/usr/bin/cc ./configure var=raboof --silent
     $ ./config.status --recheck
     running CONFIG_SHELL=/bin/sh /bin/sh ./configure var=raboof \
       CC=/usr/bin/cc  --no-create --no-recursion

7.3 Special Characters in Output Variables

Many output variables are intended to be evaluated both by make and by the shell. Some characters are expanded differently in these two contexts, so to avoid confusion these variables’ values should not contain any of the following characters:

" # $ & ' ( ) * ; < > ? [ \ ^` |

Also, these variables’ values should neither contain newlines, nor start with ‘~’, nor contain white space or ‘:’ immediately followed by ‘~’. The values can contain nonempty sequences of white space characters like tabs and spaces, but each such sequence might arbitrarily be replaced by a single space during substitution.

These restrictions apply both to the values that configure computes, and to the values set directly by the user. For example, the following invocations of configure are problematic, since they attempt to use special characters within CPPFLAGS and white space within

$(srcdir):
CPPFLAGS='-DOUCH="&\"#$*?"' '../My Source/ouch-1.0/configure'
'../My Source/ouch-1.0/configure' CPPFLAGS='-DOUCH="&\"#$*?"'

7.4 Caching Results

To avoid checking for the same features repeatedly in various configure scripts (or in repeated runs of one script), configure can optionally save the results of many checks in a cache file (see Section 7.4.2 [Cache Files], page 117). If a configure script runs with caching enabled and finds a cache file, it reads the results of previous runs from the cache and avoids rerunning those checks. As a result, configure can then run much faster than if it had to perform all of the checks every time.

[Macro] AC_CACHE_VAL (cache-id, commands-to-set-it)

Ensure that the results of the check identified by cache-id are available. If the results of the check were in the cache file that was read, and configure was not given the ‘--quiet’ or ‘--silent’ option, print a message saying that the result was cached; otherwise, run the shell commands commands-to-set-it. If the shell commands are run to determine the value, the value is saved in the cache file just before configure creates its output files. See Section 7.4.1 [Cache Variable Names], page 116, for how to choose the name of the cache-id variable.

The commands-to-set-it must have no side effects except for setting the variable cache-id, see below.

[Macro] AC_CACHE_CHECK (message, cache-id, commands-to-set-it)

A wrapper for AC_CACHE_VAL that takes care of printing the messages. This macro provides a convenient shorthand for the most common way to use these macros. It calls AC_MSG_CHECKING for message, then AC_CACHE_VAL with the cache-id and commands arguments, and AC_MSG_RESULT with cache-id.

The commands-to-set-it must have no side effects except for setting the variable cache-id, see below.

It is common to find buggy macros using AC_CACHE_VAL or AC_CACHE_CHECK, because people are tempted to call AC_DEFINE in the commands-to-set-it. Instead, the code that follows the call to AC_CACHE_VAL should call AC_DEFINE, by examining the value of the cache variable. For instance, the following macro is broken:

     AC_DEFUN([AC_SHELL_TRUE],
     [AC_CACHE_CHECK([whether true(1) works], [my_cv_shell_true_works],
                     [my_cv_shell_true_works=no
                      (true) 2>/dev/null && my_cv_shell_true_works=yes
                      if test "x$my_cv_shell_true_works" = xyes; then
                        AC_DEFINE([TRUE_WORKS], [1],
                                  [Define if `true(1)' works properly.])
                      fi])
     ])

This fails if the cache is enabled: the second time this macro is run, TRUE_WORKS will not be defined. The proper implementation is:

     AC_DEFUN([AC_SHELL_TRUE],
     [AC_CACHE_CHECK([whether true(1) works], [my_cv_shell_true_works],
                     [my_cv_shell_true_works=no
                      (true) 2>/dev/null && my_cv_shell_true_works=yes])
      if test "x$my_cv_shell_true_works" = xyes; then
        AC_DEFINE([TRUE_WORKS], [1],
                  [Define if `true(1)' works properly.])
      fi
     ])

Also, commands-to-set-it should not print any messages, for example with AC_MSG_ CHECKING; do that before calling AC_CACHE_VAL, so the messages are printed regardless of whether the results of the check are retrieved from the cache or determined by running the shell commands.

7.4.1 Cache Variable Names

The names of cache variables should have the following format:

     package-prefix_cv_value-type_specific-value_[additional-options]
for example, ‘ac_cv_header_stat_broken’ or ‘ac_cv_prog_gcc_traditional’. The parts

of the variable name are:

package-prefix An abbreviation for your package or organization; the same prefix you begin local Autoconf macros with, except lowercase by convention. For cache values used by the distributed Autoconf macros, this value is ‘ac’.

_cv_Indicates that this shell variable is a cache value. This string must be present in the variable name, including the leading underscore.

value-type A convention for classifying cache values, to produce a rational naming system. The values used in Autoconf are listed in Section 10.2 [Macro Names], page 176.

specific-value Which member of the class of cache values this test applies to. For example, which function (‘alloca’), program (‘gcc’), or output variable (‘INSTALL’). additional-options

Any particular behavior of the specific member that this test applies to. For example, ‘broken’ or ‘set’. This part of the name may be omitted if it does not apply.

The values assigned to cache variables may not contain newlines. Usually, their values are Boolean (‘yes’ or ‘no’) or the names of files or functions; so this is not an important restriction. Section B.4 [Cache Variable Index], page 357 for an index of cache variables with documented semantics.

7.4.2 Cache Files

A cache file is a shell script that caches the results of configure tests run on one system so they can be shared between configure scripts and configure runs. It is not useful on other systems. If its contents are invalid for some reason, the user may delete or edit it, or override documented cache variables on the configure command line.

By default, configure uses no cache file, to avoid problems caused by accidental use of stale cache files.

To enable caching, configure accepts ‘--config-cache’ (or ‘-C’) to cache results in the file ‘config.cache’. Alternatively, ‘--cache-file=file’ specifies that file be the cache file. The cache file is created if it does not exist already. When configure calls configure scripts in subdirectories, it uses the ‘--cache-file’ argument so that they share the same cache. See Section 4.12 [Subdirectories], page 40, for information on configuring subdirectories

with the AC_CONFIG_SUBDIRS macro.

‘config.status’ only pays attention to the cache file if it is given the ‘--recheck’ option, which makes it rerun configure.

It is wrong to try to distribute cache files for particular system types. There is too much room for error in doing that, and too much administrative overhead in maintaining them. For any features that can’t be guessed automatically, use the standard method of the canonical system type and linking files (see Chapter 14 [Manual Configuration], page 267).

The site initialization script can specify a site-wide cache file to use, instead of the usual per-program cache. In this case, the cache file gradually accumulates information whenever someone runs a new configure script. (Running configure merges the new cache results with the existing cache file.) This may cause problems, however, if the system configuration (e.g., the installed libraries or compilers) changes and the stale cache file is not deleted.

7.4.3 Cache Checkpointing

If your configure script, or a macro called from ‘configure.ac’, happens to abort the configure process, it may be useful to checkpoint the cache a few times at key points using AC_CACHE_SAVE. Doing so reduces the amount of time it takes to rerun the configure script with (hopefully) the error that caused the previous abort corrected.

[Macro] AC_CACHE_LOAD

Loads values from existing cache file, or creates a new cache file if a cache file is not found. Called automatically from AC_INIT.

[Macro] AC_CACHE_SAVE

Flushes all cached values to the cache file. Called automatically from AC_OUTPUT, but it can be quite useful to call AC_CACHE_SAVE at key points in ‘configure.ac’.

For instance:

… AC INIT, etc. …

# Checks for programs.
AC_PROG_CC
AC_PROG_AWK

… more program checks …

AC_CACHE_SAVE
# Checks for libraries.
AC_CHECK_LIB([nsl], [gethostbyname])
AC_CHECK_LIB([socket], [connect])

… more lib checks …

AC_CACHE_SAVE
# Might abort...
AM_PATH_GTK([1.0.2], [], [AC_MSG_ERROR([GTK not in path])])
AM_PATH_GTKMM([0.9.5], [], [AC_MSG_ERROR([GTK not in path])])

… AC OUTPUT, etc. …

7.5 Printing Messages

configure scripts need to give users running them several kinds of information. The following macros print messages in ways appropriate for each kind. The arguments to all of them get enclosed in shell double quotes, so the shell performs variable and back-quote substitution on them.

These macros are all wrappers around the echo shell command. They direct output to the appropriate file descriptor (see Section 9.4 [File Descriptor Macros], page 172). configure scripts should rarely need to run echo directly to print messages for the user. Using these macros makes it easy to change how and when each kind of message is printed; such changes need only be made to the macro definitions and all the callers change automatically.

To diagnose static issues, i.e., when autoconf is run, see Section 8.3.2 [Diagnostic Macros], page 138.

[Macro] AC_MSG_CHECKING (feature-description)

Notify the user that configure is checking for a particular feature. This macro prints a message that starts with ‘checking ’ and ends with ‘...’ and no newline. It must be followed by a call to AC_MSG_RESULT to print the result of the check and the newline. The feature-description should be something like ‘whether the Fortran

compiler accepts C++ comments’ or ‘for c89’.

This macro prints nothing if configure is run with the ‘--quiet’ or ‘--silent’ option.

[Macro] AC_MSG_RESULT (result-description)

Notify the user of the results of a check. result-description is almost always the value of the cache variable for the check, typically ‘yes’, ‘no’, or a file name. This macro should follow a call to AC_MSG_CHECKING, and the result-description should be the completion of the message printed by the call to AC_MSG_CHECKING. This macro prints nothing if configure is run with the ‘--quiet’ or ‘--silent’ option.

[Macro] AC_MSG_NOTICE (message)

Deliver the message to the user. It is useful mainly to print a general description of the overall purpose of a group of feature checks, e.g.,

     AC_MSG_NOTICE([checking if stack overflow is detectable])

This macro prints nothing if configure is run with the ‘--quiet’ or ‘--silent’ option.

[Macro] AC_MSG_ERROR (error-description, [exit-status = ‘$?/1’])

Notify the user of an error that prevents configure from completing. This macro prints an error message to the standard error output and exits configure with exitstatus (‘$?’ by default, except that ‘0’ is converted to ‘1’). error-description should

be something like ‘invalid value $HOME for \$HOME’.

The error-description should start with a lower-case letter, and “cannot” is preferred to “can’t”.

[Macro] AC_MSG_FAILURE (error-description, [exit-status])

This AC_MSG_ERROR wrapper notifies the user of an error that prevents configure from completing and that additional details are provided in ‘config.log’. This is typically used when abnormal results are found during a compilation.

[Macro] AC_MSG_WARN (problem-description)

Notify the configure user of a possible problem. This macro prints the message to the standard error output; configure continues running afterward, so macros that call AC_MSG_WARN should provide a default (back-up) behavior for the situations they warn about. problem-description should be something like ‘ln -s seems to make

hard links’.

8 Programming in M4

Autoconf is written on top of two layers: M4sugar, which provides convenient macros for pure M4 programming, and M4sh, which provides macros dedicated to shell script generation.

As of this version of Autoconf, these two layers still contain experimental macros, whose interface might change in the future. As a matter of fact, anything that is not documented must not be used.

8.1 M4 Quotation

The most common problem with existing macros is an improper quotation. This section, which users of Autoconf can skip, but which macro writers must read, first justifies the quotation scheme that was chosen for Autoconf and then ends with a rule of thumb. Understanding the former helps one to follow the latter.

8.1.1 Active Characters

To fully understand where proper quotation is important, you first need to know what the special characters are in Autoconf: ‘#’ introduces a comment inside which no macro expansion is performed, ‘,’ separates arguments, ‘[’ and ‘]’ are the quotes themselves^{1}, ‘(’ and ‘)’ (which M4 tries to match by pairs), and finally ‘$’ inside a macro definition.

In order to understand the delicate case of macro calls, we first have to present some obvious failures. Below they are “obvious-ified”, but when you find them in real life, they are usually in disguise.

Comments, introduced by a hash and running up to the newline, are opaque tokens to the top level: active characters are turned off, and there is no macro expansion:

     # define([def], ine)
     ⇒# define([def], ine)

Each time there can be a macro expansion, there is a quotation expansion, i.e., one level of quotes is stripped:

     int tab[10];
     ⇒int tab10;
     [int tab[10];]
     ⇒int tab[10];

Without this in mind, the reader might try hopelessly to use her macro array:

     define([array], [int tab[10];])
     array
     ⇒int tab10;
     [array]
     ⇒array

How can you correctly output the intended results^{2}?

1 By itself, M4 uses ‘`’ and ‘'’; it is the M4sugar layer that sets up the preferred quotes of ‘[’ and ‘]’.

2 Using defn.

8.1.2 One Macro Call

Let’s proceed on the interaction between active characters and macros with this small macro, which just returns its first argument:

define([car], [$1])

The two pairs of quotes above are not part of the arguments of define; rather, they are understood by the top level when it tries to find the arguments of define. Therefore, assuming car is not already defined, it is equivalent to write:

define(car, $1)

But, while it is acceptable for a ‘configure.ac’ to avoid unnecessary quotes, it is bad practice for Autoconf macros which must both be more robust and also advocate perfect style.

At the top level, there are only two possibilities: either you quote or you don’t:

car(foo, bar, baz)
⇒foo
[car(foo, bar, baz)]
⇒car(foo, bar, baz)

Let’s pay attention to the special characters:

car(#)
 error EOF in argument list

The closing parenthesis is hidden in the comment; with a hypothetical quoting, the top level understood it this way:

car([#)]

Proper quotation, of course, fixes the problem:

car([#])
⇒#

Here are more examples:

car(foo, bar)
⇒foo
car([foo, bar])
⇒foo, bar
car((foo, bar))
⇒(foo, bar)
car([(foo], [bar)])
⇒(foo
define([a], [b])

⇒

car(a)
⇒b
car([a])
⇒b
car([[a]])
⇒a
car([[[a]]])
⇒[a]

8.1.3 Quoting and Parameters

When M4 encounters ‘$’ within a macro definition, followed immediately by a character it recognizes (‘0’…‘9’, ‘#’, ‘@’, or ‘*’), it will perform M4 parameter expansion. This happens regardless of how many layers of quotes the parameter expansion is nested within, or even if it occurs in text that will be rescanned as a comment.

define([none], [$1])

⇒

define([one], [[$1]])

⇒

define([two], [[[$1]]])

⇒

define([comment], [# $1])

⇒

define([active], [ACTIVE])

⇒

none([active])
⇒ACTIVE
one([active])
⇒active
two([active])
⇒[active]
comment([active])
⇒# active

On the other hand, since autoconf generates shell code, you often want to output shell variable expansion, rather than performing M4 parameter expansion. To do this, you must use M4 quoting to separate the ‘$’ from the next character in the definition of your macro. If the macro definition occurs in single-quoted text, then insert another level of quoting; if the usage is already inside a double-quoted string, then split it into concatenated strings.

define([single], [a single-quoted $[]1 definition])

⇒

define([double], [[a double-quoted $][1 definition]])

⇒

single
⇒a single-quoted $1 definition
double
⇒a double-quoted $1 definition

Posix states that M4 implementations are free to provide implementation extensions when ‘${’ is encountered in a macro definition. Autoconf reserves the longer sequence ‘${{’ for use with planned extensions that will be available in the future GNU M4 2.0, but guarantees that all other instances of ‘${’ will be output literally. Therefore, this idiom can also be used to output shell code parameter references:

define([first], [${1}])first
⇒${1}

Posix also states that ‘$11’ should expand to the first parameter concatenated with a literal ‘1’, although some versions of GNU M4 expand the eleventh parameter instead. For portability, you should only use single-digit M4 parameter expansion.

With this in mind, we can explore the cases where macros invoke macros…

8.1.4 Quotation and Nested Macros

The examples below use the following macros:

     define([car], [$1])
     define([active], [ACT, IVE])
     define([array], [int tab[10]])

Each additional embedded macro call introduces other possible interesting quotations:

     car(active)
     ⇒ACT
     car([active])
     ⇒ACT, IVE
     car([[active]])
     ⇒active

In the first case, the top level looks for the arguments of car, and finds ‘active’. Because M4 evaluates its arguments before applying the macro, ‘active’ is expanded, which results in:

     car(ACT, IVE)
     ⇒ACT

In the second case, the top level gives ‘active’ as first and only argument of car, which results in:

     active
     ⇒ACT, IVE

i.e., the argument is evaluated after the macro that invokes it. In the third case, car receives ‘[active]’, which results in:

     [active]
     ⇒active

exactly as we already saw above.

The example above, applied to a more realistic example, gives:

     car(int tab[10];)
     ⇒int tab10;
     car([int tab[10];])
     ⇒int tab10;
     car([[int tab[10];]])
     ⇒int tab[10];

Huh? The first case is easily understood, but why is the second wrong, and the third right? To understand that, you must know that after M4 expands a macro, the resulting text is immediately subjected to macro expansion and quote removal. This means that the quote removal occurs twice—first before the argument is passed to the car macro, and second after the car macro expands to the first argument.

As the author of the Autoconf macro car, you then consider it to be incorrect that your users have to double-quote the arguments of car, so you “fix” your macro. Let’s call it qar for quoted car:

     define([qar], [[$1]])

and check that qar is properly fixed:

     qar([int tab[10];])
     ⇒int tab[10];

Ahhh! That’s much better.

But note what you’ve done: now that the result of qar is always a literal string, the only time a user can use nested macros is if she relies on an unquoted macro call:

     qar(active)
     ⇒ACT
     qar([active])
     ⇒active

leaving no way for her to reproduce what she used to do with car:

     car([active])
     ⇒ACT, IVE

Worse yet: she wants to use a macro that produces a set of cpp macros:

     define([my_includes], [#include <stdio.h>])
     car([my_includes])
     ⇒#include <stdio.h>
     qar(my_includes)
      error EOF in argument list

This macro, qar, because it double quotes its arguments, forces its users to leave their macro calls unquoted, which is dangerous. Commas and other active symbols are interpreted by M4 before they are given to the macro, often not in the way the users expect. Also, because qar behaves differently from the other macros, it’s an exception that should be avoided in Autoconf.

8.1.5 changequote is Evil

The temptation is often high to bypass proper quotation, in particular when it’s late at night. Then, many experienced Autoconf hackers finally surrender to the dark side of the force and use the ultimate weapon: changequote.

The M4 builtin changequote belongs to a set of primitives that allow one to adjust the syntax of the language to adjust it to one’s needs. For instance, by default M4 uses ‘`’ and ‘'’ as quotes, but in the context of shell programming (and actually of most programming languages), that’s about the worst choice one can make: because of strings and back-quoted expressions in shell code (such as ‘'this'’ and ‘`that`’), and because of literal characters in usual programming languages (as in ‘'0'’), there are many unbalanced ‘`’ and ‘'’. Proper M4 quotation then becomes a nightmare, if not impossible. In order to make M4 useful in such a context, its designers have equipped it with changequote, which makes it possible to choose another pair of quotes. M4sugar, M4sh, Autoconf, and Autotest all have chosen to use ‘[’ and ‘]’. Not especially because they are unlikely characters, but because they are characters unlikely to be unbalanced.

There are other magic primitives, such as changecom to specify what syntactic forms are comments (it is common to see ‘changecom(<!--, -->)’ when M4 is used to produce HTML pages), changeword and changesyntax to change other syntactic details (such as the character to denote the nth argument, ‘$’ by default, the parentheses around arguments, etc.).

These primitives are really meant to make M4 more useful for specific domains: they should be considered like command line options: ‘--quotes’, ‘--comments’, ‘--words’, and ‘--syntax’. Nevertheless, they are implemented as M4 builtins, as it makes M4 libraries self contained (no need for additional options).

There lies the problem…

The problem is that it is then tempting to use them in the middle of an M4 script, as opposed to its initialization. This, if not carefully thought out, can lead to disastrous effects: you are changing the language in the middle of the execution. Changing and restoring the syntax is often not enough: if you happened to invoke macros in between, these macros are lost, as the current syntax is probably not the one they were implemented with.

8.1.6 Quadrigraphs

When writing an Autoconf macro you may occasionally need to generate special characters that are difficult to express with the standard Autoconf quoting rules. For example, you may need to output the regular expression ‘[^[]’, which matches any character other than ‘[’. This expression contains unbalanced brackets so it cannot be put easily into an M4 macro.

Additionally, there are a few m4sugar macros (such as m4_split and m4_expand) which internally use special markers in addition to the regular quoting characters. If the arguments to these macros contain the literal strings ‘-=<{(’ or ‘)}>=-’, the macros might behave incorrectly.

You can work around these problems by using one of the following quadrigraphs:

‘@<:@’  ‘[’
‘@:>@’  ‘]’
‘@S|@’  ‘$’
‘@%:@’  ‘#’
‘@{:@’  ‘(’
‘@:}@’  ‘)’

‘@&t@’ Expands to nothing.

Quadrigraphs are replaced at a late stage of the translation process, after m4 is run, so they do not get in the way of M4 quoting. For example, the string ‘^@<:@’, independently of its quotation, appears as ‘^[’ in the output.

The empty quadrigraph can be used: − to mark trailing spaces explicitly

Trailing spaces are smashed by autom4te. This is a feature.

− to produce quadrigraphs and other strings reserved by m4sugar

For instance ‘@<@&t@:@’ produces ‘@<:@’. For a more contrived example:

    m4_define([a], [A])m4_define([b], [B])m4_define([c], [C])dnl
    m4_split([a )}>=- b -=<{( c])
    ⇒[a], [], [B], [], [c]
    m4_split([a )}@&t@>=- b -=<@&t@{( c])
    ⇒[a], [)}>=-], [b], [-=<{(], [c]

− to escape occurrences of forbidden patterns For instance you might want to mention AC_FOO in a comment, while still being sure that autom4te still catches unexpanded ‘AC_*’. Then write ‘AC@&t@_FOO’.

The name ‘@&t@’ was suggested by Paul Eggert:

I should give some credit to the ‘@&t@’ pun. The ‘&’ is my own invention, but the ‘t’ came from the source code of the ALGOL68C compiler, written by Steve Bourne (of Bourne shell fame), and which used ‘mt’ to denote the empty string. In C, it would have looked like something like:

     char const mt[] = "";

but of course the source code was written in Algol 68. I don’t know where he got ‘mt’ from: it could have been his own invention, and I suppose it could have been a common pun around the Cambridge University computer lab at the time.

8.1.7 Dealing with unbalanced parentheses

One of the pitfalls of portable shell programming is that case statements require unbalanced parentheses (see [Limitations of Shell Builtins], page 212). With syntax highlighting editors, the presence of unbalanced ‘)’ can interfere with editors that perform syntax highlighting of macro contents based on finding the matching ‘(’. Another concern is how much editing must be done when transferring code snippets between shell scripts and macro definitions. But most importantly, the presence of unbalanced parentheses can introduce expansion bugs.

For an example, here is an underquoted attempt to use the macro my_case, which happens to expand to a portable case statement:

AC_DEFUN([my_case],
[case $file_name in
  *.c) echo "C source code";;
esac])
AS_IF(:, my_case)

In the above example, the AS_IF call underquotes its arguments. As a result, the unbalanced ‘)’ generated by the premature expansion of my_case results in expanding AS_IF with a truncated parameter, and the expansion is syntactically invalid:

if :; then
  case $file_name in
  *.c
fi echo "C source code";;
esac)

If nothing else, this should emphasize the importance of the quoting arguments to macro calls. On the other hand, there are several variations for defining my_case to be more robust, even when used without proper quoting, each with some benefits and some drawbacks.

Creative literal shell comment

     AC_DEFUN([my_case],
     [case $file_name in #(
       *.c) echo "C source code";;
     esac])

This version provides balanced parentheses to several editors, and can be copied and pasted into a terminal as is. Unfortunately, it is still unbalanced as an Autoconf argument, since ‘#(’ is an M4 comment that masks the normal properties of ‘(’. Quadrigraph shell comment

     AC_DEFUN([my_case],
     [case $file_name in @%:@(
       *.c) echo "C source code";;
     esac])

This version provides balanced parentheses to even more editors, and can be used as a balanced Autoconf argument. Unfortunately, it requires some editing before it can be copied and pasted into a terminal, and the use of the quadrigraph ‘@%:@’ for ‘#’ reduces readability. Quoting just the parenthesis

     AC_DEFUN([my_case],
     [case $file_name in
       *.c[)] echo "C source code";;
     esac])

This version quotes the ‘)’, so that it can be used as a balanced Autoconf argument. As written, this is not balanced to an editor, but it can be coupled with ‘[#(]’ to meet that need, too. However, it still requires some edits before it can be copied and pasted into a terminal. Double-quoting the entire statement

     AC_DEFUN([my_case],
     [[case $file_name in #(
       *.c) echo "C source code";;
     esac]])

Since the entire macro is double-quoted, there is no problem with using this as an Autoconf argument; and since the double-quoting is over the entire statement, this code can be easily copied and pasted into a terminal. However, the double quoting prevents the expansion of any macros inside the case statement, which may cause its own set of problems.

Using AS_CASE
     AC_DEFUN([my_case],
     [AS_CASE([$file_name],
       [*.c], [echo "C source code"])])

This version avoids the balancing issue altogether, by relying on AS_CASE (see Section 9.1 [Common Shell Constructs], page 165); it also allows for the expansion of AC_REQUIRE to occur prior to the entire case statement, rather than within a branch of the case statement that might not be taken. However, the abstraction comes with a penalty that it is no longer a quick copy, paste, and edit to get back to shell code.

8.1.8 Quotation Rule Of Thumb

To conclude, the quotation rule of thumb is:

One pair of quotes per pair of parentheses.

Never over-quote, never under-quote, in particular in the definition of macros. In the few places where the macros need to use brackets (usually in C program text or regular expressions), properly quote the arguments!

It is common to read Autoconf programs with snippets like:

AC_TRY_LINK(
changequote(<<, >>)dnl
<<#include <time.h>
#ifndef tzname /* For SGI.  */
extern char *tzname[]; /* RS6000 and others reject char **tzname.  */
#endif>>,
changequote([, ])dnl
[atoi (*tzname);], ac_cv_var_tzname=yes, ac_cv_var_tzname=no)

which is incredibly useless since AC_TRY_LINK is already double quoting, so you just need:

AC_TRY_LINK(
[#include <time.h>
#ifndef tzname /* For SGI.  */
extern char *tzname[]; /* RS6000 and others reject char **tzname.  */
#endif],
            [atoi (*tzname);],
            [ac_cv_var_tzname=yes],
            [ac_cv_var_tzname=no])

The M4-fluent reader might note that these two examples are rigorously equivalent, since M4 swallows both the ‘changequote(<<, >>)’ and ‘<<’ ‘>>’ when it collects the arguments: these quotes are not part of the arguments!

Simplified, the example above is just doing this:

changequote(<<, >>)dnl
<<[]>>
changequote([, ])dnl

instead of simply:

[[]]

With macros that do not double quote their arguments (which is the rule), double-quote the (risky) literals:

AC_LINK_IFELSE([AC_LANG_PROGRAM(
[[#include <time.h>
#ifndef tzname /* For SGI.  */

     extern char *tzname[]; /* RS6000 and others reject char **tzname.  */
     #endif]],
                                     [atoi (*tzname);])],
                    [ac_cv_var_tzname=yes],
                    [ac_cv_var_tzname=no])

Please note that the macro AC_TRY_LINK is obsolete, so you really should be using AC_

LINK_IFELSE instead.

See Section 8.1.6 [Quadrigraphs], page 126, for what to do if you run into a hopeless case where quoting does not suffice.

When you create a configure script using newly written macros, examine it carefully to check whether you need to add more quotes in your macros. If one or more words have disappeared in the M4 output, you need more quotes. When in doubt, quote.

However, it’s also possible to put on too many layers of quotes. If this happens, the resulting configure script may contain unexpanded macros. The autoconf program checks for this problem by looking for the string ‘AC_’ in ‘configure’. However, this heuristic does not work in general: for example, it does not catch overquoting in AC_DEFINE descriptions.

8.2 Using autom4te

The Autoconf suite, including M4sugar, M4sh, and Autotest, in addition to Autoconf per se, heavily rely on M4. All these different uses revealed common needs factored into a layer over M4: autom4te^{3}.

autom4te is a preprocessor that is like m4. It supports M4 extensions designed for use in tools like Autoconf.

8.2.1 Invoking autom4te

The command line arguments are modeled after M4’s:

     autom4te options files

where the files are directly passed to m4. By default, GNU M4 is found during configuration, but the environment variable M4 can be set to tell autom4te where to look. In addition to the regular expansion, it handles the replacement of the quadrigraphs (see Section 8.1.6 [Quadrigraphs], page 126), and of ‘__oline__’, the current line in the output. It supports an extended syntax for the files:

‘file.m4f’

This file is an M4 frozen file. Note that all the previous files are ignored. See the option ‘--melt’ for the rationale.

‘file?’ If found in the library path, the file is included for expansion, otherwise it is ignored instead of triggering a failure.

Of course, it supports the Autoconf common subset of options:

‘--help’

‘-h’ Print a summary of the command line options and exit.

3 Yet another great name from Lars J. Aas.

‘--version’

‘-V’ Print the version number of Autoconf and exit.

‘--verbose’

‘-v’ Report processing steps.

‘--debug’

‘-d’ Don’t remove the temporary files and be even more verbose.

‘--include=dir’

‘-I dir’ Also look for input files in dir. Multiple invocations accumulate.

‘--output=file’

‘-o file’ Save output (script or trace) to file. The file ‘-’ stands for the standard output.

As an extension of m4, it includes the following options:

‘--warnings=category’
‘-W category’

Report the warnings related to category (which can actually be a comma separated list). See Section 10.3 [Reporting Messages], page 177, macro AC_ DIAGNOSE, for a comprehensive list of categories. Special values include:

‘all’ report all the warnings

‘none’ report none

‘error’ treats warnings as errors

‘no-category’

disable warnings falling into category

Warnings about ‘syntax’ are enabled by default, and the environment variable WARNINGS, a comma separated list of categories, is honored. ‘autom4te -W category’ actually behaves as if you had run:

     autom4te --warnings=syntax,$WARNINGS,category

For example, if you want to disable defaults and WARNINGS of autom4te, but enable the warnings about obsolete constructs, you would use ‘-W

none,obsolete’.

autom4te displays a back trace for errors, but not for warnings; if you want them, just pass ‘-W error’.

‘--melt’

‘-M’ Do not use frozen files. Any argument file.m4f is replaced by file.m4. This helps tracing the macros which are executed only when the files are frozen, typically m4_define. For instance, running:

     autom4te --melt 1.m4 2.m4f 3.m4 4.m4f input.m4

is roughly equivalent to running:

     m4 1.m4 2.m4 3.m4 4.m4 input.m4

while

               autom4te 1.m4 2.m4f 3.m4 4.m4f input.m4

is equivalent to:

               m4 --reload-state=4.m4f input.m4
‘--freeze’

‘-F’ Produce a frozen state file. autom4te freezing is stricter than M4’s: it must produce no warnings, and no output other than empty lines (a line with white space is not empty) and comments (starting with ‘#’). Unlike m4’s similarlynamed option, this option takes no argument:

               autom4te 1.m4 2.m4 3.m4 --freeze --output=3.m4f

corresponds to

               m4 1.m4 2.m4 3.m4 --freeze-state=3.m4f
‘--mode=octal-mode’
‘-m octal-mode’

Set the mode of the non-traces output to octal-mode; by default ‘0666’.

As another additional feature over m4, autom4te caches its results. GNU M4 is able to produce a regular output and traces at the same time. Traces are heavily used in the GNU Build System: autoheader uses them to build ‘config.h.in’, autoreconf to determine what GNU Build System components are used, automake to “parse” ‘configure.ac’ etc. To avoid recomputation, traces are cached while performing regular expansion, and conversely. This cache is (actually, the caches are) stored in the directory ‘autom4te.cache’. It can safely be removed at any moment (especially if for some reason autom4te considers it trashed).

‘--cache=directory’
‘-C directory’

Specify the name of the directory where the result should be cached. Passing an empty value disables caching. Be sure to pass a relative file name, as for the time being, global caches are not supported.

‘--no-cache’

Don’t cache the results.

‘--force’

‘-f’ If a cache is used, consider it obsolete (but update it anyway).

Because traces are so important to the GNU Build System, autom4te provides high level tracing features as compared to M4, and helps exploiting the cache:

‘--trace=macro[:format]’
‘-t macro[:format]’

Trace the invocations of macro according to the format. Multiple ‘--trace’ arguments can be used to list several macros. Multiple ‘--trace’ arguments for a single macro are not cumulative; instead, you should just make format as long as needed.

The format is a regular string, with newlines if desired, and several special escape codes. It defaults to ‘$f:$l:$n:$%’. It can use the following special escapes:

‘$$’ The character ‘$’.

‘$f’ The file name from which macro is called.

‘$l’ The line number from which macro is called.

‘$d’ The depth of the macro call. This is an M4 technical detail that you probably don’t want to know about.

‘$n’ The name of the macro.

‘$num’ The numth argument of the call to macro.

‘$@’
‘$sep@’
‘${separator}@’

All the arguments passed to macro, separated by the character sep or the string separator (‘,’ by default). Each argument is quoted, i.e., enclosed in a pair of square brackets.

‘$*’
‘$sep*’
‘${separator}*’

As above, but the arguments are not quoted.

‘$%’
‘$sep%’
‘${separator}%’

As above, but the arguments are not quoted, all new line characters in the arguments are smashed, and the default separator is ‘:’. The escape ‘$%’ produces single-line trace outputs (unless you put newlines in the ‘separator’), while ‘$@’ and ‘$*’ do not.

See Section 3.4 [autoconf Invocation], page 13, for examples of trace uses.

‘--preselect=macro’
‘-p macro’

Cache the traces of macro, but do not enable traces. This is especially important to save CPU cycles in the future. For instance, when invoked, autoconf preselects all the macros that autoheader, automake, autoreconf, etc., trace, so that running m4 is not needed to trace them: the cache suffices. This results in a huge speed-up.

Finally, autom4te introduces the concept of Autom4te libraries. They consists in a powerful yet extremely simple feature: sets of combined command line arguments:

‘--language=language’
‘-l language’

Use the language Autom4te library. Current languages include:

M4sugar create M4sugar output.

M4sh create M4sh executable shell scripts.

Autotest create Autotest executable test suites.

          Autoconf-without-aclocal-m4

create Autoconf executable configure scripts without reading

                    ‘aclocal.m4’.

Autoconf create Autoconf executable configure scripts. This language inherits all the characteristics of Autoconf-without-aclocal-m4 and additionally reads ‘aclocal.m4’.

‘--prepend-include=dir’

‘-B dir’ Prepend directory dir to the search path. This is used to include the languagespecific files before any third-party macros.

As an example, if Autoconf is installed in its default location, ‘/usr/local’, the command ‘autom4te -l m4sugar foo.m4’ is strictly equivalent to the command:

     autom4te --prepend-include /usr/local/share/autoconf \
       m4sugar/m4sugar.m4f --warnings syntax foo.m4

Recursive expansion applies here: the command ‘autom4te -l m4sh foo.m4’ is the same as

‘autom4te --language M4sugar m4sugar/m4sh.m4f foo.m4’, i.e.:
     autom4te --prepend-include /usr/local/share/autoconf \
       m4sugar/m4sugar.m4f m4sugar/m4sh.m4f --mode 777 foo.m4

The definition of the languages is stored in ‘autom4te.cfg’.

8.2.2 Customizing autom4te

One can customize autom4te via ‘~/.autom4te.cfg’ (i.e., as found in the user home directory), and ‘./.autom4te.cfg’ (i.e., as found in the directory from which autom4te is run). The order is first reading ‘autom4te.cfg’, then ‘~/.autom4te.cfg’, then ‘./.autom4te.cfg’, and finally the command line arguments.

In these text files, comments are introduced with #, and empty lines are ignored. Customization is performed on a per-language basis, wrapped in between a ‘begin-language:

"language"’, ‘end-language: "language"’ pair.

Customizing a language stands for appending options (see Section 8.2.1 [autom4te Invocation], page 130) to the current definition of the language. Options, and more generally arguments, are introduced by ‘args: arguments’. You may use the traditional shell syntax to quote the arguments.

As an example, to disable Autoconf caches (‘autom4te.cache’) globally, include the following lines in ‘~/.autom4te.cfg’:

## ------------------ ##
## User Preferences.  ##
## ------------------ ##
begin-language: "Autoconf-without-aclocal-m4"
args: --no-cache
end-language: "Autoconf-without-aclocal-m4"

8.3 Programming in M4sugar

M4 by itself provides only a small, but sufficient, set of all-purpose macros. M4sugar introduces additional generic macros. Its name was coined by Lars J. Aas: “Readability And Greater Understanding Stands 4 M4sugar”.

M4sugar reserves the macro namespace ‘^_m4_’ for internal use, and the macro namespace ‘^m4_’ for M4sugar macros. You should not define your own macros into these namespaces.

8.3.1 Redefined M4 Macros

With a few exceptions, all the M4 native macros are moved in the ‘m4_’ pseudo-namespace, e.g., M4sugar renames define as m4_define etc.

The list of macros unchanged from M4, except for their name, is: − m4 builtin − m4 changecom − m4 changequote − m4 debugfile − m4 debugmode − m4 decr − m4 define − m4 divnum − m4 errprint − m4 esyscmd − m4 eval − m4 format − m4 ifdef − m4 incr − m4 index − m4 indir − m4 len − m4 pushdef − m4 shift − m4 substr − m4 syscmd − m4 sysval − m4 traceoff − m4 traceon − m4 translit

Some M4 macros are redefined, and are slightly incompatible with their native equivalent.

[Macro] __file__
[Macro] __line__

All M4 macros starting with ‘__’ retain their original name: for example, no m4__ file__is defined.

[Macro] __oline__

This is not technically a macro, but a feature of Autom4te. The sequence __oline__ can be used similarly to the other m4sugar location macros, but rather than expanding to the location of the input file, it is translated to the line number where it appears in the output file after all other M4 expansions.

[Macro] dnl

This macro kept its original name: no m4_dnl is defined.

[Macro] m4_bpatsubst (string, regexp, [replacement])

This macro corresponds to patsubst. The name m4_patsubst is kept for future versions of M4sugar, once GNU M4 2.0 is released and supports extended regular expression syntax.

[Macro] m4_bregexp (string, regexp, [replacement])

This macro corresponds to regexp. The name m4_regexp is kept for future versions of M4sugar, once GNU M4 2.0 is released and supports extended regular expression syntax.

[Macro] m4_copy (source, dest)
[Macro] m4_copy_force (source, dest)
[Macro] m4_rename (source, dest)
[Macro] m4_rename_force (source, dest)

These macros aren’t directly builtins, but are closely related to m4_pushdef and m4_defn. m4_copy and m4_rename ensure that dest is undefined, while m4_copy_ force and m4_rename_force overwrite any existing definition. All four macros then proceed to copy the entire pushdef stack of definitions of source over to dest. m4_copy and m4_copy_force preserve the source (including in the special case where source is undefined), while m4_rename and m4_rename_force undefine the original macro name (making it an error to rename an undefined source). Note that attempting to invoke a renamed macro might not work, since the macro may have a dependence on helper macros accessed via composition of ‘$0’ but that were not also renamed; likewise, other macros may have a hard-coded dependence on source and could break if source has been deleted. On the other hand, it is always safe to rename a macro to temporarily move it out of the way, then rename it back later to restore original semantics.

[Macro] m4_defn (macro . . .)

This macro fails if macro is not defined, even when using older versions of M4 that did not warn. See m4_undefine. Unfortunately, in order to support these older versions of M4, there are some situations involving unbalanced quotes where concatenating multiple macros together will work in newer M4 but not in m4sugar; use quadrigraphs to work around this.

[Macro] m4_divert (diversion)

M4sugar relies heavily on diversions, so rather than behaving as a primitive, m4_ divert behaves like:

     m4_divert_pop()m4_divert_push([diversion])

See Section 8.3.3 [Diversion support], page 139, for more details about the use of the diversion stack. In particular, this implies that diversion should be a named diversion rather than a raw number. But be aware that it is seldom necessary to explicitly change the diversion stack, and that when done incorrectly, it can lead to syntactically invalid scripts.

[Macro] m4_dumpdef (name . . .)
[Macro] m4_dumpdefs (name . . .)

m4_dumpdef is like the M4 builtin, except that this version requires at least one argument, output always goes to standard error rather than the current debug file, no sorting is done on multiple arguments, and an error is issued if any name is undefined. m4_dumpdefs is a convenience macro that calls m4_dumpdef for all of the m4_pushdef stack of definitions, starting with the current, and silently does nothing if name is undefined. Unfortunately, due to a limitation in M4 1.4.x, any macro defined as a builtin is output as the empty string. This behavior is rectified by using M4 1.6 or newer. However, this behavior difference means that m4_dumpdef should only be used while developing m4sugar macros, and never in the final published form of a macro.

[Macro] m4_esyscmd_s (command)

Like m4_esyscmd, this macro expands to the result of running command in a shell. The difference is that any trailing newlines are removed, so that the output behaves more like shell command substitution.

[Macro] m4_exit (exit-status)

This macro corresponds to m4exit.

[Macro] m4_if (comment)
[Macro] m4_if (string-1, string-2, equal, [not-equal])
[Macro] m4_if (string-1, string-2, equal-1, string-3, string-4, equal-2,
   . . ., [not-equal])

This macro corresponds to ifelse. string-1 and string-2 are compared literally, so usually one of the two arguments is passed unquoted. See Section 8.3.4 [Conditional constructs], page 142, for more conditional idioms.

[Macro] m4_include (file)
[Macro] m4_sinclude (file)

Like the M4 builtins, but warn against multiple inclusions of file.

[Macro] m4_mkstemp (template)
[Macro] m4_maketemp (template)

Posix requires maketemp to replace the trailing ‘X’ characters in template with the process id, without regards to the existence of a file by that name, but this a security hole. When this was pointed out to the Posix folks, they agreed to invent a new macro mkstemp that always creates a uniquely named file, but not all versions of GNU M4 support the new macro. In M4sugar, m4_maketemp and m4_mkstemp are synonyms for each other, and both have the secure semantics regardless of which macro the underlying M4 provides.

[Macro] m4_popdef (macro . . .)

This macro fails if macro is not defined, even when using older versions of M4 that did not warn. See m4_undefine.

[Macro] m4_undefine (macro . . .)

This macro fails if macro is not defined, even when using older versions of M4 that did not warn. Use

     m4_ifdef([macro], [m4_undefine([macro])])

if you are not sure whether macro is defined.

[Macro] m4_undivert (diversion . . .)

Unlike the M4 builtin, at least one diversion must be specified. Also, since the M4sugar diversion stack prefers named diversions, the use of m4_undivert to include files is risky. See Section 8.3.3 [Diversion support], page 139, for more details about the use of the diversion stack. But be aware that it is seldom necessary to explicitly change the diversion stack, and that when done incorrectly, it can lead to syntactically invalid scripts.

[Macro] m4_wrap (text)
[Macro] m4_wrap_lifo (text)

These macros correspond to m4wrap. Posix requires arguments of multiple wrap calls to be reprocessed at EOF in the same order as the original calls (first-in, first-out). GNU M4 versions through 1.4.10, however, reprocess them in reverse order (last-in, first-out). Both orders are useful, therefore, you can rely on m4_wrap to provide FIFO semantics and m4_wrap_lifo for LIFO semantics, regardless of the underlying GNU M4 version. Unlike the GNU M4 builtin, these macros only recognize one argument, and avoid token pasting between consecutive invocations. On the other hand, nested calls to m4_wrap from within wrapped text work just as in the builtin.

8.3.2 Diagnostic messages from M4sugar

When macros statically diagnose abnormal situations, benign or fatal, they should report them using these macros. For issuing dynamic issues, i.e., when configure is run, see Section 7.5 [Printing Messages], page 118.

[Macro] m4_assert (expression, [exit-status = ‘1’])

Assert that the arithmetic expression evaluates to non-zero. Otherwise, issue a fatal error, and exit autom4te with exit-status.

[Macro] m4_errprintn (message)

Similar to the builtin m4_errprint, except that a newline is guaranteed after message.

[Macro] m4_fatal (message)

Report a severe error message prefixed with the current location, and have autom4te die.

[Macro] m4_location

Useful as a prefix in a message line. Short for:

          __file__:__line__
[Macro] m4_warn (category, message)

Report message as a warning (or as an error if requested by the user) if warnings of the category are turned on. If the message is emitted, it is prefixed with the current location, and followed by a call trace of all macros defined via AC_DEFUN used to get to the current expansion. You are encouraged to use standard categories, which currently include:

‘all’ messages that don’t fall into one of the following categories. Use of an empty category is equivalent.

‘cross’ related to cross compilation issues.

     ‘obsolete’

use of an obsolete construct.

‘syntax’ dubious syntactic constructs, incorrectly ordered macro calls.

8.3.3 Diversion support

M4sugar makes heavy use of diversions under the hood, because it is often the case that text that must appear early in the output is not discovered until late in the input. Additionally, some of the topological sorting algorithms used in resolving macro dependencies use diversions. However, most macros should not need to change diversions directly, but rather rely on higher-level M4sugar macros to manage diversions transparently. If you change diversions improperly, you risk generating a syntactically invalid script, because an incorrect diversion will violate assumptions made by many macros about whether prerequisite text has been previously output. In short, if you manually change the diversion, you should not expect any macros provided by the Autoconf package to work until you have restored the diversion stack back to its original state.

In the rare case that it is necessary to write a macro that explicitly outputs text to a different diversion, it is important to be aware of an M4 limitation regarding diversions: text only goes to a diversion if it is not part of argument collection. Therefore, any macro that changes the current diversion cannot be used as an unquoted argument to another macro, but must be expanded at the top level. The macro m4_expand will diagnose any attempt to change diversions, since it is generally useful only as an argument to another macro. The following example shows what happens when diversion manipulation is attempted within macro arguments:

     m4_do([normal text]
     m4_divert_push([KILL])unwanted[]m4_divert_pop([KILL])
     [m4_divert_push([KILL])discarded[]m4_divert_pop([KILL])])dnl
     ⇒normal text
     ⇒unwanted

Notice that the unquoted text unwanted is output, even though it was processed while the current diversion was KILL, because it was collected as part of the argument to m4_do. However, the text discarded disappeared as desired, because the diversion changes were single-quoted, and were not expanded until the top-level rescan of the output of m4_do.

To make diversion management easier, M4sugar uses the concept of named diversions. Rather than using diversion numbers directly, it is nicer to associate a name with each diversion. The diversion number associated with a particular diversion name is an implementation detail, and a syntax warning is issued if a diversion number is used instead of a name. In general, you should not output text to a named diversion until after calling the appropriate initialization routine for your language (m4_init, AS_INIT, AT_INIT, …), although there are some exceptions documented below.

M4sugar defines two named diversions.

KILL Text written to this diversion is discarded. This is the default diversion once M4sugar is initialized.

GROW This diversion is used behind the scenes by topological sorting macros, such as

          AC_REQUIRE.

M4sh adds several more named diversions.

BINSH This diversion is reserved for the ‘#!’ interpreter line.

HEADER-REVISION

This diversion holds text from AC_REVISION.

HEADER-COMMENT

This diversion holds comments about the purpose of a file.

HEADER-COPYRIGHT

This diversion is managed by AC_COPYRIGHT.

M4SH-SANITIZE

This diversion contains M4sh sanitization code, used to ensure M4sh is executing in a reasonable shell environment.

M4SH-INIT

This diversion contains M4sh initialization code, initializing variables that are required by other M4sh macros.

BODY This diversion contains the body of the shell code, and is the default diversion once M4sh is initialized.

Autotest inherits diversions from M4sh, and changes the default diversion from BODY back to KILL. It also adds several more named diversions, with the following subset designed for developer use.

PREPARE_TESTS

This diversion contains initialization sequences which are executed after ‘atconfig’ and ‘atlocal’, and after all command line arguments have been parsed, but prior to running any tests. It can be used to set up state that is required across all tests. This diversion will work even before AT_INIT.

Autoconf inherits diversions from M4sh, and adds the following named diversions which developers can utilize.

DEFAULTS This diversion contains shell variable assignments to set defaults that must be in place before arguments are parsed. This diversion is placed early enough in ‘configure’ that it is unsafe to expand any autoconf macros into this diversion.

HELP_ENABLE

If AC_PRESERVE_HELP_ORDER was used, then text placed in this diversion will be included as part of a quoted here-doc providing all of the ‘--help’ output of

     ‘configure’ related to options created by AC_ARG_WITH and AC_ARG_ENABLE.
INIT_PREPARE

This diversion occurs after all command line options have been parsed, but prior to the main body of the ‘configure’ script. This diversion is the last chance to insert shell code such as variable assignments or shell function declarations that will used by the expansion of other macros.

For now, the remaining named diversions of Autoconf, Autoheader, and Autotest are not documented. In other words, intentionally outputting text into an undocumented diversion is subject to breakage in a future release of Autoconf.

[Macro] m4_cleardivert (diversion . . .)

Permanently discard any text that has been diverted into diversion.

[Macro] m4_divert_once (diversion, [content])

Similar to m4_divert_text, except that content is only output to diversion if this is the first time that m4_divert_once has been called with its particular arguments.

[Macro] m4_divert_pop ([diversion])

If provided, check that the current diversion is indeed diversion. Then change to the diversion located earlier on the stack, giving an error if an attempt is made to pop beyond the initial m4sugar diversion of KILL.

[Macro] m4_divert_push (diversion)

Remember the former diversion on the diversion stack, and output subsequent text into diversion. M4sugar maintains a diversion stack, and issues an error if there is not a matching pop for every push.

[Macro] m4_divert_text (diversion, [content])

Output content and a newline into diversion, without affecting the current diversion. Shorthand for:

     m4_divert_push([diversion])content
     m4_divert_pop([diversion])dnl

One use of m4_divert_text is to develop two related macros, where macro ‘MY_A’ does the work, but adjusts what work is performed based on whether the optional macro ‘MY_B’ has also been expanded. Of course, it is possible to use AC_BEFORE within MY_A to require that ‘MY_B’ occurs first, if it occurs at all. But this imposes an ordering restriction on the user; it would be nicer if macros ‘MY_A’ and ‘MY_B’ can be invoked in either order. The trick is to let ‘MY_B’ leave a breadcrumb in an early diversion, which ‘MY_A’ can then use to determine whether ‘MY_B’ has been expanded.

     AC_DEFUN([MY_A],
     [# various actions
     if test -n "$b_was_used"; then
       # extra action
     fi])

     AC_DEFUN([MY_B],
     [AC_REQUIRE([MY_A])dnl
     m4_divert_text([INIT_PREPARE], [b_was_used=true])])
[Macro] m4_init

Initialize the M4sugar environment, setting up the default named diversion to be

KILL.

8.3.4 Conditional constructs

The following macros provide additional conditional constructs as convenience wrappers around m4_if.

[Macro] m4_bmatch (string, regex-1, value-1, [regex-2], [value-2], . . .,
   [default])

The string string is repeatedly compared against a series of regex arguments; if a match is found, the expansion is the corresponding value, otherwise, the macro moves on to the next regex. If no regex match, then the result is the optional default, or nothing.

[Macro] m4_bpatsubsts (string, regex-1, subst-1, [regex-2], [subst-2], . . .)

The string string is altered by regex-1 and subst-1, as if by:

     m4_bpatsubst([[string]], [regex], [subst])

The result of the substitution is then passed through the next set of regex and subst, and so forth. An empty subst implies deletion of any matched portions in the current string. Note that this macro over-quotes string; this behavior is intentional, so that the result of each step of the recursion remains as a quoted string. However, it means that anchors (‘^’ and ‘$’ in the regex will line up with the extra quotations, and not the characters of the original string. The overquoting is removed after the final substitution.

[Macro] m4_case (string, value-1, if-value-1, [value-2], [if-value-2], . . .,
   [default])

Test string against multiple value possibilities, resulting in the first if-value for a match, or in the optional default. This is shorthand for:

     m4_if([string], [value-1], [if-value-1],
           [string], [value-2], [if-value-2], ...,
           [default])
[Macro] m4_cond (test-1, value-1, if-value-1, [test-2], [value-2],
   [if-value-2], . . ., [default])

This macro was introduced in Autoconf 2.62. Similar to m4_if, except that each test is expanded only when it is encountered. This is useful for short-circuiting expensive tests; while m4_if requires all its strings to be expanded up front before doing comparisons, m4_cond only expands a test when all earlier tests have failed. For an example, these two sequences give the same result, but in the case where ‘$1’ does not contain a backslash, the m4_cond version only expands m4_index once, instead of five times, for faster computation if this is a common case for ‘$1’. Notice that every third argument is unquoted for m4_if, and quoted for m4_cond:

m4_if(m4_index([$1], [\]), [-1], [$2],
      m4_eval(m4_index([$1], [\\]) >= 0), [1], [$2],
      m4_eval(m4_index([$1], [\$]) >= 0), [1], [$2],
      m4_eval(m4_index([$1], [\`]) >= 0), [1], [$3],
      m4_eval(m4_index([$1], [\"]) >= 0), [1], [$3],
      [$2])
m4_cond([m4_index([$1], [\])], [-1], [$2],
        [m4_eval(m4_index([$1], [\\]) >= 0)], [1], [$2],
        [m4_eval(m4_index([$1], [\$]) >= 0)], [1], [$2],
        [m4_eval(m4_index([$1], [\`]) >= 0)], [1], [$3],
        [m4_eval(m4_index([$1], [\"]) >= 0)], [1], [$3],
        [$2])
[Macro] m4_default (expr-1, expr-2)
[Macro] m4_default_quoted (expr-1, expr-2)
[Macro] m4_default_nblank (expr-1, [expr-2])
[Macro] m4_default_nblank_quoted (expr-1, [expr-2])

If expr-1 contains text, use it. Otherwise, select expr-2. m4_default expands the result, while m4_default_quoted does not. Useful for providing a fixed default if the expression that results in expr-1 would otherwise be empty. The difference between m4_default and m4_default_nblank is whether an argument consisting of just blanks (space, tab, newline) is significant. When using the expanding versions, note that an argument may contain text but still expand to an empty string.

m4_define([active], [ACTIVE])dnl
m4_define([empty], [])dnl
m4_define([demo1], [m4_default([$1], [$2])])dnl
m4_define([demo2], [m4_default_quoted([$1], [$2])])dnl
m4_define([demo3], [m4_default_nblank([$1], [$2])])dnl
m4_define([demo4], [m4_default_nblank_quoted([$1], [$2])])dnl
demo1([active], [default])
⇒ACTIVE
demo1([], [active])
⇒ACTIVE
demo1([empty], [text])

⇒

-demo1([ ], [active])-
⇒- -
demo2([active], [default])
⇒active
demo2([], [active])
⇒active
demo2([empty], [text])
⇒empty
-demo2([ ], [active])-
⇒- -
demo3([active], [default])
⇒ACTIVE

     demo3([], [active])
     ⇒ACTIVE
     demo3([empty], [text])

⇒

     -demo3([ ], [active])-
     ⇒-ACTIVE-
     demo4([active], [default])
     ⇒active
     demo4([], [active])
     ⇒active
     demo4([empty], [text])
     ⇒empty
     -demo4([ ], [active])-
     ⇒-active-
[Macro] m4_ifblank (cond, [if-blank], [if-text])
[Macro] m4_ifnblank (cond, [if-text], [if-blank])

If cond is empty or consists only of blanks (space, tab, newline), then expand if-blank; otherwise, expand if-text. Two variants exist, in order to make it easier to select the correct logical sense when using only two parameters. Note that this is more efficient than the equivalent behavior of:

     m4_ifval(m4_normalize([cond]), if-text, if-blank)
[Macro] m4_ifndef (macro, if-not-defined, [if-defined])

This is shorthand for:

     m4_ifdef([macro], [if-defined], [if-not-defined])
[Macro] m4_ifset (macro, [if-true], [if-false])

If macro is undefined, or is defined as the empty string, expand to if-false. Otherwise, expands to if-true. Similar to:

     m4_ifval(m4_defn([macro]), [if-true], [if-false])

except that it is not an error if macro is undefined.

[Macro] m4_ifval (cond, [if-true], [if-false])

Expands to if-true if cond is not empty, otherwise to if-false. This is shorthand for:

     m4_if([cond], [], [if-true], [if-false])
[Macro] m4_ifvaln (cond, [if-true], [if-false])

Similar to m4_ifval, except guarantee that a newline is present after any non-empty expansion. Often followed by dnl.

[Macro] m4_n (text)

Expand to text, and add a newline if text is not empty. Often followed by dnl.

8.3.5 Looping constructs

The following macros are useful in implementing recursive algorithms in M4, including loop operations. An M4 list is formed by quoting a list of quoted elements; generally the lists are comma-separated, although m4_foreach_w is whitespace-separated. For example, the list ‘[[a], [b,c]]’ contains two elements: ‘[a]’ and ‘[b,c]’. It is common to see lists with unquoted elements when those elements are not likely to be macro names, as in

‘[fputc_unlocked, fgetc_unlocked]’.

Although not generally recommended, it is possible for quoted lists to have side effects; all side effects are expanded only once, and prior to visiting any list element. On the other hand, the fact that unquoted macros are expanded exactly once means that macros without side effects can be used to generate lists. For example,

m4_foreach([i], [[1], [2], [3]m4_errprintn([hi])], [i])

error hi

⇒123
m4_define([list], [[1], [2], [3]])

⇒

m4_foreach([i], [list], [i])
⇒123

[Macro] m4_argn (n, [arg]…)

Extracts argument n (larger than 0) from the remaining arguments. If there are too few arguments, the empty string is used. For any n besides 1, this is more efficient

than the similar ‘m4_car(m4_shiftn([n], [], [arg...]))’.

[Macro] m4_car (arg …)

Expands to the quoted first arg. Can be used with m4_cdr to recursively iterate through a list. Generally, when using quoted lists of quoted elements, m4_car should be called without any extra quotes.

[Macro] m4_cdr (arg …)

Expands to a quoted list of all but the first arg, or the empty string if there was only one argument. Generally, when using quoted lists of quoted elements, m4_cdr should be called without any extra quotes.

For example, this is a simple implementation of m4_map; note how each iteration checks for the end of recursion, then merely applies the first argument to the first element of the list, then repeats with the rest of the list. (The actual implementation in M4sugar is a bit more involved, to gain some speed and share code with m4_map_sep, and also to avoid expanding side effects in ‘$2’ twice).

     m4_define([m4_map], [m4_ifval([$2],
       [m4_apply([$1], m4_car($2))[]$0([$1], m4_cdr($2))])])dnl
     m4_map([ m4_eval], [[[1]], [[1+1]], [[10],[16]]])
     ⇒ 1 2 a
[Macro] m4_for (var, first, last, [step], expression)

Loop over the numeric values between first and last including bounds by increments of step. For each iteration, expand expression with the numeric value assigned to var. If step is omitted, it defaults to ‘1’ or ‘-1’ depending on the order of the limits. If given, step has to match this order. The number of iterations is determined independently from definition of var; iteration cannot be short-circuited or lengthened by modifying var from within expression.

[Macro] m4_foreach (var, list, expression)

Loop over the comma-separated M4 list list, assigning each value to var, and expand expression. The following example outputs two lines:

     m4_foreach([myvar], [[foo], [bar, baz]],
                [echo myvar
     ])dnl
     ⇒echo foo
     ⇒echo bar, baz

Note that for some forms of expression, it may be faster to use m4_map_args.

[Macro] m4_foreach_w (var, list, expression)

Loop over the white-space-separated list list, assigning each value to var, and expand expression. If var is only referenced once in expression, it is more efficient to use

m4_map_args_w.

The deprecated macro AC_FOREACH is an alias of m4_foreach_w.

[Macro] m4_map (macro, list)
[Macro] m4_mapall (macro, list)
[Macro] m4_map_sep (macro, separator, list)
[Macro] m4_mapall_sep (macro, separator, list)

Loop over the comma separated quoted list of argument descriptions in list, and invoke macro with the arguments. An argument description is in turn a commaseparated quoted list of quoted elements, suitable for m4_apply. The macros m4_ map and m4_map_sep ignore empty argument descriptions, while m4_mapall and m4_ mapall_sep invoke macro with no arguments. The macros m4_map_sep and m4_ mapall_sep additionally expand separator between invocations of macro. Note that separator is expanded, unlike in m4_join. When separating output with commas, this means that the map result can be used as a series of arguments, by using a single-quoted comma as separator, or as a single string, by using a double-quoted comma.

     m4_map([m4_count], [])

⇒

     m4_map([ m4_count], [[],
                          [[1]],
                          [[1], [2]]])
     ⇒ 1 2
     m4_mapall([ m4_count], [[],
                             [[1]],
                             [[1], [2]]])
     ⇒ 0 1 2
     m4_map_sep([m4_eval], [,], [[[1+2]],
                                 [[10], [16]]])
     ⇒3,a
     m4_map_sep([m4_echo], [,], [[[a]], [[b]]])
     ⇒a,b
     m4_count(m4_map_sep([m4_echo], [,], [[[a]], [[b]]]))

⇒2
m4_map_sep([m4_echo], [[,]], [[[a]], [[b]]])
⇒a,b
m4_count(m4_map_sep([m4_echo], [[,]], [[[a]], [[b]]]))
⇒1
[Macro] m4_map_args (macro, arg . . .)

Repeatedly invoke macro with each successive arg as its only argument. In the following example, three solutions are presented with the same expansion; the solution using m4_map_args is the most efficient.

m4_define([active], [ACTIVE])dnl
m4_foreach([var], [[plain], [active]], [ m4_echo(m4_defn([var]))])
⇒ plain active
m4_map([ m4_echo], [[[plain]], [[active]]])
⇒ plain active
m4_map_args([ m4_echo], [plain], [active])
⇒ plain active

In cases where it is useful to operate on additional parameters besides the list elements, the macro m4_curry can be used in macro to supply the argument currying necessary to generate the desired argument list. In the following example, list_add_n is more efficient than list_add_x. On the other hand, using m4_map_args_sep can be even more efficient.

m4_define([list], [[1], [2], [3]])dnl
m4_define([add], [m4_eval(([$1]) + ([$2]))])dnl
dnl list_add_n(N, ARG...)
dnl Output a list consisting of each ARG added to N
m4_define([list_add_n],
[m4_shift(m4_map_args([,m4_curry([add], [$1])], m4_shift($@)))])dnl
list_add_n([1], list)
⇒2,3,4
list_add_n([2], list)
⇒3,4,5
m4_define([list_add_x],
[m4_shift(m4_foreach([var], m4_dquote(m4_shift($@)),
  [,add([$1],m4_defn([var]))]))])dnl
list_add_x([1], list)
⇒2,3,4
[Macro] m4_map_args_pair (macro, [macro-end = ‘macro’], arg . . .)

For every pair of arguments arg, invoke macro with two arguments. If there is an odd number of arguments, invoke macro-end, which defaults to macro, with the remaining argument.

m4_map_args_pair([, m4_reverse], [], [1], [2], [3])
⇒, 2, 1, 3
m4_map_args_pair([, m4_reverse], [, m4_dquote], [1], [2], [3])
⇒, 2, 1, [3]
m4_map_args_pair([, m4_reverse], [, m4_dquote], [1], [2], [3], [4])

     ⇒, 2, 1, 4, 3

[Macro] m4_map_args_sep ([pre], [post], [sep], arg . . .)

Expand the sequence pre[arg]post for each argument, additionally expanding sep between arguments. One common use of this macro is constructing a macro call, where the opening and closing parentheses are split between pre and post; in particular,

m4_map_args([macro], [arg]) is equivalent to m4_map_args_sep([macro(], [)],

[], [arg]). This macro provides the most efficient means for iterating over an arbitrary list of arguments, particularly when repeatedly constructing a macro call with more arguments than arg.

[Macro] m4_map_args_w (string, [pre], [post], [sep])

Expand the sequence pre[word]post for each word in the whitespace-separated string, additionally expanding sep between words. This macro provides the most efficient means for iterating over a whitespace-separated string. In particular, m4_map_

args_w([string], [action(], [)]) is more efficient than m4_foreach_w([var],
[string], [action(m4_defn([var]))]).
[Macro] m4_shiftn (count, . . .)

[Macro] m4_shift2 (…)

[Macro] m4_shift3 (…)

m4_shiftn performs count iterations of m4_shift, along with validation that enough arguments were passed in to match the shift count, and that the count is positive. m4_shift2 and m4_shift3 are specializations of m4_shiftn, introduced in Autoconf 2.62, and are more efficient for two and three shifts, respectively.

[Macro] m4_stack_foreach (macro, action)
[Macro] m4_stack_foreach_lifo (macro, action)

For each of the m4_pushdef definitions of macro, expand action with the single argument of a definition of macro. m4_stack_foreach starts with the oldest definition, while m4_stack_foreach_lifo starts with the current definition. action should not push or pop definitions of macro, nor is there any guarantee that the current definition of macro matches the argument that was passed to action. The macro m4_curry can be used if action needs more than one argument, although in that case it is more efficient to use m4 stack foreach sep. Due to technical limitations, there are a few low-level m4sugar functions, such as m4_pushdef, that cannot be used as the macro argument.

     m4_pushdef([a], [1])m4_pushdef([a], [2])dnl
     m4_stack_foreach([a], [ m4_incr])
     ⇒ 2 3
     m4_stack_foreach_lifo([a], [ m4_curry([m4_substr], [abcd])])
     ⇒ cd bcd
[Macro] m4_stack_foreach_sep (macro, [pre], [post], [sep])
[Macro] m4_stack_foreach_sep_lifo (macro, [pre], [post], [sep])

Expand the sequence pre[definition]post for each m4_pushdef definition of macro, additionally expanding sep between definitions. m4_stack_foreach_sep visits the oldest definition first, while m4_stack_foreach_sep_lifo visits the current definition first. This macro provides the most efficient means for iterating

over a pushdef stack. In particular, m4_stack_foreach([macro], [action]) is
short for m4_stack_foreach_sep([macro], [action(], [)]).

8.3.6 Evaluation Macros

The following macros give some control over the order of the evaluation by adding or removing levels of quotes.

[Macro] m4_apply (macro, list)

Apply the elements of the quoted, comma-separated list as the arguments to macro. If list is empty, invoke macro without arguments. Note the difference between m4_ indir, which expects its first argument to be a macro name but can use names that are otherwise invalid, and m4_apply, where macro can contain other text, but must end in a valid macro name.

     m4_apply([m4_count], [])
     ⇒0
     m4_apply([m4_count], [[]])
     ⇒1
     m4_apply([m4_count], [[1], [2]])
     ⇒2
     m4_apply([m4_join], [[|], [1], [2]])
     ⇒1|2

[Macro] m4_count (arg, …)

This macro returns the decimal count of the number of arguments it was passed.

[Macro] m4_curry (macro, arg . . .)

This macro performs argument currying. The expansion of this macro is another macro name that expects exactly one argument; that argument is then appended to the arg list, and then macro is expanded with the resulting argument list.

     m4_curry([m4_curry], [m4_reverse], [1])([2])([3])
     ⇒3, 2, 1

Unfortunately, due to a limitation in M4 1.4.x, it is not possible to pass the definition of a builtin macro as the argument to the output of m4_curry; the empty string is used instead of the builtin token. This behavior is rectified by using M4 1.6 or newer.

[Macro] m4_do (arg, …)

This macro loops over its arguments and expands each arg in sequence. Its main use is for readability; it allows the use of indentation and fewer dnl to result in the same expansion. This macro guarantees that no expansion will be concatenated with subsequent text; to achieve full concatenation, use m4_unquote(m4_join([],

arg...)).
     m4_define([ab],[1])m4_define([bc],[2])m4_define([abc],[3])dnl
     m4_do([a],[b])c
     ⇒abc
     m4_unquote(m4_join([],[a],[b]))c
     ⇒3

     m4_define([a],[A])m4_define([b],[B])m4_define([c],[C])dnl
     m4_define([AB],[4])m4_define([BC],[5])m4_define([ABC],[6])dnl
     m4_do([a],[b])c
     ⇒ABC
     m4_unquote(m4_join([],[a],[b]))c
     ⇒3

[Macro] m4_dquote (arg, …)

Return the arguments as a quoted list of quoted arguments. Conveniently, if there is just one arg, this effectively adds a level of quoting.

[Macro] m4_dquote_elt (arg, . . .)

Return the arguments as a series of double-quoted arguments. Whereas m4_dquote returns a single argument, m4_dquote_elt returns as many arguments as it was passed.

[Macro] m4_echo (arg, …)

Return the arguments, with the same level of quoting. Other than discarding whitespace after unquoted commas, this macro is a no-op.

[Macro] m4_expand (arg)

Return the expansion of arg as a quoted string. Whereas m4_quote is designed to collect expanded text into a single argument, m4_expand is designed to perform one level of expansion on quoted text. One distinction is in the treatment of whitespace following a comma in the original arg. Any time multiple arguments are collected into one with m4_quote, the M4 argument collection rules discard the whitespace. However, with m4_expand, whitespace is preserved, even after the expansion of macros contained in arg. Additionally, m4_expand is able to expand text that would involve an unterminated comment, whereas expanding that same text as the argument to m4_quote runs into difficulty in finding the end of the argument. Since manipulating diversions during argument collection is inherently unsafe, m4_expand issues an error if arg attempts to change the current diversion (see Section 8.3.3 [Diversion support], page 139).

     m4_define([active], [ACT, IVE])dnl
     m4_define([active2], [[ACT, IVE]])dnl
     m4_quote(active, active)
     ⇒ACT,IVE,ACT,IVE
     m4_expand([active, active])
     ⇒ACT, IVE, ACT, IVE
     m4_quote(active2, active2)
     ⇒ACT, IVE,ACT, IVE
     m4_expand([active2, active2])
     ⇒ACT, IVE, ACT, IVE
     m4_expand([# m4_echo])
     ⇒# m4_echo
     m4_quote(# m4_echo)
     )
     ⇒# m4_echo)

⇒ Note that m4_expand cannot handle an arg that expands to literal unbalanced quotes, but that quadrigraphs can be used when unbalanced output is necessary. Likewise, unbalanced parentheses should be supplied with double quoting or a quadrigraph.

     m4_define([pattern], [[!@<:@]])dnl
     m4_define([bar], [BAR])dnl
     m4_expand([case $foo in
       m4_defn([pattern])@:}@ bar ;;
       *[)] blah ;;
     esac])
     ⇒case $foo in
     ⇒ [![]) BAR ;;
     ⇒ *) blah ;;
     ⇒esac

[Macro] m4_ignore (…)

This macro was introduced in Autoconf 2.62. Expands to nothing, ignoring all of its arguments. By itself, this isn’t very useful. However, it can be used to conditionally ignore an arbitrary number of arguments, by deciding which macro name to apply to a list of arguments.

     dnl foo outputs a message only if [debug] is defined.
     m4_define([foo],
     [m4_ifdef([debug],[AC_MSG_NOTICE],[m4_ignore])([debug message])])

Note that for earlier versions of Autoconf, the macro __gnu__can serve the same purpose, although it is less readable.

[Macro] m4_make_list (arg, . . .)

This macro exists to aid debugging of M4sugar algorithms. Its net effect is similar to m4_dquote—it produces a quoted list of quoted arguments, for each arg. The difference is that this version uses a comma-newline separator instead of just comma, to improve readability of the list; with the result that it is less efficient than m4_

dquote.
     m4_define([zero],[0])m4_define([one],[1])m4_define([two],[2])dnl
     m4_dquote(zero, [one], [[two]])
     ⇒[0],[one],[[two]]
     m4_make_list(zero, [one], [[two]])
     ⇒[0],
     ⇒[one],
     ⇒[[two]]
     m4_foreach([number], m4_dquote(zero, [one], [[two]]), [ number])
     ⇒ 0 1 two
     m4_foreach([number], m4_make_list(zero, [one], [[two]]), [ number])
     ⇒ 0 1 two

[Macro] m4_quote (arg, …)

Return the arguments as a single entity, i.e., wrap them into a pair of quotes. This effectively collapses multiple arguments into one, although it loses whitespace after unquoted commas in the process.

[Macro] m4_reverse (arg, . . .)

Outputs each argument with the same level of quoting, but in reverse order, and with space following each comma for readability.

     m4_define([active], [ACT,IVE])

⇒

     m4_reverse(active, [active])
     ⇒active, IVE, ACT
[Macro] m4_unquote (arg, . . .)

This macro was introduced in Autoconf 2.62. Expand each argument, separated by commas. For a single arg, this effectively removes a layer of quoting, and m4_ unquote([arg]) is more efficient than the equivalent m4_do([arg]). For multiple arguments, this results in an unquoted list of expansions. This is commonly used with m4_split, in order to convert a single quoted list into a series of quoted elements.

The following example aims at emphasizing the difference between several scenarios: not using these macros, using m4_defn, using m4_quote, using m4_dquote, and using m4_

expand.
$ cat example.m4
dnl Overquote, so that quotes are visible.
m4_define([show], [$[]1 = [$1], $[]@ = [$@]])
m4_define([a], [A])
m4_define([mkargs], [1, 2[,] 3])
m4_define([arg1], [[$1]])
m4_divert([0])dnl
show(a, b)
show([a, b])
show(m4_quote(a, b))
show(m4_dquote(a, b))
show(m4_expand([a, b]))
arg1(mkargs)
arg1([mkargs])
arg1(m4_defn([mkargs]))
arg1(m4_quote(mkargs))
arg1(m4_dquote(mkargs))
arg1(m4_expand([mkargs]))
$ autom4te -l m4sugar example.m4
$1 = A, $@ = [A],[b]
$1 = a, b, $@ = [a, b]
$1 = A,b, $@ = [A,b]
$1 = [A],[b], $@ = [[A],[b]]
$1 = A, b, $@ = [A, b]
1
mkargs
1, 2[,] 3

1,2, 3
[1],[2, 3]
1, 2, 3

8.3.7 String manipulation in M4

The following macros may be used to manipulate strings in M4. Many of the macros in this section intentionally result in quoted strings as output, rather than subjecting the arguments to further expansions. As a result, if you are manipulating text that contains active M4 characters, the arguments are passed with single quoting rather than double.

[Macro] m4_append (macro-name, string, [separator])
[Macro] m4_append_uniq (macro-name, string, [separator] [if-uniq],
   [if-duplicate])

Redefine macro-name to its former contents with separator and string added at the end. If macro-name was undefined before (but not if it was defined but empty), then no separator is added. As of Autoconf 2.62, neither string nor separator are expanded during this macro; instead, they are expanded when macro-name is invoked.

m4_append can be used to grow strings, and m4_append_uniq to grow strings without duplicating substrings. Additionally, m4_append_uniq takes two optional parameters as of Autoconf 2.62; if-uniq is expanded if string was appended, and if-duplicate is expanded if string was already present. Also, m4_append_uniq warns if separator is not empty, but occurs within string, since that can lead to duplicates.

Note that m4_append can scale linearly in the length of the final string, depending on the quality of the underlying M4 implementation, while m4_append_uniq has an inherent quadratic scaling factor. If an algorithm can tolerate duplicates in the final string, use the former for speed. If duplicates must be avoided, consider using m4_ set_add instead (see Section 8.3.9 [Set manipulation Macros], page 158).

     m4_define([active], [ACTIVE])dnl
     m4_append([sentence], [This is an])dnl
     m4_append([sentence], [ active ])dnl
     m4_append([sentence], [symbol.])dnl
     sentence
     ⇒This is an ACTIVE symbol.
     m4_undefine([active])dnl
     ⇒This is an active symbol.
     m4_append_uniq([list], [one], [, ], [new], [existing])
     ⇒new
     m4_append_uniq([list], [one], [, ], [new], [existing])
     ⇒existing
     m4_append_uniq([list], [two], [, ], [new], [existing])
     ⇒new
     m4_append_uniq([list], [three], [, ], [new], [existing])
     ⇒new
     m4_append_uniq([list], [two], [, ], [new], [existing])
     ⇒existing
     list

⇒one, two, three
m4_dquote(list)
⇒[one],[two],[three]
m4_append([list2], [one], [[, ]])dnl
m4_append_uniq([list2], [two], [[, ]])dnl
m4_append([list2], [three], [[, ]])dnl
list2
⇒one, two, three
m4_dquote(list2)
⇒[one, two, three]
[Macro] m4_append_uniq_w (macro-name, strings)

This macro was introduced in Autoconf 2.62. It is similar to m4_append_uniq, but treats strings as a whitespace separated list of words to append, and only appends unique words. macro-name is updated with a single space between new words.

m4_append_uniq_w([numbers], [1 1 2])dnl
m4_append_uniq_w([numbers], [ 2 3 ])dnl
numbers
⇒1 2 3
[Macro] m4_chomp (string)
[Macro] m4_chomp_all (string)

Output string in quotes, but without a trailing newline. The macro m4_chomp is slightly faster, and removes at most one newline; the macro m4_chomp_all removes all consecutive trailing newlines. Unlike m4_flatten, embedded newlines are left intact, and backslash does not influence the result.

[Macro] m4_combine ([separator], prefix-list, [infix], suffix-1,
[suffix-2], . . .)

This macro produces a quoted string containing the pairwise combination of every element of the quoted, comma-separated prefix-list, and every element from the suffix arguments. Each pairwise combination is joined with infix in the middle, and successive pairs are joined by separator. No expansion occurs on any of the arguments. No output occurs if either the prefix or suffix list is empty, but the lists can contain empty elements.

m4_define([a], [oops])dnl
m4_combine([, ], [[a], [b], [c]], [-], [1], [2], [3])
⇒a-1, a-2, a-3, b-1, b-2, b-3, c-1, c-2, c-3
m4_combine([, ], [[a], [b]], [-])

⇒

m4_combine([, ], [[a], [b]], [-], [])
⇒a-, b-
m4_combine([, ], [], [-], [1], [2])

⇒

m4_combine([, ], [[]], [-], [1], [2])
⇒-1, -2

[Macro] m4_escape (string)

Convert all instances of ‘[’, ‘]’, ‘#’, and ‘$’ within string into their respective quadrigraphs. The result is still a quoted string.

[Macro] m4_flatten (string)

Flatten string into a single line. Delete all backslash-newline pairs, and replace all remaining newlines with a space. The result is still a quoted string.

[Macro] m4_join ([separator], args . . .)
[Macro] m4_joinall ([separator], args . . .)

Concatenate each arg, separated by separator. joinall uses every argument, while join omits empty arguments so that there are no back-to-back separators in the output. The result is a quoted string.

     m4_define([active], [ACTIVE])dnl
     m4_join([|], [one], [], [active], [two])
     ⇒one|active|two
     m4_joinall([|], [one], [], [active], [two])
     ⇒one||active|two

Note that if all you intend to do is join args with commas between them, to form a quoted list suitable for m4_foreach, it is more efficient to use m4_dquote.

[Macro] m4_newline ([text])

This macro was introduced in Autoconf 2.62, and expands to a newline, followed by any text. It is primarily useful for maintaining macro formatting, and ensuring that M4 does not discard leading whitespace during argument collection.

[Macro] m4_normalize (string)

Remove leading and trailing spaces and tabs, sequences of backslash-then-newline, and replace multiple spaces, tabs, and newlines with a single space. This is a combination of m4_flatten and m4_strip. To determine if string consists only of bytes that would be removed by m4_normalize, you can use m4_ifblank.

[Macro] m4_re_escape (string)

Backslash-escape all characters in string that are active in regexps.

[Macro] m4_split (string, [regexp = ‘[\t ]+’])

Split string into an M4 list of elements quoted by ‘[’ and ‘]’, while keeping white space at the beginning and at the end. If regexp is given, use it instead of ‘[\t ]+’ for splitting. If string is empty, the result is an empty list.

[Macro] m4_strip (string)

Strip whitespace from string. Sequences of spaces and tabs are reduced to a single space, then leading and trailing spaces are removed. The result is still a quoted string. Note that this does not interfere with newlines; if you want newlines stripped as well, consider m4_flatten, or do it all at once with m4_normalize. To quickly test if string has only whitespace, use m4_ifblank.

[Macro] m4_text_box (message, [frame = ‘-’])

Add a text box around message, using frame as the border character above and below the message. The frame argument must be a single byte, and does not support quadrigraphs. The frame correctly accounts for the subsequent expansion of message. For example:

     m4_define([macro], [abc])dnl
     m4_text_box([macro])
     ⇒## --- ##
     ⇒## abc ##
     ⇒## --- ##

The message must contain balanced quotes and parentheses, although quadrigraphs can be used to work around this.

[Macro] m4_text_wrap (string, [prefix], [prefix1 = ‘prefix’], [width = ‘79’])

Break string into a series of whitespace-separated words, then output those words separated by spaces, and wrapping lines any time the output would exceed width columns. If given, prefix1 begins the first line, and prefix begins all wrapped lines. If prefix1 is longer than prefix, then the first line consists of just prefix1. If prefix is longer than prefix1, padding is inserted so that the first word of string begins at the same indentation as all wrapped lines. Note that using literal tab characters in any of the arguments will interfere with the calculation of width. No expansions occur on prefix, prefix1, or the words of string, although quadrigraphs are recognized. For some examples:

     m4_text_wrap([Short string */], [   ], [/* ], [20])
     ⇒/* Short string */
     m4_text_wrap([Much longer string */], [   ], [/* ], [20])
     ⇒/* Much longer
     ⇒ string */
     m4_text_wrap([Short doc.], [          ], [  --short ], [30])
     ⇒ --short Short doc.
     m4_text_wrap([Short doc.], [          ], [  --too-wide ], [30])
     ⇒ --too-wide
     ⇒  Short doc.
     m4_text_wrap([Super long documentation.], [     ],
                  [  --too-wide ], 30)
     ⇒ --too-wide
     ⇒  Super long
     ⇒  documentation.
[Macro] m4_tolower (string)
[Macro] m4_toupper (string)

Return string with letters converted to upper or lower case, respectively.

8.3.8 Arithmetic computation in M4

The following macros facilitate integer arithmetic operations. Where a parameter is documented as taking an arithmetic expression, you can use anything that can be parsed by

m4_eval.
[Macro] m4_cmp (expr-1, expr-2)

Compare the arithmetic expressions expr-1 and expr-2, and expand to ‘-1’ if expr-1 is smaller, ‘0’ if they are equal, and ‘1’ if expr-1 is larger.

[Macro] m4_list_cmp (list-1, list-2)

Compare the two M4 lists consisting of comma-separated arithmetic expressions, left to right. Expand to ‘-1’ for the first element pairing where the value from list-1 is smaller, ‘1’ where the value from list-2 is smaller, or ‘0’ if both lists have the same values. If one list is shorter than the other, the remaining elements of the longer list are compared against zero.

     m4_list_cmp([1, 0],       [1])
     ⇒0
     m4_list_cmp([1, [1 * 0]], [1, 0])
     ⇒0
     m4_list_cmp([1, 2],       [1, 0])
     ⇒1
     m4_list_cmp([1, [1+1], 3],[1, 2])
     ⇒1
     m4_list_cmp([1, 2, -3],   [1, 2])
     ⇒-1
     m4_list_cmp([1, 0],       [1, 2])
     ⇒-1
     m4_list_cmp([1],          [1, 2])
     ⇒-1

[Macro] m4_max (arg, …)

This macro was introduced in Autoconf 2.62. Expand to the decimal value of the maximum arithmetic expression among all the arguments.

[Macro] m4_min (arg, …)

This macro was introduced in Autoconf 2.62. Expand to the decimal value of the minimum arithmetic expression among all the arguments.

[Macro] m4_sign (expr)

Expand to ‘-1’ if the arithmetic expression expr is negative, ‘1’ if it is positive, and ‘0’ if it is zero.

[Macro] m4_version_compare (version-1, version-2)

This macro was introduced in Autoconf 2.53, but had a number of usability limitations that were not lifted until Autoconf 2.62. Compare the version strings version-1 and version-2, and expand to ‘-1’ if version-1 is smaller, ‘0’ if they are the same, or ‘1’ version-2 is smaller. Version strings must be a list of elements separated by ‘.’, ‘,’ or ‘-’, where each element is a number along with optional case-insensitive letters designating beta releases. The comparison stops at the leftmost element that contains a difference, although a 0 element compares equal to a missing element.

It is permissible to include commit identifiers in version, such as an abbreviated SHA1 of the commit, provided there is still a monotonically increasing prefix to allow for accurate version-based comparisons. For example, this paragraph was written when the development snapshot of autoconf claimed to be at version ‘2.61a-248-dc51’, or 248 commits after the 2.61a release, with an abbreviated commit identification of

‘dc51’.

     m4_version_compare([1.1], [2.0])
     ⇒-1
     m4_version_compare([2.0b], [2.0a])
     ⇒1
     m4_version_compare([1.1.1], [1.1.1a])
     ⇒-1
     m4_version_compare([1.2], [1.1.1a])
     ⇒1
     m4_version_compare([1.0], [1])
     ⇒0
     m4_version_compare([1.1pre], [1.1PRE])
     ⇒0
     m4_version_compare([1.1a], [1,10])
     ⇒-1
     m4_version_compare([2.61a], [2.61a-248-dc51])
     ⇒-1
     m4_version_compare([2.61b], [2.61a-248-dc51])
     ⇒1
[Macro] m4_version_prereq (version, [if-new-enough], [if-old =
   ‘m4_fatal’])

Compares version against the version of Autoconf currently running. If the running version is at version or newer, expand if-new-enough, but if version is larger than the version currently executing, expand if-old, which defaults to printing an error message and exiting m4sugar with status 63. When given only one argument, this behaves like AC_PREREQ (see Section 4.2 [Versioning], page 20). Remember that the autoconf philosophy favors feature checks over version checks.

8.3.9 Set manipulation in M4

Sometimes, it is necessary to track a set of data, where the order does not matter and where there are no duplicates in the set. The following macros facilitate set manipulations. Each set is an opaque object, which can only be accessed via these basic operations. The underlying implementation guarantees linear scaling for set creation, which is more efficient than using the quadratic m4_append_uniq. Both set names and values can be arbitrary strings, except for unbalanced quotes. This implementation ties up memory for removed elements until the next operation that must traverse all the elements of a set; and although that may slow down some operations until the memory for removed elements is pruned, it still guarantees linear performance.

[Macro] m4_set_add (set, value, [if-uniq], [if-dup])

Adds the string value as a member of set set. Expand if-uniq if the element was added, or if-dup if it was previously in the set. Operates in amortized constant time, so that set creation scales linearly.

[Macro] m4_set_add_all (set, value . . .)

Adds each value to the set set. This is slightly more efficient than repeatedly invoking

m4_set_add.

[Macro] m4_set_contains (set, value, [if-present], [if-absent])

Expands if-present if the string value is a member of set, otherwise if-absent.

m4_set_contains([a], [1], [yes], [no])
⇒no
m4_set_add([a], [1], [added], [dup])
⇒added
m4_set_add([a], [1], [added], [dup])
⇒dup
m4_set_contains([a], [1], [yes], [no])
⇒yes
m4_set_remove([a], [1], [removed], [missing])
⇒removed
m4_set_contains([a], [1], [yes], [no])
⇒no
m4_set_remove([a], [1], [removed], [missing])
⇒missing
[Macro] m4_set_contents (set, [sep])
[Macro] m4_set_dump (set, [sep])

Expands to a single string consisting of all the members of the set set, each separated by sep, which is not expanded. m4_set_contents leaves the elements in set but reclaims any memory occupied by removed elements, while m4_set_dump is a faster oneshot action that also deletes the set. No provision is made for disambiguating members that contain a non-empty sep as a substring; use m4_set_empty to distinguish between an empty set and the set containing only the empty string. The order of the output is unspecified; in the current implementation, part of the speed of m4_set_dump results from using a different output order than m4_set_contents. These macros scale linearly in the size of the set before memory pruning, and m4_set_contents([set],

[sep]) is faster than m4_joinall([sep]m4_set_listc([set])).
m4_set_add_all([a], [1], [2], [3])

⇒

m4_set_contents([a], [-])
⇒1-2-3
m4_joinall([-]m4_set_listc([a]))
⇒1-2-3
m4_set_dump([a], [-])
⇒3-2-1
m4_set_contents([a])

⇒

m4_set_add([a], [])

⇒

m4_set_contents([a], [-])

⇒

[Macro] m4_set_delete (set)

Delete all elements and memory associated with set. This is linear in the set size, and faster than removing one element at a time.

[Macro] m4_set_difference (seta, setb)
[Macro] m4_set_intersection (seta, setb)
[Macro] m4_set_union (seta, setb)

Compute the relation between seta and setb, and output the result as a list of quoted arguments without duplicates and with a leading comma. Set difference selects the elements in seta but not setb, intersection selects only elements in both sets, and union selects elements in either set. These actions are linear in the sum of the set sizes. The leading comma is necessary to distinguish between no elements and the empty string as the only element.

     m4_set_add_all([a], [1], [2], [3])

⇒

     m4_set_add_all([b], [3], [], [4])

⇒

     m4_set_difference([a], [b])
     ⇒,1,2
     m4_set_difference([b], [a])
     ⇒,,4
     m4_set_intersection([a], [b])
     ⇒,3
     m4_set_union([a], [b])
     ⇒,1,2,3,,4
[Macro] m4_set_empty (set, [if-empty], [if-elements])

Expand if-empty if the set set has no elements, otherwise expand if-elements. This macro operates in constant time. Using this macro can help disambiguate output

from m4_set_contents or m4_set_list.
[Macro] m4_set_foreach (set, variable, action)

For each element in the set set, expand action with the macro variable defined as the set element. Behavior is unspecified if action recursively lists the contents of set (although listing other sets is acceptable), or if it modifies the set in any way other than removing the element currently contained in variable. This macro is faster

than the corresponding m4_foreach([variable], m4_indir([m4_dquote]m4_set_
listc([set])), [action]), although m4_set_map might be faster still.
     m4_set_add_all([a]m4_for([i], [1], [5], [], [,i]))

⇒

     m4_set_contents([a])
     ⇒12345
     m4_set_foreach([a], [i],
       [m4_if(m4_eval(i&1), [0], [m4_set_remove([a], i, [i])])])
     ⇒24
     m4_set_contents([a])
     ⇒135
[Macro] m4_set_list (set)
[Macro] m4_set_listc (set)

Produce a list of arguments, where each argument is a quoted element from the set set. The variant m4_set_listc is unambiguous, by adding a leading comma if there are any set elements, whereas the variant m4_set_list cannot distinguish between an empty set and a set containing only the empty string. These can be directly used in macros that take multiple arguments, such as m4_join or m4_set_add_all, or wrapped by m4_dquote for macros that take a quoted list, such as m4_map or m4_foreach. Any memory occupied by removed elements is reclaimed during these macros.

     m4_set_add_all([a], [1], [2], [3])

⇒

     m4_set_list([a])
     ⇒1,2,3
     m4_set_list([b])

⇒

     m4_set_listc([b])

⇒

     m4_count(m4_set_list([b]))
     ⇒1
     m4_set_empty([b], [0], [m4_count(m4_set_list([b]))])
     ⇒0
     m4_set_add([b], [])

⇒

     m4_set_list([b])

⇒

     m4_set_listc([b])
     ⇒,
     m4_count(m4_set_list([b]))
     ⇒1
     m4_set_empty([b], [0], [m4_count(m4_set_list([b]))])
     ⇒1
[Macro] m4_set_map (set, action)

For each element in the set set, expand action with a single argument of the set element. Behavior is unspecified if action recursively lists the contents of set (although listing other sets is acceptable), or if it modifies the set in any way other than removing the element passed as an argument. This macro is faster than ei-

ther corresponding counterpart of m4_map_args([action]m4_set_listc([set]))
or m4_set_foreach([set], [var], [action(m4_defn([var]))]). It is possible to

use m4_curry if more than one argument is needed for action, although it is more efficient to use m4_set_map_sep in that case.

[Macro] m4_set_map_sep (set, [pre], [post], [sep])

For each element in the set set, expand pre[element]post, additionally expanding sep between elements. Behavior is unspecified if the expansion recursively lists the contents of set (although listing other sets is acceptable), or if it modifies the set in any way other than removing the element visited by the expansion. This macro provides the most efficient means for non-destructively visiting the elements of a set; in par-

ticular, m4_set_map([set], [action]) is equivalent to m4_set_map_sep([set],
[action(], [)]).

[Macro] m4_set_remove (set, value, [if-present], [if-absent])

If value is an element in the set set, then remove it and expand if-present. Otherwise expand if-absent. This macro operates in constant time so that multiple removals will scale linearly rather than quadratically; but when used outside of m4_set_foreach or m4_set_map, it leaves memory occupied until the set is later compacted by m4_ set_contents or m4_set_list. Several other set operations are then less efficient between the time of element removal and subsequent memory compaction, but still maintain their guaranteed scaling performance.

[Macro] m4_set_size (set)

Expand to the size of the set set. This implementation operates in constant time,

and is thus more efficient than m4_eval(m4_count(m4_set_listc([set])) - 1).

8.3.10 Forbidden Patterns

M4sugar provides a means to define suspicious patterns, patterns describing tokens which should not be found in the output. For instance, if an Autoconf ‘configure’ script includes tokens such as ‘AC_DEFINE’, or ‘dnl’, then most probably something went wrong (typically a macro was not evaluated because of overquotation).

M4sugar forbids all the tokens matching ‘^_?m4_’ and ‘^dnl$’. Additional layers, such as M4sh and Autoconf, add additional forbidden patterns to the list.

[Macro] m4_pattern_forbid (pattern)

Declare that no token matching pattern must be found in the output. Comments are not checked; this can be a problem if, for instance, you have some macro left unexpanded after an ‘#include’. No consensus is currently found in the Autoconf community, as some people consider it should be valid to name macros in comments (which doesn’t make sense to the authors of this documentation: input, such as macros, should be documented by ‘dnl’ comments; reserving ‘#’-comments to document the output).

Of course, you might encounter exceptions to these generic rules, for instance you might have to refer to ‘$m4_flags’.

[Macro] m4_pattern_allow (pattern)

Any token matching pattern is allowed, including if it matches an m4_pattern_forbid pattern.

8.4 Debugging via autom4te

At times, it is desirable to see what was happening inside m4, to see why output was not matching expectations. However, post-processing done by autom4te means that directly using the m4 builtin m4_traceon is likely to interfere with operation. Also, frequent diversion changes and the concept of forbidden tokens make it difficult to use m4_defn to generate inline comments in the final output.

There are a couple of tools to help with this. One is the use of the ‘--trace’ option provided by autom4te (as well as each of the programs that wrap autom4te, such as autoconf), in order to inspect when a macro is called and with which arguments. For example, when this paragraph was written, the autoconf version could be found by:

$ autoconf --trace=AC_INIT
configure.ac:23:AC_INIT:GNU Autoconf:2.63b.95-3963:bug-autoconf@gnu.org
$ autoconf --trace='AC_INIT:version is $2'
version is 2.63b.95-3963

Another trick is to print out the expansion of various m4 expressions to standard error or to an independent file, with no further m4 expansion, and without interfering with diversion changes or the post-processing done to standard output. m4_errprintn shows a given expression on standard error. For example, if you want to see the expansion of an autoconf primitive or of one of your autoconf macros, you can do it like this:

$ cat <<\EOF > configure.ac
AC_INIT
m4_errprintn([The definition of AC_DEFINE_UNQUOTED:])
m4_errprintn(m4_defn([AC_DEFINE_UNQUOTED]))
AC_OUTPUT
EOF
$ autoconf
 error The definition of AC_DEFINE_UNQUOTED:
 error _AC_DEFINE_Q([], $@)

9 Programming in M4sh

M4sh, pronounced “mash”, is aiming at producing portable Bourne shell scripts. This name was coined by Lars J. Aas, who notes that, according to the Webster’s Revised Unabridged Dictionary (1913):

Mash \Mash\, n. [Akin to G. meisch, maisch, meische, maische, mash, wash, and prob. to AS. miscian to mix. See “Mix”.]

1. A mass of mixed ingredients reduced to a soft pulpy state by beating or pressure…

2. A mixture of meal or bran and water fed to animals.

3. A mess; trouble. [Obs.] –Beau. & Fl.

M4sh reserves the M4 macro namespace ‘^_AS_’ for internal use, and the namespace ‘^AS_’ for M4sh macros. It also reserves the shell and environment variable namespace ‘^as_’, and the here-document delimiter namespace ‘^_AS[A-Z]’ in the output file. You should not define your own macros or output shell code that conflicts with these namespaces.

9.1 Common Shell Constructs

M4sh provides portable alternatives for some common shell constructs that unfortunately are not portable in practice.

[Macro] AS_BOX (text, [char = ‘-’])

Expand into shell code that will output text surrounded by a box with char in the top and bottom border. text should not contain a newline, but may contain shell expansions valid for unquoted here-documents. char defaults to ‘-’, but can be any character except ‘/’, ‘'’, ‘"’, ‘\’, ‘&’, or ‘`’. This is useful for outputting a comment box into log files to separate distinct phases of script operation.

[Macro] AS_CASE (word, [pattern1], [if-matched1], . . ., [default])

Expand into a shell ‘case’ statement, where word is matched against one or more patterns. if-matched is run if the corresponding pattern matched word, else default is run. Avoids several portability issues (see [Limitations of Shell Builtins], page 212).

[Macro] AS_DIRNAME (file-name)

Output the directory portion of file-name. For example, if $file is ‘/one/two/three’,

the command dir=`AS_DIRNAME(["$file"])` sets dir to ‘/one/two’.

This interface may be improved in the future to avoid forks and losing trailing newlines.

[Macro] AS_ECHO (word)

Emits word to the standard output, followed by a newline. word must be a single shell word (typically a quoted string). The bytes of word are output as-is, even if it starts with "-" or contains "\". Redirections can be placed outside the macro invocation. This is much more portable than using echo (see [Limitations of Shell Builtins], page 213).

[Macro] AS_ECHO_N (word)

Emits word to the standard output, without a following newline. word must be a single shell word (typically a quoted string) and, for portability, should not include more than one newline. The bytes of word are output as-is, even if it starts with "-" or contains "\". Redirections can be placed outside the macro invocation.

[Macro] AS_ESCAPE (string, [chars = ‘`\"$’])

Expands to string, with any characters in chars escaped with a backslash (‘\’). chars should be at most four bytes long, and only contain characters from the set ‘`\"$’; however, characters may be safely listed more than once in chars for the sake of syntax highlighting editors. The current implementation expands string after adding escapes; if string contains macro calls that in turn expand to text needing shell quoting, you

can use AS_ESCAPE(m4_dquote(m4_expand([string]))).

The default for chars (‘\"$`’) is the set of characters needing escapes when string will be used literally within double quotes. One common variant is the set of characters to protect when string will be used literally within back-ticks or an unquoted here-document (‘\$`’). Another common variant is ‘""’, which can be used to form a double-quoted string containing the same expansions that would have occurred if string were expanded in an unquoted here-document; however, when using this variant, care must be taken that string does not use double quotes within complex variable expansions (such as ‘${foo-`echo "hi"`}’) that would be broken with improper escapes. This macro is often used with AS_ECHO. For an example, observe the output generated by the shell code generated from this snippet:

     foo=bar
     AS_ECHO(["AS_ESCAPE(["$foo" = ])AS_ESCAPE(["$foo"], [""])"])
     ⇒"$foo" = "bar"
     m4_define([macro], [a, [\b]])
     AS_ECHO(["AS_ESCAPE([[macro]])"])
     ⇒macro
     AS_ECHO(["AS_ESCAPE([macro])"])
     ⇒a, b
     AS_ECHO(["AS_ESCAPE(m4_dquote(m4_expand([macro])))"])
     ⇒a, \b

To escape a string that will be placed within single quotes, use:

     m4_bpatsubst([[string]], ['], ['\\''])
[Macro] AS_EXIT ([status = ‘$?’])

Emit code to exit the shell with status, defaulting to ‘$?’. This macro works around shells that see the exit status of the command prior to exit inside a ‘trap 0’ handler (see [Limitations of Shell Builtins], page 222).

[Macro] AS_IF (test1, [run-if-true1], . . ., [run-if-false])

Run shell code test1. If test1 exits with a zero status then run shell code run-if-true1, else examine further tests. If no test exits with a zero status, run shell code run-if-false, with simplifications if either run-if-true1 or run-if-false is empty. For example,

     AS_IF([test "x$foo" = xyes], [HANDLE_FOO([yes])],
           [test "x$foo" != xno], [HANDLE_FOO([maybe])],
           [echo foo not specified])

ensures any required macros of HANDLE_FOO are expanded before the first test.

[Macro] AS_MKDIR_P (file-name)

Make the directory file-name, including intervening directories as necessary. This is equivalent to ‘mkdir -p -- file-name’, except that it is portable to older versions of mkdir that lack support for the ‘-p’ option or for the ‘--’ delimiter (see [Limitations of Usual Tools], page 232). Also, AS_MKDIR_P succeeds if file-name is a symbolic link to an existing directory, even though Posix is unclear whether ‘mkdir -p’ should succeed in that case. If creation of file-name fails, exit the script.

Also see the AC_PROG_MKDIR_P macro (see Section 5.2.1 [Particular Programs], page 45).

[Macro] AS_SET_STATUS (status)

Emit shell code to set the value of ‘$?’ to status, as efficiently as possible. However, this is not guaranteed to abort a shell running with set -e (see [Limitations of Shell Builtins], page 218). This should also be used at the end of a complex shell function instead of ‘return’ (see Section 11.11 [Shell Functions], page 209) to avoid a DJGPP shell bug.

[Macro] AS_TR_CPP (expression)

Transform expression into a valid right-hand side for a C #define. For example:

     # This outputs "#define HAVE_CHAR_P 1".
     # Notice the m4 quoting around #, to prevent an m4 comment
     type="char *"
     echo "[#]define AS_TR_CPP([HAVE_$type]) 1"
[Macro] AS_TR_SH (expression)

Transform expression into shell code that generates a valid shell variable name. The result is literal when possible at m4 time, but must be used with eval if expression causes shell indirections. For example:

     # This outputs "Have it!".
     header="sys/some file.h"
     eval AS_TR_SH([HAVE_$header])=yes
     if test "x$HAVE_sys_some_file_h" = xyes; then echo "Have it!"; fi
[Macro] AS_SET_CATFILE (var, dir, file)

Set the polymorphic shell variable var to dir/file, but optimizing the common cases (dir or file is ‘.’, file is absolute, etc.).

[Macro] AS_UNSET (var)

Unsets the shell variable var, working around bugs in older shells (see [Limitations of Shell Builtins], page 224). var can be a literal or indirect variable name.

[Macro] AS_VERSION_COMPARE (version-1, version-2, [action-if-less],
   [action-if-equal], [action-if-greater])

Compare two strings version-1 and version-2, possibly containing shell variables, as version strings, and expand action-if-less, action-if-equal, or action-if-greater depending upon the result. The algorithm to compare is similar to the one used by strverscmp in glibc (see Section “String/Array Comparison” in The GNU C Library).

9.2 Support for indirect variable names

Often, it is convenient to write a macro that will emit shell code operating on a shell variable. The simplest case is when the variable name is known. But a more powerful idiom is writing shell code that can work through an indirection, where another variable or command substitution produces the name of the variable to actually manipulate. M4sh supports the notion of polymorphic shell variables, making it easy to write a macro that can deal with either literal or indirect variable names and output shell code appropriate for both use cases. Behavior is undefined if expansion of an indirect variable does not result in a literal variable name.

[Macro] AS_LITERAL_IF (expression, [if-literal], [if-not], [if-simple-ref
   = ‘if-not’])
[Macro] AS_LITERAL_WORD_IF (expression, [if-literal], [if-not],
   [if-simple-ref = ‘if-not’])

If the expansion of expression is definitely a shell literal, expand if-literal. If the expansion of expression looks like it might contain shell indirections (such as $var or `expr`), then if-not is expanded. Sometimes, it is possible to output optimized code if expression consists only of shell variable expansions (such as ${var}), in which case if-simple-ref can be provided; but defaulting to if-not should always be safe. AS_ LITERAL_WORD_IF only expands if-literal if expression looks like a single shell word, containing no whitespace; while AS_LITERAL_IF allows whitespace in expression. In order to reduce the time spent recognizing whether an expression qualifies as a literal or a simple indirection, the implementation is somewhat conservative: expression must be a single shell word (possibly after stripping whitespace), consisting only of bytes that would have the same meaning whether unquoted or enclosed in double quotes (for example, ‘a.b’ results in if-literal, even though it is not a valid shell variable name; while both ‘'a'’ and ‘[$]’ result in if-not, because they behave differently than ‘"'a'"’ and ‘"[$]"’). This macro can be used in contexts for recognizing portable file names (such as in the implementation of AC_LIBSOURCE), or coupled with some transliterations for forming valid variable names (such as in the implementation of AS_TR_SH, which uses an additional m4_translit to convert ‘.’ to ‘_’). This example shows how to read the contents of the shell variable bar, exercising all three arguments to AS_LITERAL_IF. It results in a script that will output the line ‘hello’ three times.

     AC_DEFUN([MY_ACTION],
     [AS_LITERAL_IF([$1],
       [echo "$$1"],
       [AS_VAR_COPY([tmp], [$1])
        echo "$tmp"],

       [eval 'echo "$'"$1"\"])])
     foo=bar bar=hello
     MY_ACTION([bar])
     MY_ACTION([`echo bar`])
     MY_ACTION([$foo])
[Macro] AS_VAR_APPEND (var, text)

Emit shell code to append the shell expansion of text to the end of the current contents of the polymorphic shell variable var, taking advantage of shells that provide the ‘+=’ extension for more efficient scaling. For situations where the final contents of var are relatively short (less than 256 bytes), it is more efficient to use the simpler code sequence of var=${var}text (or

its polymorphic equivalent of AS_VAR_COPY([tmp], [var]) and AS_VAR_SET([var],

["$tmp"text])). But in the case when the script will be repeatedly appending text into var, issues of scaling start to become apparent. A naive implementation requires execution time linear to the length of the current contents of var as well as the length of text for a single append, for an overall quadratic scaling with multiple appends. This macro takes advantage of shells which provide the extension var+=text, which can provide amortized constant time for a single append, for an overall linear scaling with multiple appends. Note that unlike AS_VAR_SET, this macro requires that text be quoted properly to avoid field splitting and file name expansion.

[Macro] AS_VAR_ARITH (var, expression)

Emit shell code to compute the arithmetic expansion of expression, assigning the result as the contents of the polymorphic shell variable var. The code takes advantage of shells that provide ‘$(())’ for fewer forks, but uses expr as a fallback. Therefore, the syntax for a valid expression is rather limited: all operators must occur as separate shell arguments and with proper quoting, there is no portable equality operator, all variables containing numeric values must be expanded prior to the computation, all numeric values must be provided in decimal without leading zeroes, and the first shell argument should not be a negative number. In the following example, this snippet will print ‘(2+3)*4 == 20’.

     bar=3
     AS_VAR_ARITH([foo], [\( 2 + $bar \) \* 4])
     echo "(2+$bar)*4 == $foo"
[Macro] AS_VAR_COPY (dest, source)

Emit shell code to assign the contents of the polymorphic shell variable source to the polymorphic shell variable dest. For example, executing this M4sh snippet will output ‘bar hi’:

     foo=bar bar=hi
     AS_VAR_COPY([a], [foo])
     AS_VAR_COPY([b], [$foo])
     echo "$a $b"

When it is necessary to access the contents of an indirect variable inside a shell doublequoted context, the recommended idiom is to first copy the contents into a temporary literal shell variable.

      for header in stdint_h inttypes_h ; do
        AS_VAR_COPY([var], [ac_cv_header_$header])
        echo "$header detected: $var"
      done
[Macro] AS_VAR_IF (var, [value], [if-equal], [if-not-equal])

Output a shell conditional statement. If the contents of the polymorphic shell variable var match the string value, execute if-equal; otherwise execute if-not-equal. Avoids shell bugs if an interrupt signal arrives while a command substitution in var is being expanded.

[Macro] AS_VAR_PUSHDEF (m4-name, value)
[Macro] AS_VAR_POPDEF (m4-name)

A common M4sh idiom involves composing shell variable names from an m4 argument (for example, writing a macro that uses a cache variable). value can be an arbitrary string, which will be transliterated into a valid shell name by AS_TR_SH. In order to access the composed variable name based on value, it is easier to declare a temporary m4 macro m4-name with AS_VAR_PUSHDEF, then use that macro as the argument to subsequent AS_VAR macros as a polymorphic variable name, and finally free the temporary macro with AS_VAR_POPDEF. These macros are often followed with dnl, to avoid excess newlines in the output. Here is an involved example, that shows the power of writing macros that can handle composed shell variable names:

     m4_define([MY_CHECK_HEADER],
     [AS_VAR_PUSHDEF([my_Header], [ac_cv_header_$1])dnl
     AS_VAR_IF([my_Header], [yes], [echo "header $1 detected"])dnl
     AS_VAR_POPDEF([my_Header])dnl
     ])
     MY_CHECK_HEADER([stdint.h])
     for header in inttypes.h stdlib.h ; do
       MY_CHECK_HEADER([$header])
     done

In the above example, MY_CHECK_HEADER can operate on polymorphic variable names. In the first invocation, the m4 argument is stdint.h, which transliterates into a literal stdint_h. As a result, the temporary macro my_Header expands to the literal shell name ‘ac_cv_header_stdint_h’. In the second invocation, the m4 argument to MY_CHECK_HEADER is $header, and the temporary macro my_Header expands to the indirect shell name ‘$as_my_Header’. During the shell execution of the for loop, when ‘$header’ contains ‘inttypes.h’, then ‘$as_my_Header’ contains ‘ac_cv_header_inttypes_h’. If this script is then run on a platform where all three headers have been previously detected, the output of the script will include:

      header stdint.h detected
      header inttypes.h detected
      header stdlib.h detected
[Macro] AS_VAR_SET (var, [value])

Emit shell code to assign the contents of the polymorphic shell variable var to the shell expansion of value. value is not subject to field splitting or file name expansion, so if command substitution is used, it may be done with ‘`""`’ rather than using an intermediate variable (see Section 11.6 [Shell Substitutions], page 195). However, value does undergo rescanning for additional macro names; behavior is unspecified if late expansion results in any shell meta-characters.

[Macro] AS_VAR_SET_IF (var, [if-set], [if-undef])

Emit a shell conditional statement, which executes if-set if the polymorphic shell variable var is set to any value, and if-undef otherwise.

[Macro] AS_VAR_TEST_SET (var)

Emit a shell statement that results in a successful exit status only if the polymorphic shell variable var is set.

9.3 Initialization Macros

[Macro] AS_BOURNE_COMPATIBLE

Set up the shell to be more compatible with the Bourne shell as standardized by Posix, if possible. This may involve setting environment variables, or setting options, or similar implementation-specific actions. This macro is deprecated, since AS_INIT already invokes it.

[Macro] AS_INIT

Initialize the M4sh environment. This macro calls m4_init, then outputs the #! /bin/sh line, a notice about where the output was generated from, and code to sanitize the environment for the rest of the script. Among other initializations, this sets SHELL to the shell chosen to run the script (see [CONFIG SHELL], page 288), and LC_ALL to ensure the C locale. Finally, it changes the current diversion to BODY. AS_ INIT is called automatically by AC_INIT and AT_INIT, so shell code in ‘configure’, ‘config.status’, and ‘testsuite’ all benefit from a sanitized shell environment.

[Macro] AS_INIT_GENERATED (file, [comment])

Emit shell code to start the creation of a subsidiary shell script in file, including changing file to be executable. This macro populates the child script with information learned from the parent (thus, the emitted code is equivalent in effect, but more

efficient, than the code output by AS_INIT, AS_BOURNE_COMPATIBLE, and AS_SHELL_

SANITIZE). If present, comment is output near the beginning of the child, prior to the shell initialization code, and is subject to parameter expansion, command substitution, and backslash quote removal. The parent script should check the exit status after this macro, in case file could not be properly created (for example, if the disk was full). If successfully created, the parent script can then proceed to append additional M4sh constructs into the child script. Note that the child script starts life without a log file open, so if the parent script uses logging (see [AS MESSAGE LOG FD], page 173), you must temporarily disable any attempts to use the log file until after emitting code to open a log within the child. On the other hand, if the parent script has AS_MESSAGE_FD redirected somewhere besides ‘1’, then the child script already has code that copies stdout to that descriptor. Currently, the suggested idiom for writing a M4sh shell script from within another script is:

     AS_INIT_GENERATED([file], [[# My child script.
     ]]) || { AS_ECHO(["Failed to create child script"]); AS_EXIT; }
     m4_pushdef([AS_MESSAGE_LOG_FD])dnl
     cat >> "file" <<\__EOF__
     # Code to initialize AS_MESSAGE_LOG_FD
     m4_popdef([AS_MESSAGE_LOG_FD])dnl
     # Additional code
     __EOF__

This, however, may change in the future as the M4sh interface is stabilized further. Also, be aware that use of LINENO within the child script may report line numbers relative to their location in the parent script, even when using AS_LINENO_PREPARE, if the parent script was unable to locate a shell with working LINENO support.

[Macro] AS_LINENO_PREPARE

Find a shell that supports the special variable LINENO, which contains the number of the currently executing line. This macro is automatically invoked by AC_INIT in configure scripts.

[Macro] AS_ME_PREPARE

Set up variable as_me to be the basename of the currently executing script. This macro is automatically invoked by AC_INIT in configure scripts.

[Macro] AS_SHELL_SANITIZE

Initialize the shell suitably for configure scripts. This has the effect of AS_BOURNE_ COMPATIBLE, and sets some other environment variables for predictable results from configuration tests. For example, it sets LC_ALL to change to the default C locale. See Section 11.10 [Special Shell Variables], page 202. This macro is deprecated, since AS_INIT already invokes it.

9.4 File Descriptor Macros

The following macros define file descriptors used to output messages (or input values) from ‘configure’ scripts. For example:

echo "$wombats found" >&AS_MESSAGE_LOG_FD
echo 'Enter desired kangaroo count:' >&AS_MESSAGE_FD
read kangaroos <&AS_ORIGINAL_STDIN_FD`

However doing so is seldom needed, because Autoconf provides higher level macros as described below.

[Macro] AS_MESSAGE_FD

The file descriptor for ‘checking for...’ messages and results. By default, AS_INIT sets this to ‘1’ for standalone M4sh clients. However, AC_INIT shuffles things around to another file descriptor, in order to allow the ‘-q’ option of configure to choose whether messages should go to the script’s standard output or be discarded. If you want to display some messages, consider using one of the printing macros (see Section 7.5 [Printing Messages], page 118) instead. Copies of messages output via these macros are also recorded in ‘config.log’.

[Macro] AS_MESSAGE_LOG_FD

This must either be empty, or expand to a file descriptor for log messages. By default, AS_INIT sets this macro to the empty string for standalone M4sh clients, thus disabling logging. However, AC_INIT shuffles things around so that both configure and config.status use ‘config.log’ for log messages. Macros that run tools, like AC_COMPILE_IFELSE (see Section 6.4 [Running the Compiler], page 106), redirect all output to this descriptor. You may want to do so if you develop such a low-level macro.

[Macro] AS_ORIGINAL_STDIN_FD

This must expand to a file descriptor for the original standard input. By default, AS_ INIT sets this macro to ‘0’ for standalone M4sh clients. However, AC_INIT shuffles things around for safety. When configure runs, it may accidentally execute an interactive command that has the same name as the non-interactive meant to be used or checked. If the standard input was the terminal, such interactive programs would cause configure to stop, pending some user input. Therefore configure redirects its standard input from ‘/dev/null’ during its initialization. This is not normally a problem, since configure normally does not need user input. In the extreme case where your ‘configure’ script really needs to obtain some values from the original standard input, you can read them explicitly from AS_ORIGINAL_

STDIN_FD.

10 Writing Autoconf Macros

When you write a feature test that could be applicable to more than one software package, the best thing to do is encapsulate it in a new macro. Here are some instructions and guidelines for writing Autoconf macros.

10.1 Macro Definitions

[Macro] AC_DEFUN (name, [body])

Autoconf macros are defined using the AC_DEFUN macro, which is similar to the M4 builtin m4_define macro; this creates a macro named name and with body as its expansion. In addition to defining a macro, AC_DEFUN adds to it some code that is used to constrain the order in which macros are called, while avoiding redundant output (see Section 10.4.1 [Prerequisite Macros], page 178).

An Autoconf macro definition looks like this:

     AC_DEFUN(macro-name, macro-body)

You can refer to any arguments passed to the macro as ‘$1’, ‘$2’, etc. See Section “How to define new macros” in GNU M4, for more complete information on writing M4 macros.

Most macros fall in one of two general categories. The first category includes macros which take arguments, in order to generate output parameterized by those arguments. Macros in this category are designed to be directly expanded, often multiple times, and should not be used as the argument to AC_REQUIRE. The other category includes macros which are shorthand for a fixed block of text, and therefore do not take arguments. For this category of macros, directly expanding the macro multiple times results in redundant output, so it is more common to use the macro as the argument to AC_REQUIRE, or to declare the macro with AC_DEFUN_ONCE (see Section 10.4.3 [One-Shot Macros], page 181).

Be sure to properly quote both the macro-body and the macro-name to avoid any problems if the macro happens to have been previously defined.

Each macro should have a header comment that gives its prototype, and a brief description. When arguments have default values, display them in the prototype. For example:

     # AC_MSG_ERROR(ERROR, [EXIT-STATUS = 1])
     # --------------------------------------
     m4_define([AC_MSG_ERROR],
       [{ AS_MESSAGE([error: $1], [2])
          exit m4_default([$2], [1]); }])

Comments about the macro should be left in the header comment. Most other comments make their way into ‘configure’, so just keep using ‘#’ to introduce comments.

If you have some special comments about pure M4 code, comments that make no sense in ‘configure’ and in the header comment, then use the builtin dnl: it causes M4 to discard the text through the next newline.

Keep in mind that dnl is rarely needed to introduce comments; dnl is more useful to get rid of the newlines following macros that produce no output, such as AC_REQUIRE.

Public third-party macros need to use AC_DEFUN, and not m4_define, in order to be found by aclocal (see Section “Extending aclocal” in GNU Automake). Additionally, if it is ever determined that a macro should be made obsolete, it is easy to convert from AC_DEFUN to AU_DEFUN in order to have autoupdate assist the user in choosing a better alternative, but there is no corresponding way to make m4_define issue an upgrade notice (see [AU DEFUN], page 182).

There is another subtle, but important, difference between using m4_define and AC_ DEFUN: only the former is unaffected by AC_REQUIRE. When writing a file, it is always safe to replace a block of text with a m4_define macro that will expand to the same text. But replacing a block of text with an AC_DEFUN macro with the same content does not necessarily give the same results, because it changes the location where any embedded but unsatisfied AC_REQUIRE invocations within the block will be expanded. For an example of this, see Section 20.8 [Expanded Before Required], page 334.

10.2 Macro Names

All of the public Autoconf macros have all-uppercase names in the namespace ‘^AC_’ to prevent them from accidentally conflicting with other text; Autoconf also reserves the namespace ‘^_AC_’ for internal macros. All shell variables that they use for internal purposes have mostly-lowercase names starting with ‘ac_’. Autoconf also uses here-document delimiters in the namespace ‘^_AC[A-Z]’. During configure, files produced by Autoconf make heavy use of the file system namespace ‘^conf’.

Since Autoconf is built on top of M4sugar (see Section 8.3 [Programming in M4sugar], page 135) and M4sh (see Chapter 9 [Programming in M4sh], page 165), you must also be aware of those namespaces (‘^_?\(m4\|AS\)_’). And since ‘configure.ac’ is also designed to be scanned by Autoheader, Autoscan, Autoupdate, and Automake, you should be aware of the ‘^_?A[HNUM]_’ namespaces. In general, you should not use the namespace of a package that does not own the macro or shell code you are writing.

To ensure that your macros don’t conflict with present or future Autoconf macros, you should prefix your own macro names and any shell variables they use with some other sequence. Possibilities include your initials, or an abbreviation for the name of your organization or software package. Historically, people have not always followed the rule of using a namespace appropriate for their package, and this has made it difficult for determining the origin of a macro (and where to report bugs about that macro), as well as difficult for the true namespace owner to add new macros without interference from pre-existing uses of third-party macros. Perhaps the best example of this confusion is the AM_GNU_GETTEXT macro, which belongs, not to Automake, but to Gettext.

Most of the Autoconf macros’ names follow a structured naming convention that indicates the kind of feature check by the name. The macro names consist of several words, separated by underscores, going from most general to most specific. The names of their cache variables use the same convention (see Section 7.4.1 [Cache Variable Names], page 116, for more information on them).

The first word of the name after the namespace initials (such as ‘AC_’) usually tells the category of the feature being tested. Here are the categories used in Autoconf for specific test macros, the kind of macro that you are more likely to write. They are also used for cache variables, in all-lowercase. Use them where applicable; where they’re not, invent your own categories.

C C language builtin features.

DECL Declarations of C variables in header files.

FUNC Functions in libraries.

GROUP Posix group owners of files.

HEADER Header files.

LIB C libraries.

PROG The base names of programs.

MEMBER Members of aggregates.

SYS Operating system features.

TYPE C builtin or declared types.

VAR C variables in libraries.

After the category comes the name of the particular feature being tested. Any further words in the macro name indicate particular aspects of the feature. For example, AC_PROG_ CC_STDC checks whether the C compiler supports ISO Standard C.

An internal macro should have a name that starts with an underscore; Autoconf internals should therefore start with ‘_AC_’. Additionally, a macro that is an internal subroutine of another macro should have a name that starts with an underscore and the name of that other macro, followed by one or more words saying what the internal macro does. For

example, AC_PATH_X has internal macros _AC_PATH_X_XMKMF and _AC_PATH_X_DIRECT.

10.3 Reporting Messages

When macros statically diagnose abnormal situations, benign or fatal, it is possible to make autoconf detect the problem, and refuse to create ‘configure’ in the case of an error. The macros in this section are considered obsolescent, and new code should use M4sugar macros for this purpose, see Section 8.3.2 [Diagnostic Macros], page 138.

On the other hand, it is possible to want to detect errors when configure is run, which are dependent on the environment of the user rather than the maintainer. For dynamic diagnostics, see Section 7.5 [Printing Messages], page 118.

[Macro] AC_DIAGNOSE (category, message)

Report message as a warning (or as an error if requested by the user) if warnings of the category are turned on. This macro is obsolescent; you are encouraged to use:

          m4_warn([category], [message])

instead. See [m4 warn], page 139, for more details, including valid category names.

[Macro] AC_WARNING (message)

Report message as a syntax warning. This macro is obsolescent; you are encouraged to use:

          m4_warn([syntax], [message])

instead. See [m4 warn], page 139, for more details, as well as better finer-grained categories of warnings (not all problems have to do with syntax).

[Macro] AC_FATAL (message)

Report a severe error message, and have autoconf die. This macro is obsolescent; you are encouraged to use:

          m4_fatal([message])

instead. See [m4 fatal], page 138, for more details.

When the user runs ‘autoconf -W error’, warnings from m4_warn (including those issued through AC_DIAGNOSE and AC_WARNING) are reported as errors, see Section 3.4 [autoconf Invocation], page 13.

10.4 Dependencies Between Macros

Some Autoconf macros depend on other macros having been called first in order to work correctly. Autoconf provides a way to ensure that certain macros are called if needed and a way to warn the user if macros are called in an order that might cause incorrect operation.

10.4.1 Prerequisite Macros

A macro that you write might need to use values that have previously been computed by other macros. For example, AC_DECL_YYTEXT examines the output of flex or lex, so it depends on AC_PROG_LEX having been called first to set the shell variable LEX.

Rather than forcing the user of the macros to keep track of the dependencies between them, you can use the AC_REQUIRE macro to do it automatically. AC_REQUIRE can ensure that a macro is only called if it is needed, and only called once.

[Macro] AC_REQUIRE (macro-name)

If the M4 macro macro-name has not already been called, call it (without any arguments). Make sure to quote macro-name with square brackets. macro-name must have been defined using AC_DEFUN or else contain a call to AC_PROVIDE to indicate that it has been called.

AC_REQUIRE must be used inside a macro defined by AC_DEFUN; it must not be called from the top level. Also, it does not make sense to require a macro that takes parameters.

AC_REQUIRE is often misunderstood. It really implements dependencies between macros in the sense that if one macro depends upon another, the latter is expanded before the body of the former. To be more precise, the required macro is expanded before the outermost defined macro in the current expansion stack. In particular, ‘AC_REQUIRE([FOO])’ is not replaced with the body of FOO. For instance, this definition of macros:

     AC_DEFUN([TRAVOLTA],
     [test "$body_temperature_in_celsius" -gt "38" &&
       dance_floor=occupied])
     AC_DEFUN([NEWTON_JOHN],
     [test "x$hair_style" = xcurly &&
       dance_floor=occupied])

AC_DEFUN([RESERVE_DANCE_FLOOR],
[if date | grep '^Sat.*pm' >/dev/null 2>&1; then
  AC_REQUIRE([TRAVOLTA])
  AC_REQUIRE([NEWTON_JOHN])
fi])
with this ‘configure.ac’
AC_INIT([Dance Manager], [1.0], [bug-dance@example.org])
RESERVE_DANCE_FLOOR
if test "x$dance_floor" = xoccupied; then
  AC_MSG_ERROR([cannot pick up here, let's move])
fi

does not leave you with a better chance to meet a kindred soul at other times than Saturday night since it expands into:

test "$body_temperature_in_Celsius" -gt "38" &&
  dance_floor=occupied
test "x$hair_style" = xcurly &&
  dance_floor=occupied
fi
if date | grep '^Sat.*pm' >/dev/null 2>&1; then

fi

This behavior was chosen on purpose: (i) it prevents messages in required macros from interrupting the messages in the requiring macros; (ii) it avoids bad surprises when shell conditionals are used, as in:

if ...; then
  AC_REQUIRE([SOME_CHECK])
fi
...
SOME_CHECK

However, this implementation can lead to another class of problems. Consider the case where an outer macro first expands, then indirectly requires, an inner macro:

AC_DEFUN([TESTA], [[echo in A
if test -n "$SEEN_A" ; then echo duplicate ; fi
SEEN_A=:]])
AC_DEFUN([TESTB], [AC_REQUIRE([TESTA])[echo in B
if test -z "$SEEN_A" ; then echo bug ; fi]])
AC_DEFUN([TESTC], [AC_REQUIRE([TESTB])[echo in C]])
AC_DEFUN([OUTER], [[echo in OUTER]
TESTA
TESTC])
OUTER

Prior to Autoconf 2.64, the implementation of AC_REQUIRE recognized that TESTB needed to be hoisted prior to the expansion of OUTER, but because TESTA had already been directly expanded, it failed to hoist TESTA. Therefore, the expansion of TESTB occurs prior to its prerequisites, leading to the following output:

in B
bug
in OUTER
in A
in C

Newer Autoconf is smart enough to recognize this situation, and hoists TESTA even though it has already been expanded, but issues a syntax warning in the process. This is because the hoisted expansion of TESTA defeats the purpose of using AC_REQUIRE to avoid redundant code, and causes its own set of problems if the hoisted macro is not idempotent:

in A
in B
in OUTER
in A
duplicate
in C

The bug is not in Autoconf, but in the macro definitions. If you ever pass a particular macro name to AC_REQUIRE, then you are implying that the macro only needs to be expanded once. But to enforce this, either the macro must be declared with AC_DEFUN_ONCE (although this only helps in Autoconf 2.64 or newer), or all uses of that macro should be through AC_REQUIRE; directly expanding the macro defeats the point of using AC_REQUIRE to eliminate redundant expansion. In the example, this rule of thumb was violated because TESTB requires TESTA while OUTER directly expands it. One way of fixing the bug is to factor TESTA into two macros, the portion designed for direct and repeated use (here, named TESTA), and the portion designed for one-shot output and used only inside AC_ REQUIRE (here, named TESTA_PREREQ). Then, by fixing all clients to use the correct calling convention according to their needs:

AC_DEFUN([TESTA], [AC_REQUIRE([TESTA_PREREQ])[echo in A]])
AC_DEFUN([TESTA_PREREQ], [[echo in A_PREREQ
if test -n "$SEEN_A" ; then echo duplicate ; fi
SEEN_A=:]])
AC_DEFUN([TESTB], [AC_REQUIRE([TESTA_PREREQ])[echo in B
if test -z "$SEEN_A" ; then echo bug ; fi]])
AC_DEFUN([TESTC], [AC_REQUIRE([TESTB])[echo in C]])
AC_DEFUN([OUTER], [[echo in OUTER]
TESTA
TESTC])
OUTER

the resulting output will then obey all dependency rules and avoid any syntax warnings, whether the script is built with old or new Autoconf versions:

in A_PREREQ
in B
in OUTER
in A

     in C

The helper macros AS_IF and AS_CASE may be used to enforce expansion of required macros outside of shell conditional constructs. You are furthermore encouraged, although not required, to put all AC_REQUIRE calls at the beginning of a macro. You can use dnl to avoid the empty lines they leave.

10.4.2 Suggested Ordering

Some macros should be run before another macro if both are called, but neither requires that the other be called. For example, a macro that changes the behavior of the C compiler should be called before any macros that run the C compiler. Many of these dependencies are noted in the documentation.

Autoconf provides the AC_BEFORE macro to warn users when macros with this kind of dependency appear out of order in a ‘configure.ac’ file. The warning occurs when creating

configure from ‘configure.ac’, not when running configure.

For example, AC_PROG_CPP checks whether the C compiler can run the C preprocessor when given the ‘-E’ option. It should therefore be called after any macros that change which C compiler is being used, such as AC_PROG_CC. So AC_PROG_CC contains:

     AC_BEFORE([$0], [AC_PROG_CPP])dnl

This warns the user if a call to AC_PROG_CPP has already occurred when AC_PROG_CC is called.

[Macro] AC_BEFORE (this-macro-name, called-macro-name)

Make M4 print a warning message to the standard error output if called-macro-name has already been called. this-macro-name should be the name of the macro that is calling AC_BEFORE. The macro called-macro-name must have been defined using AC_DEFUN or else contain a call to AC_PROVIDE to indicate that it has been called.

10.4.3 One-Shot Macros

Some macros should be called only once, either because calling them multiple time is unsafe, or because it is bad style. For instance Autoconf ensures that AC_CANONICAL_BUILD and cousins (see Section 14.2 [Canonicalizing], page 268) are evaluated only once, because it makes no sense to run these expensive checks more than once. Such one-shot macros can be defined using AC_DEFUN_ONCE.

[Macro] AC_DEFUN_ONCE (macro-name, macro-body)

Declare macro macro-name like AC_DEFUN would (see Section 10.1 [Macro Definitions], page 175), but add additional logic that guarantees that only the first use of the macro (whether by direct expansion or AC_REQUIRE) causes an expansion of macro-body; the expansion will occur before the start of any enclosing macro defined by AC_DEFUN. Subsequent expansions are silently ignored. Generally, it does not make sense for macro-body to use parameters such as $1.

Prior to Autoconf 2.64, a macro defined by AC_DEFUN_ONCE would emit a warning if it was directly expanded a second time, so for portability, it is better to use AC_REQUIRE than direct invocation of macro-name inside a macro defined by AC_DEFUN (see Section 10.4.1 [Prerequisite Macros], page 178).

10.5 Obsoleting Macros

Configuration and portability technology has evolved over the years. Often better ways of solving a particular problem are developed, or ad-hoc approaches are systematized. This process has occurred in many parts of Autoconf. One result is that some of the macros are now considered obsolete; they still work, but are no longer considered the best thing to do, hence they should be replaced with more modern macros. Ideally, autoupdate should replace the old macro calls with their modern implementation.

Autoconf provides a simple means to obsolete a macro.

[Macro] AU_DEFUN (old-macro, implementation, [message])

Define old-macro as implementation. The only difference with AC_DEFUN is that the user is warned that old-macro is now obsolete. If she then uses autoupdate, the call to old-macro is replaced by the modern implementation. message should include information on what to do after running autoupdate; autoupdate prints it as a warning, and includes it in the updated

     ‘configure.ac’ file.

The details of this macro are hairy: if autoconf encounters an AU_DEFUNed macro, all macros inside its second argument are expanded as usual. However, when autoupdate is run, only M4 and M4sugar macros are expanded here, while all other macros are disabled and appear literally in the updated ‘configure.ac’.

[Macro] AU_ALIAS (old-name, new-name)

Used if the old-name is to be replaced by a call to new-macro with the same parameters. This happens for example if the macro was renamed.

10.6 Coding Style

The Autoconf macros follow a strict coding style. You are encouraged to follow this style, especially if you intend to distribute your macro, either by contributing it to Autoconf itself

or the Autoconf Macro Archive (http://www.gnu.org/software/autoconf-archive/), or

by other means.

The first requirement is to pay great attention to the quotation. For more details, see Section 3.1.2 [Autoconf Language], page 9, and Section 8.1 [M4 Quotation], page 121.

Do not try to invent new interfaces. It is likely that there is a macro in Autoconf that resembles the macro you are defining: try to stick to this existing interface (order of arguments, default values, etc.). We are conscious that some of these interfaces are not perfect; nevertheless, when harmless, homogeneity should be preferred over creativity.

Be careful about clashes both between M4 symbols and between shell variables.

If you stick to the suggested M4 naming scheme (see Section 10.2 [Macro Names], page 176), you are unlikely to generate conflicts. Nevertheless, when you need to set a special value, avoid using a regular macro name; rather, use an “impossible” name. For instance, up to version 2.13, the macro AC_SUBST used to remember what symbol macros were already defined by setting AC_SUBST_symbol, which is a regular macro name. But since there is a macro named AC_SUBST_FILE, it was just impossible to ‘AC_SUBST(FILE)’! In this case, AC_SUBST(symbol) or _AC_SUBST(symbol) should have been used (yes, with the parentheses).

No Autoconf macro should ever enter the user-variable name space; i.e., except for the variables that are the actual result of running the macro, all shell variables should start with ac_. In addition, small macros or any macro that is likely to be embedded in other macros should be careful not to use obvious names.

Do not use dnl to introduce comments: most of the comments you are likely to write are either header comments which are not output anyway, or comments that should make their way into ‘configure’. There are exceptional cases where you do want to comment special M4 constructs, in which case dnl is right, but keep in mind that it is unlikely.

M4 ignores the leading blanks and newlines before each argument. Use this feature to indent in such a way that arguments are (more or less) aligned with the opening parenthesis of the macro being called. For instance, instead of

     AC_CACHE_CHECK(for EMX OS/2 environment,
     ac_cv_emxos2,
     [AC_COMPILE_IFELSE([AC_LANG_PROGRAM(, [return __EMX__;])],
     [ac_cv_emxos2=yes], [ac_cv_emxos2=no])])

write

     AC_CACHE_CHECK([for EMX OS/2 environment], [ac_cv_emxos2],
     [AC_COMPILE_IFELSE([AC_LANG_PROGRAM([], [return __EMX__;])],
                        [ac_cv_emxos2=yes],
                        [ac_cv_emxos2=no])])

or even

     AC_CACHE_CHECK([for EMX OS/2 environment],
                    [ac_cv_emxos2],
                    [AC_COMPILE_IFELSE([AC_LANG_PROGRAM([],
                                                        [return __EMX__;])],
                                       [ac_cv_emxos2=yes],
                                       [ac_cv_emxos2=no])])

When using AC_RUN_IFELSE or any macro that cannot work when cross-compiling, provide a pessimistic value (typically ‘no’).

Feel free to use various tricks to prevent auxiliary tools, such as syntax-highlighting editors, from behaving improperly. For instance, instead of:

     m4_bpatsubst([$1], [$"])

use

     m4_bpatsubst([$1], [$""])

so that Emacsen do not open an endless “string” at the first quote. For the same reasons, avoid:

     test $[#] != 0

and use:

     test $[@%:@] != 0

Otherwise, the closing bracket would be hidden inside a ‘#’-comment, breaking the bracketmatching highlighting from Emacsen. Note the preferred style to escape from M4: ‘$[1]’, ‘$[@]’, etc. Do not escape when it is unnecessary. Common examples of useless quotation are ‘[$]$1’ (write ‘$$1’), ‘[$]var’ (use ‘$var’), etc. If you add portability issues to the picture, you’ll prefer ‘${1+"$[@]"}’ to ‘"[$]@"’, and you’ll prefer do something better than hacking Autoconf :-).

When using sed, don’t use ‘-e’ except for indenting purposes. With the s and y commands, the preferred separator is ‘/’ unless ‘/’ itself might appear in the pattern or replacement, in which case you should use ‘|’, or optionally ‘,’ if you know the pattern and replacement cannot contain a file name. If none of these characters will do, choose a printable character that cannot appear in the pattern or replacement. Characters from the set ‘"#$&'()*;<=>¿|~’ are good choices if the pattern or replacement might contain a file name, since they have special meaning to the shell and are less likely to occur in file names.

See Section 10.1 [Macro Definitions], page 175, for details on how to define a macro. If a macro doesn’t use AC_REQUIRE, is expected to never be the object of an AC_REQUIRE directive, and macros required by other macros inside arguments do not need to be expanded before this macro, then use m4_define. In case of doubt, use AC_DEFUN. Also take into account that public third-party macros need to use AC_DEFUN in order to be found by aclocal (see Section “Extending aclocal” in GNU Automake). All the AC_REQUIRE statements should be at the beginning of the macro, and each statement should be followed by

dnl.

You should not rely on the number of arguments: instead of checking whether an argument is missing, test that it is not empty. It provides both a simpler and a more predictable interface to the user, and saves room for further arguments.

Unless the macro is short, try to leave the closing ‘])’ at the beginning of a line, followed by a comment that repeats the name of the macro being defined. This introduces an additional newline in configure; normally, that is not a problem, but if you want to remove it you can use ‘[]dnl’ on the last line. You can similarly use ‘[]dnl’ after a macro call to remove its newline. ‘[]dnl’ is recommended instead of ‘dnl’ to ensure that M4 does not interpret the ‘dnl’ as being attached to the preceding text or macro output. For example, instead of:

     AC_DEFUN([AC_PATH_X],
     [AC_MSG_CHECKING([for X])
     AC_REQUIRE_CPP()

# …omitted…

       AC_MSG_RESULT([libraries $x_libraries, headers $x_includes])
     fi])

you would write:

     AC_DEFUN([AC_PATH_X],
     [AC_REQUIRE_CPP()[]dnl
     AC_MSG_CHECKING([for X])

# …omitted…

       AC_MSG_RESULT([libraries $x_libraries, headers $x_includes])
     fi[]dnl
     ])# AC_PATH_X

If the macro is long, try to split it into logical chunks. Typically, macros that check for a bug in a function and prepare its AC_LIBOBJ replacement should have an auxiliary macro to perform this setup. Do not hesitate to introduce auxiliary macros to factor your code.

In order to highlight the recommended coding style, here is a macro written the old way:

dnl Check for EMX on OS/2.
dnl _AC_EMXOS2
AC_DEFUN(_AC_EMXOS2,
[AC_CACHE_CHECK(for EMX OS/2 environment, ac_cv_emxos2,
[AC_COMPILE_IFELSE([AC_LANG_PROGRAM(, return __EMX__;)],
ac_cv_emxos2=yes, ac_cv_emxos2=no)])
test "x$ac_cv_emxos2" = xyes && EMXOS2=yes])

and the new way:

# _AC_EMXOS2
# ----------
# Check for EMX on OS/2.
m4_define([_AC_EMXOS2],
[AC_CACHE_CHECK([for EMX OS/2 environment], [ac_cv_emxos2],
[AC_COMPILE_IFELSE([AC_LANG_PROGRAM([], [return __EMX__;])],
                   [ac_cv_emxos2=yes],
                   [ac_cv_emxos2=no])])
test "x$ac_cv_emxos2" = xyes && EMXOS2=yes[]dnl
])# _AC_EMXOS2

11 Portable Shell Programming

When writing your own checks, there are some shell-script programming techniques you should avoid in order to make your code portable. The Bourne shell and upward-compatible shells like the Korn shell and Bash have evolved over the years, and many features added to the original System7 shell are now supported on all interesting porting targets. However, the following discussion between Russ Allbery and Robert Lipe is worth reading: Russ Allbery:

The GNU assumption that /bin/sh is the one and only shell leads to a permanent deadlock. Vendors don’t want to break users’ existing shell scripts, and there are some corner cases in the Bourne shell that are not completely compatible with a Posix shell. Thus, vendors who have taken this route will never (OK…“never say never”) replace the Bourne shell (as /bin/sh) with a Posix shell.

Robert Lipe:

This is exactly the problem. While most (at least most System V’s) do have a Bourne shell that accepts shell functions most vendor /bin/sh programs are not the Posix shell. So while most modern systems do have a shell somewhere that meets the Posix standard, the challenge is to find it.

For this reason, part of the job of M4sh (see Chapter 9 [Programming in M4sh], page 165) is to find such a shell. But to prevent trouble, if you’re not using M4sh you should not take advantage of features that were added after Unix version 7, circa 1977 (see Section 6.7 [Systemology], page 108); you should not use aliases, negated character classes, or even unset. # comments, while not in Unix version 7, were retrofitted in the original Bourne shell and can be assumed to be part of the least common denominator.

On the other hand, if you’re using M4sh you can assume that the shell has the features that were added in SVR2 (circa 1984), including shell functions, return, unset, and I/O redirection for builtins. For more information, refer to http://www.in-ulm.de/~mascheck/ bourne/. However, some pitfalls have to be avoided for portable use of these constructs; these will be documented in the rest of this chapter. See in particular Section 11.11 [Shell Functions], page 209 and Section 11.12 [Limitations of Shell Builtins], page 210.

Some ancient systems have quite small limits on the length of the ‘#!’ line; for instance, 32 bytes (not including the newline) on SunOS 4. However, these ancient systems are no longer of practical concern.

The set of external programs you should run in a configure script is fairly small. See Section “Utilities in Makefiles” in GNU Coding Standards, for the list. This restriction allows users to start out with a fairly small set of programs and build the rest, avoiding too many interdependencies between packages.

Some of these external utilities have a portable subset of features; see Section 11.13 [Limitations of Usual Tools], page 224.

There are other sources of documentation about shells. The specification for the Posix

Shell Command Language (http://www.opengroup.org/susv3/utilities/xcu_chap02

.html), though more generous than the restrictive shell subset described above, is fairly portable nowadays. Also please see the Shell FAQs (http://www.faqs.org/faqs/

unix-faq/shell/).

11.1 Shellology

There are several families of shells, most prominently the Bourne family and the C shell family which are deeply incompatible. If you want to write portable shell scripts, avoid members of the C shell family. The the Shell difference FAQ (http://www.faqs.org/ faqs/unix-faq/shell/shell-differences/) includes a small history of Posix shells, and a comparison between several of them.

Below we describe some of the members of the Bourne shell family.

Ash Ash is often used on GNU/Linux and BSD systems as a light-weight Bournecompatible shell. Ash 0.2 has some bugs that are fixed in the 0.3.x series, but portable shell scripts should work around them, since version 0.2 is still shipped with many GNU/Linux distributions. To be compatible with Ash 0.2: − don’t use ‘$?’ after expanding empty or unset variables, or at the start of

    an eval:
         foo=
         false
         $foo
         echo "Do not use it: $?"
         false
         eval 'echo "Do not use it: $?"'

− don’t use command substitution within variable expansion:

         cat ${FOO=`bar`}

− beware that single builtin substitutions are not performed by a subshell, hence their effect applies to the current shell! See Section 11.6 [Shell Substitutions], page 195, item “Command Substitution”.

Bash To detect whether you are running Bash, test whether BASH_VERSION is set. To require Posix compatibility, run ‘set -o posix’. See Section “Bash Posix Mode” in The GNU Bash Reference Manual, for details.

Bash 2.05 and later Versions 2.05 and later of Bash use a different format for the output of the set builtin, designed to make evaluating its output easier. However, this output is not compatible with earlier versions of Bash (or with many other shells, probably). So if you use Bash 2.05 or higher to execute configure, you’ll need to use Bash 2.05 for all other build tasks as well.

Ksh The Korn shell is compatible with the Bourne family and it mostly conforms to Posix. It has two major variants commonly called ‘ksh88’ and ‘ksh93’, named after the years of initial release. It is usually called ksh, but is called sh on some hosts if you set your path appropriately. Solaris systems have three variants: /usr/bin/ksh is ‘ksh88’; it is standard on Solaris 2.0 and later. /usr/xpg4/bin/sh is a Posix-compliant variant of ‘ksh88’; it is standard on Solaris 9 and later. /usr/dt/bin/dtksh is ‘ksh93’. Variants that are not standard may be parts of optional packages. There is no extra charge for these packages, but they are not part of a minimal OS install and therefore some installations may not have it. Starting with Tru64 Version 4.0, the Korn shell /usr/bin/ksh is also available as /usr/bin/posix/sh. If the environment variable BIN_SH is set to xpg4, subsidiary invocations of the standard shell conform to Posix.

Pdksh A public-domain clone of the Korn shell called pdksh is widely available: it has most of the ‘ksh88’ features along with a few of its own. It usually sets KSH_VERSION, except if invoked as /bin/sh on OpenBSD, and similarly to Bash you can require Posix compatibility by running ‘set -o posix’. Unfortunately, with pdksh 5.2.14 (the latest stable version as of January 2007) Posix mode is buggy and causes pdksh to depart from Posix in at least one respect, see Section 11.6 [Shell Substitutions], page 195.

Zsh To detect whether you are running zsh, test whether ZSH_VERSION is set. By default zsh is not compatible with the Bourne shell: you must execute ‘emulate sh’, and for zsh versions before 3.1.6-dev-18 you must also set NULLCMD to ‘:’. See Section “Compatibility” in The Z Shell Manual, for details. The default Mac OS X sh was originally Zsh; it was changed to Bash in Mac OS X 10.2.

11.2 Here-Documents

Don’t rely on ‘\’ being preserved just because it has no special meaning together with the next symbol. In the native sh on OpenBSD 2.7 ‘\"’ expands to ‘"’ in here-documents with unquoted delimiter. As a general rule, if ‘\\’ expands to ‘\’ use ‘\\’ to get ‘\’.

With OpenBSD 2.7’s sh

$ cat <<EOF > \" \\ > EOF " \

and with Bash:

bash-2.04$ cat <<EOF
> \" \\
> EOF
\" \

Some shells mishandle large here-documents: for example, Solaris 10 dtksh and the UnixWare 7.1.1 Posix shell, which are derived from Korn shell version M-12/28/93d, mishandle braced variable expansion that crosses a 1024- or 4096-byte buffer boundary within a here-document. Only the part of the variable name after the boundary is used. For example, ${variable} could be replaced by the expansion of ${ble}. If the end of the variable name is aligned with the block boundary, the shell reports an error, as if you used

${}. Instead of ${variable-default}, the shell may expand ${riable-default}, or even

${fault}. This bug can often be worked around by omitting the braces: $variable. The bug was fixed in ‘ksh93g’ (1998-04-30) but as of 2006 many operating systems were still shipping older versions with the bug.

Many shells (including the Bourne shell) implement here-documents inefficiently. In particular, some shells can be extremely inefficient when a single statement contains many here-documents. For instance if your ‘configure.ac’ includes something like:

     if <cross_compiling>; then
       assume this and that
     else
       check this
       check that
       check something else
       ...
       on and on forever
       ...
     fi

A shell parses the whole if/fi construct, creating temporary files for each here-document in it. Some shells create links for such here-documents on every fork, so that the clean-up code they had installed correctly removes them. It is creating the links that can take the shell forever.

Moving the tests out of the if/fi, or creating multiple if/fi constructs, would improve the performance significantly. Anyway, this kind of construct is not exactly the typical use of Autoconf. In fact, it’s even not recommended, because M4 macros can’t look into shell conditionals, so we may fail to expand a macro when it was expanded before in a conditional path, and the condition turned out to be false at runtime, and we end up not executing the macro at all.

Be careful with the use of ‘<<-’ to unindent here-documents. The behavior is only portable for stripping leading TABs, and things can silently break if an overzealous editor converts to using leading spaces (not all shells are nice enough to warn about unterminated here-documents).

     $ printf 'cat <<-x\n\t1\n\t 2\n\tx\n' | bash && echo done
     1
      2
     done
     $ printf 'cat <<-x\n 1\n  2\n x\n' | bash-3.2 && echo done
      1
       2
      x
     done

11.3 File Descriptors

Most shells, if not all (including Bash, Zsh, Ash), output traces on stderr, even for subshells. This might result in undesirable content if you meant to capture the standard-error output of the inner command:

     $ ash -x -c '(eval "echo foo >&2") 2>stderr'
     $ cat stderr

+ eval echo foo >&2
+ echo foo
foo
$ bash -x -c '(eval "echo foo >&2") 2>stderr'
$ cat stderr
+ eval 'echo foo >&2'
++ echo foo
foo
$ zsh -x -c '(eval "echo foo >&2") 2>stderr'
# Traces on startup files deleted here.
$ cat stderr
+zsh:1> eval echo foo >&2
+zsh:1> echo foo
foo

One workaround is to grep out uninteresting lines, hoping not to remove good ones.

If you intend to redirect both standard error and standard output, redirect standard output first. This works better with HP-UX, since its shell mishandles tracing if standard error is redirected first:

$ sh -x -c ': 2>err >out'
+ :
+ 2> err $ cat err
1> out

Don’t try to redirect the standard error of a command substitution. It must be done inside the command substitution. When running ‘: `cd /zorglub` 2>/dev/null’ expect the error message to escape, while ‘: `cd /zorglub 2>/dev/null`’ works properly.

On the other hand, some shells, such as Solaris or FreeBSD /bin/sh, warn about missing programs before performing redirections. Therefore, to silently check whether a program exists, it is necessary to perform redirections on a subshell:

$ /bin/sh -c 'nosuch 2>/dev/null'
nosuch: not found
$ /bin/sh -c '(nosuch) 2>/dev/null'
$ bash -c 'nosuch 2>/dev/null'

FreeBSD 6.2 sh may mix the trace output lines from the statements in a shell pipeline. It is worth noting that Zsh (but not Ash nor Bash) makes it possible in assignments

though: ‘foo=`cd /zorglub` 2>/dev/null’.

Some shells, like ash, don’t recognize bi-directional redirection (‘<>’). And even on shells that recognize it, it is not portable to use on fifos: Posix does not require read-write support for named pipes, and Cygwin does not support it:

$ mkfifo fifo
$ exec 5<>fifo
$ echo hi >&5
bash: echo: write error: Communication error on send

When catering to old systems, don’t redirect the same file descriptor several times, as you are doomed to failure under Ultrix.

ULTRIX V4.4 (Rev. 69) System #31: Thu Aug 10 19:42:23 GMT 1995
UWS V4.4 (Rev. 11)
$ eval 'echo matter >fullness' >void
illegal io
$ eval '(echo matter >fullness)' >void
illegal io
$ (eval '(echo matter >fullness)') >void
Ambiguous output redirect.

In each case the expected result is of course ‘fullness’ containing ‘matter’ and ‘void’ being empty. However, this bug is probably not of practical concern to modern platforms.

Solaris 10 sh will try to optimize away a : command in a loop after the first iteration, even if it is redirected:

$ for i in 1 2 3 ; do : >x$i; done
$ ls
x1

As a workaround, echo or eval can be used.

Don’t rely on file descriptors 0, 1, and 2 remaining closed in a subsidiary program. If any of these descriptors is closed, the operating system may open an unspecified file for the descriptor in the new process image. Posix says this may be done only if the subsidiary program is set-user-ID or set-group-ID, but HP-UX 11.23 does it even for ordinary programs.

Don’t rely on open file descriptors being open in child processes. In ksh, file descriptors above 2 which are opened using ‘exec n>file’ are closed by a subsequent ‘exec’ (such as that involved in the fork-and-exec which runs a program or script). Thus, using sh, we have:

$ cat ./descrips
#!/bin/sh -
echo hello >&5
$ exec 5>t
$ ./descrips
$ cat t
hello
$

But using ksh:

$ exec 5>t
$ ./descrips
hello
$ cat t
$

Within the process which runs the ‘descrips’ script, file descriptor 5 is closed.

Don’t rely on redirection to a closed file descriptor to cause an error. With Solaris /bin/sh, when the redirection fails, the output goes to the original file descriptor.

$ bash -c 'echo hi >&3' 3>&-; echo $?
bash: 3: Bad file descriptor

     $ /bin/sh -c 'echo hi >&3' 3>&-; echo $?
     hi
     0

DOS variants cannot rename or remove open files, such as in ‘mv foo bar >foo’ or ‘rm foo >foo’, even though this is perfectly portable among Posix hosts.

A few ancient systems reserved some file descriptors. By convention, file descriptor 3 was opened to ‘/dev/tty’ when you logged into Eighth Edition (1985) through Tenth Edition Unix (1989). File descriptor 4 had a special use on the Stardent/Kubota Titan (circa 1990), though we don’t now remember what it was. Both these systems are obsolete, so it’s now safe to treat file descriptors 3 and 4 like any other file descriptors.

11.4 File System Conventions

Autoconf uses shell-script processing extensively, so the file names that it processes should not contain characters that are special to the shell. Special characters include space, tab, newline, NUL, and the following:

     " # $ & ' ( ) * ; < = > ? [ \ ` |

Also, file names should not begin with ‘~’ or ‘-’, and should contain neither ‘-’ immediately after ‘/’ nor ‘~’ immediately after ‘:’. On Posix-like platforms, directory names should not contain ‘:’, as this runs afoul of ‘:’ used as the path separator.

These restrictions apply not only to the files that you distribute, but also to the absolute file names of your source, build, and destination directories.

On some Posix-like platforms, ‘!’ and ‘^’ are special too, so they should be avoided. Posix lets implementations treat leading ‘//’ specially, but requires leading ‘///’ and

beyond to be equivalent to ‘/’. Most Unix variants treat ‘//’ like ‘/’. However, some treat ‘//’ as a “super-root” that can provide access to files that are not otherwise reachable from ‘/’. The super-root tradition began with Apollo Domain/OS, which died out long ago, but unfortunately Cygwin has revived it.

While autoconf and friends are usually run on some Posix variety, they can be used on other systems, most notably DOS variants. This impacts several assumptions regarding file names. For example, the following code:

     case $foo_dir in
       /*) # Absolute
          ;;
       *)
          foo_dir=$dots$foo_dir ;;
     esac

fails to properly detect absolute file names on those systems, because they can use a drivespec, and usually use a backslash as directory separator. If you want to be portable to DOS variants (at the price of rejecting valid but oddball Posix file names like ‘a:\b’), you can check for absolute file names like this:

     case $foo_dir in
       [\\/]* | ?:[\\/]* ) # Absolute
          ;;

       *)
          foo_dir=$dots$foo_dir ;;
     esac

Make sure you quote the brackets if appropriate and keep the backslash as first character (see [Limitations of Shell Builtins], page 212).

Also, because the colon is used as part of a drivespec, these systems don’t use it as path separator. When creating or accessing paths, you can use the PATH_SEPARATOR output variable instead. configure sets this to the appropriate value for the build system (‘:’ or ‘;’) when it starts up.

File names need extra care as well. While DOS variants that are Posixy enough to run autoconf (such as DJGPP) are usually able to handle long file names properly, there are still limitations that can seriously break packages. Several of these issues can be easily

detected by the doschk (ftp://ftp.gnu.org/gnu/non-gnu/doschk/doschk-1.1.tar.gz)

package.

A short overview follows; problems are marked with SFN/LFN to indicate where they apply: SFN means the issues are only relevant to plain DOS, not to DOS under Microsoft Windows variants, while LFN identifies problems that exist even under Microsoft Windows variants.

No multiple dots (SFN)

DOS cannot handle multiple dots in file names. This is an especially important thing to remember when building a portable configure script, as autoconf uses a .in suffix for template files. This is perfectly OK on Posix variants:

               AC_CONFIG_HEADERS([config.h])
               AC_CONFIG_FILES([source.c foo.bar])
               AC_OUTPUT

but it causes problems on DOS, as it requires ‘config.h.in’, ‘source.c.in’ and ‘foo.bar.in’. To make your package more portable to DOS-based environments, you should use this instead:

               AC_CONFIG_HEADERS([config.h:config.hin])
               AC_CONFIG_FILES([source.c:source.cin foo.bar:foobar.in])
               AC_OUTPUT

No leading dot (SFN)

DOS cannot handle file names that start with a dot. This is usually not important for autoconf.

Case insensitivity (LFN)

DOS is case insensitive, so you cannot, for example, have both a file called ‘INSTALL’ and a directory called ‘install’. This also affects make; if there’s a file called ‘INSTALL’ in the directory, ‘make install’ does nothing (unless the ‘install’ target is marked as PHONY).

The 8+3 limit (SFN)

Because the DOS file system only stores the first 8 characters of the file name and the first 3 of the extension, those must be unique. That means that

‘foobar-part1.c’, ‘foobar-part2.c’ and ‘foobar-prettybird.c’ all resolve

to the same file name (‘FOOBAR-P.C’). The same goes for ‘foo.bar’ and

‘foo.bartender’.

The 8+3 limit is not usually a problem under Microsoft Windows, as it uses numeric tails in the short version of file names to make them unique. However, a registry setting can turn this behavior off. While this makes it possible to share file trees containing long file names between SFN and LFN environments, it also means the above problem applies there as well.

Invalid characters (LFN)

Some characters are invalid in DOS file names, and should therefore be avoided. In a LFN environment, these are ‘/’, ‘\’, ‘?’, ‘*’, ‘:’, ‘<’, ‘>’, ‘|’ and ‘"’. In a SFN environment, other characters are also invalid. These include ‘+’, ‘,’, ‘[’ and ‘]’.

Invalid names (LFN)

Some DOS file names are reserved, and cause problems if you try to use files with those names. These names include ‘CON’, ‘AUX’, ‘COM1’, ‘COM2’, ‘COM3’, ‘COM4’, ‘LPT1’, ‘LPT2’, ‘LPT3’, ‘NUL’, and ‘PRN’. File names are case insensitive, so even names like ‘aux/config.guess’ are disallowed.

11.5 Shell Pattern Matching

Nowadays portable patterns can use negated character classes like ‘[!-aeiou]’. The older syntax ‘[^-aeiou]’ is supported by some shells but not others; hence portable scripts should never use ‘^’ as the first character of a bracket pattern.

Outside the C locale, patterns like ‘[a-z]’ are problematic since they may match characters that are not lower-case letters.

11.6 Shell Substitutions

Contrary to a persistent urban legend, the Bourne shell does not systematically split variables and back-quoted expressions, in particular on the right-hand side of assignments and in the argument of case. For instance, the following code:

case "$given_srcdir" in
.)  top_srcdir="`echo "$dots" | sed 's|/$||'`" ;;
*)  top_srcdir="$dots$given_srcdir" ;;
esac

is more readable when written as:

case $given_srcdir in
.)  top_srcdir=`echo "$dots" | sed 's|/$||'` ;;
*)  top_srcdir=$dots$given_srcdir ;;
esac

and in fact it is even more portable: in the first case of the first attempt, the computation of top_srcdir is not portable, since not all shells properly understand "`..."..."...`", for example Solaris 10 ksh:

$ foo="`echo " bar" | sed 's, ,,'`"

ksh: : cannot execute
ksh: bar | sed 's, ,,': cannot execute

Posix does not specify behavior for this sequence. On the other hand, behavior for "`...\"...\"...`" is specified by Posix, but in practice, not all shells understand it the same way: pdksh 5.2.14 prints spurious quotes when in Posix mode:

$ echo "`echo \"hello\"`"
hello
$ set -o posix
$ echo "`echo \"hello\"`"
"hello"

There is just no portable way to use double-quoted strings inside double-quoted back-quoted expressions (pfew!).

$@ One of the most famous shell-portability issues is related to ‘"$@"’. When there are no positional arguments, Posix says that ‘"$@"’ is supposed to be equivalent to nothing, but the original Unix version 7 Bourne shell treated it as equivalent to ‘""’ instead, and this behavior survives in later implementations like Digital Unix 5.0. The traditional way to work around this portability problem is to use ‘${1+"$@"}’. Unfortunately this method does not work with Zsh (3.x and 4.x), which is used on Mac OS X. When emulating the Bourne shell, Zsh performs word splitting on ‘${1+"$@"}’:

     zsh $ emulate sh
     zsh $ for i in "$@"; do echo $i; done
     Hello World
     !
     zsh $ for i in ${1+"$@"}; do echo $i; done
     Hello
     World
     !

Zsh handles plain ‘"$@"’ properly, but we can’t use plain ‘"$@"’ because of the portability problems mentioned above. One workaround relies on Zsh’s “global aliases” to convert ‘${1+"$@"}’ into ‘"$@"’ by itself:

     test "${ZSH_VERSION+set}" = set && alias -g '${1+"$@"}'='"$@"'

Zsh only recognizes this alias when a shell word matches it exactly; ‘"foo"${1+"$@"}’ remains subject to word splitting. Since this case always yields at least one shell word, use plain ‘"$@"’. A more conservative workaround is to avoid ‘"$@"’ if it is possible that there may be no positional arguments. For example, instead of:

     cat conftest.c "$@"

you can use this instead:

     case $# in
     0) cat conftest.c;;
     *) cat conftest.c "$@";;
     esac

Autoconf macros often use the set command to update ‘$@’, so if you are writing shell code intended for configure you should not assume that the value of ‘$@’ persists for any length of time.

${10} The 10th, 11th, … positional parameters can be accessed only after a shift. The 7th Edition shell reported an error if given ${10}, and Solaris 10 /bin/sh still acts that way:

     $ set 1 2 3 4 5 6 7 8 9 10
     $ echo ${10}
     bad substitution
${var:-value}

Old BSD shells, including the Ultrix sh, don’t accept the colon for any shell substitution, and complain and die. Similarly for ${var:=value}, ${var:?value}, etc.

${var=literal}

Be sure to quote:

     : ${var='Some words'}

otherwise some shells, such as on Digital Unix V 5.0, die because of a “bad substitution”.

Solaris /bin/sh has a frightening bug in its interpretation of this. Imagine you need set a variable to a string containing ‘}’. This ‘}’ character confuses Solaris /bin/sh when the affected variable was already set. This bug can be exercised by running:

     $ unset foo
     $ foo=${foo='}'}
     $ echo $foo
     }
     $ foo=${foo='}'   # no error; this hints to what the bug is
     $ echo $foo
     }
     $ foo=${foo='}'}
     $ echo $foo
     }}
      ^ugh!

It seems that ‘}’ is interpreted as matching ‘${’, even though it is enclosed in single quotes. The problem doesn’t happen using double quotes.

${var=expanded-value}

On Ultrix, running

     default="yu,yaa"
     : ${var="$default"}

sets var to ‘M-yM-uM-,M-yM-aM-a’, i.e., the 8th bit of each char is set. You don’t observe the phenomenon using a simple ‘echo $var’ since apparently the shell resets the 8th bit when it expands $var. Here are two means to make this shell confess its sins:

     $ cat -v <<EOF
     $var
     EOF

and

     $ set | grep '^var=' | cat -v

One classic incarnation of this bug is:

     default="a b c"
     : ${list="$default"}
     for c in $list; do
       echo $c
     done

You’ll get ‘a b c’ on a single line. Why? Because there are no spaces in ‘$list’: there are ‘M- ’, i.e., spaces with the 8th bit set, hence no IFS splitting is performed!!! One piece of good news is that Ultrix works fine with ‘: ${list=$default}’; i.e., if you don’t quote. The bad news is then that QNX 4.25 then sets list to the last item of default! The portable way out consists in using a double assignment, to switch the 8th bit twice on Ultrix:

     list=${list="$default"}

…but beware of the ‘}’ bug from Solaris (see above). For safety, use:

     test "${var+set}" = set || var={value}
${#var}
${var%word}
${var%%word}
${var#word}
${var##word}

Posix requires support for these usages, but they do not work with many traditional shells, e.g., Solaris 10 /bin/sh. Also, pdksh 5.2.14 mishandles some word forms. For example if ‘$1’ is ‘a/b’ and ‘$2’ is ‘a’, then ‘${1#$2}’ should yield ‘/b’, but with pdksh it yields the empty string.

`commands`

Posix requires shells to trim all trailing newlines from command output before substituting it, so assignments like ‘dir=`echo "$file" | tr a A`’ do not work as expected if ‘$file’ ends in a newline. While in general it makes no sense, do not substitute a single builtin with side effects, because Ash 0.2, trying to optimize, does not fork a subshell to perform the command. For instance, if you wanted to check that cd is silent, do not use ‘test -z "`cd /`"’ because the following can happen:

     $ pwd
     /tmp

     $ test -z "`cd /`" && pwd
     /

The result of ‘foo=`exit 1`’ is left as an exercise to the reader. The MSYS shell leaves a stray byte in the expansion of a double-quoted command substitution of a native program, if the end of the substitution is not aligned with the end of the double quote. This may be worked around by inserting another pair of quotes:

     $ echo "`printf 'foo\r\n'` bar" > broken
     $ echo "`printf 'foo\r\n'`"" bar" | cmp - broken
     - broken differ: char 4, line 1

Upon interrupt or SIGTERM, some shells may abort a command substitution, replace it with a null string, and wrongly evaluate the enclosing command before entering the trap or ending the script. This can lead to spurious errors:

     $ sh -c 'if test `sleep 5; echo hi` = hi; then echo yes; fi'
     $ ^C
     sh: test: hi: unexpected operator/operand

You can avoid this by assigning the command substitution to a temporary variable:

     $ sh -c 'res=`sleep 5; echo hi`
              if test "x$res" = xhi; then echo yes; fi'
     $ ^C
$(commands)

This construct is meant to replace ‘`commands`’, and it has most of the problems listed under `commands`. This construct can be nested while this is impossible to do portably with back quotes. Unfortunately it is not yet universally supported. Most notably, even recent releases of Solaris don’t support it:

     $ showrev -c /bin/sh | grep version
     Command version: SunOS 5.10 Generic 121005-03 Oct 2006
     $ echo $(echo blah)
     syntax error: `(' unexpected

nor does IRIX 6.5’s Bourne shell:

     $ uname -a
     IRIX firebird-image 6.5 07151432 IP22
     $ echo $(echo blah)
     $(echo blah)

If you do use ‘$(commands)’, make sure that the commands do not start with a parenthesis, as that would cause confusion with a different notation ‘$((expression))’ that in modern shells is an arithmetic expression not a command. To avoid the confusion, insert a space between the two opening parentheses. Avoid commands that contain unbalanced parentheses in here-documents, comments, or case statement patterns, as many shells mishandle them. For example, Bash 3.1, ‘ksh88’, pdksh 5.2.14, and Zsh 4.2.6 all mishandle the following valid command:

echo $(case x in x) echo hello;; esac)
$((expression))

Arithmetic expansion is not portable as some shells (most notably Solaris 10 /bin/sh) don’t support it. Among shells that do support ‘$(( ))’, not all of them obey the Posix rule that octal and hexadecimal constants must be recognized:

$ bash -c 'echo $(( 010 + 0x10 ))'
24
$ zsh -c 'echo $(( 010 + 0x10 ))'
26
$ zsh -c 'emulate sh; echo $(( 010 + 0x10 ))'
24
$ pdksh -c 'echo $(( 010 + 0x10 ))'
pdksh:  010 + 0x10 : bad number `0x10'
$ pdksh -c 'echo $(( 010 ))'
10

When it is available, using arithmetic expansion provides a noticeable speedup in script execution; but testing for support requires eval to avoid syntax errors. The following construct is used by AS_VAR_ARITH to provide arithmetic computation when all arguments are provided in decimal and without a leading zero, and all operators are properly quoted and appear as distinct arguments:

if ( eval 'test $(( 1 + 1 )) = 2' ) 2>/dev/null; then
  eval 'func_arith ()
  {
    func_arith_result=$(( $* ))
  }'
else
  func_arith ()
  {
    func_arith_result=`expr "$@"`
  }
fi
func_arith 1 + 1
foo=$func_arith_result

^Always quote ‘^’, otherwise traditional shells such as /bin/sh on Solaris 10 treat this like ‘|’.

11.7 Assignments

When setting several variables in a row, be aware that the order of the evaluation is undefined. For instance ‘foo=1 foo=2; echo $foo’ gives ‘1’ with Solaris /bin/sh, but ‘2’ with Bash. You must use ‘;’ to enforce the order: ‘foo=1; foo=2; echo $foo’.

Don’t rely on the following to find ‘subdir/program’:

PATH=subdir$PATH_SEPARATOR$PATH program

as this does not work with Zsh 3.0.6. Use something like this instead:

     (PATH=subdir$PATH_SEPARATOR$PATH; export PATH; exec program)

Don’t rely on the exit status of an assignment: Ash 0.2 does not change the status and propagates that of the last statement:

     $ false || foo=bar; echo $?
     1
     $ false || foo=`:`; echo $?
     0

and to make things even worse, QNX 4.25 just sets the exit status to 0 in any case:

     $ foo=`exit 1`; echo $?
     0

To assign default values, follow this algorithm:

1. If the default value is a literal and does not contain any closing brace, use:

         : ${var='my literal'}

2. If the default value contains no closing brace, has to be expanded, and the variable

being initialized is not intended to be IFS-split (i.e., it’s not a list), then use:

         : ${var="$default"}

3. If the default value contains no closing brace, has to be expanded, and the variable

being initialized is intended to be IFS-split (i.e., it’s a list), then use:

         var=${var="$default"}

4. If the default value contains a closing brace, then use:

         test "${var+set}" = set || var="has a '}'"

In most cases ‘var=${var="$default"}’ is fine, but in case of doubt, just use the last form. See Section 11.6 [Shell Substitutions], page 195, items ‘${var:-value}’ and ‘${var=value}’ for the rationale.

11.8 Parentheses in Shell Scripts

Beware of two opening parentheses in a row, as many shell implementations treat them specially. Posix requires that the command ‘((cat))’ must behave like ‘(cat)’, but many shells, including Bash and the Korn shell, treat ‘((cat))’ as an arithmetic expression equivalent to ‘let "cat"’, and may or may not report an error when they detect that ‘cat’ is not a number. As another example, ‘pdksh’ 5.2.14 misparses the following code:

     if ((true) || false); then
       echo ok
     fi

To work around this problem, insert a space between the two opening parentheses. There is a similar problem and workaround with ‘$((’; see Section 11.6 [Shell Substitutions], page 195.

11.9 Slashes in Shell Scripts

Unpatched Tru64 5.1 sh omits the last slash of command-line arguments that contain two trailing slashes:

     $ echo / // /// //// .// //.
     / / // /// ./ //.
     $ x=//
     $ eval "echo \$x"
     /
     $ set -x
     $ echo abc | tr -t ab //
     + echo abc
     + tr -t ab /
     /bc

Unpatched Tru64 4.0 sh adds a slash after ‘"$var"’ if the variable is empty and the second double-quote is followed by a word that begins and ends with slash:

     $ sh -xc 'p=; echo "$p"/ouch/'
     p=
     + echo //ouch/
     //ouch/

However, our understanding is that patches are available, so perhaps it’s not worth worrying about working around these horrendous bugs.

11.10 Special Shell Variables

Some shell variables should not be used, since they can have a deep influence on the behavior of the shell. In order to recover a sane behavior from the shell, some variables should be unset; M4sh takes care of this and provides fallback values, whenever needed, to cater for a very old ‘/bin/sh’ that does not support unset. (see Chapter 11 [Portable Shell Programming], page 187).

As a general rule, shell variable names containing a lower-case letter are safe; you can define and use these variables without worrying about their effect on the underlying system, and without worrying about whether the shell changes them unexpectedly. (The exception is the shell variable status, as described below.)

Here is a list of names that are known to cause trouble. This list is not exhaustive, but you should be safe if you avoid the name status and names containing only upper-case letters and underscores.

? Not all shells correctly reset ‘$?’ after conditionals (see [Limitations of Shell Builtins], page 217). Not all shells manage ‘$?’ correctly in shell functions (see Section 11.11 [Shell Functions], page 209) or in traps (see [Limitations of Shell Builtins], page 222). Not all shells reset ‘$?’ to zero after an empty command.

               $ bash -c 'false; $empty; echo $?'
               0
               $ zsh -c 'false; $empty; echo $?'
               1

_Many shells reserve ‘$_’ for various purposes, e.g., the name of the last command executed.

BIN_SH In Tru64, if BIN_SH is set to xpg4, subsidiary invocations of the standard shell conform to Posix.

CDPATH When this variable is set it specifies a list of directories to search when invoking cd with a relative file name that did not start with ‘./’ or ‘../’. Posix 1003.1-2001 says that if a nonempty directory name from CDPATH is used successfully, cd prints the resulting absolute file name. Unfortunately this output can break idioms like ‘abs=`cd src && pwd`’ because abs receives the name twice. Also, many shells do not conform to this part of Posix; for example, zsh prints the result only if a directory name other than ‘.’ was chosen from CDPATH. In practice the shells that have this problem also support unset, so you can work around the problem as follows:

     (unset CDPATH) >/dev/null 2>&1 && unset CDPATH

You can also avoid output by ensuring that your directory name is absolute or

anchored at ‘./’, as in ‘abs=`cd ./src && pwd`’.

Configure scripts use M4sh, which automatically unsets CDPATH if possible, so you need not worry about this problem in those scripts.

CLICOLOR_FORCE

When this variable is set, some implementations of tools like ls attempt to add color to their output via terminal escape sequences, even when the output is not directed to a terminal, and can thus cause spurious failures in scripts. Configure scripts use M4sh, which automatically unsets this variable.

DUALCASE In the MKS shell, case statements and file name generation are case-insensitive unless DUALCASE is nonzero. Autoconf-generated scripts export this variable when they start up.

ENV
MAIL
MAILPATH
PS1
PS2

PS4 These variables should not matter for shell scripts, since they are supposed to affect only interactive shells. However, at least one shell (the pre-3.0 UWIN Korn shell) gets confused about whether it is interactive, which means that (for example) a PS1 with a side effect can unexpectedly modify ‘$?’. To work around this bug, M4sh scripts (including ‘configure’ scripts) do something like this:

     (unset ENV) >/dev/null 2>&1 && unset ENV MAIL MAILPATH
     PS1='$ '
     PS2='> '
     PS4='+ '

(actually, there is some complication due to bugs in unset; see see [Limitations of Shell Builtins], page 224).

FPATH The Korn shell uses FPATH to find shell functions, so avoid FPATH in portable scripts. FPATH is consulted after PATH, but you still need to be wary of tests that use PATH to find whether a command exists, since they might report the wrong result if FPATH is also set.

GREP_OPTIONS

When this variable is set, some implementations of grep honor these options, even if the options include direction to enable colored output via terminal escape sequences, and the result can cause spurious failures when the output is not directed to a terminal. Configure scripts use M4sh, which automatically unsets this variable.

IFS Long ago, shell scripts inherited IFS from the environment, but this caused many problems so modern shells ignore any environment settings for IFS. Don’t set the first character of IFS to backslash. Indeed, Bourne shells use the first character (backslash) when joining the components in ‘"$@"’ and some shells then reinterpret (!) the backslash escapes, so you can end up with backspace and other strange characters. The proper value for IFS (in regular code, not when performing splits) is ‘SPCTABRET’. The first character is especially important, as it is used to join the arguments in ‘$*’; however, note that traditional shells, but also bash-2.04, fail to adhere to this and join with a space anyway. M4sh guarantees that IFS will have the default value at the beginning of a script, and many macros within autoconf rely on this setting. It is okay to use blocks of shell code that temporarily change the value of IFS in order to split on another character, but remember to restore it before expanding further macros. Unsetting IFS instead of resetting it to the default sequence is not suggested, since code that tries to save and restore the variable’s value will incorrectly reset it to an empty value, thus disabling field splitting:

     unset IFS
     # default separators used for field splitting
     save_IFS=$IFS
     IFS=:
     # ...
     IFS=$save_IFS
     # no field splitting performed
LANG
LC_ALL
LC_COLLATE
LC_CTYPE
LC_MESSAGES
LC_MONETARY
LC_NUMERIC
LC_TIME

You should set all these variables to ‘C’ because so much configuration code assumes the C locale and Posix requires that locale environment variables be set to ‘C’ if the C locale is desired; ‘configure’ scripts and M4sh do that for you. Export these variables after setting them.

LANGUAGE

LANGUAGE is not specified by Posix, but it is a GNU extension that overrides LC_ALL in some cases, so you (or M4sh) should set it too.

LC_ADDRESS
LC_IDENTIFICATION
LC_MEASUREMENT
LC_NAME
LC_PAPER
LC_TELEPHONE

These locale environment variables are GNU extensions. They are treated like their Posix brethren (LC_COLLATE, etc.) as described above.

LINENO Most modern shells provide the current line number in LINENO. Its value is the line number of the beginning of the current command. M4sh, and hence Autoconf, attempts to execute configure with a shell that supports LINENO. If no such shell is available, it attempts to implement LINENO with a Sed prepass that replaces each instance of the string $LINENO (not followed by an alphanumeric character) with the line’s number. In M4sh scripts you should execute AS_LINENO_PREPARE so that these workarounds are included in your script; configure scripts do this automatically in AC_INIT. You should not rely on LINENO within eval or shell functions, as the behavior differs in practice. The presence of a quoted newline within simple commands can alter which line number is used as the starting point for $LINENO substitutions within that command. Also, the possibility of the Sed prepass means that you should not rely on $LINENO when quoted, when in here-documents, or when line continuations are used. Subshells should be OK, though. In the following example, lines 1, 9, and 14 are portable, but the other instances of $LINENO do not have deterministic values:

$ cat lineno
echo 1. $LINENO
echo "2. $LINENO
3. $LINENO"
cat <<EOF
5. $LINENO
6. $LINENO
7. \$LINENO
EOF
( echo 9. $LINENO )
eval 'echo 10. $LINENO'
eval 'echo 11. $LINENO
echo 12. $LINENO'
echo 13. '$LINENO'
echo 14. $LINENO '
15.' $LINENO
f () { echo $1 $LINENO;
echo $1 $LINENO }
f 18.
echo 19. \
$LINENO
$ bash-3.2 ./lineno
1. 1
2. 3
3. 3
5. 4
6. 4
7. $LINENO
9. 9
10. 10
11. 12
12. 13
13. $LINENO
14. 14
15. 14
18. 16
18. 17
19. 19

$ zsh-4.3.4 ./lineno
1. 1
2. 2
3. 2
5. 4
6. 4
7. $LINENO
9. 9
10. 1
11. 1
12. 2
13. $LINENO
14. 14
15. 14
18. 0
18. 1
19. 19
$ pdksh-5.2.14 ./lineno
1. 1
2. 2
3. 2
5. 4
6. 4
7. $LINENO
9. 9
10. 0
11. 0
12. 0
13. $LINENO
14. 14
15. 14
18. 16
18. 17
19. 19

$ sed '=' <lineno |
>   sed '
>     N
>     s,$,-,
>     t loop
>     :loop
>     s,^\([0-9]*\)\(.*\)[$]LINENO\([^a-zA-Z0-9_]\),\1\2\1\3,
>     t loop
>     s,-$,,
>     s,^[0-9]*\n,,
>   ' |
>   sh
1. 1
2. 2
3. 3
5. 5
6. 6
7. \7
9. 9
10. 10
11. 11
12. 12
13. 13
14. 14
15. 15
18. 16
18. 17
19. 20

In particular, note that ‘config.status’ (and any other subsidiary script created by AS_INIT_GENERATED) might report line numbers relative to the parent script as a result of the potential Sed pass.

NULLCMD When executing the command ‘>foo’, zsh executes ‘$NULLCMD >foo’ unless it is operating in Bourne shell compatibility mode and the zsh version is newer than 3.1.6-dev-18. If you are using an older zsh and forget to set NULLCMD, your script might be suspended waiting for data on its standard input.

PATH_SEPARATOR

On DJGPP systems, the PATH_SEPARATOR environment variable can be set to either ‘:’ or ‘;’ to control the path separator Bash uses to set up certain environment variables (such as PATH). You can set this variable to ‘;’ if you want configure to use ‘;’ as a separator; this might be useful if you plan to use non-Posix shells to execute files. See Section 11.4 [File System Conventions], page 193, for more information about PATH_SEPARATOR.

PWD Posix 1003.1-2001 requires that cd and pwd must update the PWD environment variable to point to the logical name of the current directory, but traditional shells do not support this. This can cause confusion if one shell instance maintains PWD but a subsidiary and different shell does not know about PWD and executes cd; in this case PWD points to the wrong directory. Use ‘`pwd`’ rather than ‘$PWD’.

RANDOM Many shells provide RANDOM, a variable that returns a different integer each time it is used. Most of the time, its value does not change when it is not used, but on IRIX 6.5 the value changes all the time. This can be observed by using set. It is common practice to use $RANDOM as part of a file name, but code shouldn’t rely on $RANDOM expanding to a nonempty string.

status This variable is an alias to ‘$?’ for zsh (at least 3.1.6), hence read-only. Do not use it.

11.11 Shell Functions

Nowadays, it is difficult to find a shell that does not support shell functions at all. However, some differences should be expected.

Inside a shell function, you should not rely on the error status of a subshell if the last command of that subshell was exit or trap, as this triggers bugs in zsh 4.x; while Autoconf tries to find a shell that does not exhibit the bug, zsh might be the only shell present on the user’s machine.

Likewise, the state of ‘$?’ is not reliable when entering a shell function. This has the effect that using a function as the first command in a trap handler can cause problems.

     $ bash -c 'foo(){ echo $?; }; trap foo 0; (exit 2); exit 2'; echo $?
     2
     2
     $ ash -c 'foo(){ echo $?; }; trap foo 0; (exit 2); exit 2'; echo $?
     0
     2

DJGPP bash 2.04 has a bug in that return from a shell function which also used a command substitution causes a segmentation fault. To work around the issue, you can use return from a subshell, or ‘AS_SET_STATUS’ as last command in the execution flow of the function (see Section 9.1 [Common Shell Constructs], page 165).

Not all shells treat shell functions as simple commands impacted by ‘set -e’, for example with Solaris 10 bin/sh:

     $ bash -c 'f(){ return 1; }; set -e; f; echo oops
     $ /bin/sh -c 'f(){ return 1; }; set -e; f; echo oops
     oops

Shell variables and functions may share the same namespace, for example with Solaris

10 /bin/sh:
     $ f () { :; }; f=; f
     f: not found

For this reason, Autoconf (actually M4sh, see Chapter 9 [Programming in M4sh], page 165) uses the prefix ‘as_fn_’ for its functions.

Handling of positional parameters and shell options varies among shells. For example, Korn shells reset and restore trace output (‘set -x’) and other options upon function entry and exit. Inside a function, IRIX sh sets ‘$0’ to the function name.

It is not portable to pass temporary environment variables to shell functions. Solaris /bin/sh does not see the variable. Meanwhile, not all shells follow the Posix rule that the assignment must affect the current environment in the same manner as special built-ins.

     $ /bin/sh -c 'func(){ echo $a;}; a=1 func; echo $a'

⇒ ⇒

     $ ash -c 'func(){ echo $a;}; a=1 func; echo $a'
     ⇒1

⇒

     $ bash -c 'set -o posix; func(){ echo $a;}; a=1 func; echo $a'
     ⇒1
     ⇒1

Some ancient Bourne shell variants with function support did not reset ‘$i, i >= 0’, upon function exit, so effectively the arguments of the script were lost after the first function invocation. It is probably not worth worrying about these shells any more.

With AIX sh, a trap on 0 installed in a shell function triggers at function exit rather than at script exit, see See [Limitations of Shell Builtins], page 222.

11.12 Limitations of Shell Builtins

No, no, we are serious: some shells do have limitations! :)

You should always keep in mind that any builtin or command may support options, and therefore differ in behavior with arguments starting with a dash. For instance, even the innocent ‘echo "$word"’ can give unexpected results when word starts with a dash. It is often possible to avoid this problem using ‘echo "x$word"’, taking the ‘x’ into account later in the pipe. Many of these limitations can be worked around using M4sh (see Chapter 9 [Programming in M4sh], page 165).

. Use . only with regular files (use ‘test -f’). Bash 2.03, for instance, chokes on ‘. /dev/null’. Remember that . uses PATH if its argument contains no slashes. Also, some shells, including bash 3.2, implicitly append the current directory to this PATH search, even though Posix forbids it. So if you want to use . on a file ‘foo’ in the current directory, you must use ‘. ./foo’. Not all shells gracefully handle syntax errors within a sourced file. On one extreme, some non-interactive shells abort the entire script. On the other, zsh 4.3.10 has a bug where it fails to react to the syntax error.

               $ echo 'fi' > syntax
               $ bash -c '. ./syntax; echo $?'
               ./syntax: line 1: syntax error near unexpected token `fi'
               ./syntax: line 1: `fi'
               1
               $ ash -c '. ./syntax; echo $?'
               ./syntax: 1: Syntax error: "fi" unexpected
               $ zsh -c '. ./syntax; echo $?'
               ./syntax:1: parse error near `fi'

! The Unix version 7 shell did not support negating the exit status of commands with !, and this feature is still absent from some shells (e.g., Solaris /bin/sh). Other shells, such as FreeBSD /bin/sh or ash, have bugs when using !:

$ sh -c '! : | :'; echo $?
1
$ ash -c '! : | :'; echo $?
0
$ sh -c '! { :; }'; echo $?
1
$ ash -c '! { :; }'; echo $?
{: not found
Syntax error: "}" unexpected
2

Shell code like this:

if ! cmp file1 file2 >/dev/null 2>&1; then
  echo files differ or trouble
fi

is therefore not portable in practice. Typically it is easy to rewrite such code, e.g.:

cmp file1 file2 >/dev/null 2>&1 ||
  echo files differ or trouble

More generally, one can always rewrite ‘! command’ as:

if command; then (exit 1); else :; fi

{...} Bash 3.2 (and earlier versions) sometimes does not properly set ‘$?’ when failing to write redirected output of a compound command. This problem is most commonly observed with ‘{...}’; it does not occur with ‘(...)’. For example:

$ bash -c '{ echo foo; } >/bad; echo $?'
bash: line 1: /bad: Permission denied
0
$ bash -c 'while :; do echo; done >/bad; echo $?'
bash: line 1: /bad: Permission denied
0

To work around the bug, prepend ‘:;’:

$ bash -c ':;{ echo foo; } >/bad; echo $?'
bash: line 1: /bad: Permission denied
1

Posix requires a syntax error if a brace list has no contents. However, not all shells obey this rule; and on shells where empty lists are permitted, the effect on ‘$?’ is inconsistent. To avoid problems, ensure that a brace list is never empty.

$ bash -c 'false; { }; echo $?' || echo $?
bash: line 1: syntax error near unexpected token `}'
bash: line 1: `false; { }; echo $?'

     $ zsh -c 'false; { }; echo $?' || echo $?
     1
     $ pdksh -c 'false; { }; echo $?' || echo $?
     0

break The use of ‘break 2’ etc. is safe.

case You don’t need to quote the argument; no splitting is performed. You don’t need the final ‘;;’, but you should use it. Posix requires support for case patterns with opening parentheses like this:

     case $file_name in
       (*.c) echo "C source code";;
     esac

but the ( in this example is not portable to many Bourne shell implementations, which is a pity for those of us using tools that rely on balanced parentheses. For instance, with Solaris /bin/sh:

     $ case foo in (foo) echo foo;; esac
      error syntax error: `(' unexpected

The leading ‘(’ can be omitted safely. Unfortunately, there are contexts where unbalanced parentheses cause other problems, such as when using a syntaxhighlighting editor that searches for the balancing counterpart, or more importantly, when using a case statement as an underquoted argument to an Autoconf macro. See Section 8.1.7 [Balancing Parentheses], page 127, for tradeoffs involved in various styles of dealing with unbalanced ‘)’. Zsh handles pattern fragments derived from parameter expansions or command substitutions as though quoted:

     $ pat=\?; case aa in ?$pat) echo match;; esac
     $ pat=\?; case a? in ?$pat) echo match;; esac
     match

Because of a bug in its fnmatch, Bash fails to properly handle backslashes in character classes:

     bash-2.02$ case /tmp in [/\\]*) echo OK;; esac
     bash-2.02$

This is extremely unfortunate, since you are likely to use this code to handle Posix or MS-DOS absolute file names. To work around this bug, always put the backslash first:

     bash-2.02$ case '\TMP' in [\\/]*) echo OK;; esac
     OK
     bash-2.02$ case /tmp in [\\/]*) echo OK;; esac
     OK

Many Bourne shells cannot handle closing brackets in character classes correctly. Some shells also have problems with backslash escaping in case you do not want to match the backslash: both a backslash and the escaped character match this pattern. To work around this, specify the character class in a variable, so that quote removal does not apply afterwards, and the special characters don’t have to be backslash-escaped:

     $ case '\' in [\<]) echo OK;; esac
     OK
     $ scanset='[<]'; case '\' in $scanset) echo OK;; esac
     $

Even with this, Solaris ksh matches a backslash if the set contains any of the characters ‘|’, ‘&’, ‘(’, or ‘)’. Conversely, Tru64 ksh (circa 2003) erroneously always matches a closing parenthesis if not specified in a character class:

     $ case foo in *\)*) echo fail ;; esac
     fail
     $ case foo in *')'*) echo fail ;; esac
     fail

Some shells, such as Ash 0.3.8, are confused by an empty case/esac:

     ash-0.3.8 $ case foo in esac;
      error Syntax error: ";" unexpected (expecting ")")

Posix requires case to give an exit status of 0 if no cases match. However, /bin/sh in Solaris 10 does not obey this rule. Meanwhile, it is unclear whether a case that matches, but contains no statements, must also change the exit status to 0. The M4sh macro AS_CASE works around these inconsistencies.

     $ bash -c 'case `false` in ?) ;; esac; echo $?'
     0
     $ /bin/sh -c 'case `false` in ?) ;; esac; echo $?'
     255

cd Posix 1003.1-2001 requires that cd must support the ‘-L’ (“logical”) and ‘-P’ (“physical”) options, with ‘-L’ being the default. However, traditional shells do not support these options, and their cd command has the ‘-P’ behavior. Portable scripts should assume neither option is supported, and should assume neither behavior is the default. This can be a bit tricky, since the Posix default behavior means that, for example, ‘ls ..’ and ‘cd ..’ may refer to different directories if the current logical directory is a symbolic link. It is safe to use cd dir if dir contains no ‘..’ components. Also, Autoconf-generated scripts check for this problem when computing variables like ac_top_srcdir (see Section 4.6 [Configuration Actions], page 23), so it is safe to cd to these variables. See See Section 11.10 [Special Shell Variables], page 202, for portability problems involving cd and the CDPATH environment variable. Also please see the discussion of the pwd command.

echo The simple echo is probably the most surprising source of portability troubles. It is not possible to use ‘echo’ portably unless both options and escape sequences are omitted. Don’t expect any option. Do not use backslashes in the arguments, as there is no consensus on their handling. For ‘echo '\n' | wc -l’, the sh of Solaris outputs 2, but Bash and Zsh (in sh emulation mode) output 1. The problem is truly echo: all the shells understand ‘'\n'’ as the string composed of a backslash and an ‘n’. Within a command substitution, ‘echo 'string\c'’ will mess up the internal state of ksh88 on AIX 6.1 so that it will print the first character ‘s’ only, followed by a newline, and then entirely drop the output of the next echo in a command substitution. Because of these problems, do not pass a string containing arbitrary characters to echo. For example, ‘echo "$foo"’ is safe only if you know that foo’s value cannot contain backslashes and cannot start with ‘-’. If this may not be true, printf is in general safer and easier to use than echo and echo -n. Thus, scripts where portability is not a major concern should use printf '%s\n' whenever echo could fail, and similarly use printf %s instead of echo -n. For portable shell scripts, instead, it is suggested to use a heredocument like this:

     cat <<EOF
     $foo
     EOF

Alternatively, M4sh provides AS_ECHO and AS_ECHO_N macros which choose between various portable implementations: ‘echo’ or ‘print’ where they work, printf if it is available, or else other creative tricks in order to work around the above problems.

eval The eval command is useful in limited circumstances, e.g., using commands

like ‘eval table_$key=\$value’ and ‘eval value=table_$key’ to simulate a

hash table when the key is known to be alphanumeric. You should also be wary of common bugs in eval implementations. In some shell implementations (e.g., older ash, OpenBSD 3.8 sh, pdksh v5.2.14 99/07/13.2, and zsh 4.2.5), the arguments of ‘eval’ are evaluated in a context where ‘$?’ is 0, so they exhibit behavior like this:

     $ false; eval 'echo $?'
     0

The correct behavior here is to output a nonzero value, but portable scripts should not rely on this. You should not rely on LINENO within eval. See Section 11.10 [Special Shell Variables], page 202. Note that, even though these bugs are easily avoided, eval is tricky to use on arbitrary arguments. It is obviously unwise to use ‘eval $cmd’ if the string value of ‘cmd’ was derived from an untrustworthy source. But even if the string value is valid, ‘eval $cmd’ might not work as intended, since it causes field splitting and file name expansion to occur twice, once for the eval and once for the command itself. It is therefore safer to use ‘eval "$cmd"’. For example, if cmd has the value ‘cat test?.c’, ‘eval $cmd’ might expand to the equivalent of ‘cat test;.c’ if there happens to be a file named ‘test;.c’ in the current directory; and this in turn mistakenly attempts to invoke cat on the file ‘test’ and then execute the command .c. To avoid this problem, use ‘eval "$cmd"’ rather than ‘eval $cmd’. However, suppose that you want to output the text of the evaluated command just before executing it. Assuming the previous example, ‘echo "Executing:

$cmd"’ outputs ‘Executing: cat test?.c’, but this output doesn’t show the user that ‘test;.c’ is the actual name of the copied file. Conversely, ‘eval "echo Executing: $cmd"’ works on this example, but it fails with ‘cmd='cat foo >bar'’, since it mistakenly replaces the contents of ‘bar’ by the string ‘cat foo’. No simple, general, and portable solution to this problem is known.

exec Posix describes several categories of shell built-ins. Special built-ins (such as exit) must impact the environment of the current shell, and need not be available through exec. All other built-ins are regular, and must not propagate variable assignments to the environment of the current shell. However, the group of regular built-ins is further distinguished by commands that do not require a PATH search (such as cd), in contrast to built-ins that are offered as a more efficient version of something that must still be found in a PATH search (such as echo). Posix is not clear on whether exec must work with the list of 17 utilities that are invoked without a PATH search, and many platforms lack an executable for some of those built-ins:

     $ sh -c 'exec cd /tmp'
     sh: line 0: exec: cd: not found

All other built-ins that provide utilities specified by Posix must have a counterpart executable that exists on PATH, although Posix allows exec to use the built-in instead of the executable. For example, contrast bash 3.2 and pdksh 5.2.14:

     $ bash -c 'pwd --version' | head -n1
     bash: line 0: pwd: --: invalid option
     pwd: usage: pwd [-LP]
     $ bash -c 'exec pwd --version' | head -n1
     pwd (GNU coreutils) 6.10
     $ pdksh -c 'exec pwd --version' | head -n1
     pdksh: pwd: --: unknown option

When it is desired to avoid a regular shell built-in, the workaround is to use some other forwarding command, such as env or nice, that will ensure a path search:

     $ pdksh -c 'exec true --version' | head -n1
     $ pdksh -c 'nice true --version' | head -n1
     true (GNU coreutils) 6.10
     $ pdksh -c 'env true --version' | head -n1
     true (GNU coreutils) 6.10

exit The default value of exit is supposed to be $?; unfortunately, some shells, such

as the DJGPP port of Bash 2.04, just perform ‘exit 0’.

     bash-2.04$ foo=`exit 1` || echo fail
     fail
     bash-2.04$ foo=`(exit 1)` || echo fail
     fail
     bash-2.04$ foo=`(exit 1); exit` || echo fail
     bash-2.04$

Using ‘exit $?’ restores the expected behavior.

Some shell scripts, such as those generated by autoconf, use a trap to clean up before exiting. If the last shell command exited with nonzero status, the trap also exits with nonzero status so that the invoker can tell that an error occurred. Unfortunately, in some shells, such as Solaris /bin/sh, an exit trap ignores the exit command’s argument. In these shells, a trap cannot determine whether it was invoked by plain exit or by exit 1. Instead of calling exit directly, use the AC_MSG_ERROR macro that has a workaround for this problem.

export The builtin export dubs a shell variable environment variable. Each update of exported variables corresponds to an update of the environment variables. Conversely, each environment variable received by the shell when it is launched should be imported as a shell variable marked as exported. Alas, many shells, such as Solaris /bin/sh, IRIX 6.3, IRIX 5.2, AIX 4.1.5, and Digital Unix 4.0, forget to export the environment variables they receive. As a result, two variables coexist: the environment variable and the shell variable. The following code demonstrates this failure:

     #!/bin/sh
     echo $FOO
     FOO=bar
     echo $FOO
     exec /bin/sh $0

when run with ‘FOO=foo’ in the environment, these shells print alternately ‘foo’ and ‘bar’, although they should print only ‘foo’ and then a sequence of ‘bar’s. Therefore you should export again each environment variable that you update; the export can occur before or after the assignment. Posix is not clear on whether the export of an undefined variable causes the variable to be defined with the value of an empty string, or merely marks any future definition of a variable by that name for export. Various shells behave differently in this regard:

     $ sh -c 'export foo; env | grep foo'
     $ ash -c 'export foo; env | grep foo'
     foo=

false Don’t expect false to exit with status 1: in native Solaris ‘/bin/false’ exits with status 255.

for To loop over positional arguments, use:

     for arg
     do
       echo "$arg"
     done

You may not leave the do on the same line as for, since some shells improperly grok:

     for arg; do
       echo "$arg"

     done

If you want to explicitly refer to the positional arguments, given the ‘$@’ bug (see Section 11.6 [Shell Substitutions], page 195), use:

     for arg in ${1+"$@"}; do
       echo "$arg"
     done

But keep in mind that Zsh, even in Bourne shell emulation mode, performs word splitting on ‘${1+"$@"}’; see Section 11.6 [Shell Substitutions], page 195, item ‘$@’, for more. In Solaris /bin/sh, when the list of arguments of a for loop starts with unquoted tokens looking like variable assignments, the loop is not executed on those tokens:

     $ /bin/sh -c 'for v in a=b c=d x e=f; do echo $v; done'
     x
     e=f

Thankfully, quoting the assignment-like tokens, or starting the list with other tokens (including unquoted variable expansion that results in an assignment-like result), avoids the problem, so it is easy to work around:

     $ /bin/sh -c 'for v in "a=b"; do echo $v; done'
     a=b
     $ /bin/sh -c 'x=a=b; for v in $x c=d; do echo $v; done'
     a=b
     c=d

if Using ‘!’ is not portable. Instead of:

     if ! cmp -s file file.new; then
       mv file.new file
     fi

use:

     if cmp -s file file.new; then :; else
       mv file.new file
     fi

Or, especially if the else branch is short, you can use ||. In M4sh, the AS_IF macro provides an easy way to write these kinds of conditionals:

     AS_IF([cmp -s file file.new], [], [mv file.new file])

This is especially useful in other M4 macros, where the then and else branches might be macro arguments. Some very old shells did not reset the exit status from an if with no else:

     $ if (exit 42); then true; fi; echo $?
     42

whereas a proper shell should have printed ‘0’. But this is no longer a portability problem; any shell that supports functions gets it correct. However, it explains why some makefiles have lengthy constructs:

     if test -f "$file"; then
       install "$file" "$dest"
     else
       :
     fi

printf A format string starting with a ‘-’ can cause problems. Bash interprets it as an option and gives an error. And ‘--’ to mark the end of options is not good in the NetBSD Almquist shell (e.g., 0.4.6) which takes that literally as the format string. Putting the ‘-’ in a ‘%c’ or ‘%s’ is probably easiest:

     printf %s -foo

Bash 2.03 mishandles an escape sequence that happens to evaluate to ‘%’:

     $ printf '\045'
     bash: printf: `%': missing format character

Large outputs may cause trouble. On Solaris 2.5.1 through 10, for example, ‘/usr/bin/printf’ is buggy, so when using /bin/sh the command ‘printf %010000x 123’ normally dumps core. Since printf is not always a shell builtin, there is a potential speed penalty for using printf '%s\n' as a replacement for an echo that does not interpret ‘\’ or leading ‘-’. With Solaris ksh, it is possible to use print -r -- for this role instead. For a discussion of portable alternatives to both printf and echo, See [Limitations of Shell Builtins], page 213.

pwd With modern shells, plain pwd outputs a “logical” directory name, some of whose components may be symbolic links. These directory names are in contrast to “physical” directory names, whose components are all directories. Posix 1003.1-2001 requires that pwd must support the ‘-L’ (“logical”) and ‘-P’ (“physical”) options, with ‘-L’ being the default. However, traditional shells do not support these options, and their pwd command has the ‘-P’ behavior. Portable scripts should assume neither option is supported, and should assume neither behavior is the default. Also, on many hosts ‘/bin/pwd’ is equivalent to ‘pwd -P’, but Posix does not require this behavior and portable scripts should not rely on it. Typically it’s best to use plain pwd. On modern hosts this outputs logical directory names, which have the following advantages:

• Logical names are what the user specified.

• Physical names may not be portable from one installation host to another due to network file system gymnastics.

• On modern hosts ‘pwd -P’ may fail due to lack of permissions to some parent directory, but plain pwd cannot fail for this reason.

Also please see the discussion of the cd command.

read No options are portable, not even support ‘-r’ (Solaris /bin/sh for example).

set With the FreeBSD 6.0 shell, the set command (without any options) does not sort its output.

The set builtin faces the usual problem with arguments starting with a dash. Modern shells such as Bash or Zsh understand ‘--’ to specify the end of the options (any argument after ‘--’ is a parameter, even ‘-x’ for instance), but many traditional shells (e.g., Solaris 10 /bin/sh) simply stop option processing as soon as a non-option argument is found. Therefore, use ‘dummy’ or simply ‘x’ to end the option processing, and use shift to pop it out:

     set x $my_list; shift

Avoid ‘set -’, e.g., ‘set - $my_list’. Posix no longer requires support for this command, and in traditional shells ‘set - $my_list’ resets the ‘-v’ and ‘-x’ options, which makes scripts harder to debug. Some nonstandard shells do not recognize more than one option (e.g., ‘set -e -x’ assigns ‘-x’ to the command line). It is better to combine them:

     set -ex

The option ‘-e’ has historically been underspecified, with enough ambiguities to cause numerous differences across various shell implementations. Perhaps

the best reference is this link (http://www.opengroup.org/austin/
mailarchives/ag-review/msg03507.html), recommending a change to Posix

2008 to match ksh88 behavior. Note that mixing set -e and shell functions is asking for surprises:

     set -e
     doit()
     {
       rm file
       echo one
     }
     doit || echo two

According to the recommendation, ‘one’ should always be output regardless of whether the rm failed, because it occurs within the body of the shell function ‘doit’ invoked on the left side of ‘||’, where the effects of ‘set -e’ are not enforced. Likewise, ‘two’ should never be printed, since the failure of rm does not abort the function, such that the status of ‘doit’ is 0. The BSD shell has had several problems with the ‘-e’ option. Older versions of the BSD shell (circa 1990) mishandled ‘&&’, ‘||’, ‘if’, and ‘case’ when ‘-e’ was in effect, causing the shell to exit unexpectedly in some cases. This was particularly a problem with makefiles, and led to circumlocutions like ‘sh -c 'test -f file || touch file'’, where the seemingly-unnecessary ‘sh -c '...'’ wrapper works around the bug (see Section 12.2 [Failure in Make Rules], page 241). Even relatively-recent versions of the BSD shell (e.g., OpenBSD 3.4) wrongly exit with ‘-e’ if the last command within a compound statement fails and is guarded by an ‘&&’ only. For example:

     #! /bin/sh
     set -e
     foo=''
     test -n "$foo" && exit 1
     echo one

if :; then
  test -n "$foo" && exit 1
  echo two
  test -n "$foo" && exit 1
fi
echo three

does not print ‘three’. One workaround is to change the last instance of

‘test -n "$foo" && exit 1’ to be ‘if test -n "$foo"; then exit 1; fi’ in-

stead. Another possibility is to warn BSD users not to use ‘sh -e’. When ‘set -e’ is in effect, a failed command substitution in Solaris /bin/sh cannot be ignored, even with ‘||’.

$ /bin/sh -c 'set -e; foo=`false` || echo foo; echo bar'
$ bash -c 'set -e; foo=`false` || echo foo; echo bar'
foo
bar

Moreover, a command substitution, successful or not, causes this shell to exit from a failing outer command even in presence of an ‘&&’ list:

$ bash -c 'set -e; false `true` && echo notreached; echo ok'
ok
$ sh -c 'set -e; false `true` && echo notreached; echo ok'
$

Portable scripts should not use ‘set -e’ if trap is used to install an exit handler. This is because Tru64/OSF 5.1 sh sometimes enters the trap handler with the exit status of the command prior to the one that triggered the errexit handler:

$ sh -ec 'trap '\''echo $?'\'' 0; false'
0
$ sh -c 'set -e; trap '\''echo $?'\'' 0; false'
1

Thus, when writing a script in M4sh, rather than trying to rely on ‘set -e’, it is better to append ‘|| AS_EXIT’ to any statement where it is desirable to abort on failure. Job control is not provided by all shells, so the use of ‘set -m’ or ‘set -b’ must be done with care. When using zsh in native mode, asynchronous notification (‘set -b’) is enabled by default, and using ‘emulate sh’ to switch to Posix mode does not clear this setting (although asynchronous notification has no impact unless job monitoring is also enabled). Also, zsh 4.3.10 and earlier have a bug where job control can be manipulated in interactive shells, but not in subshells or scripts. Furthermore, some shells, like pdksh, fail to treat subshells as interactive, even though the parent shell was.

$ echo $ZSH_VERSION
4.3.10
$ set -m; echo $?
0
$ zsh -c 'set -m; echo $?'
set: can't change option: -m

     $ (set -m); echo $?
     set: can't change option: -m
     1
     $ pdksh -ci 'echo $-; (echo $-)'
     cim
     c

shift Not only is shifting a bad idea when there is nothing left to shift, but in addition it is not portable: the shell of MIPS RISC/OS 4.52 refuses to do it.

Don’t use ‘shift 2’ etc.; while it in the SVR1 shell (1983), it is also absent in many pre-Posix shells.

source This command is not portable, as Posix does not require it; use . instead.

test The test program is the way to perform many file and string tests. It is often invoked by the alternate name ‘[’, but using that name in Autoconf code is asking for trouble since it is an M4 quote character.

The ‘-a’, ‘-o’, ‘(’, and ‘)’ operands are not present in all implementations, and have been marked obsolete by Posix 2008. This is because there are inherent ambiguities in using them. For example, ‘test "$1" -a "$2"’ looks like a binary operator to check whether two strings are both non-empty, but if ‘$1’ is the literal ‘!’, then some implementations of test treat it as a negation of the unary operator ‘-a’.

Thus, portable uses of test should never have more than four arguments, and scripts should use shell constructs like ‘&&’ and ‘||’ instead. If you combine ‘&&’ and ‘||’ in the same statement, keep in mind that they have equal precedence, so it is often better to parenthesize even when this is redundant. For example:

      # Not portable:
      test "X$a" = "X$b" -a \
        '(' "X$c" != "X$d" -o "X$e" = "X$f" ')'
      # Portable:
      test "X$a" = "X$b" &&
        { test "X$c" != "X$d" || test "X$e" = "X$f"; }

test does not process options like most other commands do; for example, it does not recognize the ‘--’ argument as marking the end of options.

It is safe to use ‘!’ as a test operator. For example, ‘if test ! -d foo; ...’ is portable even though ‘if ! test -d foo; ...’ is not.

test (files)

To enable configure scripts to support cross-compilation, they shouldn’t do anything that tests features of the build system instead of the host system. But occasionally you may find it necessary to check whether some arbitrary file exists. To do so, use ‘test -f’ or ‘test -r’. Do not use ‘test -x’, because 4.3BSD does not have it. Do not use ‘test -e’ either, because Solaris /bin/sh lacks it. To test for symbolic links on systems that have them, use ‘test -h’ rather than ‘test -L’; either form conforms to Posix 1003.1-2001, but older shells like Solaris 8 /bin/sh support only ‘-h’. test (strings)

Posix says that ‘test "string"’ succeeds if string is not null, but this usage is not portable to traditional platforms like Solaris 10 /bin/sh, which mishandle strings like ‘!’ and ‘-n’.

Posix also says that ‘test ! "string"’, ‘test -n "string"’ and ‘test -z

"string"’ work with any string, but many shells (such as Solaris, AIX 3.2, UNICOS 10.0.0.6, Digital Unix 4, etc.) get confused if string looks like an operator:

     $ test -n =
     test: argument expected
     $ test ! -n
     test: argument expected

Similarly, Posix says that both ‘test "string1" = "string2"’ and ‘test "string1" != "string2"’ work for any pairs of strings, but in practice this is not true for troublesome strings that look like operators or parentheses, or that begin with ‘-’. It is best to protect such strings with a leading ‘X’, e.g., ‘test "Xstring" !=

X’ rather than ‘test -n "string"’ or ‘test ! "string"’.

It is common to find variations of the following idiom:

     test -n "`echo $ac_feature | sed 's/[-a-zA-Z0-9_]//g'`" &&
       action

to take an action when a token matches a given pattern. Such constructs should be avoided by using:

     case $ac_feature in
       *[!-a-zA-Z0-9_]*) action;;
     esac

If the pattern is a complicated regular expression that cannot be expressed as a shell pattern, use something like this instead:

     expr "X$ac_feature" : 'X.*[^-a-zA-Z0-9_]' >/dev/null &&
       action
‘expr "Xfoo" : "Xbar"’ is more robust than ‘echo "Xfoo" | grep "^Xbar"’,

because it avoids problems when ‘foo’ contains backslashes.

trap It is safe to trap at least the signals 1, 2, 13, and 15. You can also trap 0, i.e., have the trap run when the script ends (either via an explicit exit, or the end of the script). The trap for 0 should be installed outside of a shell function, or AIX 5.3 /bin/sh will invoke the trap at the end of this function. Posix says that ‘trap - 1 2 13 15’ resets the traps for the specified signals to their default values, but many common shells (e.g., Solaris /bin/sh) misinterpret this and attempt to execute a “command” named - when the specified conditions arise. Posix 2008 also added a requirement to support ‘trap 1 2 13 15’ to reset traps, as this is supported by a larger set of shells, but there are still shells like dash that mistakenly try to execute 1 instead of resetting the traps. Therefore, there is no portable workaround, except for ‘trap - 0’, for which ‘trap '' 0’ is a portable substitute.

Although Posix is not absolutely clear on this point, it is widely admitted that when entering the trap ‘$?’ should be set to the exit status of the last command run before the trap. The ambiguity can be summarized as: “when the trap is launched by an exit, what is the last command run: that before exit, or exit itself?” Bash considers exit to be the last command, while Zsh and Solaris /bin/sh consider that when the trap is run it is still in the exit, hence it is the previous exit status that the trap receives:

$ cat trap.sh
trap 'echo $?' 0
(exit 42); exit 0
$ zsh trap.sh
42
$ bash trap.sh
0

The portable solution is then simple: when you want to ‘exit 42’, run ‘(exit 42); exit 42’, the first exit being used to set the exit status to 42 for Zsh, and the second to trigger the trap and pass 42 as exit status for Bash. In M4sh, this is covered by using AS_EXIT. The shell in FreeBSD 4.0 has the following bug: ‘$?’ is reset to 0 by empty lines if the code is inside trap.

$ trap 'false
echo $?' 0
$ exit
0

Fortunately, this bug only affects trap. Several shells fail to execute an exit trap that is defined inside a subshell, when the last command of that subshell is not a builtin. A workaround is to use ‘exit $?’ as the shell builtin.

$ bash -c '(trap "echo hi" 0; /bin/true)'
hi
$ /bin/sh -c '(trap "echo hi" 0; /bin/true)'
$ /bin/sh -c '(trap "echo hi" 0; /bin/true; exit $?)'
hi

Likewise, older implementations of bash failed to preserve ‘$?’ across an exit trap consisting of a single cleanup command.

$ bash -c 'trap "/bin/true" 0; exit 2'; echo $?
2
$ bash-2.05b -c 'trap "/bin/true" 0; exit 2'; echo $?
0
$ bash-2.05b -c 'trap ":; /bin/true" 0; exit 2'; echo $?
2

true Don’t worry: as far as we know true is portable. Nevertheless, it’s not always a builtin (e.g., Bash 1.x), and the portable shell community tends to prefer using :. This has a funny side effect: when asked whether false is more portable than true Alexandre Oliva answered:

In a sense, yes, because if it doesn’t exist, the shell will produce an exit status of failure, which is correct for false, but not for true.

unset In some nonconforming shells (e.g., Bash 2.05a), unset FOO fails when FOO is

not set. You can use

      FOO=; unset FOO

if you are not sure that FOO is set. A few ancient shells lack unset entirely. For some variables such as PS1, you can use a neutralizing value instead:

      PS1='$ '

Usually, shells that do not support unset need less effort to make the environment sane, so for example is not a problem if you cannot unset CDPATH on those shells. However, Bash 2.01 mishandles unset MAIL in some cases and dumps core. So, you should do something like

      ( (unset MAIL) || exit 1) >/dev/null 2>&1 && unset MAIL || :

See Section 11.10 [Special Shell Variables], page 202, for some neutralizing values. Also, see [Limitations of Builtins], page 216, for the case of environment variables.

wait The exit status of wait is not always reliable.

11.13 Limitations of Usual Tools

The small set of tools you can expect to find on any machine can still include some limitations you should be aware of.

awk Don’t leave white space before the opening parenthesis in a user function call.

Posix does not allow this and GNU Awk rejects it:

     $ gawk 'function die () { print "Aaaaarg!"  }
             BEGIN { die () }'
     gawk: cmd. line:2:         BEGIN { die () }
     gawk: cmd. line:2:                      ^parse error
     $ gawk 'function die () { print "Aaaaarg!"  }
             BEGIN { die() }'
     Aaaaarg!

Posix says that if a program contains only ‘BEGIN’ actions, and contains no instances of getline, then the program merely executes the actions without reading input. However, traditional Awk implementations (such as Solaris 10 awk) read and discard input in this case. Portable scripts can redirect input from ‘/dev/null’ to work around the problem. For example:

     awk 'BEGIN {print "hello world"}' </dev/null

Posix says that in an ‘END’ action, ‘$NF’ (and presumably, ‘$1’) retain their value from the last record read, if no intervening ‘getline’ occurred. However, some implementations (such as Solaris 10 ‘/usr/bin/awk’, ‘nawk’, or Darwin ‘awk’) reset these variables. A workaround is to use an intermediate variable prior to the ‘END’ block. For example:

$ cat end.awk
{ tmp = $1 }
END { print "a", $1, $NF, "b", tmp }
$ echo 1 | awk -f end.awk
a   b 1
$ echo 1 | gawk -f end.awk
a 1 1 b 1

If you want your program to be deterministic, don’t depend on for on arrays:

$ cat for.awk
END {
  arr["foo"] = 1
  arr["bar"] = 1
  for (i in arr)
    print i
}
$ gawk -f for.awk </dev/null
foo
bar
$ nawk -f for.awk </dev/null
bar
foo

Some Awk implementations, such as HP-UX 11.0’s native one, mishandle anchors:

$ echo xfoo | $AWK '/foo|^bar/ { print }'
$ echo bar | $AWK '/foo|^bar/ { print }'
bar
$ echo xfoo | $AWK '/^bar|foo/ { print }'
xfoo
$ echo bar | $AWK '/^bar|foo/ { print }'
bar

Either do not depend on such patterns (i.e., use ‘/^(.*foo|bar)/’, or use a simple test to reject such implementations. On ‘ia64-hp-hpux11.23’, Awk mishandles printf conversions after %u:

$ awk 'BEGIN { printf "%u %d\n", 0, -1 }'
0 0

AIX version 5.2 has an arbitrary limit of 399 on the length of regular expressions and literal strings in an Awk program. Traditional Awk implementations derived from Unix version 7, such as Solaris /bin/awk, have many limitations and do not conform to Posix. Nowadays AC_ PROG_AWK (see Section 5.2.1 [Particular Programs], page 45) finds you an Awk that doesn’t have these problems, but if for some reason you prefer not to use AC_PROG_AWK you may need to address them. Traditional Awk does not support multidimensional arrays or user-defined functions.

Traditional Awk does not support the ‘-v’ option. You can use assignments after the program instead, e.g., $AWK '{print v $1}' v=x; however, don’t forget that such assignments are not evaluated until they are encountered (e.g., after any BEGIN action). Traditional Awk does not support the keywords delete or do. Traditional Awk does not support the expressions a?b:c, !a, a^b, or a^=b. Traditional Awk does not support the predefined CONVFMT variable. Traditional Awk supports only the predefined functions exp, index, int,

length, log, split, sprintf, sqrt, and substr.

Traditional Awk getline is not at all compatible with Posix; avoid it. Traditional Awk has for (i in a) ... but no other uses of the in keyword. For example, it lacks if (i in a) .... In code portable to both traditional and modern Awk, FS must be a string containing just one ordinary character, and similarly for the field-separator argument to split. Traditional Awk has a limit of 99 fields in a record. Since some Awk implementations, like Tru64’s, split the input even if you don’t refer to any field in the script, to circumvent this problem, set ‘FS’ to an unusual character and use

split.

Traditional Awk has a limit of at most 99 bytes in a number formatted by OFMT; for example, OFMT="%.300e"; print 0.1; typically dumps core. The original version of Awk had a limit of at most 99 bytes per split field, 99 bytes per substr substring, and 99 bytes per run of non-special characters in a printf format, but these bugs have been fixed on all practical hosts that we know of. HP-UX 11.00 and IRIX 6.5 Awk require that input files have a line length of at most 3070 bytes.

basename Not all hosts have a working basename. You can use expr instead.

cat Don’t rely on any option.

cc The command ‘cc -c foo.c’ traditionally produces an object file named ‘foo.o’. Most compilers allow ‘-c’ to be combined with ‘-o’ to specify a different object file name, but Posix does not require this combination and a few compilers lack support for it. See Section 5.10.3 [C Compiler], page 81, for how GNU Make tests for this feature with AC_PROG_CC_C_O. When a compilation such as ‘cc -o foo foo.c’ fails, some compilers (such as CDS on Reliant Unix) leave a ‘foo.o’. HP-UX cc doesn’t accept ‘.S’ files to preprocess and assemble. ‘cc -c foo.S’ appears to succeed, but in fact does nothing. The default executable, produced by ‘cc foo.c’, can be

• ‘a.out’ — usual Posix convention.

• ‘b.out’ — i960 compilers (including gcc).

• ‘a.exe’ — DJGPP port of gcc.

• ‘a_out.exe’ — GNV cc wrapper for DEC C on OpenVMS.

• ‘foo.exe’ — various MS-DOS compilers.

The C compiler’s traditional name is cc, but other names like gcc are common. Posix 1003.1-2001 specifies the name c99, but older Posix editions specified c89 and anyway these standard names are rarely used in practice. Typically the C compiler is invoked from makefiles that use ‘$(CC)’, so the value of the ‘CC’ make variable selects the compiler name.

chgrp

chown It is not portable to change a file’s group to a group that the owner does not belong to.

chmod Avoid usages like ‘chmod -w file’; use ‘chmod a-w file’ instead, for two reasons. First, plain ‘-w’ does not necessarily make the file unwritable, since it does not affect mode bits that correspond to bits in the file mode creation mask. Second, Posix says that the ‘-w’ might be interpreted as an implementation-specific option, not as a mode; Posix suggests using ‘chmod -- -w file’ to avoid this confusion, but unfortunately ‘--’ does not work on some older hosts.

cmp cmp performs a raw data comparison of two files, while diff compares two text files. Therefore, if you might compare DOS files, even if only checking whether two files are different, use diff to avoid spurious differences due to differences of newline encoding.

cp Avoid the ‘-r’ option, since Posix 1003.1-2004 marks it as obsolescent and its behavior on special files is implementation-defined. Use ‘-R’ instead. On GNU hosts the two options are equivalent, but on Solaris hosts (for example) cp -r reads from pipes instead of replicating them. Some cp implementations (e.g., BSD/OS 4.2) do not allow trailing slashes at the end of nonexistent destination directories. To avoid this problem, omit the trailing slashes. For example, use ‘cp -R source /tmp/newdir’ rather than ‘cp

-R source /tmp/newdir/’ if ‘/tmp/newdir’ does not exist.

The ancient SunOS 4 cp does not support ‘-f’, although its mv does. Traditionally, file timestamps had 1-second resolution, and ‘cp -p’ copied the timestamps exactly. However, many modern file systems have timestamps with 1-nanosecond resolution. Unfortunately, ‘cp -p’ implementations truncate timestamps when copying files, so this can result in the destination file appearing to be older than the source. The exact amount of truncation depends on the resolution of the system calls that cp uses; traditionally this was utime, which has 1-second resolution, but some newer cp implementations use utimes, which has 1-microsecond resolution. These newer implementations include GNU Core Utilities 5.0.91 or later, and Solaris 8 (sparc) patch 109933-02 or later. Unfortunately as of January 2006 there is still no system call to set timestamps to the full nanosecond resolution. Bob Proulx notes that ‘cp -p’ always tries to copy ownerships. But whether it actually does copy ownerships or not is a system dependent policy decision implemented by the kernel. If the kernel allows it then it happens. If the kernel does not allow it then it does not happen. It is not something cp itself has control over. In Unix System V any user can chown files to any other user, and System V also has a non-sticky ‘/tmp’. That probably derives from the heritage of System V in a business environment without hostile users. BSD changed this to be a more secure model where only root can chown files and a sticky ‘/tmp’ is used. That undoubtedly derives from the heritage of BSD in a campus environment. GNU/Linux and Solaris by default follow BSD, but can be configured to allow a System V style chown. On the other hand, HP-UX follows System V, but can be configured to use the modern security model and disallow chown. Since it is an administrator-configurable parameter you can’t use the name of the kernel as an indicator of the behavior.

date Some versions of date do not recognize special ‘%’ directives, and unfortunately,

instead of complaining, they just pass them through, and exit with success:

     $ uname -a
     OSF1 medusa.sis.pasteur.fr V5.1 732 alpha
     $ date "+%s"
     %s

diff Option ‘-u’ is nonportable.

Some implementations, such as Tru64’s, fail when comparing to ‘/dev/null’. Use an empty file instead.

dirname Not all hosts have a working dirname, and you should instead use AS_DIRNAME

(see Chapter 9 [Programming in M4sh], page 165). For example:

     dir=`dirname "$file"`       # This is not portable.
     dir=`AS_DIRNAME(["$file"])` # This is more portable.

egrep Posix 1003.1-2001 no longer requires egrep, but many hosts do not yet support the Posix replacement grep -E. Also, some traditional implementations do not work on long input lines. To work around these problems, invoke AC_PROG_

EGREP and then use $EGREP.

Portable extended regular expressions should use ‘\’ only to escape characters in the string ‘$()*+.?[\^{|’. For example, ‘\}’ is not portable, even though it typically matches ‘}’. The empty alternative is not portable. Use ‘?’ instead. For instance with Digital Unix v5.0:

     > printf "foo\n|foo\n" | $EGREP '^(|foo|bar)$'
     |foo
     > printf "bar\nbar|\n" | $EGREP '^(foo|bar|)$'
     bar|
     > printf "foo\nfoo|\n|bar\nbar\n" | $EGREP '^(foo||bar)$'
     foo
     |bar

$EGREP also suffers the limitations of grep (see [Limitations of Usual Tools], page 231). expr Not all implementations obey the Posix rule that ‘--’ separates options from

arguments; likewise, not all implementations provide the extension to Posix that the first argument can be treated as part of a valid expression rather than an invalid option if it begins with ‘-’. When performing arithmetic, use ‘expr 0 + $var’ if ‘$var’ might be a negative number, to keep expr from interpreting it as an option. No expr keyword starts with ‘X’, so use ‘expr X"word" : 'Xregex'’ to keep expr from misinterpreting word.

Don’t use length, substr, match and index.

expr (‘|’) You can use ‘|’. Although Posix does require that ‘expr ''’ return the empty string, it does not specify the result when you ‘|’ together the empty string (or zero) with the empty string. For example:

     expr '' \| ''

Posix 1003.2-1992 returns the empty string for this case, but traditional Unix returns ‘0’ (Solaris is one such example). In Posix 1003.1-2001, the specification was changed to match traditional Unix’s behavior (which is bizarre, but it’s too late to fix this). Please note that the same problem does arise when the empty string results from a computation, as in:

     expr bar : foo \| foo : bar

Avoid this portability problem by avoiding the empty string.

expr (‘:’) Portable expr regular expressions should use ‘\’ to escape only characters in the string ‘$()*.0123456789[\^n{}’. For example, alternation, ‘\|’, is common but Posix does not require its support, so it should be avoided in portable scripts. Similarly, ‘\+’ and ‘\?’ should be avoided. Portable expr regular expressions should not begin with ‘^’. Patterns are automatically anchored so leading ‘^’ is not needed anyway. On the other hand, the behavior of the ‘$’ anchor is not portable on multi-line strings. Posix is ambiguous whether the anchor applies to each line, as was done in older versions of GNU Coreutils, or whether it applies only to the end of the overall string, as in Coreutils 6.0 and most other implementations.

     $ baz='foo
     > bar'
     $ expr "X$baz" : 'X\(foo\)$'
     $ expr-5.97 "X$baz" : 'X\(foo\)$'
     foo

The Posix standard is ambiguous as to whether ‘expr 'a' : '\(b\)'’ outputs ‘0’ or the empty string. In practice, it outputs the empty string on most platforms, but portable scripts should not assume this. For instance, the QNX 4.25 native expr returns ‘0’. One might think that a way to get a uniform behavior would be to use the empty string as a default value:

     expr a : '\(b\)' \| ''

Unfortunately this behaves exactly as the original expression; see the expr (‘|’) entry for more information. Some ancient expr implementations (e.g., SunOS 4 expr and Solaris 8 /usr/ucb/expr) have a silly length limit that causes expr to fail if the matched substring is longer than 120 bytes. In this case, you might want to fall back on ‘echo|sed’ if expr fails. Nowadays this is of practical importance only for the rare installer who mistakenly puts ‘/usr/ucb’ before ‘/usr/bin’

in PATH.

On Mac OS X 10.4, expr mishandles the pattern ‘[^-]’ in some cases. For example, the command

     expr Xpowerpc-apple-darwin8.1.0 : 'X[^-]*-[^-]*-\(.*\)'

outputs ‘apple-darwin8.1.0’ rather than the correct ‘darwin8.1.0’. This particular case can be worked around by substituting ‘[^--]’ for ‘[^-]’. Don’t leave, there is some more! The QNX 4.25 expr, in addition of preferring ‘0’ to the empty string, has a funny behavior in its exit status: it’s always 1 when parentheses are used!

     $ val=`expr 'a' : 'a'`; echo "$?: $val"
     0: 1
     $ val=`expr 'a' : 'b'`; echo "$?: $val"
     1: 0
     $ val=`expr 'a' : '\(a\)'`; echo "?: $val"
     1: a
     $ val=`expr 'a' : '\(b\)'`; echo "?: $val"
     1: 0

In practice this can be a big problem if you are ready to catch failures of expr programs with some other method (such as using sed), since you may get twice the result. For instance

     $ expr 'a' : '\(a\)' || echo 'a' | sed 's/^\(a\)$/\1/'

outputs ‘a’ on most hosts, but ‘aa’ on QNX 4.25. A simple workaround consists of testing expr and using a variable set to expr or to false according to the result. Tru64 expr incorrectly treats the result as a number, if it can be interpreted that way:

     $ expr 00001 : '.*\(...\)'
     1

On HP-UX 11, expr only supports a single sub-expression.

     $ expr 'Xfoo' : 'X\(f\(oo\)*\)$'
     expr: More than one '\(' was used.

fgrep Posix 1003.1-2001 no longer requires fgrep, but many hosts do not yet support the Posix replacement grep -F. Also, some traditional implementations do not work on long input lines. To work around these problems, invoke AC_PROG_

FGREP and then use $FGREP.

Tru64/OSF 5.1 fgrep does not match an empty pattern. find The option ‘-maxdepth’ seems to be GNU specific. Tru64 v5.1, NetBSD 1.5

and Solaris find commands do not understand it. The replacement of ‘{}’ is guaranteed only if the argument is exactly {}, not if it’s only a part of an argument. For instance on DU, and HP-UX 10.20 and HP-UX 11:

     $ touch foo
     $ find . -name foo -exec echo "{}-{}" \;
     {}-{}

while GNU find reports ‘./foo-./foo’.

grep Portable scripts can rely on the grep options ‘-c’, ‘-l’, ‘-n’, and ‘-v’, but should avoid other options. For example, don’t use ‘-w’, as Posix does not require it and Irix 6.5.16m’s grep does not support it. Also, portable scripts should not combine ‘-c’ with ‘-l’, as Posix does not allow this. Some of the options required by Posix are not portable in practice. Don’t use ‘grep -q’ to suppress output, because many grep implementations (e.g., Solaris) do not support ‘-q’. Don’t use ‘grep -s’ to suppress output either, because Posix says ‘-s’ does not suppress output, only some error messages; also, the ‘-s’ option of traditional grep behaved like ‘-q’ does in most modern implementations. Instead, redirect the standard output and standard error (in case the file doesn’t exist) of grep to ‘/dev/null’. Check the exit status of grep to determine whether it found a match. The QNX4 implementation fails to count lines with grep -c '$', but works with grep -c '^'. Other alternatives for counting lines are to use sed -n '$='

or wc -l.

Some traditional grep implementations do not work on long input lines. On AIX the default grep silently truncates long lines on the input before matching. Also, many implementations do not support multiple regexps with ‘-e’: they either reject ‘-e’ entirely (e.g., Solaris) or honor only the last pattern (e.g., IRIX 6.5 and NeXT). To work around these problems, invoke AC_PROG_GREP and then use $GREP. Another possible workaround for the multiple ‘-e’ problem is to separate the patterns by newlines, for example:

     grep 'foo
     bar' in.txt

except that this fails with traditional grep implementations and with OpenBSD

3.8 grep.

Traditional grep implementations (e.g., Solaris) do not support the ‘-E’ or ‘-F’ options. To work around these problems, invoke AC_PROG_EGREP and then use $EGREP, and similarly for AC_PROG_FGREP and $FGREP. Even if you are willing to require support for Posix grep, your script should not use both ‘-E’ and ‘-F’, since Posix does not allow this combination. Portable grep regular expressions should use ‘\’ only to escape characters in the string ‘$()*.0123456789[\^{}’. For example, alternation, ‘\|’, is common but Posix does not require its support in basic regular expressions, so it should be avoided in portable scripts. Solaris and HP-UX grep do not support it. Similarly, the following escape sequences should also be avoided: ‘\<’, ‘\>’, ‘\+’, ‘\?’, ‘\`’, ‘\'’, ‘\B’, ‘\b’, ‘\S’, ‘\s’, ‘\W’, and ‘\w’.

Posix does not specify the behavior of grep on binary files. An example where this matters is using BSD grep to search text that includes embedded ANSI escape sequences for colored output to terminals (‘\033[m’ is the sequence to restore normal output); the behavior depends on whether input is seekable:

     $ printf 'esc\033[mape\n' > sample
     $ grep . sample
     Binary file sample matches
     $ cat sample | grep .
     escape

join Solaris 8 join has bugs when the second operand is standard input, and when standard input is a pipe. For example, the following shell script causes Solaris 8 join to loop forever:

     cat >file <<'EOF'
     1 x
     2 y
     EOF
     cat file | join file -

Use ‘join - file’ instead.

ln Don’t rely on ln having a ‘-f’ option. Symbolic links are not available on old systems; use ‘$(LN_S)’ as a portable substitute.

For versions of the DJGPP before 2.04, ln emulates symbolic links to executables by generating a stub that in turn calls the real program. This feature also works with nonexistent files like in the Posix spec. So ‘ln -s file link’ generates ‘link.exe’, which attempts to call ‘file.exe’ if run. But this feature only works for executables, so ‘cp -p’ is used instead for these systems. DJGPP versions 2.04 and later have full support for symbolic links.

ls The portable options are ‘-acdilrtu’. Current practice is for ‘-l’ to output both owner and group, even though ancient versions of ls omitted the group.

On ancient hosts, ‘ls foo’ sent the diagnostic ‘foo not found’ to standard output if ‘foo’ did not exist. Hence a shell command like ‘sources=`ls *.c 2>/dev/null`’ did not always work, since it was equivalent to ‘sources='*.c not found'’ in the absence of ‘.c’ files. This is no longer a practical problem, since current ls implementations send diagnostics to standard error.

The behavior of ls on a directory that is being concurrently modified is not always predictable, because of a data race where cached information returned by readdir does not match the current directory state. In fact, MacOS 10.5 has an intermittent bug where readdir, and thus ls, sometimes lists a file more than once if other files were added or removed from the directory immediately prior to the ls call. Since ls already sorts its output, the duplicate entries can be avoided by piping the results through uniq. mkdir No mkdir option is portable to older systems. Instead of ‘mkdir -p file-

name’, you should use AS_MKDIR_P(file-name) (see Chapter 9 [Programming in M4sh], page 165) or AC_PROG_MKDIR_P (see Section 5.2.1 [Particular Programs], page 45). Combining the ‘-m’ and ‘-p’ options, as in ‘mkdir -m go-w -p dir’, often leads to trouble. FreeBSD mkdir incorrectly attempts to change the permissions of dir even if it already exists. HP-UX 11.23 and IRIX 6.5 mkdir often assign the wrong permissions to any newly-created parents of dir. Posix does not clearly specify whether ‘mkdir -p foo’ should succeed when ‘foo’ is a symbolic link to an already-existing directory. The GNU Core Utilities 5.1.0 mkdir succeeds, but Solaris mkdir fails. Traditional mkdir -p implementations suffer from race conditions. For example, if you invoke mkdir -p a/b and mkdir -p a/c at the same time, both processes might detect that ‘a’ is missing, one might create ‘a’, then the other might try to create ‘a’ and fail with a File exists diagnostic. The GNU Core Utilities (‘fileutils’ version 4.1), FreeBSD 5.0, NetBSD 2.0.2, and OpenBSD 2.4 are known to be race-free when two processes invoke mkdir -p simultaneously, but earlier versions are vulnerable. Solaris mkdir is still vulnerable as of Solaris 10, and other traditional Unix systems are probably vulnerable too. This possible race is harmful in parallel builds when several Make rules call mkdir -p to construct directories. You may use install-sh -d as a safe replacement, provided this script is recent enough; the copy shipped with Autoconf 2.60 and Automake 1.10 is OK, but copies from older versions are vulnerable.

mkfifo

mknod The GNU Coding Standards state that mknod is safe to use on platforms where it has been tested to exist; but it is generally portable only for creating named FIFOs, since device numbers are platform-specific. Autotest uses mkfifo to implement parallel testsuites. Posix states that behavior is unspecified when opening a named FIFO for both reading and writing; on at least Cygwin, this results in failure on any attempt to read or write to that file descriptor.

mktemp Shell scripts can use temporary files safely with mktemp, but it does not exist on all systems. A portable way to create a safe temporary file name is to create a temporary directory with mode 700 and use a file inside this directory. Both methods prevent attackers from gaining control, though mktemp is far less likely to fail gratuitously under attack. Here is sample code to create a new temporary directory safely:

     # Create a temporary directory $tmp in $TMPDIR (default /tmp).
     # Use mktemp if possible; otherwise fall back on mkdir,
     # with $RANDOM to make collisions less likely.
     : ${TMPDIR=/tmp}
     {
       tmp=`
         (umask 077 && mktemp -d "$TMPDIR/fooXXXXXX") 2>/dev/null
       ` &&
       test -n "$tmp" && test -d "$tmp"

     } || {
       tmp=$TMPDIR/foo$$-$RANDOM
       (umask 077 && mkdir "$tmp")
     } || exit $?

mv The only portable options are ‘-f’ and ‘-i’.

Moving individual files between file systems is portable (it was in Unix version 6), but it is not always atomic: when doing ‘mv new existing’, there’s a critical section where neither the old nor the new version of ‘existing’ actually exists. On some systems moving files from ‘/tmp’ can sometimes cause undesirable (but perfectly valid) warnings, even if you created these files. This is because ‘/tmp’ belongs to a group that ordinary users are not members of, and files created in ‘/tmp’ inherit the group of ‘/tmp’. When the file is copied, mv issues a diagnostic without failing:

      $ touch /tmp/foo
      $ mv /tmp/foo .
       error mv: ./foo: set owner/group (was: 100/0): Operation not permitted
      $ echo $?
      0
      $ ls foo
      foo

This annoying behavior conforms to Posix, unfortunately. Moving directories across mount points is not portable, use cp and rm. DOS variants cannot rename or remove open files, and do not support commands like ‘mv foo bar >foo’, even though this is perfectly portable among Posix hosts.

od

In Mac OS X 10.3, od does not support the standard Posix options ‘-A’, ‘-j’, ‘-N’, or ‘-t’, or the XSI option ‘-s’. The only supported Posix option is ‘-v’, and the only supported XSI options are those in ‘-bcdox’. The BSD hexdump program can be used instead. This problem no longer exists in Mac OS X 10.4.3.

rm The ‘-f’ and ‘-r’ options are portable.

It is not portable to invoke rm without operands. For example, on many systems ‘rm -f -r’ (with no other arguments) silently succeeds without doing anything, but it fails with a diagnostic on NetBSD 2.0.2. A file might not be removed even if its parent directory is writable and searchable. Many Posix hosts cannot remove a mount point, a named stream, a working directory, or a last link to a file that is being executed. DOS variants cannot rename or remove open files, and do not support commands like ‘rm foo >foo’, even though this is perfectly portable among Posix hosts.

rmdir Just as with rm, some platforms refuse to remove a working directory.

sed Patterns should not include the separator (unless escaped), even as part of a character class. In conformance with Posix, the Cray sed rejects ‘s/[^/]*$//’: use ‘s%[^/]*$%%’. Even when escaped, patterns should not include separators that are also used as sed metacharacters. For example, GNU sed 4.0.9 rejects ‘s,x\{1\,\},,’, while sed 4.1 strips the backslash before the comma before evaluating the basic regular expression. Avoid empty patterns within parentheses (i.e., ‘\(\)’). Posix does not require support for empty patterns, and Unicos 9 sed rejects them. Unicos 9 sed loops endlessly on patterns like ‘.*\n.*’. Sed scripts should not use branch labels longer than 7 characters and should not contain comments. HP-UX sed has a limit of 99 commands (not counting ‘:’ commands) and 48 labels, which can not be circumvented by using more than one script file. It can execute up to 19 reads with the ‘r’ command per cycle. Solaris /usr/ucb/sed rejects usages that exceed a limit of about 6000 bytes for the internal representation of commands. Avoid redundant ‘;’, as some sed implementations, such as NetBSD 1.4.2’s, incorrectly try to interpret the second ‘;’ as a command:

     $ echo a | sed 's/x/x/;;s/x/x/'
     sed: 1: "s/x/x/;;s/x/x/": invalid command code ;

Input should not have unreasonably long lines, since some sed implementations have an input buffer limited to 4000 bytes. Likewise, not all sed implementations can handle embedded NUL or a missing trailing newline. Portable sed regular expressions should use ‘\’ only to escape characters in the string ‘$()*.0123456789[\^n{}’. For example, alternation, ‘\|’, is common but Posix does not require its support, so it should be avoided in portable scripts. Solaris sed does not support alternation; e.g., ‘sed '/a\|b/d'’ deletes only lines that contain the literal string ‘a|b’. Similarly, ‘\+’ and ‘\?’ should be avoided. Anchors (‘^’ and ‘$’) inside groups are not portable. Nested parentheses in patterns (e.g., ‘\(\(a*\)b*)\)’) are quite portable to current hosts, but was not supported by some ancient sed implementations like SVR3. Some sed implementations, e.g., Solaris, restrict the special role of the asterisk ‘*’ to one-character regular expressions and back-references, and the special role of interval expressions ‘\{m\}’, ‘\{m,\}’, or ‘\{m,n\}’ to one-character regular expressions. This may lead to unexpected behavior:

     $ echo '1*23*4' | /usr/bin/sed 's/\(.\)*/x/g'
     x2x4
     $ echo '1*23*4' | /usr/xpg4/bin/sed 's/\(.\)*/x/g'
     x

The ‘-e’ option is mostly portable. However, its argument cannot start with ‘a’, ‘c’, or ‘i’, as this runs afoul of a Tru64 5.1 bug. Also, its argument cannot be empty, as this fails on AIX 5.3. Some people prefer to use ‘-e’:

     sed -e 'command-1' \
         -e 'command-2'

as opposed to the equivalent:

     sed '
       command-1
       command-2
     '

The following usage is sometimes equivalent:

     sed 'command-1;command-2'

but Posix says that this use of a semicolon has undefined effect if command-1’s verb is ‘{’, ‘a’, ‘b’, ‘c’, ‘i’, ‘r’, ‘t’, ‘w’, ‘:’, or ‘#’, so you should use semicolon only with simple scripts that do not use these verbs. Posix up to the 2008 revision requires the argument of the ‘-e’ option to be a syntactically complete script. GNU sed allows to pass multiple script fragments, each as argument of a separate ‘-e’ option, that are then combined, with newlines between the fragments, and a future Posix revision may allow this as well. This approach is not portable with script fragments ending in backslash; for example, the sed programs on Solaris 10, HP-UX 11, and AIX don’t allow splitting in this case:

     $ echo a | sed -n -e 'i\
     0'
     0
     $ echo a | sed -n -e 'i\' -e 0
     Unrecognized command: 0

In practice, however, this technique of joining fragments through ‘-e’ works for multiple sed functions within ‘{’ and ‘}’, even if that is not specified by Posix:

     $ echo a | sed -n -e '/a/{' -e s/a/b/ -e p -e '}'
     b

Commands inside { } brackets are further restricted. Posix 2008 says that they cannot be preceded by addresses, ‘!’, or ‘;’, and that each command must be followed immediately by a newline, without any intervening blanks or semicolons. The closing bracket must be alone on a line, other than white space preceding or following it. However, a future version of Posix may standardize the use of addresses within brackets. Contrary to yet another urban legend, you may portably use ‘&’ in the replacement part of the s command to mean “what was matched”. All descendants of Unix version 7 sed (at least; we don’t have first hand experience with older sed implementations) have supported it. Posix requires that you must not have any white space between ‘!’ and the following command. It is OK to have blanks between the address and the ‘!’. For instance, on Solaris:

     $ echo "foo" | sed -n '/bar/ ! p'
      error Unrecognized command: /bar/ ! p
     $ echo "foo" | sed -n '/bar/! p'
      error Unrecognized command: /bar/! p
     $ echo "foo" | sed -n '/bar/ !p'
     foo

Posix also says that you should not combine ‘!’ and ‘;’. If you use ‘!’, it is best to put it on a command that is delimited by newlines rather than ‘;’. Also note that Posix requires that the ‘b’, ‘t’, ‘r’, and ‘w’ commands be followed by exactly one space before their argument. On the other hand, no white space is allowed between ‘:’ and the subsequent label name. If a sed script is specified on the command line and ends in an ‘a’, ‘c’, or ‘i’ command, the last line of inserted text should be followed by a newline. Otherwise some sed implementations (e.g., OpenBSD 3.9) do not append a newline to the inserted text. Many sed implementations (e.g., MacOS X 10.4, OpenBSD 3.9, Solaris 10 /usr/ucb/sed) strip leading white space from the text of ‘a’, ‘c’, and ‘i’ commands. Prepend a backslash to work around this incompatibility with Posix:

     $ echo flushleft | sed 'a\
     >    indented
     > '
     flushleft
     indented
     $ echo foo | sed 'a\
     > \   indented
     > '
     flushleft
        indented

Posix requires that with an empty regular expression, the last non-empty regular expression from either an address specification or substitution command is applied. However, busybox 1.6.1 complains when using a substitution command with a replacement containing a back-reference to an empty regular expression; the workaround is repeating the regular expression.

     $ echo abc | busybox sed '/a\(b\)c/ s//\1/'
     sed: No previous regexp.
     $ echo abc | busybox sed '/a\(b\)c/ s/a\(b\)c/\1/'
     b

sed (‘t’) Some old systems have sed that “forget” to reset their ‘t’ flag when starting a new cycle. For instance on MIPS RISC/OS, and on IRIX 5.3, if you run the following sed script (the line numbers are not actual part of the texts):

     s/keep me/kept/g  # a
     t end             # b
     s/.*/deleted/g    # c
     :end              # d

on

     delete me         # 1
     delete me         # 2
     keep me           # 3
     delete me         # 4

you get

     deleted
     delete me
     kept
     deleted

instead of

     deleted
     deleted
     kept
     deleted

Why? When processing line 1, (c) matches, therefore sets the ‘t’ flag, and the output is produced. When processing line 2, the ‘t’ flag is still set (this is the bug). Command (a) fails to match, but sed is not supposed to clear the ‘t’ flag when a substitution fails. Command (b) sees that the flag is set, therefore it clears it, and jumps to (d), hence you get ‘delete me’ instead of ‘deleted’. When processing line (3), ‘t’ is clear, (a) matches, so the flag is set, hence (b) clears the flags and jumps. Finally, since the flag is clear, line 4 is processed properly. There are two things one should remember about ‘t’ in sed. Firstly, always remember that ‘t’ jumps if some substitution succeeded, not only the immediately preceding substitution. Therefore, always use a fake ‘t clear’ followed by a ‘:clear’ on the next line, to reset the ‘t’ flag where needed. Secondly, you cannot rely on sed to clear the flag at each new cycle. One portable implementation of the script above is:

     t clear
     :clear
     s/keep me/kept/g
     t end
     s/.*/deleted/g
     :end

sleep Using sleep is generally portable. However, remember that adding a sleep to work around timestamp issues, with a minimum granularity of one second, doesn’t scale well for parallel builds on modern machines with sub-second process completion.

sort Remember that sort order is influenced by the current locale. Inside ‘configure’, the C locale is in effect, but in Makefile snippets, you may need

to specify LC_ALL=C sort.

tar There are multiple file formats for tar; if you use Automake, the macro AM_ INIT_AUTOMAKE has some options controlling which level of portability to use.

touch If you specify the desired timestamp (e.g., with the ‘-r’ option), touch typically uses the utime or utimes system call, which can result in the same kind of timestamp truncation problems that ‘cp -p’ has. On ancient BSD systems, touch or any command that results in an empty file does not update the timestamps, so use a command like echo as a workaround.

Also, GNU touch 3.16r (and presumably all before that) fails to work on SunOS 4.1.3 when the empty file is on an NFS-mounted 4.2 volume. However, these problems are no longer of practical concern.

tr Not all versions of tr handle all backslash character escapes. For example, Solaris 10 /usr/ucb/tr falls over, even though Solaris contains more modern tr in other locations. Using octal escapes is more portable for carriage returns, since ‘\015’ is the same for both ASCII and EBCDIC, and since use of literal carriage returns in scripts causes a number of other problems. But for other characters, like newline, using octal escapes ties the operation to ASCII, so it is better to use literal characters.

$ { echo moon; echo light; } | /usr/ucb/tr -d '\n' ; echo
moo
light
$ { echo moon; echo light; } | /usr/bin/tr -d '\n' ; echo
moonlight
$ { echo moon; echo light; } | /usr/ucb/tr -d '\012' ; echo
moonlight
$ nl='
'; { echo moon; echo light; } | /usr/ucb/tr -d "$nl" ; echo
moonlight

Not all versions of tr recognize ranges of characters: at least Solaris /usr/bin/tr still fails to do so. But you can use /usr/xpg4/bin/tr instead.

$ echo "Hazy Fantazy" | LC_ALL=C /usr/bin/tr a-z A-Z
HAZy FAntAZy
$ echo "Hazy Fantazy" | LC_ALL=C /usr/xpg4/bin/tr a-z A-Z
HAZY FANTAZY

When providing two arguments, be sure the second string is at least as long as the first.

$ echo abc | /usr/xpg4/bin/tr bc d
adc
$ echo abc | coreutils/tr bc d
add

Posix requires tr to operate on binary files. But at least Solaris /usr/ucb/tr and /usr/bin/tr silently discard NUL in the input prior to doing any translation. When using tr to process a binary file that may contain NUL bytes, it is

necessary to use /usr/xpg4/bin/tr instead, or /usr/xpg6/bin/tr if that is

available.

$ printf 'a\0b' | /usr/ucb/tr x x | od -An -tx1
 61 62
$ printf 'a\0b' | /usr/bin/tr x x | od -An -tx1
 61 62
$ printf 'a\0b' | /usr/xpg4/bin/tr x x | od -An -tx1
 61 00 62

Solaris /usr/ucb/tr additionally fails to handle ‘\0’ as the octal escape for

NUL.

$ printf 'abc' | /usr/ucb/tr 'bc' '\0d' | od -An -tx1
 61 62 63
$ printf 'abc' | /usr/bin/tr 'bc' '\0d' | od -An -tx1
 61 00 64
$ printf 'abc' | /usr/xpg4/bin/tr 'bc' '\0d' | od -An -tx1
 61 00 64

12 Portable Make Programming

Writing portable makefiles is an art. Since a makefile’s commands are executed by the shell, you must consider the shell portability issues already mentioned. However, other issues are specific to make itself.

12.1 $< in Ordinary Make Rules

Posix says that the ‘$<’ construct in makefiles can be used only in inference rules and in the ‘.DEFAULT’ rule; its meaning in ordinary rules is unspecified. Solaris make for instance replaces it with the empty string. OpenBSD (3.0 and later) make diagnoses these uses and errors out.

12.2 Failure in Make Rules

Posix 2008 requires that make must invoke each command with the equivalent of a ‘sh -e -c’ subshell, which causes the subshell to exit immediately if a subsidiary simple-command fails, although not all make implementations have historically followed this rule. For example, the command ‘touch T; rm -f U’ may attempt to remove ‘U’ even if the touch fails, although this is not permitted with Posix make. One way to work around failures in simple commands is to reword them so that they always succeed, e.g., ‘touch T || :; rm -f U’. However, even this approach can run into common bugs in BSD implementations of the ‘-e’ option of sh and set (see [Limitations of Shell Builtins], page 218), so if you are worried about porting to buggy BSD shells it may be simpler to migrate complicated make actions into separate scripts.

12.3 Special Characters in Make Macro Names

Posix limits macro names to nonempty strings containing only ASCII letters and digits, ‘.’, and ‘_’. Many make implementations allow a wider variety of characters, but portable makefiles should avoid them. It is portable to start a name with a special character, e.g.,

‘$(.FOO)’.

Some ancient make implementations don’t support leading underscores in macro names. An example is NEWS-OS 4.2R.

     $ cat Makefile
     _am_include = #
     _am_quote =
     all:; @echo this is test
     $ make
     Make: Must be a separator on rules line 2.  Stop.
     $ cat Makefile2
     am_include = #
     am_quote =
     all:; @echo this is test
     $ make -f Makefile2
     this is test

However, this problem is no longer of practical concern.

12.4 Backslash-Newline-Newline in Make Macro Values

On some versions of HP-UX, make reads multiple newlines following a backslash, continuing to the next non-empty line. For example,

     FOO = one \
     BAR = two
     test:
             : FOO is "$(FOO)"
             : BAR is "$(BAR)"

shows FOO equal to one BAR = two. Other implementations sensibly let a backslash continue only to the immediately following line.

12.5 Backslash-Newline in Make Comments

According to Posix, Make comments start with # and continue until an unescaped newline is reached.

     $ cat Makefile
     # A = foo \
           bar \
           baz
     all:
             @echo ok
     $ make   # GNU make
     ok

However this is not always the case. Some implementations discard everything from # through the end of the line, ignoring any trailing backslash.

     $ pmake  # BSD make
     "Makefile", line 3: Need an operator
     Fatal errors encountered -- cannot continue

Therefore, if you want to comment out a multi-line definition, prefix each line with #, not only the first.

     # A = foo \
     #     bar \
     #     baz

12.6 Long Lines in Makefiles

Tru64 5.1’s make has been reported to crash when given a makefile with lines longer than around 20 kB. Earlier versions are reported to exit with Line too long diagnostics.

12.7 make macro=value and Submakes

A command-line variable definition such as foo=bar overrides any definition of foo in a makefile. Some make implementations (such as GNU make) propagate this override to subsidiary invocations of make. Some other implementations do not pass the substitution along to submakes.

$ cat Makefile
foo = foo
one:
        @echo $(foo)
        $(MAKE) two
two:
        @echo $(foo)
$ make foo=bar            # GNU make 3.79.1
bar
make two
make[1]: Entering directory `/home/adl'
bar
make[1]: Leaving directory `/home/adl'
$ pmake foo=bar           # BSD make
bar
pmake two
foo

You have a few possibilities if you do want the foo=bar override to propagate to submakes. One is to use the ‘-e’ option, which causes all environment variables to have precedence over the makefile macro definitions, and declare foo as an environment variable:

$ env foo=bar make -e

The ‘-e’ option is propagated to submakes automatically, and since the environment is inherited between make invocations, the foo macro is overridden in submakes as expected.

This syntax (foo=bar make -e) is portable only when used outside of a makefile, for instance from a script or from the command line. When run inside a make rule, GNU make 3.80 and prior versions forget to propagate the ‘-e’ option to submakes.

Moreover, using ‘-e’ could have unexpected side effects if your environment contains some other macros usually defined by the makefile. (See also the note about make -e and SHELL below.)

If you can foresee all macros that a user might want to override, then you can propagate them to submakes manually, from your makefile:

foo = foo
one:
        @echo $(foo)
        $(MAKE) foo=$(foo) two
two:
        @echo $(foo)

Another way to propagate a variable to submakes in a portable way is to expand an extra variable in every invocation of ‘$(MAKE)’ within your makefile:

foo = foo
one:
        @echo $(foo)
        $(MAKE) $(SUBMAKEFLAGS) two

     two:
             @echo $(foo)

Users must be aware that this technique is in use to take advantage of it, e.g. with make foo=bar SUBMAKEFLAGS='foo=bar', but it allows any macro to be overridden. Makefiles generated by automake use this technique, expanding $(AM_MAKEFLAGS) on the command lines of submakes (see Section “Automake” in GNU Automake).

12.8 The Make Macro MAKEFLAGS

Posix requires make to use MAKEFLAGS to affect the current and recursive invocations of make, but allows implementations several formats for the variable. It is tricky to parse $MAKEFLAGS to determine whether ‘-s’ for silent execution or ‘-k’ for continued execution are in effect. For example, you cannot assume that the first space-separated word in $MAKEFLAGS contains single-letter options, since in the Cygwin version of GNU make it is either ‘--unix’ or ‘--win32’ with the second word containing single-letter options.

     $ cat Makefile
     all:
             @echo MAKEFLAGS = $(MAKEFLAGS)
     $ make
     MAKEFLAGS = --unix
     $ make -k
     MAKEFLAGS = --unix -k

12.9 The Make Macro SHELL

Posix-compliant make internally uses the $(SHELL) macro to spawn shell processes and execute Make rules. This is a builtin macro supplied by make, but it can be modified by a makefile or by a command-line argument.

Not all make implementations define this SHELL macro. Tru64 make is an example; this implementation always uses /bin/sh. So it’s a good idea to always define SHELL in your makefiles. If you use Autoconf, do

     SHELL = @SHELL@

If you use Automake, this is done for you.

Do not force SHELL = /bin/sh because that is not correct everywhere. Remember, ‘/bin/sh’ is not Posix compliant on many systems, such as FreeBSD 4, NetBSD 3, AIX 3, Solaris 10, or Tru64. Additionally, DJGPP lacks /bin/sh, and when its GNU make port sees such a setting it enters a special emulation mode where features like pipes and redirections are emulated on top of DOS’s command.com. Unfortunately this emulation is incomplete; for instance it does not handle command substitutions. Using @SHELL@ means that your makefile will benefit from the same improved shell, such as bash or ksh, that was discovered during configure, so that you aren’t fighting two different sets of shell bugs between the two contexts.

Posix-compliant make should never acquire the value of $(SHELL) from the environment, even when make -e is used (otherwise, think about what would happen to your rules if

SHELL=/bin/tcsh).

However not all make implementations have this exception. For instance it’s not surprising that Tru64 make doesn’t protect SHELL, since it doesn’t use it.

$ cat Makefile
SHELL = /bin/sh
FOO = foo
all:
        @echo $(SHELL)
        @echo $(FOO)
$ env SHELL=/bin/tcsh FOO=bar make -e   # Tru64 Make
/bin/tcsh
bar
$ env SHELL=/bin/tcsh FOO=bar gmake -e  # GNU make
/bin/sh
bar

Conversely, make is not supposed to export any changes to the macro SHELL to child processes. Again, many implementations break this rule:

$ cat Makefile
all:
        @echo $(SHELL)
        @printenv SHELL
$ env SHELL=sh make -e SHELL=/bin/ksh   # BSD Make, GNU make 3.80
/bin/ksh
/bin/ksh
$ env SHELL=sh gmake -e SHELL=/bin/ksh  # GNU make 3.81
/bin/ksh
sh

12.10 Parallel Make

Support for parallel execution in make implementation varies. Generally, using GNU make is your best bet. When NetBSD make is invoked with ‘-jN’, it will reuse the same shell for multiple commands within one recipe. This can have unexpected consequences.^{1} For example, change of directories or variables persist between commands:

all:
        @var=value; cd /; pwd; echo $$var; echo $$$$
        @pwd; echo $$var; echo $$$$

may output the following with make -j1:

--- all ---
/
value
32235
/
value

1 Note that GNU make has heuristics to avoid spawning a shell at all if the command is deemed safe to

be executed directly.

     32235

while without ‘-j1’, or with ‘-B’, the output looks less surprising:

     /
     value
     32238
     /tmp
     32239

Another consequence of this is that, if one command in a recipe uses exit 0 to indicate a successful exit, the shell will be gone and the remaining commands of this recipe will not be executed.

The above example also shows additional status output NetBSD make produces in parallel mode for targets being updated.

Furthermore, parallel NetBSD make will route standard error from commands that it spawns into its own standard output, and may remove leading whitespace from output lines.

You can avoid these issues by using the ‘-B’ option to enable compatibility semantics. However, that will effectively also disable all parallelism as that will cause prerequisites to be updated in the order they are listed in a rule.

12.11 Comments in Make Rules

Never put comments in a rule.

Some make treat anything starting with a tab as a command for the current rule, even if the tab is immediately followed by a #. The make from Tru64 Unix V5.1 is one of them. The following makefile runs # foo through the shell.

     all:
             # foo

As a workaround, you can use the : no-op command with a string argument that gets ignored:

     all:
             : "foo"

12.12 Newlines in Make Rules

In shell scripts, newlines can be used inside string literals. But in the shell statements of ‘Makefile’ rules, this is not possible: A newline not preceded by a backslash is a separator between shell statements. Whereas a newline that is preceded by a backslash becomes part of the shell statement according to POSIX, but gets replaced, together with the backslash that precedes it, by a space in GNU make 3.80 and older. So, how can a newline be used in a string literal?

The trick is to set up a shell variable that contains a newline:

     nlinit=`echo 'nl="'; echo '"'`; eval "$$nlinit"

For example, in order to create a multiline ‘sed’ expression that inserts a blank line after every line of a file, this code can be used:

     nlinit=`echo 'nl="'; echo '"'`; eval "$$nlinit"; \
     sed -e "s/\$$/\\$${nl}/" < input > output

12.13 The ‘obj/’ Subdirectory and Make

Never name one of your subdirectories ‘obj/’ if you don’t like surprises.

If an ‘obj/’ directory exists, BSD make enters it before reading the makefile. Hence the makefile in the current directory is not read.

$ cat Makefile
all:
        echo Hello
$ cat obj/Makefile
all:
        echo World
$ make      # GNU make
echo Hello
Hello
$ pmake     # BSD make
echo World
World

12.14 Exit Status of make -k

Do not rely on the exit status of make -k. Some implementations reflect whether they encountered an error in their exit status; other implementations always succeed.

$ cat Makefile
all:
        false
$ make -k; echo exit status: $?    # GNU make
false
make: *** [all] Error 1
exit status: 2
$ pmake -k; echo exit status: $?   # BSD make
false
*** Error code 1 (continuing)
exit status: 0

12.15 VPATH and Make

Posix does not specify the semantics of VPATH. Typically, make supports VPATH, but its implementation is not consistent.

Autoconf and Automake support makefiles whose usages of VPATH are portable to recentenough popular implementations of make, but to keep the resulting makefiles portable, a package’s makefile prototypes must take the following issues into account. These issues are complicated and are often poorly understood, and installers who use VPATH should expect to find many bugs in this area. If you use VPATH, the simplest way to avoid these portability bugs is to stick with GNU make, since it is the most commonly-used make among Autoconf users.

Here are some known issues with some VPATH implementations.

12.15.1 Variables listed in VPATH

Do not set VPATH to the value of another variable, for example ‘VPATH = $(srcdir)’, because some ancient versions of make do not do variable substitutions on the value of VPATH. For example, use this

     srcdir = @srcdir@
     VPATH = @srcdir@

rather than ‘VPATH = $(srcdir)’. Note that with GNU Automake, there is no need to set this yourself.

12.15.2 VPATH and Double-colon Rules

With ancient versions of Sun make, any assignment to VPATH causes make to execute only the first set of double-colon rules. However, this problem is no longer of practical concern.

12.15.3 $< Not Supported in Explicit Rules

Using $< in explicit rules is not portable. The prerequisite file must be named explicitly in the rule. If you want to find the prerequisite via a VPATH search, you have to code the whole thing manually. See Section 4.8.4 [Build Directories], page 33.

12.15.4 Automatic Rule Rewriting

Some make implementations, such as Solaris and Tru64, search for prerequisites in VPATH and then rewrite each occurrence as a plain word in the rule. For instance:

     # This isn't portable to GNU make.
     VPATH = ../pkg/src
     f.c: if.c
             cp if.c f.c
executes cp ../pkg/src/if.c f.c if ‘if.c’ is found in ‘../pkg/src’.

However, this rule leads to real problems in practice. For example, if the source directory contains an ordinary file named ‘test’ that is used in a dependency, Solaris make rewrites

commands like ‘if test -r foo; ...’ to ‘if ../pkg/src/test -r foo; ...’, which is typ-

ically undesirable. To avoid this problem, portable makefiles should never mention a source file whose name is that of a shell keyword like ‘until’ or a shell command like cat or gcc

or test.

Because of these problems GNU make and many other make implementations do not rewrite commands, so portable makefiles should search VPATH manually. It is tempting to write this:

      # This isn't portable to Solaris make.
      VPATH = ../pkg/src
      f.c: if.c
              cp `test -f if.c || echo $(VPATH)/`if.c f.c

However, the “prerequisite rewriting” still applies here. So if ‘if.c’ is in ‘../pkg/src’, Solaris and Tru64 make execute

      cp `test -f ../pkg/src/if.c || echo ../pkg/src/`if.c f.c

which reduces to

     cp if.c f.c

and thus fails. Oops.

A simple workaround, and good practice anyway, is to use ‘$?’ and ‘$@’ when possible:

      VPATH = ../pkg/src
      f.c: if.c
              cp $? $@

but this does not generalize well to commands with multiple prerequisites. A more general workaround is to rewrite the rule so that the prerequisite ‘if.c’ never appears as a plain word. For example, these three rules would be safe, assuming ‘if.c’ is in ‘../pkg/src’ and the other files are in the working directory:

      VPATH = ../pkg/src
      f.c: if.c f1.c
              cat `test -f ./if.c || echo $(VPATH)/`if.c f1.c >$@
      g.c: if.c g1.c
              cat `test -f 'if.c' || echo $(VPATH)/`if.c g1.c >$@
      h.c: if.c h1.c
              cat `test -f "if.c" || echo $(VPATH)/`if.c h1.c >$@

Things get worse when your prerequisites are in a macro.

     VPATH = ../pkg/src
     HEADERS = f.h g.h h.h
     install-HEADERS: $(HEADERS)
             for i in $(HEADERS); do \
               $(INSTALL) -m 644 \
                 `test -f $$i || echo $(VPATH)/`$$i \
                 $(DESTDIR)$(includedir)/$$i; \
             done

The above install-HEADERS rule is not Solaris-proof because for i in $(HEADERS); is expanded to for i in f.h g.h h.h; where f.h and g.h are plain words and are hence subject to VPATH adjustments.

If the three files are in ‘../pkg/src’, the rule is run as:

     for i in ../pkg/src/f.h ../pkg/src/g.h h.h; do \
       install -m 644 \
          `test -f $i || echo ../pkg/src/`$i \
          /usr/local/include/$i; \
     done

where the two first install calls fail. For instance, consider the f.h installation:

     install -m 644 \
       `test -f ../pkg/src/f.h || \
         echo ../pkg/src/ \
       `../pkg/src/f.h \
       /usr/local/include/../pkg/src/f.h;

It reduces to:

     install -m 644 \
       ../pkg/src/f.h \
       /usr/local/include/../pkg/src/f.h;

Note that the manual VPATH search did not cause any problems here; however this command installs ‘f.h’ in an incorrect directory.

Trying to quote $(HEADERS) in some way, as we did for foo.c a few makefiles ago, does not help:

     install-HEADERS: $(HEADERS)
             headers='$(HEADERS)'; \
             for i in $$headers; do \
               $(INSTALL) -m 644 \
                 `test -f $$i || echo $(VPATH)/`$$i \
                 $(DESTDIR)$(includedir)/$$i; \
             done
   Now, headers='$(HEADERS)' macro-expands to:
     headers='f.h g.h h.h'

but g.h is still a plain word. (As an aside, the idiom headers='$(HEADERS)'; for i in $$headers; is a good idea if $(HEADERS) can be empty, because some shells diagnose a syntax error on for i in;.)

One workaround is to strip this unwanted ‘../pkg/src/’ prefix manually:

     VPATH = ../pkg/src
     HEADERS = f.h g.h h.h
     install-HEADERS: $(HEADERS)
             headers='$(HEADERS)'; \
             for i in $$headers; do \
               i=`expr "$$i" : '$(VPATH)/\(.*\)'`;
               $(INSTALL) -m 644 \
                 `test -f $$i || echo $(VPATH)/`$$i \
                 $(DESTDIR)$(includedir)/$$i; \
             done

Automake does something similar. However the above hack works only if the files listed in HEADERS are in the current directory or a subdirectory; they should not be in an enclosing directory. If we had HEADERS = ../f.h, the above fragment would fail in a VPATH build with Tru64 make. The reason is that not only does Tru64 make rewrite dependencies, but it also simplifies them. Hence ../f.h becomes ../pkg/f.h instead of ../pkg/src/../f.h. This obviously defeats any attempt to strip a leading ‘../pkg/src/’ component.

The following example makes the behavior of Tru64 make more apparent.

     $ cat Makefile
     VPATH = sub
     all: ../foo
             echo ../foo
     $ ls
     Makefile foo
     $ make
     echo foo
     foo

Dependency ‘../foo’ was found in ‘sub/../foo’, but Tru64 make simplified it as ‘foo’. (Note that the ‘sub/’ directory does not even exist, this just means that the simplification occurred before the file was checked for.)

For the record here is how SunOS 4 make behaves on this example.

      $ make
      make: Fatal error: Don't know how to make target `../foo'
      $ mkdir sub
      $ make
      echo sub/../foo
      sub/../foo

12.15.5 Tru64 make Creates Prerequisite Directories Magically

When a prerequisite is a subdirectory of VPATH, Tru64 make creates it in the current directory.

     $ mkdir -p foo/bar build
     $ cd build
     $ cat >Makefile <<END
     VPATH = ..
     all: foo/bar
     END
     $ make
     mkdir foo
     mkdir foo/bar

This can yield unexpected results if a rule uses a manual VPATH search as presented before.

     VPATH = ..
     all : foo/bar
             command `test -d foo/bar || echo ../`foo/bar

The above command is run on the empty ‘foo/bar’ directory that was created in the current directory.

12.15.6 Make Target Lookup

GNU make uses a complex algorithm to decide when it should use files found via a VPATH search. See Section “How Directory Searches are Performed” in The GNU Make Manual.

If a target needs to be rebuilt, GNU make discards the file name found during the VPATH search for this target, and builds the file locally using the file name given in the makefile. If a target does not need to be rebuilt, GNU make uses the file name found during the VPATH search.

Other make implementations, like NetBSD make, are easier to describe: the file name found during the VPATH search is used whether the target needs to be rebuilt or not. Therefore new files are created locally, but existing files are updated at their VPATH location.

OpenBSD and FreeBSD make, however, never perform a VPATH search for a dependency that has an explicit rule. This is extremely annoying.

When attempting a VPATH build for an autoconfiscated package (e.g., mkdir build && cd build && ../configure), this means GNU make builds everything locally in the ‘build’ directory, while BSD make builds new files locally and updates existing files in the source directory.

     $ cat Makefile
     VPATH = ..
     all: foo.x bar.x

foo.x bar.x: newer.x
        @echo Building $@
$ touch ../bar.x
$ touch ../newer.x
$ make        # GNU make
Building foo.x
Building bar.x
$ pmake       # NetBSD make
Building foo.x
Building ../bar.x
$ fmake       # FreeBSD make, OpenBSD make
Building foo.x
Building bar.x
$ tmake       # Tru64 make
Building foo.x
Building bar.x
$ touch ../bar.x
$ make        # GNU make
Building foo.x
$ pmake       # NetBSD make
Building foo.x
$ fmake       # FreeBSD make, OpenBSD make
Building foo.x
Building bar.x
$ tmake       # Tru64 make
Building foo.x
Building bar.x

Note how NetBSD make updates ‘../bar.x’ in its VPATH location, and how FreeBSD, OpenBSD, and Tru64 make always update ‘bar.x’, even when ‘../bar.x’ is up to date.

Another point worth mentioning is that once GNU make has decided to ignore a VPATH file name (e.g., it ignored ‘../bar.x’ in the above example) it continues to ignore it when the target occurs as a prerequisite of another rule.

The following example shows that GNU make does not look up ‘bar.x’ in VPATH before performing the .x.y rule, because it ignored the VPATH result of ‘bar.x’ while running the

bar.x: newer.x rule.
$ cat Makefile
VPATH = ..
all: bar.y
bar.x: newer.x
        @echo Building $@
.SUFFIXES: .x .y
.x.y:
        cp $< $@
$ touch ../bar.x
$ touch ../newer.x
$ make        # GNU make

Building bar.x
cp bar.x bar.y
cp: cannot stat `bar.x': No such file or directory
make: *** [bar.y] Error 1
$ pmake       # NetBSD make
Building ../bar.x
cp ../bar.x bar.y
$ rm bar.y
$ fmake       # FreeBSD make, OpenBSD make
echo Building bar.x
cp bar.x bar.y
cp: cannot stat `bar.x': No such file or directory
*** Error code 1
$ tmake       # Tru64 make
Building bar.x
cp: bar.x: No such file or directory
*** Exit 1

Note that if you drop away the command from the bar.x: newer.x rule, GNU make magically starts to work: it knows that bar.x hasn’t been updated, therefore it doesn’t discard the result from VPATH (‘../bar.x’) in succeeding uses. Tru64 also works, but FreeBSD and OpenBSD still don’t.

$ cat Makefile
VPATH = ..
all: bar.y
bar.x: newer.x
.SUFFIXES: .x .y
.x.y:
        cp $< $@
$ touch ../bar.x
$ touch ../newer.x
$ make        # GNU make
cp ../bar.x bar.y
$ rm bar.y
$ pmake       # NetBSD make
cp ../bar.x bar.y
$ rm bar.y
$ fmake       # FreeBSD make, OpenBSD make
cp bar.x bar.y
cp: cannot stat `bar.x': No such file or directory
*** Error code 1
$ tmake       # Tru64 make
cp ../bar.x bar.y

It seems the sole solution that would please every make implementation is to never rely on VPATH searches for targets. In other words, VPATH should be reserved to unbuilt sources.

12.16 Single Suffix Rules and Separated Dependencies

A Single Suffix Rule is basically a usual suffix (inference) rule (‘.from.to:’), but which destination suffix is empty (‘.from:’).

Separated dependencies simply refers to listing the prerequisite of a target, without defining a rule. Usually one can list on the one hand side, the rules, and on the other hand side, the dependencies.

Solaris make does not support separated dependencies for targets defined by single suffix rules:

$ cat Makefile
.SUFFIXES: .in
foo: foo.in
.in:
        cp $< $@
$ touch foo.in
$ make
$ ls
Makefile  foo.in

while GNU Make does:

$ gmake
cp foo.in foo
$ ls
Makefile  foo       foo.in

Note it works without the ‘foo: foo.in’ dependency.

$ cat Makefile
.SUFFIXES: .in
.in:
        cp $< $@
$ make foo
cp foo.in foo

and it works with double suffix inference rules:

$ cat Makefile
foo.out: foo.in
.SUFFIXES: .in .out
.in.out:
        cp $< $@
$ make
cp foo.in foo.out

As a result, in such a case, you have to write target rules.

12.17 Timestamp Resolution and Make

Traditionally, file timestamps had 1-second resolution, and make used those timestamps to determine whether one file was newer than the other. However, many modern file systems have timestamps with 1-nanosecond resolution. Some make implementations look at the entire timestamp; others ignore the fractional part, which can lead to incorrect results.

Normally this is not a problem, but in some extreme cases you may need to use tricks like ‘sleep 1’ to work around timestamp truncation bugs.

Commands like ‘cp -p’ and ‘touch -r’ typically do not copy file timestamps to their full resolutions (see [Limitations of Usual Tools], page 238). Hence you should be wary of rules like this:

     dest: src
             cp -p src dest

as ‘dest’ often appears to be older than ‘src’ after the timestamp is truncated, and this can cause make to do needless rework the next time it is invoked. To work around this problem, you can use a timestamp file, e.g.:

     dest-stamp: src
             cp -p src dest
             date >dest-stamp

13 Portable C and C++ Programming

C and C++ programs often use low-level features of the underlying system, and therefore are often more difficult to make portable to other platforms.

Several standards have been developed to help make your programs more portable. If you write programs with these standards in mind, you can have greater confidence that your programs work on a wide variety of systems. See Section “Language Standards Supported by GCC” in Using the GNU Compiler Collection (GCC), for a list of C-related standards. Many programs also assume the Posix standard (http://www.opengroup.org/susv3).

Some old code is written to be portable to K&R C, which predates any C standard. K&R C compilers are no longer of practical interest, though, and the rest of section assumes at least C89, the first C standard.

Program portability is a huge topic, and this section can only briefly introduce common pitfalls. See Section “Portability between System Types” in GNU Coding Standards, for more information.

13.1 Varieties of Unportability

Autoconf tests and ordinary programs often need to test what is allowed on a system, and therefore they may need to deliberately exceed the boundaries of what the standards allow, if only to see whether an optional feature is present. When you write such a program, you should keep in mind the difference between constraints, unspecified behavior, and undefined behavior.

In C, a constraint is a rule that the compiler must enforce. An example constraint is that C programs must not declare a bit-field with negative width. Tests can therefore reliably assume that programs with negative-width bit-fields are rejected by a compiler that conforms to the standard.

Unspecified behavior is valid behavior, where the standard allows multiple possibilities. For example, the order of evaluation of function arguments is unspecified. Some unspecified behavior is implementation-defined, i.e., documented by the implementation, but since Autoconf tests cannot read the documentation they cannot distinguish between implementation-defined and other unspecified behavior. It is common for Autoconf tests to probe implementations to determine otherwise-unspecified behavior.

Undefined behavior is invalid behavior, where the standard allows the implementation to do anything it pleases. For example, dereferencing a null pointer leads to undefined behavior. If possible, test programs should avoid undefined behavior, since a program with undefined behavior might succeed on a test that should fail.

The above rules apply to programs that are intended to conform to the standard. However, strictly-conforming programs are quite rare, since the standards are so limiting. A major goal of Autoconf is to support programs that use implementation features not described by the standard, and it is fairly common for test programs to violate the above rules, if the programs work well enough in practice.

13.2 Integer Overflow

In practice many portable C programs assume that signed integer overflow wraps around reliably using two’s complement arithmetic. Yet the C standard says that program behavior is undefined on overflow, and in a few cases C programs do not work on some modern implementations because their overflows do not wrap around as their authors expected. Conversely, in signed integer remainder, the C standard requires overflow behavior that is commonly not implemented.

13.2.1 Basics of Integer Overflow

In languages like C, unsigned integer overflow reliably wraps around; e.g., UINT_MAX + 1 yields zero. This is guaranteed by the C standard and is portable in practice, unless you specify aggressive, nonstandard optimization options suitable only for special applications.

In contrast, the C standard says that signed integer overflow leads to undefined behavior where a program can do anything, including dumping core or overrunning a buffer. The misbehavior can even precede the overflow. Such an overflow can occur during addition, subtraction, multiplication, division, and left shift.

Despite this requirement of the standard, many C programs and Autoconf tests assume that signed integer overflow silently wraps around modulo a power of two, using two’s complement arithmetic, so long as you cast the resulting value to a signed integer type or store it into a signed integer variable. If you use conservative optimization flags, such programs are generally portable to the vast majority of modern platforms, with a few exceptions discussed later.

For historical reasons the C standard also allows implementations with ones’ complement or signed magnitude arithmetic, but it is safe to assume two’s complement nowadays.

Also, overflow can occur when converting an out-of-range value to a signed integer type. Here a standard implementation must define what happens, but this might include raising an exception. In practice all known implementations support silent wraparound in this case, so you need not worry about other possibilities.

13.2.2 Examples of Code Assuming Wraparound Overflow

There has long been a tension between what the C standard requires for signed integer overflow, and what C programs commonly assume. The standard allows aggressive optimizations based on assumptions that overflow never occurs, but many practical C programs rely on overflow wrapping around. These programs do not conform to the standard, but they commonly work in practice because compiler writers are understandably reluctant to implement optimizations that would break many programs, unless perhaps a user specifies aggressive optimization.

The C Standard says that if a program has signed integer overflow its behavior is undefined, and the undefined behavior can even precede the overflow. To take an extreme example:

     if (password == expected_password)
       allow_superuser_privileges ();
     else if (counter++ == INT_MAX)
       abort ();
     else

       printf ("%d password mismatches\n", counter);

If the int variable counter equals INT_MAX, counter++ must overflow and the behavior is undefined, so the C standard allows the compiler to optimize away the test against INT_ MAX and the abort call. Worse, if an earlier bug in the program lets the compiler deduce that counter == INT_MAX or that counter previously overflowed, the C standard allows the compiler to optimize away the password test and generate code that allows superuser privileges unconditionally.

Despite this requirement by the standard, it has long been common for C code to assume wraparound arithmetic after signed overflow, and all known practical C implementations support some C idioms that assume wraparound signed arithmetic, even if the idioms do not conform strictly to the standard. If your code looks like the following examples it will almost surely work with real-world compilers.

Here is an example derived from the 7th Edition Unix implementation of atoi (1979-01-10):

     char *p;
     int f, n;
     ...
     while (*p >= '0' && *p <= '9')
       n = n * 10 + *p++ - '0';
     return (f ? -n : n);

Even if the input string is in range, on most modern machines this has signed overflow when computing the most negative integer (the -n overflows) or a value near an extreme integer (the first + overflows).

Here is another example, derived from the 7th Edition implementation of rand (1979-01-10). Here the programmer expects both multiplication and addition to wrap on overflow:

     static long int randx = 1;
     ...
     randx = randx * 1103515245 + 12345;
     return (randx >> 16) & 077777;

In the following example, derived from the GNU C Library 2.5 implementation of mktime (2006-09-09), the code assumes wraparound arithmetic in + to detect signed overflow:

     time_t t, t1, t2;
     int sec_requested, sec_adjustment;
     ...
     t1 = t + sec_requested;
     t2 = t1 + sec_adjustment;
     if (((t1 < t) != (sec_requested < 0))
         | ((t2 < t1) != (sec_adjustment < 0)))
       return -1;

If your code looks like these examples, it is probably safe even though it does not strictly conform to the C standard. This might lead one to believe that one can generally assume wraparound on overflow, but that is not always true, as can be seen in the next section.

13.2.3 Optimizations That Break Wraparound Arithmetic

Compilers sometimes generate code that is incompatible with wraparound integer arithmetic. A simple example is an algebraic simplification: a compiler might translate (i * 2000) / 1000 to i * 2 because it assumes that i * 2000 does not overflow. The translation is not equivalent to the original when overflow occurs: e.g., in the typical case of 32-bit signed two’s complement wraparound int, if i has type int and value 1073742, the original expression returns −2147483 but the optimized version returns the mathematically correct value 2147484.

More subtly, loop induction optimizations often exploit the undefined behavior of signed overflow. Consider the following contrived function sumc:

     int
     sumc (int lo, int hi)
     {
       int sum = 0;
       int i;
       for (i = lo; i <= hi; i++)
         sum ^= i * 53;
       return sum;
     }

To avoid multiplying by 53 each time through the loop, an optimizing compiler might internally transform sumc to the equivalent of the following:

     int
     transformed_sumc (int lo, int hi)
     {
       int sum = 0;
       int hic = hi * 53;
       int ic;
       for (ic = lo * 53; ic <= hic; ic += 53)
         sum ^= ic;
       return sum;
     }

This transformation is allowed by the C standard, but it is invalid for wraparound arithmetic when INT_MAX / 53 < hi, because then the overflow in computing expressions like hi * 53 can cause the expression i <= hi to yield a different value from the transformed expression

ic <= hic.

For this reason, compilers that use loop induction and similar techniques often do not support reliable wraparound arithmetic when a loop induction variable like ic is involved. Since loop induction variables are generated by the compiler, and are not visible in the source code, it is not always trivial to say whether the problem affects your code.

Hardly any code actually depends on wraparound arithmetic in cases like these, so in practice these loop induction optimizations are almost always useful. However, edge cases in this area can cause problems. For example:

     int j;
     for (j = 1; 0 < j; j *= 2)
       test (j);

Here, the loop attempts to iterate through all powers of 2 that int can represent, but the C standard allows a compiler to optimize away the comparison and generate an infinite loop, under the argument that behavior is undefined on overflow. As of this writing this optimization is not done by any production version of GCC with ‘-O2’, but it might be performed by other compilers, or by more aggressive GCC optimization options, and the GCC developers have not decided whether it will continue to work with GCC and ‘-O2’.

13.2.4 Practical Advice for Signed Overflow Issues

Ideally the safest approach is to avoid signed integer overflow entirely. For example, instead of multiplying two signed integers, you can convert them to unsigned integers, multiply the unsigned values, then test whether the result is in signed range.

Rewriting code in this way will be inconvenient, though, particularly if the signed values might be negative. Also, it may hurt performance. Using unsigned arithmetic to check for overflow is particularly painful to do portably and efficiently when dealing with an integer type like uid_t whose width and signedness vary from platform to platform.

Furthermore, many C applications pervasively assume wraparound behavior and typically it is not easy to find and remove all these assumptions. Hence it is often useful to maintain nonstandard code that assumes wraparound on overflow, instead of rewriting the code. The rest of this section attempts to give practical advice for this situation.

If your code wants to detect signed integer overflow in sum = a + b, it is generally safe to use an expression like (sum < a) != (b < 0).

If your code uses a signed loop index, make sure that the index cannot overflow, along with all signed expressions derived from the index. Here is a contrived example of problematic code with two instances of overflow.

     for (i = INT_MAX - 10; i <= INT_MAX; i++)
       if (i + 1 < 0)
         {
           report_overflow ();
           break;
         }

Because of the two overflows, a compiler might optimize away or transform the two comparisons in a way that is incompatible with the wraparound assumption.

If your code uses an expression like (i * 2000) / 1000 and you actually want the multiplication to wrap around on overflow, use unsigned arithmetic to do it, e.g., ((int) (i *

2000u)) / 1000.

If your code assumes wraparound behavior and you want to insulate it against any GCC optimizations that would fail to support that behavior, you should use GCC’s ‘-fwrapv’ option, which causes signed overflow to wrap around reliably (except for division and remainder, as discussed in the next section).

If you need to port to platforms where signed integer overflow does not reliably wrap around (e.g., due to hardware overflow checking, or to highly aggressive optimizations), you should consider debugging with GCC’s ‘-ftrapv’ option, which causes signed overflow to raise an exception.

13.2.5 Signed Integer Division and Integer Overflow

Overflow in signed integer division is not always harmless: for example, on CPUs of the i386 family, dividing INT_MIN by -1 yields a SIGFPE signal which by default terminates the program. Worse, taking the remainder of these two values typically yields the same signal on these CPUs, even though the C standard requires INT_MIN % -1 to yield zero because the expression does not overflow.

13.3 Preprocessor Arithmetic

In C99, preprocessor arithmetic, used for #if expressions, must be evaluated as if all signed values are of type intmax_t and all unsigned values of type uintmax_t. Many compilers are buggy in this area, though. For example, as of 2007, Sun C mishandles #if LLONG_ MIN < 0 on a platform with 32-bit long int and 64-bit long long int. Also, some older preprocessors mishandle constants ending in LL. To work around these problems, you can compute the value of expressions like LONG_MAX < LLONG_MAX at configure-time rather than at #if-time.

13.4 Properties of Null Pointers

Most modern hosts reliably fail when you attempt to dereference a null pointer.

On almost all modern hosts, null pointers use an all-bits-zero internal representation, so you can reliably use memset with 0 to set all the pointers in an array to null values.

If p is a null pointer to an object type, the C expression p + 0 always evaluates to p on modern hosts, even though the standard says that it has undefined behavior.

13.5 Buffer Overruns and Subscript Errors

Buffer overruns and subscript errors are the most common dangerous errors in C programs. They result in undefined behavior because storing outside an array typically modifies storage that is used by some other object, and most modern systems lack runtime checks to catch these errors. Programs should not rely on buffer overruns being caught.

There is one exception to the usual rule that a portable program cannot address outside an array. In C, it is valid to compute the address just past an object, e.g., &a[N] where a has N elements, so long as you do not dereference the resulting pointer. But it is not valid to compute the address just before an object, e.g., &a[-1]; nor is it valid to compute two past the end, e.g., &a[N+1]. On most platforms &a[-1] < &a[0] && &a[N] < &a[N+1], but this is not reliable in general, and it is usually easy enough to avoid the potential portability problem, e.g., by allocating an extra unused array element at the start or end.

Valgrind (http://valgrind.org/) can catch many overruns. GCC users might also consider using the ‘-fmudflap’ option to catch overruns.

Buffer overruns are usually caused by off-by-one errors, but there are more subtle ways to get them.

Using int values to index into an array or compute array sizes causes problems on typical 64-bit hosts where an array index might be 2^{3}1 or larger. Index values of type size_t avoid this problem, but cannot be negative. Index values of type ptrdiff_t are signed, and are wide enough in practice.

If you add or multiply two numbers to calculate an array size, e.g., malloc (x * sizeof y + z), havoc ensues if the addition or multiplication overflows.

Many implementations of the alloca function silently misbehave and can generate buffer overflows if given sizes that are too large. The size limits are implementation dependent, but are at least 4000 bytes on all platforms that we know about.

The standard functions asctime, asctime_r, ctime, ctime_r, and gets are prone to buffer overflows, and portable code should not use them unless the inputs are known to be within certain limits. The time-related functions can overflow their buffers if given timestamps out of range (e.g., a year less than -999 or greater than 9999). Time-related buffer overflows cannot happen with recent-enough versions of the GNU C library, but are possible with other implementations. The gets function is the worst, since it almost invariably overflows its buffer when presented with an input line larger than the buffer.

13.6 Volatile Objects

The keyword volatile is often misunderstood in portable code. Its use inhibits some memory-access optimizations, but programmers often wish that it had a different meaning than it actually does.

volatile was designed for code that accesses special objects like memory-mapped device registers whose contents spontaneously change. Such code is inherently low-level, and it is difficult to specify portably what volatile means in these cases. The C standard says, “What constitutes an access to an object that has volatile-qualified type is implementationdefined,” so in theory each implementation is supposed to fill in the gap by documenting what volatile means for that implementation. In practice, though, this documentation is usually absent or incomplete.

One area of confusion is the distinction between objects defined with volatile types, and volatile lvalues. From the C standard’s point of view, an object defined with a volatile type has externally visible behavior. You can think of such objects as having little oscilloscope probes attached to them, so that the user can observe some properties of accesses to them, just as the user can observe data written to output files. However, the standard does not make it clear whether users can observe accesses by volatile lvalues to ordinary objects. For example:

     /* Declare and access a volatile object.
        Accesses to X are "visible" to users.  */
     static int volatile x;
     x = 1;
     /* Access two ordinary objects via a volatile lvalue.
        It's not clear whether accesses to *P are "visible".  */
     int y;
     int *z = malloc (sizeof (int));
     int volatile *p;
     p = &y;
     *p = 1;
     p = z;
     *p = 1;

Programmers often wish that volatile meant “Perform the memory access here and now, without merging several memory accesses, without changing the memory word size, and without reordering.” But the C standard does not require this. For objects defined with a volatile type, accesses must be done before the next sequence point; but otherwise merging, reordering, and word-size change is allowed. Worse, it is not clear from the standard whether volatile lvalues provide more guarantees in general than nonvolatile lvalues, if the underlying objects are ordinary.

Even when accessing objects defined with a volatile type, the C standard allows only extremely limited signal handlers: the behavior is undefined if a signal handler reads any nonlocal object, or writes to any nonlocal object whose type is not sig_atomic_t volatile, or calls any standard library function other than abort, signal, and (if C99) _Exit. Hence C compilers need not worry about a signal handler disturbing ordinary computation, unless the computation accesses a sig_atomic_t volatile lvalue that is not a local variable. (There is an obscure exception for accesses via a pointer to a volatile character, since it may point into part of a sig_atomic_t volatile object.) Posix adds to the list of library functions callable from a portable signal handler, but otherwise is like the C standard in this area.

Some C implementations allow memory-access optimizations within each translation unit, such that actual behavior agrees with the behavior required by the standard only when calling a function in some other translation unit, and a signal handler acts like it was called from a different translation unit. The C standard hints that in these implementations, objects referred to by signal handlers “would require explicit specification of volatile storage, as well as other implementation-defined restrictions.” But unfortunately even for this special case these other restrictions are often not documented well. See Section “When is a Volatile Object Accessed?” in Using the GNU Compiler Collection (GCC), for some restrictions imposed by GCC. See Section “Defining Signal Handlers” in The GNU C Library, for some restrictions imposed by the GNU C library. Restrictions differ on other platforms.

If possible, it is best to use a signal handler that fits within the limits imposed by the C and Posix standards.

If this is not practical, you can try the following rules of thumb. A signal handler should access only volatile lvalues, preferably lvalues that refer to objects defined with a volatile type, and should not assume that the accessed objects have an internally consistent state if they are larger than a machine word. Furthermore, installers should employ compilers and compiler options that are commonly used for building operating system kernels, because kernels often need more from volatile than the C Standard requires, and installers who compile an application in a similar environment can sometimes benefit from the extra constraints imposed by kernels on compilers. Admittedly we are handwaving somewhat here, as there are few guarantees in this area; the rules of thumb may help to fix some bugs but there is a good chance that they will not fix them all.

For volatile, C++ has the same problems that C does. Multithreaded applications have even more problems with volatile, but they are beyond the scope of this section.

The bottom line is that using volatile typically hurts performance but should not hurt correctness. In some cases its use does help correctness, but these cases are often so poorly understood that all too often adding volatile to a data structure merely alleviates some symptoms of a bug while not fixing the bug in general.

13.7 Floating Point Portability

Almost all modern systems use IEEE-754 floating point, and it is safe to assume IEEE-754 in most portable code these days. For more information, please see David Goldberg’s classic paper What Every Computer Scientist Should Know About Floating-Point Arithmetic

(http://www.validlab.com/goldberg/paper.pdf).

13.8 Exiting Portably

A C or C++ program can exit with status N by returning N from the main function. Portable programs are supposed to exit either with status 0 or EXIT_SUCCESS to succeed, or with status EXIT_FAILURE to fail, but in practice it is portable to fail by exiting with status 1, and test programs that assume Posix can fail by exiting with status values from 1 through 255. Programs on SunOS 2.0 (1985) through 3.5.2 (1988) incorrectly exited with zero status when main returned nonzero, but ancient systems like these are no longer of practical concern.

A program can also exit with status N by passing N to the exit function, and a program can fail by calling the abort function. If a program is specialized to just some platforms, it can fail by calling functions specific to those platforms, e.g., _exit (Posix) and _Exit (C99). However, like other functions, an exit function should be declared, typically by including a header. For example, if a C program calls exit, it should include ‘stdlib.h’ either directly or via the default includes (see Section 5.1.2 [Default Includes], page 43).

A program can fail due to undefined behavior such as dereferencing a null pointer, but this is not recommended as undefined behavior allows an implementation to do whatever it pleases and this includes exiting successfully.

14 Manual Configuration

A few kinds of features can’t be guessed automatically by running test programs. For example, the details of the object-file format, or special options that need to be passed to the compiler or linker. You can check for such features using ad-hoc means, such as having configure check the output of the uname program, or looking for libraries that are unique to particular systems. However, Autoconf provides a uniform method for handling unguessable features.

14.1 Specifying target triplets

Autoconf-generated configure scripts can make decisions based on a canonical name for the system type, or target triplet, which has the form: ‘cpu-vendor-os’, where os can be

‘system’ or ‘kernel-system’

configure can usually guess the canonical name for the type of system it’s running on. To do so it runs a script called config.guess, which infers the name using the uname command or symbols predefined by the C preprocessor.

Alternately, the user can specify the system type with command line arguments to configure (see Section 16.7 [System Type], page 284. Doing so is necessary when crosscompiling. In the most complex case of cross-compiling, three system types are involved. The options to specify them are:

‘--build=build-type’

the type of system on which the package is being configured and compiled. It defaults to the result of running config.guess.

‘--host=host-type’

the type of system on which the package runs. By default it is the same as the build machine. Specifying it enables the cross-compilation mode.

‘--target=target-type’

the type of system for which any compiler tools in the package produce code (rarely needed). By default, it is the same as host.

If you mean to override the result of config.guess, use ‘--build’, not ‘--host’, since the latter enables cross-compilation. For historical reasons, whenever you specify ‘--host’, be sure to specify ‘--build’ too; this will be fixed in the future. So, to enter cross-compilation mode, use a command like this

     ./configure --build=i686-pc-linux-gnu --host=m68k-coff

Note that if you do not specify ‘--host’, configure fails if it can’t run the code generated by the specified compiler. For example, configuring as follows fails:

     ./configure CC=m68k-coff-gcc

When cross-compiling, configure will warn about any tools (compilers, linkers, assemblers) whose name is not prefixed with the host type. This is an aid to users performing cross-compilation. Continuing the example above, if a cross-compiler named cc is used with a native pkg-config, then libraries found by pkg-config will likely cause subtle build failures; but using the names m68k-coff-cc and m68k-coff-pkg-config avoids any confusion. Avoiding the warning is as simple as creating the correct symlinks naming the cross tools.

configure recognizes short aliases for many system types; for example, ‘decstation’ can be used instead of ‘mips-dec-ultrix4.2’. configure runs a script called config.sub to canonicalize system type aliases.

This section deliberately omits the description of the obsolete interface; see Section 18.6.3 [Hosts and Cross-Compilation], page 310.

14.2 Getting the Canonical System Type

The following macros make the system type available to configure scripts.

The variables ‘build_alias’, ‘host_alias’, and ‘target_alias’ are always exactly the arguments of ‘--build’, ‘--host’, and ‘--target’; in particular, they are left empty if the user did not use them, even if the corresponding AC_CANONICAL macro was run. Any configure script may use these variables anywhere. These are the variables that should be used when in interaction with the user.

If you need to recognize some special environments based on their system type, run the following macros to get canonical system names. These variables are not set before the macro call.

If you use these macros, you must distribute config.guess and config.sub along with your source code. See Section 4.5 [Output], page 22, for information about the AC_CONFIG_ AUX_DIR macro which you can use to control in which directory configure looks for those scripts.

[Macro] AC_CANONICAL_BUILD

Compute the canonical build-system type variable, build, and its three individual

     parts build_cpu, build_vendor, and build_os.

If ‘--build’ was specified, then build is the canonicalization of build_alias by config.sub, otherwise it is determined by the shell script config.guess.

[Macro] AC_CANONICAL_HOST

Compute the canonical host-system type variable, host, and its three individual parts

     host_cpu, host_vendor, and host_os.

If ‘--host’ was specified, then host is the canonicalization of host_alias by config.sub, otherwise it defaults to build.

[Macro] AC_CANONICAL_TARGET

Compute the canonical target-system type variable, target, and its three individual

     parts target_cpu, target_vendor, and target_os.

If ‘--target’ was specified, then target is the canonicalization of target_alias by config.sub, otherwise it defaults to host.

Note that there can be artifacts due to the backward compatibility code. See See Section 18.6.3 [Hosts and Cross-Compilation], page 310, for more.

14.3 Using the System Type

In ‘configure.ac’ the system type is generally used by one or more case statements to select system-specifics. Shell wildcards can be used to match a group of system types.

For example, an extra assembler code object file could be chosen, giving access to a CPU cycle counter register. $(CYCLE_OBJ) in the following would be used in a makefile to add the object to a program or library.

     AS_CASE([$host],
       [alpha*-*-*], [CYCLE_OBJ=rpcc.o],
       [i?86-*-*],   [CYCLE_OBJ=rdtsc.o],
       [CYCLE_OBJ=""]
     )
     AC_SUBST([CYCLE_OBJ])

AC_CONFIG_LINKS (see Section 4.11 [Configuration Links], page 40) is another good way to select variant source files, for example optimized code for some CPUs. The configured CPU type doesn’t always indicate exact CPU types, so some runtime capability checks may be necessary too.

     case $host in
       alpha*-*-*)   AC_CONFIG_LINKS([dither.c:alpha/dither.c]) ;;
       powerpc*-*-*) AC_CONFIG_LINKS([dither.c:powerpc/dither.c]) ;;
       *-*-*)        AC_CONFIG_LINKS([dither.c:generic/dither.c]) ;;
     esac

The host system type can also be used to find cross-compilation tools with AC_CHECK_ TOOL (see Section 5.2.2 [Generic Programs], page 48).

The above examples all show ‘$host’, since this is where the code is going to run. Only rarely is it necessary to test ‘$build’ (which is where the build is being done).

Whenever you’re tempted to use ‘$host’ it’s worth considering whether some sort of probe would be better. New system types come along periodically or previously missing features are added. Well-written probes can adapt themselves to such things, but hardcoded lists of names can’t. Here are some guidelines,

• Availability of libraries and library functions should always be checked by probing.

• Variant behavior of system calls is best identified with runtime tests if possible, but bug workarounds or obscure difficulties might have to be driven from ‘$host’.

• Assembler code is inevitably highly CPU-specific and is best selected according to

    ‘$host_cpu’.

• Assembler variations like underscore prefix on globals or ELF versus COFF type directives are however best determined by probing, perhaps even examining the compiler output.

‘$target’ is for use by a package creating a compiler or similar. For ordinary packages it’s meaningless and should not be used. It indicates what the created compiler should generate code for, if it can cross-compile. ‘$target’ generally selects various hard-coded CPU and system conventions, since usually the compiler or tools under construction themselves determine how the target works.

15 Site Configuration

configure scripts support several kinds of local configuration decisions. There are ways for users to specify where external software packages are, include or exclude optional features, install programs under modified names, and set default values for configure options.

15.1 Controlling Help Output

Users consult ‘configure --help’ to learn of configuration decisions specific to your package. By default, configure breaks this output into sections for each type of option; within each section, help strings appear in the order ‘configure.ac’ defines them:

     Optional Features:
       ...
       --enable-bar            include bar
     Optional Packages:
       ...
       --with-foo              use foo
[Macro] AC_PRESERVE_HELP_ORDER

Request an alternate ‘--help’ format, in which options of all types appear together, in the order defined. Call this macro before any AC_ARG_ENABLE or AC_ARG_WITH.

          Optional Features and Packages:
            ...
            --enable-bar            include bar
            --with-foo              use foo

15.2 Working With External Software

Some packages require, or can optionally use, other software packages that are already installed. The user can give configure command line options to specify which such external software to use. The options have one of these forms:

     --with-package[=arg]
     --without-package

For example, ‘--with-gnu-ld’ means work with the GNU linker instead of some other linker. ‘--with-x’ means work with The X Window System.

The user can give an argument by following the package name with ‘=’ and the argument. Giving an argument of ‘no’ is for packages that are used by default; it says to not use the package. An argument that is neither ‘yes’ nor ‘no’ could include a name or number of a version of the other package, to specify more precisely which other package this program is supposed to work with. If no argument is given, it defaults to ‘yes’. ‘--without-package’

is equivalent to ‘--with-package=no’.

Normally configure scripts complain about ‘--with-package’ options that they do not support. See Section 15.5 [Option Checking], page 275, for details, and for how to override the defaults.

For each external software package that may be used, ‘configure.ac’ should call AC_ ARG_WITH to detect whether the configure user asked to use it. Whether each package is used or not by default, and which arguments are valid, is up to you.

[Macro] AC_ARG_WITH (package, help-string, [action-if-given],
   [action-if-not-given])
If the user gave configure the option ‘--with-package’ or ‘--without-package’,

run shell commands action-if-given. If neither option was given, run shell commands action-if-not-given. The name package indicates another software package that this program should work with. It should consist only of alphanumeric characters, dashes, plus signs, and dots.

The option’s argument is available to the shell commands action-if-given in the shell variable withval, which is actually just the value of the shell variable named with_ package, with any non-alphanumeric characters in package changed into ‘_’. You may use that variable instead, if you wish.

The argument help-string is a description of the option that looks like this:

       --with-readline         support fancy command line editing

help-string may be more than one line long, if more detail is needed. Just make sure the columns line up in ‘configure --help’. Avoid tabs in the help string. The easiest way to provide the proper leading whitespace is to format your help-string with the macro AS_HELP_STRING (see Section 15.4 [Pretty Help Strings], page 274).

The following example shows how to use the AC_ARG_WITH macro in a common situation. You want to let the user decide whether to enable support for an external library (e.g., the readline library); if the user specified neither ‘--with-readline’ nor ‘--without-readline’, you want to enable support for readline only if the library is available on the system.

     AC_ARG_WITH([readline],
       [AS_HELP_STRING([--with-readline],
         [support fancy command line editing @<:@default=check@:>@])],
       [],
       [with_readline=check])
     LIBREADLINE=
     AS_IF([test "x$with_readline" != xno],
       [AC_CHECK_LIB([readline], [main],
         [AC_SUBST([LIBREADLINE], ["-lreadline -lncurses"])
          AC_DEFINE([HAVE_LIBREADLINE], [1],
                    [Define if you have libreadline])
         ],
         [if test "x$with_readline" != xcheck; then
            AC_MSG_FAILURE(
              [--with-readline was given, but test for readline failed])
          fi
         ], -lncurses)])

The next example shows how to use AC_ARG_WITH to give the user the possibility to enable support for the readline library, in case it is still experimental and not well tested, and is therefore disabled by default.

AC_ARG_WITH([readline],
[AS_HELP_STRING([--with-readline],
[enable experimental support for readline])],
[],
[with_readline=no])
LIBREADLINE=
AS_IF([test "x$with_readline" != xno],
[AC_CHECK_LIB([readline], [main],
[AC_SUBST([LIBREADLINE], ["-lreadline -lncurses"])
 AC_DEFINE([HAVE_LIBREADLINE], [1],
           [Define if you have libreadline])
],
[AC_MSG_FAILURE(
   [--with-readline was given, but test for readline failed])],
[-lncurses])])

The last example shows how to use AC_ARG_WITH to give the user the possibility to disable support for the readline library, given that it is an important feature and that it should be enabled by default.

AC_ARG_WITH([readline],
[AS_HELP_STRING([--without-readline],
[disable support for readline])],
[],
[with_readline=yes])
LIBREADLINE=
AS_IF([test "x$with_readline" != xno],
[AC_CHECK_LIB([readline], [main],
[AC_SUBST([LIBREADLINE], ["-lreadline -lncurses"])
 AC_DEFINE([HAVE_LIBREADLINE], [1],
           [Define if you have libreadline])
],
[AC_MSG_FAILURE(
   [readline test failed (--without-readline to disable)])],
[-lncurses])])

These three examples can be easily adapted to the case where AC_ARG_ENABLE should be preferred to AC_ARG_WITH (see Section 15.3 [Package Options], page 273).

15.3 Choosing Package Options

If a software package has optional compile-time features, the user can give configure command line options to specify whether to compile them. The options have one of these forms:

--enable-feature[=arg]
--disable-feature

These options allow users to choose which optional features to build and install. ‘--enable-feature’ options should never make a feature behave differently or cause one feature to replace another. They should only cause parts of the program to be built rather than left out.

The user can give an argument by following the feature name with ‘=’ and the argument. Giving an argument of ‘no’ requests that the feature not be made available. A feature with an argument looks like ‘--enable-debug=stabs’. If no argument is given, it defaults to

‘yes’. ‘--disable-feature’ is equivalent to ‘--enable-feature=no’.

Normally configure scripts complain about ‘--enable-package’ options that they do not support. See Section 15.5 [Option Checking], page 275, for details, and for how to override the defaults.

For each optional feature, ‘configure.ac’ should call AC_ARG_ENABLE to detect whether the configure user asked to include it. Whether each feature is included or not by default, and which arguments are valid, is up to you.

[Macro] AC_ARG_ENABLE (feature, help-string, [action-if-given],
   [action-if-not-given])
If the user gave configure the option ‘--enable-feature’ or ‘--disable-feature’,

run shell commands action-if-given. If neither option was given, run shell commands action-if-not-given. The name feature indicates an optional user-level facility. It should consist only of alphanumeric characters, dashes, plus signs, and dots. The option’s argument is available to the shell commands action-if-given in the shell variable enableval, which is actually just the value of the shell variable named enable_feature, with any non-alphanumeric characters in feature changed into ‘_’. You may use that variable instead, if you wish. The help-string argument is like that of AC_ARG_WITH (see Section 15.2 [External Software], page 271). You should format your help-string with the macro AS_HELP_STRING (see Section 15.4 [Pretty Help Strings], page 274). See the examples suggested with the definition of AC_ARG_WITH (see Section 15.2 [External Software], page 271) to get an idea of possible applications of AC_ARG_

ENABLE.

15.4 Making Your Help Strings Look Pretty

Properly formatting the ‘help strings’ which are used in AC_ARG_WITH (see Section 15.2 [External Software], page 271) and AC_ARG_ENABLE (see Section 15.3 [Package Options], page 273) can be challenging. Specifically, you want your own ‘help strings’ to line up in the appropriate columns of ‘configure --help’ just like the standard Autoconf ‘help strings’ do. This is the purpose of the AS_HELP_STRING macro.

[Macro] AS_HELP_STRING (left-hand-side, right-hand-side [indent-column
   = ‘26’], [wrap-column = ‘79’])

Expands into a help string that looks pretty when the user executes ‘configure --help’. It is typically used in AC_ARG_WITH (see Section 15.2 [External Software], page 271) or AC_ARG_ENABLE (see Section 15.3 [Package Options], page 273). The following example makes this clearer.

     AC_ARG_WITH([foo],
       [AS_HELP_STRING([--with-foo],
          [use foo (default is no)])],
       [use_foo=$withval],
       [use_foo=no])

Then the last few lines of ‘configure --help’ appear like this:

     --enable and --with options recognized:
       --with-foo              use foo (default is no)

Macro expansion is performed on the first argument. However, the second argument of AS_HELP_STRING is treated as a whitespace separated list of text to be reformatted, and is not subject to macro expansion. Since it is not expanded, it should not be double quoted. See Section 3.1.2 [Autoconf Language], page 9, for a more detailed explanation.

The AS_HELP_STRING macro is particularly helpful when the left-hand-side and/or right-hand-side are composed of macro arguments, as shown in the following example. Be aware that left-hand-side may not expand to unbalanced quotes, although quadrigraphs can be used.

     AC_DEFUN([MY_ARG_WITH],
       [AC_ARG_WITH(m4_translit([[$1]], [_], [-]),
          [AS_HELP_STRING([--with-m4_translit([$1], [_], [-])],
                          [use $1 (default is $2)])],
          [use_[]$1=$withval],
          [use_[]$1=$2])])
     MY_ARG_WITH([a_b], [no])

Here, the last few lines of ‘configure --help’ will include:

     --enable and --with options recognized:
       --with-a-b              use a_b (default is no)

The parameters indent-column and wrap-column were introduced in Autoconf 2.62. Generally, they should not be specified; they exist for fine-tuning of the wrapping.

     AS_HELP_STRING([--option], [description of option])
     ⇒ --option                description of option
     AS_HELP_STRING([--option], [description of option], [15], [30])
     ⇒ --option     description of
     ⇒  option

15.5 Controlling Checking of configure Options

The configure script checks its command-line options against a list of known options, like ‘--help’ or ‘--config-cache’. An unknown option ordinarily indicates a mistake by the user and configure halts with an error. However, by default unknown ‘--with-package’ and ‘--enable-feature’ options elicit only a warning, to support configuring entire source trees.

Source trees often contain multiple packages with a top-level configure script that uses the AC_CONFIG_SUBDIRS macro (see Section 4.12 [Subdirectories], page 40). Because the packages generally support different ‘--with-package’ and ‘--enable-feature’ options, the GNU Coding Standards say they must accept unrecognized options without halting. Even a warning message is undesirable here, so AC_CONFIG_SUBDIRS automatically disables the warnings.

This default behavior may be modified in two ways. First, the installer can invoke

configure --disable-option-checking to disable these warnings, or invoke configure

--enable-option-checking=fatal options to turn them into fatal errors, respectively. Second, the maintainer can use AC_DISABLE_OPTION_CHECKING.

[Macro] AC_DISABLE_OPTION_CHECKING

By default, disable warnings related to any unrecognized ‘--with-package’ or

     ‘--enable-feature’ options. This is implied by AC_CONFIG_SUBDIRS.

The installer can override this behavior by passing ‘--enable-option-checking’ (en-

     able warnings) or ‘--enable-option-checking=fatal’ (enable errors) to configure.

15.6 Configuring Site Details

Some software packages require complex site-specific information. Some examples are host names to use for certain services, company names, and email addresses to contact. Since some configuration scripts generated by Metaconfig ask for such information interactively, people sometimes wonder how to get that information in Autoconf-generated configuration scripts, which aren’t interactive.

Such site configuration information should be put in a file that is edited only by users, not by programs. The location of the file can either be based on the prefix variable, or be a standard location such as the user’s home directory. It could even be specified by an environment variable. The programs should examine that file at runtime, rather than at compile time. Runtime configuration is more convenient for users and makes the configuration process simpler than getting the information while configuring. See Section “Variables for Installation Directories” in GNU Coding Standards, for more information on where to put data files.

15.7 Transforming Program Names When Installing

Autoconf supports changing the names of programs when installing them. In order to use these transformations, ‘configure.ac’ must call the macro AC_ARG_PROGRAM.

[Macro] AC_ARG_PROGRAM

Place in output variable program_transform_name a sequence of sed commands for changing the names of installed programs.

If any of the options described below are given to configure, program names are transformed accordingly. Otherwise, if AC_CANONICAL_TARGET has been called and a ‘--target’ value is given, the target type followed by a dash is used as a prefix. Otherwise, no program name transformation is done.

15.7.1 Transformation Options

You can specify name transformations by giving configure these command line options:

‘--program-prefix=prefix’

prepend prefix to the names;

‘--program-suffix=suffix’

append suffix to the names;

‘--program-transform-name=expression’

perform sed substitution expression on the names.

15.7.2 Transformation Examples

These transformations are useful with programs that can be part of a cross-compilation development environment. For example, a cross-assembler running on a Sun 4 configured with ‘--target=i960-vxworks’ is normally installed as ‘i960-vxworks-as’, rather than ‘as’, which could be confused with a native Sun 4 assembler.

You can force a program name to begin with ‘g’, if you don’t want GNU programs installed on your system to shadow other programs with the same name. For example, if you configure GNU diff with ‘--program-prefix=g’, then when you run ‘make install’

it is installed as ‘/usr/local/bin/gdiff’.

As a more sophisticated example, you could use

     --program-transform-name='s/^/g/; s/^gg/g/; s/^gless/less/'

to prepend ‘g’ to most of the program names in a source tree, excepting those like gdb that already have one and those like less and lesskey that aren’t GNU programs. (That is assuming that you have a source tree containing those programs that is set up to use this feature.)

One way to install multiple versions of some programs simultaneously is to append a version number to the name of one or both. For example, if you want to keep Autoconf version 1 around for awhile, you can configure Autoconf version 2 using ‘--program-suffix=2’

to install the programs as ‘/usr/local/bin/autoconf2’, ‘/usr/local/bin/autoheader2’,

etc. Nevertheless, pay attention that only the binaries are renamed, therefore you’d have problems with the library files which might overlap.

15.7.3 Transformation Rules

Here is how to use the variable program_transform_name in a ‘Makefile.in’:
     PROGRAMS = cp ls rm
     transform = @program_transform_name@
     install:
             for p in $(PROGRAMS); do \
               $(INSTALL_PROGRAM) $$p $(DESTDIR)$(bindir)/`echo $$p | \
                                                   sed '$(transform)'`; \
             done
     uninstall:
             for p in $(PROGRAMS); do \

               rm -f $(DESTDIR)$(bindir)/`echo $$p | sed '$(transform)'`; \
             done

It is guaranteed that program_transform_name is never empty, and that there are no useless separators. Therefore you may safely embed program_transform_name within a sed program using ‘;’:

     transform = @program_transform_name@
     transform_exe = s/$(EXEEXT)$$//;$(transform);s/$$/$(EXEEXT)/

Whether to do the transformations on documentation files (Texinfo or man) is a tricky question; there seems to be no perfect answer, due to the several reasons for name transforming. Documentation is not usually particular to a specific architecture, and Texinfo files do not conflict with system documentation. But they might conflict with earlier versions of the same files, and man pages sometimes do conflict with system documentation. As a compromise, it is probably best to do name transformations on man pages but not on Texinfo manuals.

15.8 Setting Site Defaults

Autoconf-generated configure scripts allow your site to provide default values for some configuration values. You do this by creating site- and system-wide initialization files.

If the environment variable CONFIG_SITE is set, configure uses its value as the name of a shell script to read; it is recommended that this be an absolute file name. Otherwise, it reads

the shell script ‘prefix/share/config.site’ if it exists, then ‘prefix/etc/config.site’

if it exists. Thus, settings in machine-specific files override those in machine-independent ones in case of conflict.

Site files can be arbitrary shell scripts, but only certain kinds of code are really appropriate to be in them. Because configure reads any cache file after it has read any site files, a site file can define a default cache file to be shared between all Autoconf-generated configure scripts run on that system (see Section 7.4.2 [Cache Files], page 117). If you set a default cache file in a site file, it is a good idea to also set the output variable CC in that site file, because the cache file is only valid for a particular compiler, but many systems have several available.

You can examine or override the value set by a command line option to configure in a site file; options set shell variables that have the same names as the options, with any dashes turned into underscores. The exceptions are that ‘--without-’ and ‘--disable-’ options are like giving the corresponding ‘--with-’ or ‘--enable-’ option and the value ‘no’.

Thus, ‘--cache-file=localcache’ sets the variable cache_file to the value ‘localcache’;
‘--enable-warnings=no’ or ‘--disable-warnings’ sets the variable enable_warnings to

the value ‘no’; ‘--prefix=/usr’ sets the variable prefix to the value ‘/usr’; etc.

Site files are also good places to set default values for other output variables, such as CFLAGS, if you need to give them non-default values: anything you would normally do, repetitively, on the command line. If you use non-default values for prefix or exec prefix (wherever you locate the site file), you can set them in the site file if you specify it with the CONFIG_SITE environment variable.

You can set some cache values in the site file itself. Doing this is useful if you are cross-compiling, where it is impossible to check features that require running a test program. You could “prime the cache” by setting those values correctly for that system in ‘prefix/etc/config.site’. To find out the names of the cache variables you need to set, see the documentation of the respective Autoconf macro. If the variables or their semantics are undocumented, you may need to look for shell variables with ‘_cv_’ in their names in the affected configure scripts, or in the Autoconf M4 source code for those macros; but in that case, their name or semantics may change in a future Autoconf version.

The cache file is careful to not override any variables set in the site files. Similarly, you should not override command-line options in the site files. Your code should check that variables such as prefix and cache_file have their default values (as set near the top of configure) before changing them.

Here is a sample file ‘/usr/share/local/gnu/share/config.site’. The command
‘configure --prefix=/usr/share/local/gnu’ would read this file (if CONFIG_SITE is not

set to a different file).

# /usr/share/local/gnu/share/config.site for configure
#
# Change some defaults.
test "$prefix" = NONE && prefix=/usr/share/local/gnu
test "$exec_prefix" = NONE && exec_prefix=/usr/local/gnu
test "$sharedstatedir" = '${prefix}/com' && sharedstatedir=/var
test "$localstatedir" = '${prefix}/var' && localstatedir=/var
# Give Autoconf 2.x generated configure scripts a shared default
# cache file for feature test results, architecture-specific.
if test "$cache_file" = /dev/null; then
  cache_file="$prefix/var/config.cache"
  # A cache file is only valid for one C compiler.
  CC=gcc
fi

Another use of ‘config.site’ is for priming the directory variables in a manner consistent with the Filesystem Hierarchy Standard (FHS). Once the following file is installed

at ‘/usr/share/config.site’, a user can execute simply ./configure --prefix=/usr to

get all the directories chosen in the locations recommended by FHS.

# /usr/share/config.site for FHS defaults when installing below /usr,
# and the respective settings were not changed on the command line.
if test "$prefix" = /usr; then
  test "$sysconfdir" = '${prefix}/etc' && sysconfdir=/etc
  test "$sharedstatedir" = '${prefix}/com' && sharedstatedir=/var
  test "$localstatedir" = '${prefix}/var' && localstatedir=/var
fi

Likewise, on platforms where 64-bit libraries are built by default, then installed in

‘/usr/local/lib64’ instead of ‘/usr/local/lib’, it is appropriate to install ‘/usr/local/
share/config.site’:
# /usr/local/share/config.site for platforms that prefer
# the directory /usr/local/lib64 over /usr/local/lib.
test "$libdir" = '${exec_prefix}/lib' && libdir='${exec_prefix}/lib64'

16 Running configure Scripts

Below are instructions on how to configure a package that uses a configure script, suitable for inclusion as an ‘INSTALL’ file in the package. A plain-text version of ‘INSTALL’ which you may use comes with Autoconf.

16.1 Basic Installation

Briefly, the shell commands ‘./configure; make; make install’ should configure, build, and install this package. The following more-detailed instructions are generic; see the ‘README’ file for instructions specific to this package. More recommendations for GNU packages can be found in Section “Makefile Conventions” in GNU Coding Standards.

The configure shell script attempts to guess correct values for various system-dependent variables used during compilation. It uses those values to create a ‘Makefile’ in each directory of the package. It may also create one or more ‘.h’ files containing systemdependent definitions. Finally, it creates a shell script ‘config.status’ that you can run in the future to recreate the current configuration, and a file ‘config.log’ containing compiler output (useful mainly for debugging configure).

It can also use an optional file (typically called ‘config.cache’ and enabled with ‘--cache-file=config.cache’ or simply ‘-C’) that saves the results of its tests to speed up reconfiguring. Caching is disabled by default to prevent problems with accidental use of stale cache files.

If you need to do unusual things to compile the package, please try to figure out how configure could check whether to do them, and mail diffs or instructions to the address given in the ‘README’ so they can be considered for the next release. If you are using the cache, and at some point ‘config.cache’ contains results you don’t want to keep, you may remove or edit it.

The file ‘configure.ac’ (or ‘configure.in’) is used to create ‘configure’ by a program called autoconf. You need ‘configure.ac’ if you want to change it or regenerate ‘configure’ using a newer version of autoconf.

The simplest way to compile this package is:

1. cd to the directory containing the package’s source code and type ‘./configure’ to configure the package for your system. Running configure might take a while. While running, it prints some messages telling which features it is checking for.

2. Type ‘make’ to compile the package.

3. Optionally, type ‘make check’ to run any self-tests that come with the package, generally using the just-built uninstalled binaries.

4. Type ‘make install’ to install the programs and any data files and documentation. When installing into a prefix owned by root, it is recommended that the package be configured and built as a regular user, and only the ‘make install’ phase executed with root privileges.

5. Optionally, type ‘make installcheck’ to repeat any self-tests, but this time using the binaries in their final installed location. This target does not install anything.

Running this target as a regular user, particularly if the prior ‘make install’ required root privileges, verifies that the installation completed correctly.

6. You can remove the program binaries and object files from the source code directory by typing ‘make clean’. To also remove the files that configure created (so you can compile the package for a different kind of computer), type ‘make distclean’. There is also a ‘make maintainer-clean’ target, but that is intended mainly for the package’s developers. If you use it, you may have to get all sorts of other programs in order to regenerate files that came with the distribution.

7. Often, you can also type ‘make uninstall’ to remove the installed files again. In practice, not all packages have tested that uninstallation works correctly, even though it is required by the GNU Coding Standards.

8. Some packages, particularly those that use Automake, provide ‘make distcheck’, which can by used by developers to test that all other targets like ‘make install’ and ‘make uninstall’ work correctly. This target is generally not run by end users.

16.2 Compilers and Options

Some systems require unusual options for compilation or linking that the configure script does not know about. Run ‘./configure --help’ for details on some of the pertinent environment variables.

You can give configure initial values for configuration parameters by setting variables in the command line or in the environment. Here is an example:

     ./configure CC=c99 CFLAGS=-g LIBS=-lposix

See Section 16.9 [Defining Variables], page 285, for more details.

16.3 Compiling For Multiple Architectures

You can compile the package for more than one kind of computer at the same time, by placing the object files for each architecture in their own directory. To do this, you can use GNU make. cd to the directory where you want the object files and executables to go and run the configure script. configure automatically checks for the source code in the directory that configure is in and in ‘..’. This is known as a VPATH build.

With a non-GNU make, it is safer to compile the package for one architecture at a time in the source code directory. After you have installed the package for one architecture, use ‘make distclean’ before reconfiguring for another architecture.

On MacOS X 10.5 and later systems, you can create libraries and executables that work on multiple system types—known as fat or universal binaries—by specifying multiple ‘-arch’ options to the compiler but only a single ‘-arch’ option to the preprocessor. Like this:

     ./configure CC="gcc -arch i386 -arch x86_64 -arch ppc -arch ppc64" \
                 CXX="g++ -arch i386 -arch x86_64 -arch ppc -arch ppc64" \
                 CPP="gcc -E" CXXCPP="g++ -E"

This is not guaranteed to produce working output in all cases, you may have to build one architecture at a time and combine the results using the lipo tool if you have problems.

16.4 Installation Names

By default, ‘make install’ installs the package’s commands under ‘/usr/local/bin’, include files under ‘/usr/local/include’, etc. You can specify an installation prefix other than ‘/usr/local’ by giving configure the option ‘--prefix=prefix’, where prefix must be an absolute file name.

You can specify separate installation prefixes for architecture-specific files and architecture-independent files. If you pass the option ‘--exec-prefix=prefix’ to configure, the package uses prefix as the prefix for installing programs and libraries. Documentation and other data files still use the regular prefix.

In addition, if you use an unusual directory layout you can give options like ‘--bindir=dir’ to specify different values for particular kinds of files. Run ‘configure --help’ for a list of the directories you can set and what kinds of files go in them. In general, the default for these options is expressed in terms of ‘${prefix}’, so that specifying just ‘--prefix’ will affect all of the other directory specifications that were not explicitly provided.

The most portable way to affect installation locations is to pass the correct locations to configure; however, many packages provide one or both of the following shortcuts of passing variable assignments to the ‘make install’ command line to change installation locations without having to reconfigure or recompile.

The first method involves providing an override variable for each affected directory. For example, ‘make install prefix=/alternate/directory’ will choose an alternate location for all directory configuration variables that were expressed in terms of ‘${prefix}’. Any directories that were specified during configure, but not in terms of ‘${prefix}’, must each be overridden at install time for the entire installation to be relocated. The approach of makefile variable overrides for each directory variable is required by the GNU Coding Standards, and ideally causes no recompilation. However, some platforms have known limitations with the semantics of shared libraries that end up requiring recompilation when using this method, particularly noticeable in packages that use GNU Libtool.

The second method involves providing the ‘DESTDIR’ variable. For example, ‘make

install DESTDIR=/alternate/directory’ will prepend ‘/alternate/directory’ before

all installation names. The approach of ‘DESTDIR’ overrides is not required by the GNU Coding Standards, and does not work on platforms that have drive letters. On the other hand, it does better at avoiding recompilation issues, and works well even when some directory options were not specified in terms of ‘${prefix}’ at configure time.

16.5 Optional Features

If the package supports it, you can cause programs to be installed with an extra prefix or suffix on their names by giving configure the option ‘--program-prefix=PREFIX’ or

‘--program-suffix=SUFFIX’.

Some packages pay attention to ‘--enable-feature’ options to configure, where feature indicates an optional part of the package. They may also pay attention to ‘--with-package’ options, where package is something like ‘gnu-as’ or ‘x’ (for the X Window System). The ‘README’ should mention any ‘--enable-’ and ‘--with-’ options that the package recognizes.

For packages that use the X Window System, configure can usually find the X include and library files automatically, but if it doesn’t, you can use the configure options

‘--x-includes=dir’ and ‘--x-libraries=dir’ to specify their locations.

Some packages offer the ability to configure how verbose the execution of make will be. For these packages, running ‘./configure --enable-silent-rules’ sets the default to minimal output, which can be overridden with make V=1; while running ‘./configure --disable-silent-rules’ sets the default to verbose, which can be overridden with make

V=0.

16.6 Particular systems

On HP-UX, the default C compiler is not ANSI C compatible. If GNU CC is not installed, it is recommended to use the following options in order to use an ANSI C compiler:

     ./configure CC="cc -Ae -D_XOPEN_SOURCE=500"

and if that doesn’t work, install pre-built binaries of GCC for HP-UX.

On OSF/1 a.k.a. Tru64, some versions of the default C compiler cannot parse its <wchar.h> header file. The option ‘-nodtk’ can be used as a workaround. If GNU CC is not installed, it is therefore recommended to try

     ./configure CC="cc"

and if that doesn’t work, try

     ./configure CC="cc -nodtk"

On Solaris, don’t put /usr/ucb early in your PATH. This directory contains several dysfunctional programs; working variants of these programs are available in /usr/bin. So, if you need /usr/ucb in your PATH, put it after /usr/bin.

On Haiku, software installed for all users goes in ‘/boot/common’, not ‘/usr/local’. It is recommended to use the following options:

     ./configure --prefix=/boot/common

16.7 Specifying the System Type

There may be some features configure cannot figure out automatically, but needs to determine by the type of machine the package will run on. Usually, assuming the package is built to be run on the same architectures, configure can figure that out, but if it prints a message saying it cannot guess the machine type, give it the ‘--build=type’ option. type can either be a short name for the system type, such as ‘sun4’, or a canonical name which has the form:

     cpu-company-system

where system can have one of these forms:

     os
     kernel-os

See the file ‘config.sub’ for the possible values of each field. If ‘config.sub’ isn’t included in this package, then this package doesn’t need to know the machine type.

If you are building compiler tools for cross-compiling, you should use the option ‘--target=type’ to select the type of system they will produce code for.

If you want to use a cross compiler, that generates code for a platform different from the build platform, you should specify the host platform (i.e., that on which the generated programs will eventually be run) with ‘--host=type’.

16.8 Sharing Defaults

If you want to set default values for configure scripts to share, you can create a site shell script called ‘config.site’ that gives default values for variables like CC, cache_

file, and prefix. configure looks for ‘prefix/share/config.site’ if it exists, then

‘prefix/etc/config.site’ if it exists. Or, you can set the CONFIG_SITE environment variable to the location of the site script. A warning: not all configure scripts look for a site script.

16.9 Defining Variables

Variables not defined in a site shell script can be set in the environment passed to configure. However, some packages may run configure again during the build, and the customized values of these variables may be lost. In order to avoid this problem, you should set them in the configure command line, using ‘VAR=value’. For example:

     ./configure CC=/usr/local2/bin/gcc

causes the specified gcc to be used as the C compiler (unless it is overridden in the site shell script).

Unfortunately, this technique does not work for CONFIG_SHELL due to an Autoconf bug. Until the bug is fixed you can use this workaround:

     CONFIG_SHELL=/bin/bash /bin/bash ./configure CONFIG_SHELL=/bin/bash

16.10 configure Invocation

configure recognizes the following options to control how it operates.

‘--help’

‘-h’ Print a summary of all of the options to configure, and exit.

‘--help=short’
‘--help=recursive’

Print a summary of the options unique to this package’s configure, and exit. The short variant lists options used only in the top level, while the recursive variant lists options also present in any nested packages.

‘--version’

‘-V’ Print the version of Autoconf used to generate the configure script, and exit.

‘--cache-file=file’

Enable the cache: use and save the results of the tests in file, traditionally ‘config.cache’. file defaults to ‘/dev/null’ to disable caching.

‘--config-cache’
‘-C’  Alias for ‘--cache-file=config.cache’.

‘--quiet’
‘--silent’

‘-q’ Do not print messages saying which checks are being made. To suppress all normal output, redirect it to ‘/dev/null’ (any error messages will still be shown).

‘--srcdir=dir’

Look for the package’s source code in directory dir. Usually configure can determine that directory automatically.

‘--prefix=dir’

Use dir as the installation prefix. Section 16.4 [Installation Names], page 283 for more details, including other options available for fine-tuning the installation locations.

‘--no-create’

‘-n’ Run the configure checks, but stop before creating any output files.

configure also accepts some other, not widely useful, options. Run ‘configure --help’ for more details.

17 config.status Invocation

The configure script creates a file named ‘config.status’, which actually configures, instantiates, the template files. It also records the configuration options that were specified when the package was last configured in case reconfiguring is needed.

Synopsis:

     ./config.status [option]... [tag]...

It configures each tag; if none are specified, all the templates are instantiated. A tag refers to a file or other tag associated with a configuration action, as specified by an AC_ CONFIG_ITEMS macro (see Section 4.6 [Configuration Actions], page 23). The files must be specified without their dependencies, as in

     ./config.status foobar

not

     ./config.status foobar:foo.in:bar.in

The supported options are:

‘--help’

‘-h’ Print a summary of the command line options, the list of the template files, and exit.

‘--version’

‘-V’ Print the version number of Autoconf and the configuration settings, and exit.

‘--config’

Print the configuration settings in reusable way, quoted for the shell, and exit. For example, for a debugging build that otherwise reuses the configuration from a different build directory build-dir of a package in src-dir, you could use the following:

               args=`build-dir/config.status --config`
               eval src-dir/configure "$args" CFLAGS=-g --srcdir=src-dir

Note that it may be necessary to override a ‘--srcdir’ setting that was saved in the configuration, if the arguments are used in a different build directory.

‘--silent’
‘--quiet’

‘-q’ Do not print progress messages.

‘--debug’

‘-d’ Don’t remove the temporary files.

‘--file=file[:template]’

Require that file be instantiated as if ‘AC_CONFIG_FILES(file:template)’ was used. Both file and template may be ‘-’ in which case the standard output and/or standard input, respectively, is used. If a template file name is relative, it is first looked for in the build tree, and then in the source tree. See Section 4.6 [Configuration Actions], page 23, for more details. This option and the following ones provide one way for separately distributed packages to share the values computed by configure. Doing so can be useful if some of the packages need a superset of the features that one of them, perhaps a common library, does. These options allow a ‘config.status’ file to create files other than the ones that its ‘configure.ac’ specifies, so it can be used for a different package, or for extracting a subset of values. For example,

     echo '@CC@' | ./config.status --file=-

provides the value of @CC@ on standard output.

‘--header=file[:template]’

Same as ‘--file’ above, but with ‘AC_CONFIG_HEADERS’.

‘--recheck’

Ask ‘config.status’ to update itself and exit (no instantiation). This option is useful if you change configure, so that the results of some tests might be different from the previous run. The ‘--recheck’ option reruns configure with the same arguments you used before, plus the ‘--no-create’ option, which prevents configure from running ‘config.status’ and creating ‘Makefile’ and other files, and the ‘--no-recursion’ option, which prevents configure from running other configure scripts in subdirectories. (This is so other Make rules can run ‘config.status’ when it changes; see Section 4.8.5 [Automatic Remaking], page 34, for an example).

‘config.status’ checks several optional environment variables that can alter its behavior:

[Variable] CONFIG_SHELL

The shell with which to run configure for the ‘--recheck’ option. It must be Bournecompatible. The default is a shell that supports LINENO if available, and ‘/bin/sh’ otherwise. Invoking configure by hand bypasses this setting, so you may need to use

a command like ‘CONFIG_SHELL=/bin/bash /bin/bash ./configure’ to insure that

the same shell is used everywhere. The absolute name of the shell should be passed.

[Variable] CONFIG_STATUS

The file name to use for the shell script that records the configuration. The default is ‘./config.status’. This variable is useful when one package uses parts of another and the configure scripts shouldn’t be merged because they are maintained separately.

You can use ‘./config.status’ in your makefiles. For example, in the dependencies given above (see Section 4.8.5 [Automatic Remaking], page 34), ‘config.status’ is run twice when ‘configure.ac’ has changed. If that bothers you, you can make each run only regenerate the files for that rule:

config.h: stamp-h
stamp-h: config.h.in config.status
   ./config.status config.h
   echo > stamp-h
Makefile: Makefile.in config.status
   ./config.status Makefile

Chapter 17: config.status Invocation 289

The calling convention of ‘config.status’ has changed; see Section 18.1 [Obsolete config.status Use], page 291, for details.

18 Obsolete Constructs

Autoconf changes, and throughout the years some constructs have been obsoleted. Most of the changes involve the macros, but in some cases the tools themselves, or even some concepts, are now considered obsolete.

You may completely skip this chapter if you are new to Autoconf. Its intention is mainly to help maintainers updating their packages by understanding how to move to more modern constructs.

18.1 Obsolete ‘config.status’ Invocation

‘config.status’ now supports arguments to specify the files to instantiate; see Chapter 17 [config.status Invocation], page 287, for more details. Before, environment variables had to be used.

[Variable] CONFIG_COMMANDS

The tags of the commands to execute. The default is the arguments given to AC_

     OUTPUT and AC_CONFIG_COMMANDS in ‘configure.ac’.
[Variable] CONFIG_FILES

The files in which to perform ‘@variable@’ substitutions. The default is the argu-

     ments given to AC_OUTPUT and AC_CONFIG_FILES in ‘configure.ac’.
[Variable] CONFIG_HEADERS

The files in which to substitute C #define statements. The default is the arguments given to AC_CONFIG_HEADERS; if that macro was not called, ‘config.status’ ignores this variable.

[Variable] CONFIG_LINKS

The symbolic links to establish. The default is the arguments given to AC_CONFIG_ LINKS; if that macro was not called, ‘config.status’ ignores this variable.

In Chapter 17 [config.status Invocation], page 287, using this old interface, the example would be:

     config.h: stamp-h
     stamp-h: config.h.in config.status
             CONFIG_COMMANDS= CONFIG_LINKS= CONFIG_FILES= \
               CONFIG_HEADERS=config.h ./config.status
             echo > stamp-h
     Makefile: Makefile.in config.status
             CONFIG_COMMANDS= CONFIG_LINKS= CONFIG_HEADERS= \
               CONFIG_FILES=Makefile ./config.status
(If ‘configure.ac’ does not call AC_CONFIG_HEADERS, there is no need to set CONFIG_

HEADERS in the make rules. Equally for CONFIG_COMMANDS, etc.)

18.2 ‘acconfig.h’

In order to produce ‘config.h.in’, autoheader needs to build or to find templates for each symbol. Modern releases of Autoconf use AH_VERBATIM and AH_TEMPLATE (see Section 4.9.3 [Autoheader Macros], page 38), but in older releases a file, ‘acconfig.h’, contained the list of needed templates. autoheader copied comments and #define and #undef statements from ‘acconfig.h’ in the current directory, if present. This file used to be mandatory if you AC_DEFINE any additional symbols.

Modern releases of Autoconf also provide AH_TOP and AH_BOTTOM if you need to prepend/append some information to ‘config.h.in’. Ancient versions of Autoconf had a similar feature: if ‘./acconfig.h’ contains the string ‘@TOP@’, autoheader copies the lines before the line containing ‘@TOP@’ into the top of the file that it generates. Similarly, if ‘./acconfig.h’ contains the string ‘@BOTTOM@’, autoheader copies the lines after that line to the end of the file it generates. Either or both of those strings may be omitted. An even older alternate way to produce the same effect in ancient versions of Autoconf is to create the files ‘file.top’ (typically ‘config.h.top’) and/or ‘file.bot’ in the current directory. If they exist, autoheader copies them to the beginning and end, respectively, of its output.

In former versions of Autoconf, the files used in preparing a software package for distribution were:

     configure.ac --.   .------> autoconf* -----> configure
                    +---+
     [aclocal.m4] --+   `---.
     [acsite.m4] ---'       |
                            +--> [autoheader*] -> [config.h.in]
     [acconfig.h] ----.     |
                      +-----'
     [config.h.top] --+
     [config.h.bot] --'

Using only the AH_macros, ‘configure.ac’ should be self-contained, and should not depend upon ‘acconfig.h’ etc.

18.3 Using autoupdate to Modernize ‘configure.ac’

The autoupdate program updates a ‘configure.ac’ file that calls Autoconf macros by their old names to use the current macro names. In version 2 of Autoconf, most of the macros were renamed to use a more uniform and descriptive naming scheme. See Section 10.2 [Macro Names], page 176, for a description of the new scheme. Although the old names still work (see Section 18.4 [Obsolete Macros], page 293, for a list of the old macros and the corresponding new names), you can make your ‘configure.ac’ files more readable and make it easier to use the current Autoconf documentation if you update them to use the new macro names.

If given no arguments, autoupdate updates ‘configure.ac’, backing up the original version with the suffix ‘~’ (or the value of the environment variable SIMPLE_BACKUP_SUFFIX, if that is set). If you give autoupdate an argument, it reads that file instead of ‘configure.ac’ and writes the updated file to the standard output.

autoupdate accepts the following options:

‘--help’

‘-h’ Print a summary of the command line options and exit.

‘--version’

‘-V’ Print the version number of Autoconf and exit.

‘--verbose’

‘-v’ Report processing steps.

‘--debug’

‘-d’ Don’t remove the temporary files.

‘--force’

‘-f’ Force the update even if the file has not changed. Disregard the cache.

‘--include=dir’

‘-I dir’ Also look for input files in dir. Multiple invocations accumulate. Directories are browsed from last to first.

‘--prepend-include=dir’

‘-B dir’ Prepend directory dir to the search path. This is used to include the languagespecific files before any third-party macros.

18.4 Obsolete Macros

Several macros are obsoleted in Autoconf, for various reasons (typically they failed to quote properly, couldn’t be extended for more recent issues, etc.). They are still supported, but deprecated: their use should be avoided.

During the jump from Autoconf version 1 to version 2, most of the macros were renamed to use a more uniform and descriptive naming scheme, but their signature did not change. See Section 10.2 [Macro Names], page 176, for a description of the new naming scheme. Below, if there is just the mapping from old names to new names for these macros, the reader is invited to refer to the definition of the new macro for the signature and the description.

[Macro] AC_AIX

This macro is a platform-specific subset of AC_USE_SYSTEM_EXTENSIONS (see [AC USE SYSTEM EXTENSIONS], page 95).

[Macro] AC_ALLOCA

Replaced by AC_FUNC_ALLOCA (see [AC FUNC ALLOCA], page 55).

[Macro] AC_ARG_ARRAY

Removed because of limited usefulness.

[Macro] AC_C_CROSS

This macro is obsolete; it does nothing.

[Macro] AC_C_LONG_DOUBLE

If the C compiler supports a working long double type with more range or precision

     than the double type, define HAVE_LONG_DOUBLE.
     You should use AC_TYPE_LONG_DOUBLE or AC_TYPE_LONG_DOUBLE_WIDER instead. See

Section 5.9.1 [Particular Types], page 76.

[Macro] AC_CANONICAL_SYSTEM

Determine the system type and set output variables to the names of the canonical system types. See Section 14.2 [Canonicalizing], page 268, for details about the variables this macro sets.

The user is encouraged to use either AC_CANONICAL_BUILD, or AC_CANONICAL_HOST,
or AC_CANONICAL_TARGET, depending on the needs. Using AC_CANONICAL_TARGET is

enough to run the two other macros (see Section 14.2 [Canonicalizing], page 268).

[Macro] AC_CHAR_UNSIGNED

Replaced by AC_C_CHAR_UNSIGNED (see [AC C CHAR UNSIGNED], page 85).

[Macro] AC_CHECK_TYPE (type, default)

Autoconf, up to 2.13, used to provide this version of AC_CHECK_TYPE, deprecated because of its flaws. First, although it is a member of the CHECK clan, it does more than just checking. Secondly, missing types are defined using #define, not typedef, and this can lead to problems in the case of pointer types. This use of AC_CHECK_TYPE is obsolete and discouraged; see Section 5.9.2 [Generic Types], page 78, for the description of the current macro. If the type type is not defined, define it to be the C (or C++) builtin type default,

e.g., ‘short int’ or ‘unsigned int’.

This macro is equivalent to:

     AC_CHECK_TYPE([type], [],
       [AC_DEFINE_UNQUOTED([type], [default],
          [Define to `default'
           if <sys/types.h> does not define.])])

In order to keep backward compatibility, the two versions of AC_CHECK_TYPE are implemented, selected using these heuristics:

1. If there are three or four arguments, the modern version is used.

2. If the second argument appears to be a C or C++ type, then the obsolete version is used. This happens if the argument is a C or C++ builtin type or a C identifier ending in ‘_t’, optionally followed by one of ‘[(* ’ and then by a string of zero or more characters taken from the set ‘[]()* _a-zA-Z0-9’.

3. If the second argument is spelled with the alphabet of valid C and C++ types, the user is warned and the modern version is used.

4. Otherwise, the modern version is used.

You are encouraged either to use a valid builtin type, or to use the equivalent modern code (see above), or better yet, to use AC_CHECK_TYPES together with

     #ifndef HAVE_LOFF_T
     typedef loff_t off_t;
     #endif
[Macro] AC_CHECKING (feature-description)

Same as

     AC_MSG_NOTICE([checking feature-description...]

See [AC MSG NOTICE], page 119.

[Macro] AC_COMPILE_CHECK (echo-text, includes, function-body,
   action-if-true, [action-if-false])

This is an obsolete version of AC_TRY_COMPILE itself replaced by AC_COMPILE_IFELSE (see Section 6.4 [Running the Compiler], page 106), with the addition that it prints ‘checking for echo-text’ to the standard output first, if echo-text is non-empty. Use AC_MSG_CHECKING and AC_MSG_RESULT instead to print messages (see Section 7.5 [Printing Messages], page 118).

[Macro] AC_CONST

Replaced by AC_C_CONST (see [AC C CONST], page 84).

[Macro] AC_CROSS_CHECK

Same as AC_C_CROSS, which is obsolete too, and does nothing :-).

[Macro] AC_CYGWIN

Check for the Cygwin environment in which case the shell variable CYGWIN is set to ‘yes’. Don’t use this macro, the dignified means to check the nature of the host is using AC_CANONICAL_HOST (see Section 14.2 [Canonicalizing], page 268). As a matter of fact this macro is defined as:

     AC_REQUIRE([AC_CANONICAL_HOST])[]dnl
     case $host_os in
       *cygwin* ) CYGWIN=yes;;
              * ) CYGWIN=no;;
     esac

Beware that the variable CYGWIN has a special meaning when running Cygwin, and should not be changed. That’s yet another reason not to use this macro.

[Macro] AC_DECL_SYS_SIGLIST

Same as:

     AC_CHECK_DECLS([sys_siglist], [], [],
     [#include <signal.h>
     /* NetBSD declares sys_siglist in unistd.h.  */
     #ifdef HAVE_UNISTD_H
     # include <unistd.h>
     #endif
     ])

See [AC CHECK DECLS], page 73.

[Macro] AC_DECL_YYTEXT

Does nothing, now integrated in AC_PROG_LEX (see [AC PROG LEX], page 46).

[Macro] AC_DIR_HEADER

Like calling AC_FUNC_CLOSEDIR_VOID (see [AC FUNC CLOSEDIR VOID], page 56) and AC_HEADER_DIRENT (see [AC HEADER DIRENT], page 66), but defines a different set of C preprocessor macros to indicate which header file is found:

Header  Old Symbol New Symbol

‘dirent.h’  DIRENT       HAVE_DIRENT_H
‘sys/ndir.h’ SYSNDIR      HAVE_SYS_NDIR_H

‘sys/dir.h’ SYSDIR       HAVE_SYS_DIR_H
‘ndir.h’  NDIR         HAVE_NDIR_H

[Macro] AC_DYNIX_SEQ

If on DYNIX/ptx, add ‘-lseq’ to output variable LIBS. This macro used to be defined as

     AC_CHECK_LIB([seq], [getmntent], [LIBS="-lseq $LIBS"])

now it is just AC_FUNC_GETMNTENT (see [AC FUNC GETMNTENT], page 58).

[Macro] AC_EXEEXT

Defined the output variable EXEEXT based on the output of the compiler, which is now done automatically. Typically set to empty string if Posix and ‘.exe’ if a DOS variant.

[Macro] AC_EMXOS2

Similar to AC_CYGWIN but checks for the EMX environment on OS/2 and sets EMXOS2. Don’t use this macro, the dignified means to check the nature of the host is using AC_CANONICAL_HOST (see Section 14.2 [Canonicalizing], page 268).

[Macro] AC_ENABLE (feature, action-if-given, [action-if-not-given])

This is an obsolete version of AC_ARG_ENABLE that does not support providing a help string (see [AC ARG ENABLE], page 274).

[Macro] AC_ERROR

Replaced by AC_MSG_ERROR (see [AC MSG ERROR], page 119).

[Macro] AC_FIND_X

Replaced by AC_PATH_X (see [AC PATH X], page 94).

[Macro] AC_FIND_XTRA

Replaced by AC_PATH_XTRA (see [AC PATH XTRA], page 94).

[Macro] AC_FOREACH

Replaced by m4_foreach_w (see [m4 foreach w], page 146).

[Macro] AC_FUNC_CHECK

Replaced by AC_CHECK_FUNC (see [AC CHECK FUNC], page 62).

[Macro] AC_FUNC_SETVBUF_REVERSED

Do nothing. Formerly, this macro checked whether setvbuf takes the buffering type as its second argument and the buffer pointer as the third, instead of the other way around, and defined SETVBUF_REVERSED. However, the last systems to have the problem were those based on SVR2, which became obsolete in 1987, and the macro is no longer needed.

[Macro] AC_FUNC_WAIT3

If wait3 is found and fills in the contents of its third argument (a ‘struct rusage *’), which HP-UX does not do, define HAVE_WAIT3. These days portable programs should use waitpid, not wait3, as wait3 has been removed from Posix.

[Macro] AC_GCC_TRADITIONAL

Replaced by AC_PROG_GCC_TRADITIONAL (see [AC PROG GCC TRADITIONAL], page 86).

[Macro] AC_GETGROUPS_T

Replaced by AC_TYPE_GETGROUPS (see [AC TYPE GETGROUPS], page 76).

[Macro] AC_GETLOADAVG

Replaced by AC_FUNC_GETLOADAVG (see [AC FUNC GETLOADAVG], page 58).

[Macro] AC_GNU_SOURCE

This macro is a platform-specific subset of AC_USE_SYSTEM_EXTENSIONS (see [AC USE SYSTEM EXTENSIONS], page 95).

[Macro] AC_HAVE_FUNCS

Replaced by AC_CHECK_FUNCS (see [AC CHECK FUNCS], page 63).

[Macro] AC_HAVE_HEADERS

Replaced by AC_CHECK_HEADERS (see [AC CHECK HEADERS], page 71).

[Macro] AC_HAVE_LIBRARY (library, [action-if-found],
   [action-if-not-found], [other-libraries])

This macro is equivalent to calling AC_CHECK_LIB with a function argument of main. In addition, library can be written as any of ‘foo’, ‘-lfoo’, or ‘libfoo.a’. In all of those cases, the compiler is passed ‘-lfoo’. However, library cannot be a shell variable; it must be a literal name. See [AC CHECK LIB], page 51.

[Macro] AC_HAVE_POUNDBANG

Replaced by AC_SYS_INTERPRETER (see [AC SYS INTERPRETER], page 94).

[Macro] AC_HEADER_CHECK

Replaced by AC_CHECK_HEADER (see [AC CHECK HEADER], page 71).

[Macro] AC_HEADER_EGREP

Replaced by AC_EGREP_HEADER (see [AC EGREP HEADER], page 105).

[Macro] AC_HELP_STRING

Replaced by AS_HELP_STRING (see [AS HELP STRING], page 274).

[Macro] AC_INIT (unique-file-in-source-dir)

Formerly AC_INIT used to have a single argument, and was equivalent to:

     AC_INIT
     AC_CONFIG_SRCDIR(unique-file-in-source-dir)

See [AC INIT], page 19 and [AC CONFIG SRCDIR], page 21.

[Macro] AC_INLINE

Replaced by AC_C_INLINE (see [AC C INLINE], page 85).

[Macro] AC_INT_16_BITS
If the C type int is 16 bits wide, define INT_16_BITS. Use ‘AC_CHECK_SIZEOF(int)’

instead (see [AC CHECK SIZEOF], page 79).

[Macro] AC_IRIX_SUN

If on IRIX (Silicon Graphics Unix), add ‘-lsun’ to output LIBS. If you were using it to get getmntent, use AC_FUNC_GETMNTENT instead. If you used it for the NIS versions of the password and group functions, use ‘AC_CHECK_LIB(sun, getpwnam)’. Up to Autoconf 2.13, it used to be

     AC_CHECK_LIB([sun], [getmntent], [LIBS="-lsun $LIBS"])

now it is defined as

     AC_FUNC_GETMNTENT
     AC_CHECK_LIB([sun], [getpwnam])

See [AC FUNC GETMNTENT], page 58 and [AC CHECK LIB], page 51.

[Macro] AC_ISC_POSIX

This macro adds ‘-lcposix’ to output variable LIBS if necessary for Posix facilities. Sun dropped support for the obsolete INTERACTIVE Systems Corporation Unix on 2006-07-23. New programs need not use this macro. It is implemented as AC_SEARCH_ LIBS([strerror], [cposix]) (see [AC SEARCH LIBS], page 52).

[Macro] AC_LANG_C

Same as ‘AC_LANG([C])’ (see [AC LANG], page 99).

[Macro] AC_LANG_CPLUSPLUS

Same as ‘AC_LANG([C++])’ (see [AC LANG], page 99).

[Macro] AC_LANG_FORTRAN77

Same as ‘AC_LANG([Fortran 77])’ (see [AC LANG], page 99).

[Macro] AC_LANG_RESTORE

Select the language that is saved on the top of the stack, as set by AC_LANG_SAVE, remove it from the stack, and call AC_LANG(language). See Section 6.1 [Language Choice], page 99, for the preferred way to change languages.

[Macro] AC_LANG_SAVE

Remember the current language (as set by AC_LANG) on a stack. The current language does not change. AC_LANG_PUSH is preferred (see [AC LANG PUSH], page 100).

[Macro] AC_LINK_FILES (source . . ., dest . . .)

This is an obsolete version of AC_CONFIG_LINKS (see [AC CONFIG LINKS], page 40. An updated version of:

     AC_LINK_FILES(config/$machine.h config/$obj_format.h,
                   host.h            object.h)

is:

     AC_CONFIG_LINKS([host.h:config/$machine.h
                     object.h:config/$obj_format.h])
[Macro] AC_LN_S

Replaced by AC_PROG_LN_S (see [AC PROG LN S], page 47).

[Macro] AC_LONG_64_BITS

Define LONG_64_BITS if the C type long int is 64 bits wide. Use the generic macro ‘AC_CHECK_SIZEOF([long int])’ instead (see [AC CHECK SIZEOF], page 79).

[Macro] AC_LONG_DOUBLE

If the C compiler supports a working long double type with more range or precision

than the double type, define HAVE_LONG_DOUBLE.
You should use AC_TYPE_LONG_DOUBLE or AC_TYPE_LONG_DOUBLE_WIDER instead. See

Section 5.9.1 [Particular Types], page 76.

[Macro] AC_LONG_FILE_NAMES

Replaced by

     AC_SYS_LONG_FILE_NAMES

See [AC SYS LONG FILE NAMES], page 95.

[Macro] AC_MAJOR_HEADER

Replaced by AC_HEADER_MAJOR (see [AC HEADER MAJOR], page 67).

[Macro] AC_MEMORY_H

Used to define NEED_MEMORY_H if the mem functions were defined in ‘memory.h’. Today it is equivalent to ‘AC_CHECK_HEADERS([memory.h])’ (see [AC CHECK HEADERS], page 71). Adjust your code to depend upon HAVE_MEMORY_H, not NEED_MEMORY_H; see Section 5.1.1 [Standard Symbols], page 43.

[Macro] AC_MINGW32

Similar to AC_CYGWIN but checks for the MinGW compiler environment and sets MINGW32. Don’t use this macro, the dignified means to check the nature of the host is using AC_CANONICAL_HOST (see Section 14.2 [Canonicalizing], page 268).

[Macro] AC_MINIX

This macro is a platform-specific subset of AC_USE_SYSTEM_EXTENSIONS (see [AC USE SYSTEM EXTENSIONS], page 95).

[Macro] AC_MINUS_C_MINUS_O

Replaced by AC_PROG_CC_C_O (see [AC PROG CC C O], page 82).

[Macro] AC_MMAP

Replaced by AC_FUNC_MMAP (see [AC FUNC MMAP], page 60).

[Macro] AC_MODE_T

Replaced by AC_TYPE_MODE_T (see [AC TYPE MODE T], page 77).

[Macro] AC_OBJEXT

Defined the output variable OBJEXT based on the output of the compiler, after .c files have been excluded. Typically set to ‘o’ if Posix, ‘obj’ if a DOS variant. Now the compiler checking macros handle this automatically.

[Macro] AC_OBSOLETE (this-macro-name, [suggestion])

Make M4 print a message to the standard error output warning that this-macro-name is obsolete, and giving the file and line number where it was called. this-macro-name should be the name of the macro that is calling AC_OBSOLETE. If suggestion is given, it is printed at the end of the warning message; for example, it can be a suggestion for what to use instead of this-macro-name. For instance

     AC_OBSOLETE([$0], [; use AC_CHECK_HEADERS(unistd.h) instead])dnl

You are encouraged to use AU_DEFUN instead, since it gives better services to the user (see [AU DEFUN], page 182).

[Macro] AC_OFF_T

Replaced by AC_TYPE_OFF_T (see [AC TYPE OFF T], page 77).

[Macro] AC_OUTPUT ([file]. . ., [extra-cmds], [init-cmds])

The use of AC_OUTPUT with arguments is deprecated. This obsoleted interface is equivalent to:

     AC_CONFIG_FILES(file...)
     AC_CONFIG_COMMANDS([default],
                        extra-cmds, init-cmds)
     AC_OUTPUT

See [AC CONFIG FILES], page 25, [AC CONFIG COMMANDS], page 39, and [AC OUTPUT], page 22.

[Macro] AC_OUTPUT_COMMANDS (extra-cmds, [init-cmds])

Specify additional shell commands to run at the end of ‘config.status’, and shell commands to initialize any variables from configure. This macro may be called multiple times. It is obsolete, replaced by AC_CONFIG_COMMANDS (see [AC CONFIG COMMANDS], page 39). Here is an unrealistic example:

     fubar=27
     AC_OUTPUT_COMMANDS([echo this is extra $fubar, and so on.],
                        [fubar=$fubar])
     AC_OUTPUT_COMMANDS([echo this is another, extra, bit],
                        [echo init bit])

Aside from the fact that AC_CONFIG_COMMANDS requires an additional key, an important difference is that AC_OUTPUT_COMMANDS is quoting its arguments twice, unlike

AC_CONFIG_COMMANDS. This means that AC_CONFIG_COMMANDS can safely be given

macro calls as arguments:

     AC_CONFIG_COMMANDS(foo, [my_FOO()])

Conversely, where one level of quoting was enough for literal strings with AC_OUTPUT_ COMMANDS, you need two with AC_CONFIG_COMMANDS. The following lines are equivalent:

     AC_OUTPUT_COMMANDS([echo "Square brackets: []"])
     AC_CONFIG_COMMANDS([default], [[echo "Square brackets: []"]])
[Macro] AC_PID_T

Replaced by AC_TYPE_PID_T (see [AC TYPE PID T], page 77).

[Macro] AC_PREFIX

Replaced by AC_PREFIX_PROGRAM (see [AC PREFIX PROGRAM], page 41).

[Macro] AC_PROGRAMS_CHECK

Replaced by AC_CHECK_PROGS (see [AC CHECK PROGS], page 48).

[Macro] AC_PROGRAMS_PATH

Replaced by AC_PATH_PROGS (see [AC PATH PROGS], page 50).

[Macro] AC_PROGRAM_CHECK

Replaced by AC_CHECK_PROG (see [AC CHECK PROG], page 48).

[Macro] AC_PROGRAM_EGREP

Replaced by AC_EGREP_CPP (see [AC EGREP CPP], page 105).

[Macro] AC_PROGRAM_PATH

Replaced by AC_PATH_PROG (see [AC PATH PROG], page 50).

[Macro] AC_REMOTE_TAPE

Removed because of limited usefulness.

[Macro] AC_RESTARTABLE_SYSCALLS

This macro was renamed AC_SYS_RESTARTABLE_SYSCALLS. However, these days portable programs should use sigaction with SA_RESTART if they want restartable system calls. They should not rely on HAVE_RESTARTABLE_SYSCALLS, since nowadays whether a system call is restartable is a dynamic issue, not a configuration-time issue.

[Macro] AC_RETSIGTYPE

Replaced by AC_TYPE_SIGNAL (see [AC TYPE SIGNAL], page 304), which itself is obsolete when assuming C89 or better.

[Macro] AC_RSH

Removed because of limited usefulness.

[Macro] AC_SCO_INTL

If on SCO Unix, add ‘-lintl’ to output variable LIBS. This macro used to do this:

     AC_CHECK_LIB([intl], [strftime], [LIBS="-lintl $LIBS"])

Now it just calls AC_FUNC_STRFTIME instead (see [AC FUNC STRFTIME], page 61).

[Macro] AC_SETVBUF_REVERSED

Replaced by

     AC_FUNC_SETVBUF_REVERSED

See [AC FUNC SETVBUF REVERSED], page 296.

[Macro] AC_SET_MAKE

Replaced by AC_PROG_MAKE_SET (see [AC PROG MAKE SET], page 22).

[Macro] AC_SIZEOF_TYPE

Replaced by AC_CHECK_SIZEOF (see [AC CHECK SIZEOF], page 79).

[Macro] AC_SIZE_T

Replaced by AC_TYPE_SIZE_T (see [AC TYPE SIZE T], page 77).

[Macro] AC_STAT_MACROS_BROKEN

Replaced by AC_HEADER_STAT (see [AC HEADER STAT], page 68).

[Macro] AC_STDC_HEADERS

Replaced by AC_HEADER_STDC (see [AC HEADER STDC], page 68).

[Macro] AC_STRCOLL

Replaced by AC_FUNC_STRCOLL (see [AC FUNC STRCOLL], page 61).

[Macro] AC_STRUCT_ST_BLKSIZE
If struct stat contains an st_blksize member, define HAVE_STRUCT_STAT_ST_

BLKSIZE. The former name, HAVE_ST_BLKSIZE is to be avoided, as its support will cease in the future. This macro is obsoleted, and should be replaced by

     AC_CHECK_MEMBERS([struct stat.st_blksize])

See [AC CHECK MEMBERS], page 75.

[Macro] AC_STRUCT_ST_RDEV
If struct stat contains an st_rdev member, define HAVE_STRUCT_STAT_ST_RDEV.

The former name for this macro, HAVE_ST_RDEV, is to be avoided as it will cease to be supported in the future. Actually, even the new macro is obsolete and should be replaced by:

     AC_CHECK_MEMBERS([struct stat.st_rdev])

See [AC CHECK MEMBERS], page 75.

[Macro] AC_ST_BLKSIZE

Replaced by AC_CHECK_MEMBERS (see [AC CHECK MEMBERS], page 75).

[Macro] AC_ST_BLOCKS

Replaced by AC_STRUCT_ST_BLOCKS (see [AC STRUCT ST BLOCKS], page 74).

[Macro] AC_ST_RDEV

Replaced by AC_CHECK_MEMBERS (see [AC CHECK MEMBERS], page 75).

[Macro] AC_SYS_RESTARTABLE_SYSCALLS

If the system automatically restarts a system call that is interrupted by a signal, define HAVE_RESTARTABLE_SYSCALLS. This macro does not check whether system calls are restarted in general—it checks whether a signal handler installed with signal (but not sigaction) causes system calls to be restarted. It does not check whether system calls can be restarted when interrupted by signals that have no handler. These days portable programs should use sigaction with SA_RESTART if they want restartable system calls. They should not rely on HAVE_RESTARTABLE_SYSCALLS, since nowadays whether a system call is restartable is a dynamic issue, not a configurationtime issue.

[Macro] AC_SYS_SIGLIST_DECLARED

This macro was renamed AC_DECL_SYS_SIGLIST. However, even that name is obsolete, as the same functionality is now acheived via AC_CHECK_DECLS (see [AC CHECK DECLS], page 73).

[Macro] AC_TEST_CPP

This macro was renamed AC_TRY_CPP, which in turn was replaced by AC_PREPROC_ IFELSE (see [AC PREPROC IFELSE], page 104).

[Macro] AC_TEST_PROGRAM

This macro was renamed AC_TRY_RUN, which in turn was replaced by AC_RUN_IFELSE (see [AC RUN IFELSE], page 107).

[Macro] AC_TIMEZONE

Replaced by AC_STRUCT_TIMEZONE (see [AC STRUCT TIMEZONE], page 75).

[Macro] AC_TIME_WITH_SYS_TIME

Replaced by AC_HEADER_TIME (see [AC HEADER TIME], page 70).

[Macro] AC_TRY_COMPILE (includes, function-body, [action-if-true],
   [action-if-false])

Same as:

     AC_COMPILE_IFELSE(
       [AC_LANG_PROGRAM([[includes]],
          [[function-body]])],
       [action-if-true],
       [action-if-false])

See Section 6.4 [Running the Compiler], page 106. This macro double quotes both includes and function-body. For C and C++, includes is any #include statements needed by the code in functionbody (includes is ignored if the currently selected language is Fortran or Fortran 77). The compiler and compilation flags are determined by the current language (see Section 6.1 [Language Choice], page 99).

[Macro] AC_TRY_CPP (input, [action-if-true], [action-if-false])

Same as:

     AC_PREPROC_IFELSE(
       [AC_LANG_SOURCE([[input]])],
       [action-if-true],
       [action-if-false])

See Section 6.3 [Running the Preprocessor], page 104. This macro double quotes the input.

[Macro] AC_TRY_LINK (includes, function-body, [action-if-true],
   [action-if-false])

Same as:

     AC_LINK_IFELSE(
       [AC_LANG_PROGRAM([[includes]],
          [[function-body]])],
       [action-if-true],
       [action-if-false])

See Section 6.4 [Running the Compiler], page 106.

This macro double quotes both includes and function-body. Depending on the current language (see Section 6.1 [Language Choice], page 99), create a test program to see whether a function whose body consists of functionbody can be compiled and linked. If the file compiles and links successfully, run shell commands action-if-found, otherwise run action-if-not-found. This macro double quotes both includes and function-body. For C and C++, includes is any #include statements needed by the code in functionbody (includes is ignored if the currently selected language is Fortran or Fortran 77). The compiler and compilation flags are determined by the current language (see Section 6.1 [Language Choice], page 99), and in addition LDFLAGS and LIBS are used for linking.

[Macro] AC_TRY_LINK_FUNC (function, [action-if-found],
   [action-if-not-found])

This macro is equivalent to

     AC_LINK_IFELSE([AC_LANG_CALL([], [function])],
       [action-if-found], [action-if-not-found])

See [AC LINK IFELSE], page 106.

[Macro] AC_TRY_RUN (program, [action-if-true], [action-if-false],
   [action-if-cross-compiling])

Same as:

     AC_RUN_IFELSE(
       [AC_LANG_SOURCE([[program]])],
       [action-if-true],
       [action-if-false],
       [action-if-cross-compiling])

See Section 6.6 [Runtime], page 107.

[Macro] AC_TYPE_SIGNAL

If ‘signal.h’ declares signal as returning a pointer to a function returning void, define RETSIGTYPE to be void; otherwise, define it to be int. These days, it is portable to assume C89, and that signal handlers return void, without needing to use this macro or RETSIGTYPE. When targetting older K&R C, it is possible to define signal handlers as returning type RETSIGTYPE, and omit a return statement:

     RETSIGTYPE
     hup_handler ()
     {
     ...
     }
[Macro] AC_UID_T

Replaced by AC_TYPE_UID_T (see [AC TYPE UID T], page 77).

[Macro] AC_UNISTD_H

Same as ‘AC_CHECK_HEADERS([unistd.h])’ (see [AC CHECK HEADERS], page 71).

[Macro] AC_USG

Define USG if the BSD string functions are defined in ‘strings.h’. You should no longer depend upon USG, but on HAVE_STRING_H; see Section 5.1.1 [Standard Symbols], page 43.

[Macro] AC_UTIME_NULL

Replaced by AC_FUNC_UTIME_NULL (see [AC FUNC UTIME NULL], page 62).

[Macro] AC_VALIDATE_CACHED_SYSTEM_TUPLE ([cmd])

If the cache file is inconsistent with the current host, target and build system types, it used to execute cmd or print a default error message. This is now handled by default.

[Macro] AC_VERBOSE (result-description)

Replaced by AC_MSG_RESULT (see [AC MSG RESULT], page 118).

[Macro] AC_VFORK

Replaced by AC_FUNC_FORK (see [AC FUNC FORK], page 57).

[Macro] AC_VPRINTF

Replaced by AC_FUNC_VPRINTF (see [AC FUNC VPRINTF], page 62).

[Macro] AC_WAIT3

This macro was renamed AC_FUNC_WAIT3. However, these days portable programs should use waitpid, not wait3, as wait3 has been removed from Posix.

[Macro] AC_WARN

Replaced by AC_MSG_WARN (see [AC MSG WARN], page 119).

[Macro] AC_WITH (package, action-if-given, [action-if-not-given])

This is an obsolete version of AC_ARG_WITH that does not support providing a help string (see [AC ARG WITH], page 272).

[Macro] AC_WORDS_BIGENDIAN

Replaced by AC_C_BIGENDIAN (see [AC C BIGENDIAN], page 84).

[Macro] AC_XENIX_DIR

This macro used to add ‘-lx’ to output variable LIBS if on Xenix. Also, if ‘dirent.h’ is being checked for, added ‘-ldir’ to LIBS. Now it is merely an alias of AC_HEADER_ DIRENT instead, plus some code to detect whether running XENIX on which you should not depend:

     AC_MSG_CHECKING([for Xenix])
     AC_EGREP_CPP([yes],
     [#if defined M_XENIX && !defined M_UNIX
       yes
     #endif],
                  [AC_MSG_RESULT([yes]); XENIX=yes],
                  [AC_MSG_RESULT([no]); XENIX=])

Don’t use this macro, the dignified means to check the nature of the host is using AC_CANONICAL_HOST (see Section 14.2 [Canonicalizing], page 268).

[Macro] AC_YYTEXT_POINTER

This macro was renamed AC_DECL_YYTEXT, which in turn was integrated into AC_ PROG_LEX (see [AC PROG LEX], page 46).

18.5 Upgrading From Version 1

Autoconf version 2 is mostly backward compatible with version 1. However, it introduces better ways to do some things, and doesn’t support some of the ugly things in version 1. So, depending on how sophisticated your ‘configure.ac’ files are, you might have to do some manual work in order to upgrade to version 2. This chapter points out some problems to watch for when upgrading. Also, perhaps your configure scripts could benefit from some of the new features in version 2; the changes are summarized in the file ‘NEWS’ in the Autoconf distribution.

18.5.1 Changed File Names

If you have an ‘aclocal.m4’ installed with Autoconf (as opposed to in a particular package’s source directory), you must rename it to ‘acsite.m4’. See Section 3.4 [autoconf Invocation], page 13.

If you distribute ‘install.sh’ with your package, rename it to ‘install-sh’ so make builtin rules don’t inadvertently create a file called ‘install’ from it. AC_PROG_INSTALL looks for the script under both names, but it is best to use the new name.

If you were using ‘config.h.top’, ‘config.h.bot’, or ‘acconfig.h’, you still can, but you have less clutter if you use the AH_macros. See Section 4.9.3 [Autoheader Macros], page 38.

18.5.2 Changed Makefiles

Add ‘@CFLAGS@’, ‘@CPPFLAGS@’, and ‘@LDFLAGS@’ in your ‘Makefile.in’ files, so they can

take advantage of the values of those variables in the environment when configure is run. Doing this isn’t necessary, but it’s a convenience for users.

Also add ‘@configure_input@’ in a comment to each input file for AC_OUTPUT, so that the output files contain a comment saying they were produced by configure. Automatically selecting the right comment syntax for all the kinds of files that people call AC_OUTPUT on became too much work.

Add ‘config.log’ and ‘config.cache’ to the list of files you remove in distclean targets.

If you have the following in ‘Makefile.in’:

     prefix = /usr/local
     exec_prefix = $(prefix)

you must change it to:

     prefix = @prefix@
     exec_prefix = @exec_prefix@

The old behavior of replacing those variables without ‘@’ characters around them has been removed.

18.5.3 Changed Macros

Many of the macros were renamed in Autoconf version 2. You can still use the old names, but the new ones are clearer, and it’s easier to find the documentation for them. See Section 18.4 [Obsolete Macros], page 293, for a table showing the new names for the old macros. Use the autoupdate program to convert your ‘configure.ac’ to using the new macro names. See Section 18.3 [autoupdate Invocation], page 292.

Some macros have been superseded by similar ones that do the job better, but are not call-compatible. If you get warnings about calling obsolete macros while running autoconf, you may safely ignore them, but your configure script generally works better if you follow the advice that is printed about what to replace the obsolete macros with. In particular, the mechanism for reporting the results of tests has changed. If you were using echo or AC_ VERBOSE (perhaps via AC_COMPILE_CHECK), your configure script’s output looks better if you switch to AC_MSG_CHECKING and AC_MSG_RESULT. See Section 7.5 [Printing Messages], page 118. Those macros work best in conjunction with cache variables. See Section 7.4 [Caching Results], page 115.

18.5.4 Changed Results

If you were checking the results of previous tests by examining the shell variable DEFS, you need to switch to checking the values of the cache variables for those tests. DEFS no longer exists while configure is running; it is only created when generating output files. This difference from version 1 is because properly quoting the contents of that variable turned out to be too cumbersome and inefficient to do every time AC_DEFINE is called. See Section 7.4.1 [Cache Variable Names], page 116.

For example, here is a ‘configure.ac’ fragment written for Autoconf version 1:

     AC_HAVE_FUNCS(syslog)
     case "$DEFS" in
     *-DHAVE_SYSLOG*) ;;
     *) # syslog is not in the default libraries.  See if it's in some other.
       saved_LIBS="$LIBS"
       for lib in bsd socket inet; do
         AC_CHECKING(for syslog in -l$lib)
         LIBS="-l$lib $saved_LIBS"
         AC_HAVE_FUNCS(syslog)
         case "$DEFS" in
         *-DHAVE_SYSLOG*) break ;;
         *) ;;
         esac
         LIBS="$saved_LIBS"
       done ;;
     esac

Here is a way to write it for version 2:

     AC_CHECK_FUNCS([syslog])
     if test "x$ac_cv_func_syslog" = xno; then
       # syslog is not in the default libraries.  See if it's in some other.
       for lib in bsd socket inet; do

         AC_CHECK_LIB([$lib], [syslog], [AC_DEFINE([HAVE_SYSLOG])
           LIBS="-l$lib $LIBS"; break])
       done
     fi

If you were working around bugs in AC_DEFINE_UNQUOTED by adding backslashes before quotes, you need to remove them. It now works predictably, and does not treat quotes (except back quotes) specially. See Section 7.2 [Setting Output Variables], page 112.

All of the Boolean shell variables set by Autoconf macros now use ‘yes’ for the true value. Most of them use ‘no’ for false, though for backward compatibility some use the empty string instead. If you were relying on a shell variable being set to something like 1 or ‘t’ for true, you need to change your tests.

18.5.5 Changed Macro Writing

When defining your own macros, you should now use AC_DEFUN instead of define. AC_DEFUN automatically calls AC_PROVIDE and ensures that macros called via AC_REQUIRE do not interrupt other macros, to prevent nested ‘checking...’ messages on the screen. There’s no actual harm in continuing to use the older way, but it’s less convenient and attractive. See Section 10.1 [Macro Definitions], page 175.

You probably looked at the macros that came with Autoconf as a guide for how to do things. It would be a good idea to take a look at the new versions of them, as the style is somewhat improved and they take advantage of some new features.

If you were doing tricky things with undocumented Autoconf internals (macros, variables, diversions), check whether you need to change anything to account for changes that have been made. Perhaps you can even use an officially supported technique in version 2 instead of kludging. Or perhaps not.

To speed up your locally written feature tests, add caching to them. See whether any of your tests are of general enough usefulness to encapsulate them into macros that you can share.

18.6 Upgrading From Version 2.13

The introduction of the previous section (see Section 18.5 [Autoconf 1], page 306) perfectly suits this section…

Autoconf version 2.50 is mostly backward compatible with version 2.13. However, it introduces better ways to do some things, and doesn’t support some of the ugly things in version 2.13. So, depending on how sophisticated your ‘configure.ac’ files are, you might have to do some manual work in order to upgrade to version 2.50. This chapter points out some problems to watch for when upgrading. Also, perhaps your configure scripts could benefit from some of the new features in version 2.50; the changes are summarized in the file ‘NEWS’ in the Autoconf distribution.

18.6.1 Changed Quotation

The most important changes are invisible to you: the implementation of most macros have completely changed. This allowed more factorization of the code, better error messages, a higher uniformity of the user’s interface etc. Unfortunately, as a side effect, some construct which used to (miraculously) work might break starting with Autoconf 2.50. The most common culprit is bad quotation.

For instance, in the following example, the message is not properly quoted:

AC_INIT
AC_CHECK_HEADERS(foo.h, ,
  AC_MSG_ERROR(cannot find foo.h, bailing out))
AC_OUTPUT

Autoconf 2.13 simply ignores it:

$ autoconf-2.13; ./configure --silent
creating cache ./config.cache
configure: error: cannot find foo.h
$

while Autoconf 2.50 produces a broken ‘configure’:

$ autoconf-2.50; ./configure --silent
configure: error: cannot find foo.h
./configure: exit: bad non-numeric arg `bailing'
./configure: exit: bad non-numeric arg `bailing'
$

The message needs to be quoted, and the AC_MSG_ERROR invocation too!

AC_INIT([Example], [1.0], [bug-example@example.org])
AC_CHECK_HEADERS([foo.h], [],
  [AC_MSG_ERROR([cannot find foo.h, bailing out])])
AC_OUTPUT

Many many (and many more) Autoconf macros were lacking proper quotation, including no less than… AC_DEFUN itself!

$ cat configure.in
AC_DEFUN([AC_PROG_INSTALL],
[# My own much better version
])
AC_INIT
AC_PROG_INSTALL
AC_OUTPUT
$ autoconf-2.13
autoconf: Undefined macros:
***BUG in Autoconf--please report*** AC_FD_MSG
***BUG in Autoconf--please report*** AC_EPI
configure.in:1:AC_DEFUN([AC_PROG_INSTALL],
configure.in:5:AC_PROG_INSTALL
$ autoconf-2.50
$

18.6.2 New Macros

While Autoconf was relatively dormant in the late 1990s, Automake provided Autoconflike macros for a while. Starting with Autoconf 2.50 in 2001, Autoconf provided versions of these macros, integrated in the AC_namespace, instead of AM_. But in order to ease the upgrading via autoupdate, bindings to such AM_macros are provided.

Unfortunately older versions of Automake (e.g., Automake 1.4) did not quote the names of these macros. Therefore, when m4 finds something like ‘AC_DEFUN(AM_TYPE_PTRDIFF_T, ...)’ in ‘aclocal.m4’, AM_TYPE_PTRDIFF_T is expanded, replaced with its Autoconf definition.

Fortunately Autoconf catches pre-AC_INIT expansions, and complains, in its own words:

$ cat configure.ac
AC_INIT([Example], [1.0], [bug-example@example.org])
AM_TYPE_PTRDIFF_T
$ aclocal-1.4
$ autoconf
aclocal.m4:17: error: m4_defn: undefined macro: _m4_divert_diversion
aclocal.m4:17: the top level
autom4te: m4 failed with exit status: 1
$

Modern versions of Automake no longer define most of these macros, and properly quote the names of the remaining macros. If you must use an old Automake, do not depend upon macros from Automake as it is simply not its job to provide macros (but the one it requires itself):

$ cat configure.ac
AC_INIT([Example], [1.0], [bug-example@example.org])
AM_TYPE_PTRDIFF_T
$ rm aclocal.m4
$ autoupdate
autoupdate: `configure.ac' is updated
$ cat configure.ac
AC_INIT([Example], [1.0], [bug-example@example.org])
AC_CHECK_TYPES([ptrdiff_t])
$ aclocal-1.4
$ autoconf
$

18.6.3 Hosts and Cross-Compilation

Based on the experience of compiler writers, and after long public debates, many aspects of the cross-compilation chain have changed: − the relationship between the build, host, and target architecture types, − the command line interface for specifying them to configure, − the variables defined in configure, − the enabling of cross-compilation mode.

The relationship between build, host, and target have been cleaned up: the chain of default is now simply: target defaults to host, host to build, and build to the result of config.guess. Nevertheless, in order to ease the transition from 2.13 to 2.50, the following transition scheme is implemented. Do not rely on it, as it will be completely disabled in a couple of releases (we cannot keep it, as it proves to cause more problems than it cures).

They all default to the result of running config.guess, unless you specify either ‘--build’ or ‘--host’. In this case, the default becomes the system type you specified. If you specify both, and they’re different, configure enters cross compilation mode, so it doesn’t run any tests that require execution.

Hint: if you mean to override the result of config.guess, prefer ‘--build’ over ‘--host’. In the future, ‘--host’ will not override the name of the build system type. Whenever you specify ‘--host’, be sure to specify ‘--build’ too.

For backward compatibility, configure accepts a system type as an option by itself. Such an option overrides the defaults for build, host, and target system types. The following configure statement configures a cross toolchain that runs on NetBSD/alpha but generates code for GNU Hurd/sparc, which is also the build platform.

     ./configure --host=alpha-netbsd sparc-gnu

In Autoconf 2.13 and before, the variables build, host, and target had a different semantics before and after the invocation of AC_CANONICAL_BUILD etc. Now, the argument of ‘--build’ is strictly copied into build_alias, and is left empty otherwise. After the AC_CANONICAL_BUILD, build is set to the canonicalized build type. To ease the transition, before, its contents is the same as that of build_alias. Do not rely on this broken feature.

For consistency with the backward compatibility scheme exposed above, when ‘--host’ is specified but ‘--build’ isn’t, the build system is assumed to be the same as ‘--host’, and ‘build_alias’ is set to that value. Eventually, this historically incorrect behavior will go away.

The former scheme to enable cross-compilation proved to cause more harm than good, in particular, it used to be triggered too easily, leaving regular end users puzzled in front of cryptic error messages. configure could even enter cross-compilation mode only because the compiler was not functional. This is mainly because configure used to try to detect cross-compilation, instead of waiting for an explicit flag from the user.

Now, configure enters cross-compilation mode if and only if ‘--host’ is passed.

That’s the short documentation. To ease the transition between 2.13 and its successors, a more complicated scheme is implemented. Do not rely on the following, as it will be removed in the near future.

If you specify ‘--host’, but not ‘--build’, when configure performs the first compiler test it tries to run an executable produced by the compiler. If the execution fails, it enters cross-compilation mode. This is fragile. Moreover, by the time the compiler test is performed, it may be too late to modify the build-system type: other tests may have already been performed. Therefore, whenever you specify ‘--host’, be sure to specify ‘--build’ too.

     ./configure --build=i686-pc-linux-gnu --host=m68k-coff

enters cross-compilation mode. The former interface, which consisted in setting the compiler to a cross-compiler without informing configure is obsolete. For instance, configure fails if it can’t run the code generated by the specified compiler if you configure as follows:

     ./configure CC=m68k-coff-gcc
18.6.4 AC_LIBOBJ vs. LIBOBJS

Up to Autoconf 2.13, the replacement of functions was triggered via the variable LIBOBJS. Since Autoconf 2.50, the macro AC_LIBOBJ should be used instead (see Section 5.5.3 [Generic Functions], page 62). Starting at Autoconf 2.53, the use of LIBOBJS is an error.

This change is mandated by the unification of the GNU Build System components. In particular, the various fragile techniques used to parse a ‘configure.ac’ are all replaced with the use of traces. As a consequence, any action must be traceable, which obsoletes critical variable assignments. Fortunately, LIBOBJS was the only problem, and it can even be handled gracefully (read, “without your having to change something”).

There were two typical uses of LIBOBJS: asking for a replacement function, and adjusting LIBOBJS for Automake and/or Libtool.

As for function replacement, the fix is immediate: use AC_LIBOBJ. For instance:

     LIBOBJS="$LIBOBJS fnmatch.o"
     LIBOBJS="$LIBOBJS malloc.$ac_objext"

should be replaced with:

     AC_LIBOBJ([fnmatch])
     AC_LIBOBJ([malloc])

When used with Automake 1.10 or newer, a suitable value for LIBOBJDIR is set so that the LIBOBJS and LTLIBOBJS can be referenced from any ‘Makefile.am’. Even without Automake, arranging for LIBOBJDIR to be set correctly enables referencing LIBOBJS and LTLIBOBJS in another directory. The LIBOBJDIR feature is experimental.

18.6.5 AC_ACT_IFELSE vs. AC_TRY_ACT
Since Autoconf 2.50, internal codes uses AC_PREPROC_IFELSE, AC_COMPILE_IFELSE, AC_
LINK_IFELSE, and AC_RUN_IFELSE on one hand and AC_LANG_SOURCES, and AC_LANG_

PROGRAM on the other hand instead of the deprecated AC_TRY_CPP, AC_TRY_COMPILE, AC_ TRY_LINK, and AC_TRY_RUN. The motivations where: − a more consistent interface: AC_TRY_COMPILE etc. were double quoting their arguments; − the combinatoric explosion is solved by decomposing on the one hand the generation

of sources, and on the other hand executing the program; − this scheme helps supporting more languages than plain C and C++.

In addition to the change of syntax, the philosophy has changed too: while emphasis was put on speed at the expense of accuracy, today’s Autoconf promotes accuracy of the testing framework at, ahem…, the expense of speed.

As a perfect example of what is not to be done, here is how to find out whether a header file contains a particular declaration, such as a typedef, a structure, a structure member, or a function. Use AC_EGREP_HEADER instead of running grep directly on the header file; on some systems the symbol might be defined in another header file that the file you are checking includes.

As a (bad) example, here is how you should not check for C preprocessor symbols, either defined by header files or predefined by the C preprocessor: using AC_EGREP_CPP:

AC_EGREP_CPP(yes,
[#ifdef _AIX
  yes
#endif
], is_aix=yes, is_aix=no)

The above example, properly written would (i) use AC_LANG_PROGRAM, and (ii) run the compiler:

AC_COMPILE_IFELSE([AC_LANG_PROGRAM(
[[#ifndef _AIX
 error: This isn't AIX!
#endif
]])],
                   [is_aix=yes],
                   [is_aix=no])

19 Generating Test Suites with Autotest

N.B.: This section describes a feature which is still

stabilizing. Although we believe that Autotest is useful as-is, this documentation describes an interface which might change in the future: do not depend upon Autotest without subscribing to the Autoconf mailing lists.

It is paradoxical that portable projects depend on nonportable tools to run their test suite. Autoconf by itself is the paragon of this problem: although it aims at perfectly portability, up to 2.13 its test suite was using DejaGNU, a rich and complex testing framework, but which is far from being standard on Posix systems. Worse yet, it was likely to be missing on the most fragile platforms, the very platforms that are most likely to torture Autoconf and exhibit deficiencies.

To circumvent this problem, many package maintainers have developed their own testing framework, based on simple shell scripts whose sole outputs are exit status values describing whether the test succeeded. Most of these tests share common patterns, and this can result in lots of duplicated code and tedious maintenance.

Following exactly the same reasoning that yielded to the inception of Autoconf, Autotest provides a test suite generation framework, based on M4 macros building a portable shell script. The suite itself is equipped with automatic logging and tracing facilities which greatly diminish the interaction with bug reporters, and simple timing reports.

Autoconf itself has been using Autotest for years, and we do attest that it has considerably improved the strength of the test suite and the quality of bug reports. Other projects are known to use some generation of Autotest, such as Bison, Free Recode, Free Wdiff, GNU Tar, each of them with different needs, and this usage has validated Autotest as a general testing framework.

Nonetheless, compared to DejaGNU, Autotest is inadequate for interactive tool testing, which is probably its main limitation.

19.1 Using an Autotest Test Suite

19.1.1 testsuite Scripts

Generating testing or validation suites using Autotest is rather easy. The whole validation suite is held in a file to be processed through autom4te, itself using GNU M4 under the hood, to produce a stand-alone Bourne shell script which then gets distributed. Neither autom4te nor GNU M4 are needed at the installer’s end.

Each test of the validation suite should be part of some test group. A test group is a sequence of interwoven tests that ought to be executed together, usually because one test in the group creates data files than a later test in the same group needs to read. Complex test groups make later debugging more tedious. It is much better to keep only a few tests per test group. Ideally there is only one test per test group.

For all but the simplest packages, some file such as ‘testsuite.at’ does not fully hold all test sources, as these are often easier to maintain in separate files. Each of these separate files holds a single test group, or a sequence of test groups all addressing some common functionality in the package. In such cases, ‘testsuite.at’ merely initializes the validation suite, and sometimes does elementary health checking, before listing include statements for all other test files. The special file ‘package.m4’, containing the identification of the package, is automatically included if found.

A convenient alternative consists in moving all the global issues (local Autotest macros, elementary health checking, and AT_INIT invocation) into the file local.at, and making ‘testsuite.at’ be a simple list of m4_include of sub test suites. In such case, generating the whole test suite or pieces of it is only a matter of choosing the autom4te command line arguments.

The validation scripts that Autotest produces are by convention called testsuite. When run, testsuite executes each test group in turn, producing only one summary line per test to say if that particular test succeeded or failed. At end of all tests, summarizing counters get printed. One debugging directory is left for each test group which failed, if any: such directories are named ‘testsuite.dir/nn’, where nn is the sequence number of the test group, and they include:

• a debugging script named ‘run’ which reruns the test in debug mode (see Section 19.3 [testsuite Invocation], page 323). The automatic generation of debugging scripts has the purpose of easing the chase for bugs.

• all the files created with AT_DATA

• all the Erlang source code files created with AT_CHECK_EUNIT

• a log of the run, named ‘testsuite.log’

In the ideal situation, none of the tests fail, and consequently no debugging directory is left behind for validation.

It often happens in practice that individual tests in the validation suite need to get information coming out of the configuration process. Some of this information, common for all validation suites, is provided through the file ‘atconfig’, automatically created by AC_CONFIG_TESTDIR. For configuration informations which your testing environment specifically needs, you might prepare an optional file named ‘atlocal.in’, instantiated by AC_CONFIG_FILES. The configuration process produces ‘atconfig’ and ‘atlocal’ out of these two input files, and these two produced files are automatically read by the ‘testsuite’ script.

Here is a diagram showing the relationship between files. Files used in preparing a software package for distribution:

                     [package.m4] -->.
                                      \
     subfile-1.at ->.  [local.at] ---->+
         ...         \                  \
     subfile-i.at ---->-- testsuite.at -->-- autom4te* -->testsuite
         ...         /
     subfile-n.at ->'

Files used in configuring a software package:

                                          .--> atconfig
                                         /
     [atlocal.in] -->  config.status* --<
                                         \

                                `--> [atlocal]

Files created during test suite execution:

atconfig -->.                    .--> testsuite.log
        \                  /
         >-- testsuite* --<
        /                  \
[atlocal] ->'                    `--> [testsuite.dir]

19.1.2 Autotest Logs

When run, the test suite creates a log file named after itself, e.g., a test suite named testsuite creates ‘testsuite.log’. It contains a lot of information, usually more than maintainers actually need, but therefore most of the time it contains all that is needed:

command line arguments A bad but unfortunately widespread habit consists of setting environment variables before the command, such as in ‘CC=my-home-grown-cc ./testsuite’. The test suite does not know this change, hence (i) it cannot report it to you, and (ii) it cannot preserve the value of CC for subsequent runs. Autoconf faced exactly the same problem, and solved it by asking users to pass the variable definitions as command line arguments. Autotest requires this rule, too, but has no means to enforce it; the log then contains a trace of the variables that were changed by the user.

‘ChangeLog’ excerpts The topmost lines of all the ‘ChangeLog’ files found in the source hierarchy. This is especially useful when bugs are reported against development versions of the package, since the version string does not provide sufficient information to know the exact state of the sources the user compiled. Of course, this relies on the use of a ‘ChangeLog’.

build machine Running a test suite in a cross-compile environment is not an easy task, since it would mean having the test suite run on a machine build, while running programs on a machine host. It is much simpler to run both the test suite and the programs on host, but then, from the point of view of the test suite, there remains a single environment, host = build. The log contains relevant information on the state of the build machine, including some important environment variables.

tested programs The absolute file name and answers to ‘--version’ of the tested programs (see Section 19.2 [Writing Testsuites], page 318, AT_TESTED).

configuration log The contents of ‘config.log’, as created by configure, are appended. It contains the configuration flags and a detailed report on the configuration itself.

19.2 Writing ‘testsuite.at’

The ‘testsuite.at’ is a Bourne shell script making use of special Autotest M4 macros. It often contains a call to AT_INIT near its beginning followed by one call to m4_include per source file for tests. Each such included file, or the remainder of ‘testsuite.at’ if include files are not used, contain a sequence of test groups. Each test group begins with a call to AT_SETUP, then an arbitrary number of shell commands or calls to AT_CHECK, and then completes with a call to AT_CLEANUP. Multiple test groups can be categorized by a call to

AT_BANNER.

All of the public Autotest macros have all-uppercase names in the namespace ‘^AT_’ to prevent them from accidentally conflicting with other text; Autoconf also reserves the namespace ‘^_AT_’ for internal macros. All shell variables used in the testsuite for internal purposes have mostly-lowercase names starting with ‘at_’. Autotest also uses heredocument delimiters in the namespace ‘^_AT[A-Z]’, and makes use of the file system namespace ‘^at-’.

Since Autoconf is built on top of M4sugar (see Section 8.3 [Programming in M4sugar], page 135) and M4sh (see Chapter 9 [Programming in M4sh], page 165), you must also be aware of those namespaces (‘^_?\(m4\|AS\)_’). In general, you should not use the namespace of a package that does not own the macro or shell code you are writing.

[Macro] AT_INIT ([name])

Initialize Autotest. Giving a name to the test suite is encouraged if your package includes several test suites. Before this macro is called, AT_PACKAGE_STRING and AT_ PACKAGE_BUGREPORT must be defined, which are used to display information about the testsuite to the user. Typically, these macros are provided by a file ‘package.m4’ built by make (see Section 19.4 [Making testsuite Scripts], page 325), in order to inherit the package name, version, and bug reporting address from ‘configure.ac’.

[Macro] AT_COPYRIGHT (copyright-notice)

State that, in addition to the Free Software Foundation’s copyright on the Autotest macros, parts of your test suite are covered by copyright-notice. The copyright-notice shows up in both the head of testsuite and in ‘testsuite

--version’.
[Macro] AT_ARG_OPTION (options, help-text, [action-if-given],
   [action-if-not-given])

Accept options from the space-separated list options, a list that has leading dashes removed from the options. Long options will be prefixed with ‘--’, single-character options with ‘-’. The first word in this list is the primary option, any others are assumed to be short-hand aliases. The variable associated with it is at_arg_option, with any dashes in option replaced with underscores. If the user passes ‘--option’ to the testsuite, the variable will be set to ‘:’. If the user does not pass the option, or passes ‘--no-option’, then the variable will be set

to ‘false’.

action-if-given is run each time the option is encountered; here, the variable at_ optarg will be set to ‘:’ or ‘false’ as appropriate. at_optarg is actually just a copy

of at_arg_option.

action-if-not-given will be run once after option parsing is complete and if no option from options was used. help-text is added to the end of the list of options shown in testsuite --help (see [AS HELP STRING], page 274). It it recommended that you use a package-specific prefix to options names in order to avoid clashes with future Autotest built-in options.

[Macro] AT_ARG_OPTION_ARG (options, help-text, [action-if-given],
   [action-if-not-given])

Accept options with arguments from the space-separated list options, a list that has leading dashes removed from the options. Long options will be prefixed with ‘--’, single-character options with ‘-’. The first word in this list is the primary option, any others are assumed to be short-hand aliases. The variable associated with it is at_arg_option, with any dashes in option replaced with underscores. If the user passes ‘--option=arg’ or ‘--option arg’ to the testsuite, the variable will be set to ‘arg’. action-if-given is run each time the option is encountered; here, the variable at_ optarg will be set to ‘arg’. at_optarg is actually just a copy of at_arg_option. action-if-not-given will be run once after option parsing is complete and if no option from options was used. help-text is added to the end of the list of options shown in testsuite --help (see [AS HELP STRING], page 274). It it recommended that you use a package-specific prefix to options names in order to avoid clashes with future Autotest built-in options.

[Macro] AT_COLOR_TESTS

Enable colored test results by default when the output is connected to a terminal.

[Macro] AT_TESTED (executables)

Log the file name and answer to ‘--version’ of each program in space-separated list executables. Several invocations register new executables, in other words, don’t fear registering one program several times. Autotest test suites rely on PATH to find the tested program. This avoids the need to generate absolute names of the various tools, and makes it possible to test installed programs. Therefore, knowing which programs are being exercised is crucial to understanding problems in the test suite itself, or its occasional misuses. It is a good idea to also subscribe foreign programs you depend upon, to avoid incompatible diagnostics.

[Macro] AT_BANNER (test-category-name)

This macro identifies the start of a category of related test groups. When the resulting ‘testsuite’ is invoked with more than one test group to run, its output will include a banner containing test-category-name prior to any tests run from that category. The banner should be no more than about 40 or 50 characters. A blank banner will not print, effectively ending a category and letting subsequent test groups behave as though they are uncategorized when run in isolation.

[Macro] AT_SETUP (test-group-name)

This macro starts a group of related tests, all to be executed in the same subshell. It accepts a single argument, which holds a few words (no more than about 30 or 40 characters) quickly describing the purpose of the test group being started. testgroup-name must not expand to unbalanced quotes, although quadrigraphs can be used.

[Macro] AT_KEYWORDS (keywords)

Associate the space-separated list of keywords to the enclosing test group. This makes it possible to run “slices” of the test suite. For instance, if some of your test groups exercise some ‘foo’ feature, then using ‘AT_KEYWORDS(foo)’ lets you run ‘./testsuite -k foo’ to run exclusively these test groups. The test-group-name of the test group is automatically recorded to AT_KEYWORDS. Several invocations within a test group accumulate new keywords. In other words, don’t fear registering the same keyword several times in a test group.

[Macro] AT_CAPTURE_FILE (file)

If the current test group fails, log the contents of file. Several identical calls within one test group have no additional effect.

[Macro] AT_FAIL_IF (shell-condition)

Make the test group fail and skip the rest of its execution, if shell-condition is true. shell-condition is a shell expression such as a test command. Tests before AT_FAIL_IF will be executed and may still cause the test group to be skipped. You can instantiate this macro many times from within the same test group. You should use this macro only for very simple failure conditions. If the shell-condition could emit any kind of output you should instead use AT_CHECK like

     AT_CHECK([if shell-condition; then exit 99; fi])

so that such output is properly recorded in the ‘testsuite.log’ file.

[Macro] AT_SKIP_IF (shell-condition)

Determine whether the test should be skipped because it requires features that are unsupported on the machine under test. shell-condition is a shell expression such as a test command. Tests before AT_SKIP_IF will be executed and may still cause the test group to fail. You can instantiate this macro many times from within the same test group. You should use this macro only for very simple skip conditions. If the shell-condition could emit any kind of output you should instead use AT_CHECK like

     AT_CHECK([if shell-condition; then exit 77; fi])

so that such output is properly recorded in the ‘testsuite.log’ file.

[Macro] AT_XFAIL_IF (shell-condition)

Determine whether the test is expected to fail because it is a known bug (for unsupported features, you should skip the test). shell-condition is a shell expression such as a test command; you can instantiate this macro many times from within the same test group, and one of the conditions is enough to turn the test into an expected failure.

[Macro] AT_CLEANUP

End the current test group.

[Macro] AT_DATA (file, contents)

Initialize an input data file with given contents. Of course, the contents have to be properly quoted between square brackets to protect against included commas or spurious M4 expansion. The contents must end with an end of line. file must be a single shell word that expands into a single file name.

[Macro] AT_CHECK (commands, [status = ‘0’], [stdout], [stderr],
   [run-if-fail], [run-if-pass])
[Macro] AT_CHECK_UNQUOTED (commands, [status = ‘0’], [stdout], [stderr],
   [run-if-fail], [run-if-pass])

Execute a test by performing given shell commands. commands is output as-is, so shell expansions are honored. These commands should normally exit with status, while producing expected stdout and stderr contents. If commands exit with unexpected status 77, then the rest of the test group is skipped. If commands exit with unexpected status 99, then the test group is immediately failed. Otherwise, if this test fails, run shell commands run-if-fail or, if this test passes, run shell commands run-if-pass. This macro must be invoked in between AT_SETUP and AT_CLEANUP. If status is the literal ‘ignore’, then the corresponding exit status is not checked, except for the special cases of 77 (skip) and 99 (hard failure). The existence of hard failures allows one to mark a test as an expected failure with AT_XFAIL_IF because a feature has not yet been implemented, but to still distinguish between gracefully handling the missing feature and dumping core. A hard failure also inhibits post-test actions in run-if-fail. If the value of the stdout or stderr parameter is one of the literals in the following table, then the test treats the output according to the rules of that literal. Otherwise, the value of the parameter is treated as text that must exactly match the output given by commands on standard output and standard error (including an empty parameter for no output); any differences are captured in the testsuite log and the test is failed (unless an unexpected exit status of 77 skipped the test instead). The difference between AT_CHECK and AT_CHECK_UNQUOTED is that only the latter performs shell variable expansion (‘$’), command substitution (‘`’), and backslash escaping (‘\’) on comparison text given in the stdout and stderr arguments; if the text includes a trailing newline, this would be the same as if it were specified via an unquoted here-document. (However, there is no difference in the interpretation of commands).

‘ignore’ The content of the output is ignored, but still captured in the test group log (if the testsuite is run with option ‘-v’, the test group log is displayed as the test is run; if the test group later fails, the test group log is also copied into the overall testsuite log). This action is valid for both stdout and stderr.

‘ignore-nolog’

The content of the output is ignored, and nothing is captured in the log files. If commands are likely to produce binary output (including long lines) or large amounts of output, then logging the output can make it harder to locate details related to subsequent tests within the group, and could potentially corrupt terminal display of a user running testsuite

          -v.

‘stdout’ For the stdout parameter, capture the content of standard output to both the file ‘stdout’ and the test group log. Subsequent commands in the test group can then post-process the file. This action is often used when it is desired to use grep to look for a substring in the output, or when the output must be post-processed to normalize error messages into a common form.

‘stderr’ Like ‘stdout’, except that it only works for the stderr parameter, and the standard error capture file will be named ‘stderr’.

‘stdout-nolog’
‘stderr-nolog’

Like ‘stdout’ or ‘stderr’, except that the captured output is not duplicated into the test group log. This action is particularly useful for an intermediate check that produces large amounts of data, which will be followed by another check that filters down to the relevant data, as it makes it easier to locate details in the log.

‘expout’ For the stdout parameter, compare standard output contents with the previously created file ‘expout’, and list any differences in the testsuite log.

‘experr’ Like ‘expout’, except that it only works for the stderr parameter, and the standard error contents are compared with ‘experr’.

[Macro] AT_CHECK_EUNIT (module, test-spec, [erlflags], [run-if-fail],
   [run-if-pass])

Initialize and execute an Erlang module named module that performs tests following the test-spec EUnit test specification. test-spec must be a valid EUnit test specification, as defined in the EUnit Reference Manual (http://erlang.org/doc/apps/ eunit/index.html). erlflags are optional command-line options passed to the Erlang interpreter to execute the test Erlang module. Typically, erlflags defines at least the paths to directories containing the compiled Erlang modules under test, as ‘-pa

path1 path2 ...’.

For example, the unit tests associated with Erlang module ‘testme’, which compiled code is in subdirectory ‘src’, can be performed with:

     AT_CHECK_EUNIT([testme_testsuite], [{module, testme}],
                    [-pa "${abs_top_builddir}/src"])

This macro must be invoked in between AT_SETUP and AT_CLEANUP. Variables ERL, ERLC, and (optionally) ERLCFLAGS must be defined as the path of the Erlang interpreter, the path of the Erlang compiler, and the command-line flags to pass to the compiler, respectively. Those variables should be configured

in ‘configure.ac’ using the AC_ERLANG_PATH_ERL and AC_ERLANG_PATH_ERLC

macros, and the configured values of those variables are automatically defined in the testsuite. If ERL or ERLC is not defined, the test group is skipped.

If the EUnit library cannot be found, i.e. if module eunit cannot be loaded, the test group is skipped. Otherwise, if test-spec is an invalid EUnit test specification, the test group fails. Otherwise, if the EUnit test passes, shell commands run-if-pass are executed or, if the EUnit test fails, shell commands run-if-fail are executed and the test group fails. Only the generated test Erlang module is automatically compiled and executed. If test-spec involves testing other Erlang modules, e.g. module ‘testme’ in the example above, those modules must be already compiled. If the testsuite is run in verbose mode, with option ‘--verbose’, EUnit is also run in verbose mode to output more details about individual unit tests.

19.3 Running testsuite Scripts

Autotest test suites support the following options:

‘--help’

‘-h’ Display the list of options and exit successfully.

‘--version’

‘-V’ Display the version of the test suite and exit successfully.

‘--directory=dir’

‘-C dir’ Change the current directory to dir before creating any files. Useful for running the testsuite in a subdirectory from a top-level Makefile.

‘--jobs[=n]’

‘-j[n]’ Run n tests in parallel, if possible. If n is not given, run all given tests in parallel. Note that there should be no space before the argument to ‘-j’, as ‘-j number’ denotes the separate arguments ‘-j’ and ‘number’, see below. In parallel mode, the standard input device of the testsuite script is not available to commands inside a test group. Furthermore, banner lines are not printed, and the summary line for each test group is output after the test group completes. Summary lines may appear unordered. If verbose and trace output are enabled (see below), they may appear intermixed from concurrently running tests. Parallel mode requires the mkfifo command to work, and will be silently disabled otherwise.

‘--clean’

‘-c’ Remove all the files the test suite might have created and exit. Meant for clean Make targets.

‘--list’

‘-l’ List all the tests (or only the selection), including their possible keywords.

By default all tests are performed (or described with ‘--list’) silently in the default environment, but the environment, set of tests, and verbosity level can be tuned:

‘variable=value’

Set the environment variable to value. Use this rather than ‘FOO=foo ./testsuite’ as debugging scripts would then run in a different environment.

The variable AUTOTEST_PATH specifies the testing path to prepend to PATH. Relative directory names (not starting with ‘/’) are considered to be relative to the top level of the package being built. All directories are made absolute, first starting from the top level build tree, then from the source tree. For instance

‘./testsuite AUTOTEST_PATH=tests:bin’ for a ‘/src/foo-1.0’ source pack-
age built in ‘/tmp/foo’ results in ‘/tmp/foo/tests:/tmp/foo/bin’ and then
‘/src/foo-1.0/tests:/src/foo-1.0/bin’ being prepended to PATH.
‘number’
‘number-number’
‘number-’

‘-number’ Add the corresponding test groups, with obvious semantics, to the selection.

‘--keywords=keywords’
‘-k keywords’

Add to the selection the test groups with title or keywords (arguments to AT_ SETUP or AT_KEYWORDS) that match all keywords of the comma separated list keywords, case-insensitively. Use ‘!’ immediately before the keyword to invert the selection for this keyword. By default, the keywords match whole words; enclose them in ‘.*’ to also match parts of words. For example, running

     ./testsuite -k 'autoupdate,.*FUNC.*'

selects all tests tagged ‘autoupdate’ and with tags containing ‘FUNC’ (as in

‘AC_CHECK_FUNC’, ‘AC_FUNC_ALLOCA’, etc.), while
     ./testsuite -k '!autoupdate' -k '.*FUNC.*'

selects all tests not tagged ‘autoupdate’ or with tags containing ‘FUNC’.

‘--errexit’

‘-e’ If any test fails, immediately abort testing. This implies ‘--debug’: post test group clean up, and top-level logging are inhibited. This option is meant for the full test suite, it is not really useful for generated debugging scripts. If the testsuite is run in parallel mode using ‘--jobs’, then concurrently running tests will finish before exiting.

‘--verbose’

‘-v’ Force more verbosity in the detailed output of what is being done. This is the default for debugging scripts.

‘--color’
‘--color[=never|auto|always]’

Enable colored test results. Without an argument, or with ‘always’, test results will be colored. With ‘never’, color mode is turned off. Otherwise, if either the macro AT_COLOR_TESTS is used by the testsuite author, or the argument ‘auto’ is given, then test results are colored if standard output is connected to a terminal.

‘--debug’

‘-d’ Do not remove the files after a test group was performed—but they are still removed before, therefore using this option is sane when running several test groups. Create debugging scripts. Do not overwrite the top-level log (in order to preserve a supposedly existing full log file). This is the default for debugging scripts, but it can also be useful to debug the testsuite itself.

‘--recheck’

Add to the selection all test groups that failed or passed unexpectedly during the last non-debugging test run.

‘--trace’

‘-x’ Trigger shell tracing of the test groups.

Besides these options accepted by every Autotest testsuite, the testsuite author might have added package-specific options via the AT_ARG_OPTION and AT_ARG_OPTION_ARG macros (see Section 19.2 [Writing Testsuites], page 318); refer to testsuite --help and the package documentation for details.

19.4 Making testsuite Scripts

For putting Autotest into movement, you need some configuration and makefile machinery. We recommend, at least if your package uses deep or shallow hierarchies, that you use ‘tests/’ as the name of the directory holding all your tests and their makefile. Here is a check list of things to do. − Make sure to create the file ‘package.m4’, which defines the identity of the package.

It must define AT_PACKAGE_STRING, the full signature of the package, and AT_PACKAGE_ BUGREPORT, the address to which bug reports should be sent. For sake of completeness,

    we suggest that you also define AT_PACKAGE_NAME, AT_PACKAGE_TARNAME, AT_PACKAGE_

VERSION, and AT_PACKAGE_URL. See Section 4.1 [Initializing configure], page 19, for a description of these variables. Be sure to distribute ‘package.m4’ and to put it into the source hierarchy: the test suite ought to be shipped! See below for an example ‘Makefile’ excerpt.

 − Invoke AC_CONFIG_TESTDIR.
   [Macro] AC_CONFIG_TESTDIR (directory, [test-path = ‘directory’])

An Autotest test suite is to be configured in directory. This macro requires the in-

         stantiation of ‘directory/atconfig’ from ‘directory/atconfig.in’, and sets

the default AUTOTEST_PATH to test-path (see Section 19.3 [testsuite Invocation], page 323).

− Still within ‘configure.ac’, as appropriate, ensure that some AC_CONFIG_FILES com-

mand includes substitution for ‘tests/atlocal’. − The appropriate ‘Makefile’ should be modified so the validation in your package is

triggered by ‘make check’. An example is provided below.

With Automake, here is a minimal example for inclusion in ‘tests/Makefile.am’, in order to link ‘make check’ with a validation suite.

     # The `:;' works around a Bash 3.2 bug when the output is not writeable.
     $(srcdir)/package.m4: $(top_srcdir)/configure.ac
             :;{ \
               echo '# Signature of the current package.' && \
               echo 'm4_define([AT_PACKAGE_NAME],' && \

          echo '  [$(PACKAGE_NAME)])' && \
          echo 'm4_define([AT_PACKAGE_TARNAME],' && \
          echo '  [$(PACKAGE_TARNAME)])' && \
          echo 'm4_define([AT_PACKAGE_VERSION],' && \
          echo '  [$(PACKAGE_VERSION)])' && \
          echo 'm4_define([AT_PACKAGE_STRING],' && \
          echo '  [$(PACKAGE_STRING)])' && \
          echo 'm4_define([AT_PACKAGE_BUGREPORT],' && \
          echo '  [$(PACKAGE_BUGREPORT)])'; \
          echo 'm4_define([AT_PACKAGE_URL],' && \
          echo '  [$(PACKAGE_URL)])'; \
        } >'$(srcdir)/package.m4'
EXTRA_DIST = testsuite.at $(srcdir)/package.m4 $(TESTSUITE) atlocal.in
TESTSUITE = $(srcdir)/testsuite
check-local: atconfig atlocal $(TESTSUITE)
        $(SHELL) '$(TESTSUITE)' $(TESTSUITEFLAGS)
installcheck-local: atconfig atlocal $(TESTSUITE)
        $(SHELL) '$(TESTSUITE)' AUTOTEST_PATH='$(bindir)' \
          $(TESTSUITEFLAGS)
clean-local:
        test ! -f '$(TESTSUITE)' || \
         $(SHELL) '$(TESTSUITE)' --clean
AUTOM4TE = $(SHELL) $(srcdir)/build-aux/missing --run autom4te
AUTOTEST = $(AUTOM4TE) --language=autotest
$(TESTSUITE): $(srcdir)/testsuite.at $(srcdir)/package.m4
        $(AUTOTEST) -I '$(srcdir)' -o $@.tmp $@.at
        mv $@.tmp $@

Note that the built testsuite is distributed; this is necessary because users might not have Autoconf installed, and thus would not be able to rebuild it. Likewise, the use of ‘missing’ provides the user with a nicer error message if they modify a source file to the testsuite, and accidentally trigger the rebuild rules.

You might want to list explicitly the dependencies, i.e., the list of the files ‘testsuite.at’ includes.

If you don’t use Automake, you should include the above example in ‘tests/ Makefile.in’, along with additional lines inspired from the following:

subdir = tests
PACKAGE_NAME = @PACKAGE_NAME@
PACKAGE_TARNAME = @PACKAGE_TARNAME@
PACKAGE_VERSION = @PACKAGE_VERSION@
PACKAGE_STRING = @PACKAGE_STRING@
PACKAGE_BUGREPORT = @PACKAGE_BUGREPORT@

     PACKAGE_URL = @PACKAGE_URL@
     atconfig: $(top_builddir)/config.status
             cd $(top_builddir) && \
                $(SHELL) ./config.status $(subdir)/$@
     atlocal: $(srcdir)/atlocal.in $(top_builddir)/config.status
             cd $(top_builddir) && \
                $(SHELL) ./config.status $(subdir)/$@

and manage to have $(EXTRA_DIST) distributed. You will also want to distribute the file ‘build-aux/missing’ from the Automake project; a copy of this file resides in the Autoconf source tree.

With all this in place, and if you have not initialized ‘TESTSUITEFLAGS’ within your makefile, you can fine-tune test suite execution with this variable, for example:

     make check TESTSUITEFLAGS='-v -d -x 75 -k AC_PROG_CC CFLAGS=-g'

20 Frequent Autoconf Questions, with answers

Several questions about Autoconf come up occasionally. Here some of them are addressed.

20.1 Distributing configure Scripts

What are the restrictions on distributing configure

scripts that Autoconf generates? How does that affect my programs that use them?

There are no restrictions on how the configuration scripts that Autoconf produces may be distributed or used. In Autoconf version 1, they were covered by the GNU General Public License. We still encourage software authors to distribute their work under terms like those of the GPL, but doing so is not required to use Autoconf.

Of the other files that might be used with configure, ‘config.h.in’ is under whatever copyright you use for your ‘configure.ac’. ‘config.sub’ and ‘config.guess’ have an exception to the GPL when they are used with an Autoconf-generated configure script, which permits you to distribute them under the same terms as the rest of your package. ‘install-sh’ is from the X Consortium and is not copyrighted.

20.2 Why Require GNU M4?

Why does Autoconf require GNU M4?

Many M4 implementations have hard-coded limitations on the size and number of macros that Autoconf exceeds. They also lack several builtin macros that it would be difficult to get along without in a sophisticated application like Autoconf, including:

     m4_builtin
     m4_indir
     m4_bpatsubst
     __file__
     __line__

Autoconf requires version 1.4.6 or later of GNU M4.

Since only software maintainers need to use Autoconf, and since GNU M4 is simple to configure and install, it seems reasonable to require GNU M4 to be installed also. Many maintainers of GNU and other free software already have most of the GNU utilities installed, since they prefer them.

20.3 How Can I Bootstrap?

If Autoconf requires GNU M4 and GNU M4 has an Autoconf configure script, how do I bootstrap? It seems like a chicken and egg problem!

This is a misunderstanding. Although GNU M4 does come with a configure script produced by Autoconf, Autoconf is not required in order to run the script and install GNU M4. Autoconf is only required if you want to change the M4 configure script, which few people have to do (mainly its maintainer).

20.4 Why Not Imake?

Why not use Imake instead of configure scripts? Several people have written addressing this question, so I include adaptations of their

explanations here.

The following answer is based on one written by Richard Pixley:

Autoconf generated scripts frequently work on machines that it has never been set up to handle before. That is, it does a good job of inferring a configuration for a new system. Imake cannot do this. Imake uses a common database of host specific data. For X11, this makes sense because the distribution is made as a collection of tools, by one central authority who has control over the database. GNU tools are not released this way. Each GNU tool has a maintainer; these maintainers are scattered across the world. Using a common database would be a maintenance nightmare. Autoconf may appear to be this kind of database, but in fact it is not. Instead of listing host dependencies, it lists program requirements. If you view the GNU suite as a collection of native tools, then the problems are similar. But the GNU development tools can be configured as cross tools in almost any host+target permutation. All of these configurations can be installed concurrently. They can even be configured to share host independent files across hosts. Imake doesn’t address these issues. Imake templates are a form of standardization. The GNU coding standards address the same issues without necessarily imposing the same restrictions.

Here is some further explanation, written by Per Bothner:

One of the advantages of Imake is that it easy to generate large makefiles using the ‘#include’ and macro mechanisms of cpp. However, cpp is not programmable: it has limited conditional facilities, and no looping. And cpp cannot inspect its environment. All of these problems are solved by using sh instead of cpp. The shell is fully programmable, has macro substitution, can execute (or source) other shell scripts, and can inspect its environment.

Paul Eggert elaborates more:

With Autoconf, installers need not assume that Imake itself is already installed and working well. This may not seem like much of an advantage to people who are accustomed to Imake. But on many hosts Imake is not installed or the default installation is not working well, and requiring Imake to install a package hinders the acceptance of that package on those hosts. For example, the Imake template and configuration files might not be installed properly on a host, or the Imake build procedure might wrongly assume that all source files are in one big directory tree, or the Imake configuration might assume one compiler whereas the package or the installer needs to use another, or there might be a version mismatch between the Imake expected by the package and the Imake supported by the host. These problems are much rarer with Autoconf, where each package comes with its own independent configuration processor.

Also, Imake often suffers from unexpected interactions between make and the installer’s C preprocessor. The fundamental problem here is that the C preprocessor was designed to preprocess C programs, not makefiles. This is much less of a problem with Autoconf, which uses the general-purpose preprocessor M4, and where the package’s author (rather than the installer) does the preprocessing in a standard way.

Finally, Mark Eichin notes:

Imake isn’t all that extensible, either. In order to add new features to Imake, you need to provide your own project template, and duplicate most of the features of the existing one. This means that for a sophisticated project, using the vendor-provided Imake templates fails to provide any leverage—since they don’t cover anything that your own project needs (unless it is an X11 program). On the other side, though: The one advantage that Imake has over configure: ‘Imakefile’ files tend to be much shorter (likewise, less redundant) than ‘Makefile.in’ files. There is a fix to this, however—at least for the Kerberos V5 tree, we’ve modified things to call in common ‘post.in’ and ‘pre.in’ makefile fragments for the entire tree. This means that a lot of common things don’t have to be duplicated, even though they normally are in configure setups.

20.5 How Do I #define Installation Directories?

My program needs library files, installed in datadir and similar. If I use

     AC_DEFINE_UNQUOTED([DATADIR], [$datadir],
       [Define to the read-only architecture-independent
        data directory.])

I get

     #define DATADIR "${prefix}/share"

As already explained, this behavior is on purpose, mandated by the GNU Coding Standards, see Section 4.8.2 [Installation Directory Variables], page 29. There are several means to achieve a similar goal: − Do not use AC_DEFINE but use your makefile to pass the actual value of datadir via

compilation flags. See Section 4.8.2 [Installation Directory Variables], page 29, for the details.

− This solution can be simplified when compiling a program: you may either extend the

CPPFLAGS:
    CPPFLAGS = -DDATADIR='"$(datadir)"' @CPPFLAGS@

If you are using Automake, you should use AM_CPPFLAGS instead:

    AM_CPPFLAGS = -DDATADIR='"$(datadir)"'

Alternatively, create a dedicated header file:

         DISTCLEANFILES = myprog-paths.h
         myprog-paths.h: Makefile
                 echo '#define DATADIR "$(datadir)"' >$@

− Use AC_DEFINE but have configure compute the literal value of datadir and others. Many people have wrapped macros to automate this task; for an example, see the macro

    AC_DEFINE_DIR from the Autoconf Macro Archive (http://www.gnu.org/software/
    autoconf-archive/).

This solution does not conform to the GNU Coding Standards.

− Note that all the previous solutions hard wire the absolute name of these directories in the executables, which is not a good property. You may try to compute the names relative to prefix, and try to find prefix at runtime, this way your package is relocatable.

20.6 What is ‘autom4te.cache’?

What is this directory ‘autom4te.cache’? Can I safely remove it?

In the GNU Build System, ‘configure.ac’ plays a central role and is read by many tools:

autoconf to create ‘configure’, autoheader to create ‘config.h.in’, automake to create
‘Makefile.in’, autoscan to check the completeness of ‘configure.ac’, autoreconf to

check the GNU Build System components that are used. To “read ‘configure.ac’” actually means to compile it with M4, which can be a long process for complex ‘configure.ac’.

This is why all these tools, instead of running directly M4, invoke autom4te (see Section 8.2.1 [autom4te Invocation], page 130) which, while answering to a specific demand, stores additional information in ‘autom4te.cache’ for future runs. For instance, if you run autoconf, behind the scenes, autom4te also stores information for the other tools, so that when you invoke autoheader or automake etc., reprocessing ‘configure.ac’ is not needed. The speed up is frequently 30%, and is increasing with the size of ‘configure.ac’.

But it is and remains being simply a cache: you can safely remove it.

Can I permanently get rid of it?

The creation of this cache can be disabled from ‘~/.autom4te.cfg’, see Section 8.2.2 [Customizing autom4te], page 134, for more details. You should be aware that disabling the cache slows down the Autoconf test suite by 40%. The more GNU Build System components are used, the more the cache is useful; for instance running ‘autoreconf -f’ on the Core Utilities is twice slower without the cache although ‘--force’ implies that the cache is not fully exploited, and eight times slower than without ‘--force’.

20.7 Header Present But Cannot Be Compiled

The most important guideline to bear in mind when checking for features is to mimic as much as possible the intended use. Unfortunately, old versions of AC_CHECK_HEADER and AC_CHECK_HEADERS failed to follow this idea, and called the preprocessor, instead of the compiler, to check for headers. As a result, incompatibilities between headers went unnoticed during configuration, and maintainers finally had to deal with this issue elsewhere.

Chapter 20: Frequent Autoconf Questions, with answers 333

The transition began with Autoconf 2.56. As of Autoconf 2.64 both checks are performed, and configure complains loudly if the compiler and the preprocessor do not agree. However, only the compiler result is considered.

Consider the following example:

$ cat number.h
typedef int number;
$ cat pi.h
const number pi = 3;
$ cat configure.ac
AC_INIT([Example], [1.0], [bug-example@example.org])
AC_CHECK_HEADERS([pi.h])
$ autoconf -Wall
$ ./configure
checking for gcc... gcc
checking for C compiler default output file name... a.out
checking whether the C compiler works... yes
checking whether we are cross compiling... no
checking for suffix of executables...
checking for suffix of object files... o
checking whether we are using the GNU C compiler... yes
checking whether gcc accepts -g... yes
checking for gcc option to accept ISO C89... none needed
checking how to run the C preprocessor... gcc -E
checking for grep that handles long lines and -e... grep
checking for egrep... grep -E
checking for ANSI C header files... yes
checking for sys/types.h... yes
checking for sys/stat.h... yes
checking for stdlib.h... yes
checking for string.h... yes
checking for memory.h... yes
checking for strings.h... yes
checking for inttypes.h... yes
checking for stdint.h... yes
checking for unistd.h... yes
checking pi.h usability... no
checking pi.h presence... yes
configure: WARNING: pi.h: present but cannot be compiled
configure: WARNING: pi.h:     check for missing prerequisite headers?
configure: WARNING: pi.h: see the Autoconf documentation
configure: WARNING: pi.h:     section "Present But Cannot Be Compiled"
configure: WARNING: pi.h: proceeding with the compiler's result
configure: WARNING:     ## -------------------------------------- ##
configure: WARNING:     ## Report this to bug-example@example.org ##
configure: WARNING:     ## -------------------------------------- ##
checking for pi.h... yes

The proper way the handle this case is using the fourth argument (see Section 5.6.3 [Generic Headers], page 71):

$ cat configure.ac
AC_INIT([Example], [1.0], [bug-example@example.org])
AC_CHECK_HEADERS([number.h pi.h], [], [],
[[#ifdef HAVE_NUMBER_H
# include <number.h>
#endif

]])
$ autoconf -Wall
$ ./configure
checking for gcc... gcc
checking for C compiler default output... a.out
checking whether the C compiler works... yes
checking whether we are cross compiling... no
checking for suffix of executables...
checking for suffix of object files... o
checking whether we are using the GNU C compiler... yes
checking whether gcc accepts -g... yes
checking for gcc option to accept ANSI C... none needed
checking for number.h... yes
checking for pi.h... yes

See Section 5.6.2 [Particular Headers], page 66, for a list of headers with their prerequisites.

20.8 Expanded Before Required

Older versions of Autoconf silently built files with incorrect ordering between dependent macros if an outer macro first expanded, then later indirectly required, an inner macro. Starting with Autoconf 2.64, this situation no longer generates out-of-order code, but results in duplicate output and a syntax warning:

$ cat configure.ac
⇒AC_DEFUN([TESTA], [[echo in A
⇒if test -n "$SEEN_A" ; then echo duplicate ; fi
⇒SEEN_A=:]])
⇒AC_DEFUN([TESTB], [AC_REQUIRE([TESTA])[echo in B
⇒if test -z "$SEEN_A" ; then echo bug ; fi]])
⇒AC_DEFUN([TESTC], [AC_REQUIRE([TESTB])[echo in C]])
⇒AC_DEFUN([OUTER], [[echo in OUTER]
⇒TESTA
⇒TESTC])
⇒AC_INIT
⇒OUTER
⇒AC_OUTPUT
$ autoconf
⇒configure.ac:11: warning: AC_REQUIRE:
⇒ `TESTA' was expanded before it was required
⇒configure.ac:4: TESTB is expanded from...
⇒configure.ac:6: TESTC is expanded from...
⇒configure.ac:7: OUTER is expanded from...
⇒configure.ac:11: the top level

To avoid this warning, decide what purpose the macro in question serves. If it only needs to be expanded once (for example, if it provides initialization text used by later macros), then the simplest fix is to change the macro to be declared with AC_DEFUN_ONCE (see Section 10.4.3 [One-Shot Macros], page 181), although this only works in Autoconf 2.64 and newer. A more portable fix is to change all instances of direct calls to instead go through AC_REQUIRE (see Section 10.4.1 [Prerequisite Macros], page 178). If, instead, the macro is parameterized by arguments or by the current definition of other macros in the m4 environment, then the macro should always be directly expanded instead of required.

For another case study, consider this example trimmed down from an actual package. Originally, the package contained shell code and multiple macro invocations at the top level

of ‘configure.ac’:
     AC_DEFUN([FOO], [AC_COMPILE_IFELSE([...])])
     foobar=
     AC_PROG_CC
     FOO

but that was getting complex, so the author wanted to offload some of the text into a new macro in another file included via ‘aclocal.m4’. The naïve approach merely wraps the text in a new macro:

     AC_DEFUN([FOO], [AC_COMPILE_IFELSE([...])])
     AC_DEFUN([BAR], [
     foobar=
     AC_PROG_CC
     FOO
     ])
     BAR

With older versions of Autoconf, the setting of ‘foobar=’ occurs before the single compiler check, as the author intended. But with Autoconf 2.64, this issues the “expanded before it was required” warning for AC_PROG_CC, and outputs two copies of the compiler check, one before ‘foobar=’, and one after. To understand why this is happening, remember that the

use of AC_COMPILE_IFELSE includes a call to AC_REQUIRE([AC_PROG_CC]) under the hood.

According to the documented semantics of AC_REQUIRE, this means that AC_PROG_CC must occur before the body of the outermost AC_DEFUN, which in this case is BAR, thus preceeding the use of ‘foobar=’. The older versions of Autoconf were broken with regards to the rules of AC_REQUIRE, which explains why the code changed from one over to two copies of AC_ PROG_CC when upgrading autoconf. In other words, the author was unknowingly relying on a bug exploit to get the desired results, and that exploit broke once the bug was fixed.

So, what recourse does the author have, to restore their intended semantics of setting ‘foobar=’ prior to a single compiler check, regardless of whether Autoconf 2.63 or 2.64 is used? One idea is to remember that only AC_DEFUN is impacted by AC_REQUIRE; there is always the possibility of using the lower-level m4_define:

     AC_DEFUN([FOO], [AC_COMPILE_IFELSE([...])])
     m4_define([BAR], [
     foobar=
     AC_PROG_CC
     FOO
     ])
     BAR

This works great if everything is in the same file. However, it does not help in the case where the author wants to have aclocal find the definition of BAR from its own file, since aclocal requires the use of AC_DEFUN. In this case, a better fix is to recognize that if BAR also uses AC_REQUIRE, then there will no longer be direct expansion prior to a subsequent require. Then, by creating yet another helper macro, the author can once again guarantee a single invocation of AC_PROG_CC, which will still occur after foobar=. The author can also use AC_BEFORE to make sure no other macro appearing before BAR has triggered an unwanted expansion of AC_PROG_CC.

     AC_DEFUN([FOO], [AC_COMPILE_IFELSE([...])])
     AC_DEFUN([BEFORE_CC], [
     foobar=
     ])
     AC_DEFUN([BAR], [
     AC_BEFORE([$0], [AC_PROG_CC])dnl
     AC_REQUIRE([BEFORE_CC])dnl
     AC_REQUIRE([AC_PROG_CC])dnl
     FOO
     ])
     BAR

20.9 Debugging configure scripts

While in general, configure scripts generated by Autoconf strive to be fairly portable to various systems, compilers, shells, and other tools, it may still be necessary to debug a failing test, broken script or makefile, or fix or override an incomplete, faulty, or erroneous test, especially during macro development. Failures can occur at all levels, in M4 syntax or semantics, shell script issues, or due to bugs in the test or the tools invoked by configure. Together with the rather arcane error message that m4 and make may produce when their input contains syntax errors, this can make debugging rather painful.

Nevertheless, here is a list of hints and strategies that may help:

• When autoconf fails, common causes for error include:

• mismatched or unbalanced parentheses or braces (see Section 8.1.7 [Balancing Parentheses], page 127),

• under- or overquoted macro arguments (see Section 3.1.2 [Autoconf Language], page 9, see Section 8.1.3 [Quoting and Parameters], page 123, see Section 8.1.4 [Quotation and Nested Macros], page 124),

• spaces between macro name and opening parenthesis (see Section 3.1.2 [Autoconf Language], page 9).

Typically, it helps to go back to the last working version of the input and compare the differences for each of these errors. Another possibility is to sprinkle pairs of m4_ traceon and m4_traceoff judiciously in the code, either without a parameter or listing some macro names and watch m4 expand its input verbosely (see Section 8.4 [Debugging via autom4te], page 162).

• Sometimes autoconf succeeds but the generated configure script has invalid shell syntax. You can detect this case by running ‘bash -n configure’ or ‘sh -n configure’. If this command fails, the same tips apply, as if autoconf had failed.

• Debugging configure script execution may be done by sprinkling pairs of set -x and set +x into the shell script before and after the region that contains a bug. Running

the whole script with ‘shell ./configure -vx 2>&1 | tee log-file’ with a decent

shell may work, but produces lots of output. Here, it can help to search for markers like ‘checking for’ a particular test in the log-file.

• When configure tests produce invalid results for your system, it may be necessary to

override them:

• For programs, tools or libraries variables, preprocessor, compiler, or linker flags, it is often sufficient to override them at make run time with some care (see Section 12.7 [Macros and Submakes], page 242). Since this normally won’t cause configure to be run again with these changed settings, it may fail if the changed variable would have caused different test results from configure, so this may work only for simple differences.

• Most tests which produce their result in a substituted variable allow to override the test by setting the variable on the configure command line (see Section 16.2 [Compilers and Options], page 282, see Section 16.9 [Defining Variables], page 285, see Section 16.6 [Particular Systems], page 284).

• Many tests store their result in a cache variable (see Section 7.4 [Caching Results], page 115). This lets you override them either on the configure command line as above, or through a primed cache or site file (see Section 7.4.2 [Cache Files], page 117, see Section 15.8 [Site Defaults], page 278). The name of a cache variable is documented with a test macro or may be inferred from Section 7.4.1 [Cache Variable Names], page 116; the precise semantics of undocumented variables are often internal details, subject to change.

• Alternatively, configure may produce invalid results because of uncaught programming errors, in your package or in an upstream library package. For example, when AC_ CHECK_LIB fails to find a library with a specified function, always check ‘config.log’. This will reveal the exact error that produced the failing result: the library linked by AC_CHECK_LIB probably has a fatal bug.

Conversely, as macro author, you can make it easier for users of your macro:

• by minimizing dependencies between tests and between test results as far as possible,

• by using make variables to factorize and allow override of settings at make run time,

• by honoring the GNU Coding Standards and not overriding flags reserved for the user except temporarily during configure tests,

• by not requiring users of your macro to use the cache variables. Instead, expose the result of the test via run-if-true and run-if-false parameters. If the result is not a boolean, then provide it through documented shell variables.

21 History of Autoconf

You may be wondering, Why was Autoconf originally written? How did it get into its present form? (Why does it look like gorilla spit?) If you’re not wondering, then this chapter contains no information useful to you, and you might as well skip it. If you are wondering, then let there be light…

21.1 Genesis

In June 1991 I was maintaining many of the GNU utilities for the Free Software Foundation. As they were ported to more platforms and more programs were added, the number of ‘-D’ options that users had to select in the makefile (around 20) became burdensome. Especially for me—I had to test each new release on a bunch of different systems. So I wrote a little shell script to guess some of the correct settings for the fileutils package, and released it as part of fileutils 2.0. That configure script worked well enough that the next month I adapted it (by hand) to create similar configure scripts for several other GNU utilities packages. Brian Berliner also adapted one of my scripts for his CVS revision control system.

Later that summer, I learned that Richard Stallman and Richard Pixley were developing similar scripts to use in the GNU compiler tools; so I adapted my configure scripts to support their evolving interface: using the file name ‘Makefile.in’ as the templates; adding ‘+srcdir’, the first option (of many); and creating ‘config.status’ files.

21.2 Exodus

As I got feedback from users, I incorporated many improvements, using Emacs to search and replace, cut and paste, similar changes in each of the scripts. As I adapted more GNU utilities packages to use configure scripts, updating them all by hand became impractical. Rich Murphey, the maintainer of the GNU graphics utilities, sent me mail saying that the configure scripts were great, and asking if I had a tool for generating them that I could send him. No, I thought, but I should! So I started to work out how to generate them. And the journey from the slavery of hand-written configure scripts to the abundance and ease of Autoconf began.

Cygnus configure, which was being developed at around that time, is table driven; it is meant to deal mainly with a discrete number of system types with a small number of mainly unguessable features (such as details of the object file format). The automatic configuration system that Brian Fox had developed for Bash takes a similar approach. For general use, it seems to me a hopeless cause to try to maintain an up-to-date database of which features each variant of each operating system has. It’s easier and more reliable to check for most features on the fly—especially on hybrid systems that people have hacked on locally or that have patches from vendors installed.

I considered using an architecture similar to that of Cygnus configure, where there is a single configure script that reads pieces of ‘configure.in’ when run. But I didn’t want to have to distribute all of the feature tests with every package, so I settled on having a different configure made from each ‘configure.in’ by a preprocessor. That approach also offered more control and flexibility.

I looked briefly into using the Metaconfig package, by Larry Wall, Harlan Stenn, and Raphael Manfredi, but I decided not to for several reasons. The Configure scripts it produces are interactive, which I find quite inconvenient; I didn’t like the ways it checked for some features (such as library functions); I didn’t know that it was still being maintained, and the Configure scripts I had seen didn’t work on many modern systems (such as System V R4 and NeXT); it wasn’t flexible in what it could do in response to a feature’s presence or absence; I found it confusing to learn; and it was too big and complex for my needs (I didn’t realize then how much Autoconf would eventually have to grow).

I considered using Perl to generate my style of configure scripts, but decided that M4 was better suited to the job of simple textual substitutions: it gets in the way less, because output is implicit. Plus, everyone already has it. (Initially I didn’t rely on the GNU extensions to M4.) Also, some of my friends at the University of Maryland had recently been putting M4 front ends on several programs, including tvtwm, and I was interested in trying out a new language.

21.3 Leviticus

Since my configure scripts determine the system’s capabilities automatically, with no interactive user intervention, I decided to call the program that generates them Autoconfig. But with a version number tacked on, that name would be too long for old Unix file systems, so I shortened it to Autoconf.

In the fall of 1991 I called together a group of fellow questers after the Holy Grail of portability (er, that is, alpha testers) to give me feedback as I encapsulated pieces of my handwritten scripts in M4 macros and continued to add features and improve the techniques used in the checks. Prominent among the testers were François Pinard, who came up with the idea of making an Autoconf shell script to run M4 and check for unresolved macro calls; Richard Pixley, who suggested running the compiler instead of searching the file system to find include files and symbols, for more accurate results; Karl Berry, who got Autoconf to configure TEX and added the macro index to the documentation; and Ian Lance Taylor, who added support for creating a C header file as an alternative to putting ‘-D’ options in a makefile, so he could use Autoconf for his UUCP package. The alpha testers cheerfully adjusted their files again and again as the names and calling conventions of the Autoconf macros changed from release to release. They all contributed many specific checks, great ideas, and bug fixes.

21.4 Numbers

In July 1992, after months of alpha testing, I released Autoconf 1.0, and converted many GNU packages to use it. I was surprised by how positive the reaction to it was. More people started using it than I could keep track of, including people working on software that wasn’t part of the GNU Project (such as TCL, FSP, and Kerberos V5). Autoconf continued to improve rapidly, as many people using the configure scripts reported problems they encountered.

Autoconf turned out to be a good torture test for M4 implementations. Unix M4 started to dump core because of the length of the macros that Autoconf defined, and several bugs showed up in GNU M4 as well. Eventually, we realized that we needed to use some features that only GNU M4 has. 4.3BSD M4, in particular, has an impoverished set of builtin macros; the System V version is better, but still doesn’t provide everything we need.

More development occurred as people put Autoconf under more stresses (and to uses I hadn’t anticipated). Karl Berry added checks for X11. david zuhn contributed C++ support. François Pinard made it diagnose invalid arguments. Jim Blandy bravely coerced it into configuring GNU Emacs, laying the groundwork for several later improvements. Roland McGrath got it to configure the GNU C Library, wrote the autoheader script to automate the creation of C header file templates, and added a ‘--verbose’ option to configure. Noah Friedman added the ‘--autoconf-dir’ option and AC_MACRODIR environment variable. (He also coined the term autoconfiscate to mean “adapt a software package to use Autoconf”.) Roland and Noah improved the quoting protection in AC_DEFINE and fixed many bugs, especially when I got sick of dealing with portability problems from February through June, 1993.

21.5 Deuteronomy

A long wish list for major features had accumulated, and the effect of several years of patching by various people had left some residual cruft. In April 1994, while working for Cygnus Support, I began a major revision of Autoconf. I added most of the features of the Cygnus configure that Autoconf had lacked, largely by adapting the relevant parts of Cygnus configure with the help of david zuhn and Ken Raeburn. These features include support for using ‘config.sub’, ‘config.guess’, ‘--host’, and ‘--target’; making links to files; and running configure scripts in subdirectories. Adding these features enabled Ken to convert GNU as, and Rob Savoye to convert DejaGNU, to using Autoconf.

I added more features in response to other peoples’ requests. Many people had asked for configure scripts to share the results of the checks between runs, because (particularly when configuring a large source tree, like Cygnus does) they were frustratingly slow. Mike Haertel suggested adding site-specific initialization scripts. People distributing software that had to unpack on MS-DOS asked for a way to override the ‘.in’ extension on the file names, which produced file names like ‘config.h.in’ containing two dots. Jim Avera did an extensive examination of the problems with quoting in AC_DEFINE and AC_SUBST; his insights led to significant improvements. Richard Stallman asked that compiler output be sent to ‘config.log’ instead of ‘/dev/null’, to help people debug the Emacs configure script.

I made some other changes because of my dissatisfaction with the quality of the program. I made the messages showing results of the checks less ambiguous, always printing a result. I regularized the names of the macros and cleaned up coding style inconsistencies. I added some auxiliary utilities that I had developed to help convert source code packages to use Autoconf. With the help of François Pinard, I made the macros not interrupt each others’ messages. (That feature revealed some performance bottlenecks in GNU M4, which he hastily corrected!) I reorganized the documentation around problems people want to solve. And I began a test suite, because experience had shown that Autoconf has a pronounced tendency to regress when we change it.

Again, several alpha testers gave invaluable feedback, especially François Pinard, Jim Meyering, Karl Berry, Rob Savoye, Ken Raeburn, and Mark Eichin.

Finally, version 2.0 was ready. And there was much rejoicing. (And I have free time again. I think. Yeah, right.)

Appendix A GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

The “publisher” means any person or entity that distributes copies of the Document to the public.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Appendix B: Indices 351

Appendix B Indices

B.1 Environment Variable Index

This is an alphabetical list of the environment variables that might influence Autoconf checks.

_. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202

B

BIN_SH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202

C

CC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 CDPATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 CFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26, 82 CLICOLOR_FORCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 CONFIG_COMMANDS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 CONFIG_FILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 CONFIG_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 CONFIG_LINKS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 CONFIG_SHELL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 CONFIG_SITE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 CONFIG_STATUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 CPP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 CPPFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 CXX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 CXXCPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 CXXFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27, 87 CYGWIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295

D

DUALCASE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203

E

ENV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 ERL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 ERLC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 ERLCFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27, 88

F

F77. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 FC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 FCFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27, 89 FFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27, 89 FPATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203

G

GREP_OPTIONS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204

I

IFS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204

L

LANG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 LANGUAGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 LC_ADDRESS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 LC_ALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 204 LC_COLLATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 LC_CTYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 LC_IDENTIFICATION . . . . . . . . . . . . . . . . . . . . . . . . . . 205 LC_MEASUREMENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 LC_MESSAGES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 LC_MONETARY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 LC_NAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 LC_NUMERIC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 LC_PAPER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 LC_TELEPHONE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 LC_TIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204 LDFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 LIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 LINENO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 205

M

M4. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 MAIL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 MAILPATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203

N

NULLCMD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208

O

OBJC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 OBJCFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28, 87 OBJCPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 OBJCXX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 OBJCXXCPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 OBJCXXFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28, 88

352 Autoconf

P

PATH_SEPARATOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 PS1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 PS2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 PS4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 PWD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208

R

RANDOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209

S

SHELL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 SIMPLE_BACKUP_SUFFIX . . . . . . . . . . . . . . . . . . . . . . . 292 status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209

W

WARNINGS . . . . . . . . . . . . . . . . . . . . . . . . . . 14, 17, 38, 131

X

XMKMF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

Y

YACC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 YFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

B.2 Output Variable Index

This is an alphabetical list of the variables that Autoconf can substitute into files that it creates, typically one or more makefiles. See Section 7.2 [Setting Output Variables], page 112, for more information on how this is done.

A

abs_builddir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 abs_srcdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 abs_top_builddir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 abs_top_srcdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 ALLOCA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 AWK. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

B

bindir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 build . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 build_alias . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 build_cpu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 build_os . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 build_vendor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 builddir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

C

CC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82, 86, 95 CFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26, 82 configure_input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 CPP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 CPPFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 cross_compiling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 CXX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 CXXCPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 CXXFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27, 87

D

datadir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 datarootdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 DEFS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 docdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 dvidir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

E

ECHO_C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ECHO_N . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 ECHO_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 EGREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 ERL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89, 100, 106 ERLANG_ERTS_VER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 ERLANG_INSTALL_LIB_DIR . . . . . . . . . . . . . . . . . . 32, 97

ERLANG_INSTALL_LIB_DIR_library . . . . . . . . . 32, 97

ERLANG_LIB_DIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 ERLANG_LIB_DIR_library . . . . . . . . . . . . . . . . . . . . . . 96 ERLANG_LIB_VER_library . . . . . . . . . . . . . . . . . . . . . . 96 ERLANG_ROOT_DIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 ERLC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88, 100 ERLCFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27, 88, 100 exec_prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 EXEEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79, 296

F

F77. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 FC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 FCFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27, 89 FCLIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 FFLAGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27, 89 FGREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 FLIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

G

GETGROUPS_LIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 GETLOADAVG_LIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 GREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

H

host . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 host_alias . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 host_cpu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 host_os . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 host_vendor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 htmldir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

I

includedir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 infodir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 INSTALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 INSTALL_DATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 INSTALL_PROGRAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 INSTALL_SCRIPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

K

KMEM_GROUP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

L

LDFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 LEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 LEX_OUTPUT_ROOT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 LEXLIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 libdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 libexecdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 LIBOBJDIR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 LIBOBJS. . . . . . . . . . . . . . . . . . . . . . . . . 58, 60, 63, 64, 74 LIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28, 298, 301, 305 LN_S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 localedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 localstatedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

M

mandir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 MKDIR_P . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

N

NEED_SETGID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

O

OBJC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 OBJCFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28, 87 OBJCPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 OBJCXX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 OBJCXXCPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 OBJCXXFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28, 88 OBJEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79, 299 oldincludedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 OPENMP_CFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 OPENMP_CXXFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 OPENMP_FCFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 OPENMP_FFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

P

PACKAGE_BUGREPORT. . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PACKAGE_NAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PACKAGE_STRING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PACKAGE_TARNAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PACKAGE_URL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PACKAGE_VERSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 pdfdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 POW_LIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 program_transform_name. . . . . . . . . . . . . . . . . . . . . 276 psdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

R

RANLIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

S

sbindir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 SED. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 SET_MAKE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 sharedstatedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 srcdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 subdirs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 sysconfdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

T

target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 target_alias . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 target_cpu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 target_os. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 target_vendor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 top_build_prefix. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 top_builddir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 top_srcdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

X

X_CFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 X_EXTRA_LIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 X_LIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 X_PRE_LIBS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

Y

YACC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

B.3 Preprocessor Symbol Index

This is an alphabetical list of the C preprocessor symbols that the Autoconf macros define. To work with Autoconf, C source code needs to use these names in #if or #ifdef directives.

__CHAR_UNSIGNED__. . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 __EXTENSIONS__. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 __PROTOTYPES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 _ALL_SOURCE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95, 293 _FILE_OFFSET_BITS. . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 _GNU_SOURCE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95, 297 _LARGE_FILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 _LARGEFILE_SOURCE. . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 _MINIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95, 299 _OPENMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 _POSIX_1_SOURCE . . . . . . . . . . . . . . . . . . . . . . . . . 95, 299

_POSIX_PTHREAD_SEMANTICS . . . . . . . . . . . . . . . . . . . 95

_POSIX_SOURCE . . . . . . . . . . . . . . . . . . . . . . . . . . . 95, 299 _POSIX_VERSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 _TANDEM_SOURCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

A

ALIGNOF_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

C

C_ALLOCA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 C_GETLOADAVG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 CLOSEDIR_VOID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 const . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 CXX_NO_MINUS_C_MINUS_O . . . . . . . . . . . . . . . . . . . . . . 87

D

DGUX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 DIRENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295

F

F77_DUMMY_MAIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 F77_FUNC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 F77_FUNC_. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 F77_MAIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 F77_NO_MINUS_C_MINUS_O . . . . . . . . . . . . . . . . . . . . . . 90 FC_FUNC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 FC_FUNC_. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 FC_MAIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 FC_NO_MINUS_C_MINUS_O . . . . . . . . . . . . . . . . . . . . . . . 90 FLEXIBLE_ARRAY_MEMBER . . . . . . . . . . . . . . . . . . . . . . . 86

G

GETGROUPS_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 GETLOADAVG_PRIVILEGED . . . . . . . . . . . . . . . . . . . . . . . 58 GETPGRP_VOID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 gid_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 GWINSZ_IN_SYS_IOCTL . . . . . . . . . . . . . . . . . . . . . . . . . 71

H

HAVE__BOOL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 HAVE_aggregate_member . . . . . . . . . . . . . . . . . . . . . . 75 HAVE_ALLOCA_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 HAVE_C_BACKSLASH_A . . . . . . . . . . . . . . . . . . . . . . . . . . 84 HAVE_C_VARARRAYS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 HAVE_CHOWN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 HAVE_CONFIG_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 HAVE_DECL_STRERROR_R . . . . . . . . . . . . . . . . . . . . . . . . 61 HAVE_DECL_symbol . . . . . . . . . . . . . . . . . . . . . . . . . 73, 74 HAVE_DECL_TZNAME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 HAVE_DIRENT_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 HAVE_DOPRNT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 HAVE_FSEEKO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 HAVE_function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63, 64 HAVE_GETGROUPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 HAVE_GETMNTENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 HAVE_header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71, 72 HAVE_INT16_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 HAVE_INT32_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 HAVE_INT64_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 HAVE_INT8_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 HAVE_INTMAX_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 HAVE_INTPTR_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 HAVE_LONG_DOUBLE . . . . . . . . . . . . . . . . . . . . . . . . 76, 293 HAVE_LONG_DOUBLE_WIDER . . . . . . . . . . . . . . . . . . . . . . 77 HAVE_LONG_FILE_NAMES . . . . . . . . . . . . . . . . . . . . . . . . 95 HAVE_LONG_LONG_INT . . . . . . . . . . . . . . . . . . . . . . . . . . 77

HAVE_LSTAT_EMPTY_STRING_BUG . . . . . . . . . . . . . . . . 61

HAVE_MALLOC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 HAVE_MBRTOWC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 HAVE_MMAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 HAVE_NDIR_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 HAVE_NLIST_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 HAVE_OBSTACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 HAVE_REALLOC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 HAVE_RESOLV_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

HAVE_RESTARTABLE_SYSCALLS . . . . . . . . . . . . . . . . . 302

HAVE_ST_BLKSIZE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 HAVE_ST_BLOCKS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 HAVE_ST_RDEV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302

HAVE_STAT_EMPTY_STRING_BUG . . . . . . . . . . . . . . . . . 61

HAVE_STDBOOL_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 HAVE_STRCOLL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 HAVE_STRERROR_R . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 HAVE_STRFTIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 HAVE_STRINGIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 HAVE_STRNLEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 HAVE_STRTOLD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

HAVE_STRUCT_DIRENT_D_INO . . . . . . . . . . . . . . . . . . . 74
HAVE_STRUCT_DIRENT_D_TYPE . . . . . . . . . . . . . . . . . . 74
HAVE_STRUCT_STAT_ST_BLKSIZE . . . . . . . . . . . . . . . 302
HAVE_STRUCT_STAT_ST_BLOCKS . . . . . . . . . . . . . . . . . 74
HAVE_STRUCT_STAT_ST_RDEV . . . . . . . . . . . . . . . . . . 302

HAVE_STRUCT_TM_TM_ZONE . . . . . . . . . . . . . . . . . . . . . . 75 HAVE_SYS_DIR_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 HAVE_SYS_NDIR_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 HAVE_SYS_WAIT_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 HAVE_TM_ZONE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 HAVE_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 HAVE_TYPEOF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 HAVE_TZNAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 HAVE_UINT16_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 HAVE_UINT32_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 HAVE_UINT64_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 HAVE_UINT8_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 HAVE_UINTMAX_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 HAVE_UINTPTR_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

HAVE_UNSIGNED_LONG_LONG_INT . . . . . . . . . . . . . . . . 78

HAVE_UTIME_NULL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 HAVE_VFORK_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 HAVE_VPRINTF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 HAVE_WAIT3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 HAVE_WORKING_FORK. . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 HAVE_WORKING_VFORK . . . . . . . . . . . . . . . . . . . . . . . . . . 57

I

inline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 INT_16_BITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 int16_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 int32_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 int64_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 int8_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 intmax_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 intptr_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

L

LONG_64_BITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299

LSTAT_FOLLOWS_SLASHED_SYMLINK . . . . . . . . . . . . . . 59

M

MAJOR_IN_MKDEV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 MAJOR_IN_SYSMACROS . . . . . . . . . . . . . . . . . . . . . . . . . . 67 malloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 mbstate_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 mode_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

N

NDEBUG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 NDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 NEED_MEMORY_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 NEED_SETGID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 NLIST_NAME_UNION. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 NO_MINUS_C_MINUS_O . . . . . . . . . . . . . . . . . . . . . . . . . . 82

O

off_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

P

PACKAGE_BUGREPORT. . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PACKAGE_NAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PACKAGE_STRING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PACKAGE_TARNAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PACKAGE_URL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PACKAGE_VERSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PARAMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 pid_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 PROTOTYPES. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

R

realloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 restrict . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 RETSIGTYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304

S

SELECT_TYPE_ARG1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 SELECT_TYPE_ARG234 . . . . . . . . . . . . . . . . . . . . . . . . . . 60 SELECT_TYPE_ARG5. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 SETPGRP_VOID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 SETVBUF_REVERSED. . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 size_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 SIZEOF_type-or-expr . . . . . . . . . . . . . . . . . . . . . . . . . 79 ssize_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 STAT_MACROS_BROKEN . . . . . . . . . . . . . . . . . . . . . . . . . . 68 STDC_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 STRERROR_R_CHAR_P. . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 SVR4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 SYS_SIGLIST_DECLARED . . . . . . . . . . . . . . . . . . . . . . . 295 SYSDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 SYSNDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295

T

TIME_WITH_SYS_TIME . . . . . . . . . . . . . . . . . . . . . . . . . . 70 TM_IN_SYS_TIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 typeof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

U

uid_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 uint16_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 uint32_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 uint64_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 uint8_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 uintmax_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 uintptr_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 UMAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 UMAX4_3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 USG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305

V

variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111, 112 vfork . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 volatile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

W

WORDS_BIGENDIAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

X

X_DISPLAY_MISSING. . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

Y

YYTEXT_POINTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

Appendix B: Indices 357

B.4 Cache Variable Index

This is an alphabetical list of documented cache variables used by macros defined in Autoconf. Autoconf macros may use additional cache variables internally. To make the list easier to use, the variables are listed without their preceding ‘ac_cv_’.

A

alignof_type-or-expr . . . . . . . . . . . . . . . . . . . . . . . . 80

C

c_const . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 c_int16_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 c_int32_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 c_int64_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 c_int8_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 c_restrict. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 c_uint16_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 c_uint32_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 c_uint64_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 c_uint8_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

F

file_file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 func_chown_works. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 func_closedir_void . . . . . . . . . . . . . . . . . . . . . . . . . . 56 func_fnmatch_gnu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 func_fnmatch_works . . . . . . . . . . . . . . . . . . . . . . . 57, 62 func_function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 func_getgroups_works . . . . . . . . . . . . . . . . . . . . . . . . 58 func_getpgrp_void. . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

func_lstat_dereferences_slashed_symlink . . 59
func_lstat_empty_string_bug . . . . . . . . . . . . . . . . 61

func_malloc_0_nonnull . . . . . . . . . . . . . . . . . . . . . . . 59 func_mbrtowc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 func_memcmp_working . . . . . . . . . . . . . . . . . . . . . . . . . 60 func_mmap_fixed_mapped . . . . . . . . . . . . . . . . . . . . . . 60 func_obstack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 func_pow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 func_realloc_0_nonnull . . . . . . . . . . . . . . . . . . . . . . 60 func_setpgrp_void. . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

func_stat_empty_string_bug . . . . . . . . . . . . . . . . . 61

func_strcoll_works . . . . . . . . . . . . . . . . . . . . . . . . . . 61 func_strerror_r_char_p . . . . . . . . . . . . . . . . . . . . . . 61 func_strnlen_working . . . . . . . . . . . . . . . . . . . . . . . . 62 func_strtod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 func_strtold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 func_utime_null . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 func_working_mktime . . . . . . . . . . . . . . . . . . . . . . . . . 60

H

have_decl_symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 header_header-file . . . . . . . . . . . . . . . . . . . . . . . . . . 71 header_stdbool_h. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 header_stdc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 header_sys_wait_h. . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 header_time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

L

lib_error_at_line. . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 lib_library_function . . . . . . . . . . . . . . . . . . . . . . . . 51

M

                                                member_aggregate_member . . . . . . . . . . . . . . . . . . . . 75
                                                member_struct_stat_st_blocks . . . . . . . . . . . . . . . 74

P

path_install . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 path_mkdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 path_SED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 path_variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 prog_AWK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 prog_cc_c89 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82, 83 prog_cc_c99 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 prog_cc_compiler_c_o . . . . . . . . . . . . . . . . . . . . . . . . 82 prog_cc_stdc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 prog_EGREP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 prog_FGREP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 prog_GREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 prog_LEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 prog_variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 prog_YACC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

S

search_function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 search_getmntent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 sizeof_type-or-expr . . . . . . . . . . . . . . . . . . . . . . . . . 79 sys_posix_termios. . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

T

type_getgroups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 type_long_double. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 type_long_double_wider . . . . . . . . . . . . . . . . . . . . . . 77 type_long_long_int . . . . . . . . . . . . . . . . . . . . . . . . . . 77 type_mbstate_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 type_mode_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 type_off_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 type_pid_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 type_size_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 type_ssize_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 type_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 type_uid_t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

                                                type_unsigned_long_long_int . . . . . . . . . . . . . . . . 78

B.5 Autoconf Macro Index

This is an alphabetical list of the Autoconf macros. To make the list easier to use, the macros are listed without their preceding ‘AC_’.

A

ACT_IFELSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 AH_BOTTOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 AH_HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 AH_TEMPLATE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 AH_TOP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 AH_VERBATIM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 AIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 ALLOCA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 ARG_ARRAY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 ARG_ENABLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 ARG_PROGRAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 ARG_VAR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 ARG_WITH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 AU_ALIAS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 AU_DEFUN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 AUTOCONF_VERSION. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

B

BEFORE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181

C

C_BACKSLASH_A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 C_BIGENDIAN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 C_CHAR_UNSIGNED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 C_CONST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 C_CROSS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293

C_FLEXIBLE_ARRAY_MEMBER. . . . . . . . . . . . . . . . . . . . . 86

C_INLINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 C_LONG_DOUBLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 C_PROTOTYPES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 C_RESTRICT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 C_STRINGIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 C_TYPEOF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 C_VARARRAYS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 C_VOLATILE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 CACHE_CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 CACHE_LOAD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 CACHE_SAVE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 CACHE_VAL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 CANONICAL_BUILD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 CANONICAL_HOST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 CANONICAL_SYSTEM. . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 CANONICAL_TARGET. . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 CHAR_UNSIGNED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 CHECK_ALIGNOF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 CHECK_DECL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 CHECK_DECLS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 CHECK_DECLS_ONCE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 CHECK_FILE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 CHECK_FILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 CHECK_FUNC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 CHECK_FUNCS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 CHECK_FUNCS_ONCE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 CHECK_HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 CHECK_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 CHECK_HEADERS_ONCE . . . . . . . . . . . . . . . . . . . . . . . . . . 72 CHECK_LIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 CHECK_MEMBER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 CHECK_MEMBERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 CHECK_PROG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 CHECK_PROGS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 CHECK_SIZEOF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 CHECK_TARGET_TOOL. . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 CHECK_TARGET_TOOLS . . . . . . . . . . . . . . . . . . . . . . . . . . 49 CHECK_TOOL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 CHECK_TOOLS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 CHECK_TYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78, 294 CHECK_TYPES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 CHECKING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 COMPILE_CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 COMPILE_IFELSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 COMPUTE_INT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 CONFIG_AUX_DIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 CONFIG_COMMANDS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 CONFIG_COMMANDS_POST . . . . . . . . . . . . . . . . . . . . . . . . 40 CONFIG_COMMANDS_PRE . . . . . . . . . . . . . . . . . . . . . . . . . 40 CONFIG_FILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 CONFIG_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 CONFIG_ITEMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 CONFIG_LIBOBJ_DIR. . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 CONFIG_LINKS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 CONFIG_MACRO_DIR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 CONFIG_SRCDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 CONFIG_SUBDIRS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 CONFIG_TESTDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 CONST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 COPYRIGHT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 CROSS_CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 CYGWIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295

D

DATAROOTDIR_CHECKED . . . . . . . . . . . . . . . . . . . . . . . . . 33 DECL_SYS_SIGLIST. . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 DECL_YYTEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 DEFINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 DEFINE_UNQUOTED. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 DEFUN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 DEFUN_ONCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 DIAGNOSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 DIR_HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295

DISABLE_OPTION_CHECKING . . . . . . . . . . . . . . . . . . . 276

DYNIX_SEQ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296

E

EGREP_CPP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 EGREP_HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 EMXOS2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 ENABLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 ERLANG_CHECK_LIB. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 ERLANG_NEED_ERL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 ERLANG_NEED_ERLC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 ERLANG_PATH_ERL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 ERLANG_PATH_ERLC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 ERLANG_SUBST_ERTS_VER . . . . . . . . . . . . . . . . . . . . . . . 96

ERLANG_SUBST_INSTALL_LIB_DIR . . . . . . . . . . . 32, 97
ERLANG_SUBST_INSTALL_LIB_SUBDIR . . . . . . . . 32, 97

ERLANG_SUBST_LIB_DIR . . . . . . . . . . . . . . . . . . . . . . . . 96 ERLANG_SUBST_ROOT_DIR . . . . . . . . . . . . . . . . . . . . . . . 96 ERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 EXEEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296

F

F77_DUMMY_MAIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 F77_FUNC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 F77_LIBRARY_LDFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . 90 F77_MAIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 F77_WRAPPERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 FATAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 FC_FIXEDFORM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 FC_FREEFORM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 FC_FUNC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 FC_LIBRARY_LDFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . 90 FC_MAIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 FC_SRCEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 FC_WRAPPERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 FIND_X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 FIND_XTRA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 FOREACH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 FUNC_ALLOCA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 FUNC_CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 FUNC_CHOWN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 FUNC_CLOSEDIR_VOID . . . . . . . . . . . . . . . . . . . . . . . . . . 56 FUNC_ERROR_AT_LINE . . . . . . . . . . . . . . . . . . . . . . . . . . 57 FUNC_FNMATCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 FUNC_FNMATCH_GNU. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 FUNC_FORK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 FUNC_FSEEKO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 FUNC_GETGROUPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 FUNC_GETLOADAVG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 FUNC_GETMNTENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 FUNC_GETPGRP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 FUNC_LSTAT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

FUNC_LSTAT_FOLLOWS_SLASHED_SYMLINK . . . . . . . . 59

FUNC_MALLOC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 FUNC_MBRTOWC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 FUNC_MEMCMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 FUNC_MKTIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 FUNC_MMAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 FUNC_OBSTACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 FUNC_REALLOC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 FUNC_SELECT_ARGTYPES . . . . . . . . . . . . . . . . . . . . . . . . 60 FUNC_SETPGRP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 FUNC_SETVBUF_REVERSED . . . . . . . . . . . . . . . . . . . . . . 296 FUNC_STAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 FUNC_STRCOLL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 FUNC_STRERROR_R . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 FUNC_STRFTIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 FUNC_STRNLEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 FUNC_STRTOD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 FUNC_STRTOLD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 FUNC_UTIME_NULL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 FUNC_VPRINTF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 FUNC_WAIT3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296

G

GCC_TRADITIONAL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 GETGROUPS_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 GETLOADAVG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 GNU_SOURCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297

H

HAVE_FUNCS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 HAVE_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 HAVE_LIBRARY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 HAVE_POUNDBANG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 HEADER_ASSERT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 HEADER_CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 HEADER_DIRENT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 HEADER_EGREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 HEADER_MAJOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 HEADER_RESOLV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 HEADER_STAT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 HEADER_STDBOOL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 HEADER_STDC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 HEADER_SYS_WAIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 HEADER_TIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 HEADER_TIOCGWINSZ. . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 HELP_STRING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297

I

INCLUDES_DEFAULT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 INIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 297 INLINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 INT_16_BITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 IRIX_SUN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 ISC_POSIX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298

L

LANG_ASSERT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 LANG_C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 LANG_CALL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 LANG_CONFTEST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 LANG_CPLUSPLUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 LANG_FORTRAN77 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 LANG_FUNC_LINK_TRY . . . . . . . . . . . . . . . . . . . . . . . . . 104 LANG_POP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 LANG_PROGRAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 LANG_PUSH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 LANG_RESTORE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 LANG_SAVE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 LANG_SOURCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 LANG_WERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 LIBOBJ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 LIBSOURCE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 LIBSOURCES. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 LINK_FILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 LINK_IFELSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 LN_S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 LONG_64_BITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 LONG_DOUBLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 LONG_FILE_NAMES. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299

M

MAJOR_HEADER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 MEMORY_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 MINGW32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 MINIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 MINUS_C_MINUS_O. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 MMAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 MODE_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 MSG_CHECKING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 MSG_ERROR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 MSG_FAILURE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 MSG_NOTICE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 MSG_RESULT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 MSG_WARN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

O

OBJEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 OBSOLETE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 299 OFF_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 OPENMP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 OUTPUT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22, 300 OUTPUT_COMMANDS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300

P

PACKAGE_BUGREPORT. . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PACKAGE_NAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PACKAGE_STRING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PACKAGE_TARNAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PACKAGE_URL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PACKAGE_VERSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 PATH_PROG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 PATH_PROGS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

PATH_PROGS_FEATURE_CHECK . . . . . . . . . . . . . . . . . . . 50

PATH_TARGET_TOOL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 PATH_TOOL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 PATH_X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 PATH_XTRA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 PID_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 PREFIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 PREFIX_DEFAULT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 PREFIX_PROGRAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 PREPROC_IFELSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 PREREQ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 PRESERVE_HELP_ORDER . . . . . . . . . . . . . . . . . . . . . . . . 271 PROG_AWK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 PROG_CC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 PROG_CC_C_O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 PROG_CC_C89 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 PROG_CC_C99 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 PROG_CC_STDC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 PROG_CPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 PROG_CPP_WERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 PROG_CXX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 PROG_CXX_C_O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 PROG_CXXCPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 PROG_EGREP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 PROG_F77 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 PROG_F77_C_O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 PROG_FC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 PROG_FC_C_O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 PROG_FGREP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 PROG_GCC_TRADITIONAL . . . . . . . . . . . . . . . . . . . . . . . . 86 PROG_GREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 PROG_INSTALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 PROG_LEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 PROG_LN_S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 PROG_MAKE_SET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 PROG_MKDIR_P . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 PROG_OBJC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 PROG_OBJCPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 PROG_OBJCXX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 PROG_OBJCXXCPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 PROG_RANLIB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 PROG_SED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 PROG_YACC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 PROGRAM_CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 PROGRAM_EGREP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 PROGRAM_PATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 PROGRAMS_CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 PROGRAMS_PATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301

R

REMOTE_TAPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 REPLACE_FNMATCH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 REPLACE_FUNCS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 REQUIRE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 REQUIRE_AUX_FILE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 REQUIRE_CPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 RESTARTABLE_SYSCALLS . . . . . . . . . . . . . . . . . . . . . . . 301 RETSIGTYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 REVISION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 RSH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 RUN_IFELSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107

S

SCO_INTL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 SEARCH_LIBS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 SET_MAKE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 SETVBUF_REVERSED. . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 SIZE_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 SIZEOF_TYPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 ST_BLKSIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 ST_BLOCKS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 ST_RDEV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 STAT_MACROS_BROKEN . . . . . . . . . . . . . . . . . . . . . . . . . 302 STDC_HEADERS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 STRCOLL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 STRUCT_DIRENT_D_INO . . . . . . . . . . . . . . . . . . . . . . . . . 74 STRUCT_DIRENT_D_TYPE . . . . . . . . . . . . . . . . . . . . . . . . 74 STRUCT_ST_BLKSIZE . . . . . . . . . . . . . . . . . . . . . . . . . . 302 STRUCT_ST_BLOCKS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 STRUCT_ST_RDEV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 STRUCT_TIMEZONE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 STRUCT_TM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 SUBST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 SUBST_FILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 SYS_INTERPRETER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 SYS_LARGEFILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 SYS_LONG_FILE_NAMES . . . . . . . . . . . . . . . . . . . . . . . . . 95 SYS_POSIX_TERMIOS. . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

SYS_RESTARTABLE_SYSCALLS . . . . . . . . . . . . . . . . . . 302

SYS_SIGLIST_DECLARED . . . . . . . . . . . . . . . . . . . . . . . 302

T

TEST_CPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 TEST_PROGRAM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 TIME_WITH_SYS_TIME . . . . . . . . . . . . . . . . . . . . . . . . . 303 TIMEZONE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 TRY_ACT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 TRY_COMPILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 TRY_CPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 TRY_LINK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 TRY_LINK_FUNC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 TRY_RUN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 TYPE_GETGROUPS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 TYPE_INT16_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 TYPE_INT32_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 TYPE_INT64_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 TYPE_INT8_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 TYPE_INTMAX_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 TYPE_INTPTR_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 TYPE_LONG_DOUBLE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 TYPE_LONG_DOUBLE_WIDER . . . . . . . . . . . . . . . . . . . . . . 77 TYPE_LONG_LONG_INT . . . . . . . . . . . . . . . . . . . . . . . . . . 77 TYPE_MBSTATE_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 TYPE_MODE_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 TYPE_OFF_T. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 TYPE_PID_T. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 TYPE_SIGNAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 TYPE_SIZE_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 TYPE_SSIZE_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 TYPE_UID_T. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 TYPE_UINT16_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 TYPE_UINT32_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 TYPE_UINT64_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 TYPE_UINT8_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 TYPE_UINTMAX_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 TYPE_UINTPTR_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

                                                TYPE_UNSIGNED_LONG_LONG_INT . . . . . . . . . . . . . . . . 78

U

UID_T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 UNISTD_H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 USE_SYSTEM_EXTENSIONS . . . . . . . . . . . . . . . . . . . . . . . 95 USG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 UTIME_NULL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305

V

                                                VALIDATE_CACHED_SYSTEM_TUPLE . . . . . . . . . . . . . . 305

VERBOSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 VFORK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 VPRINTF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305

W

WAIT3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 WARN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 WARNING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 WITH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305 WORDS_BIGENDIAN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305

X

XENIX_DIR. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 305

Y

YYTEXT_POINTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306

B.6 M4 Macro Index

This is an alphabetical list of the M4, M4sugar, and M4sh macros. To make the list easier to use, the macros are listed without their preceding ‘m4_’ or ‘AS_’.

__file__. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136  __line__. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136

__oline__. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136

A

append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 append_uniq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 append_uniq_w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 apply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 argn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 assert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138

B

bmatch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 BOURNE_COMPATIBLE . . . . . . . . . . . . . . . . . . . . . . . . . . 171 BOX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 bpatsubst. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 bpatsubsts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 bregexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 builtin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

C

car . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 CASE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 cdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 changecom. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 changequote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 chomp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 chomp_all. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 cleardivert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 cmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 combine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 cond . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 copy_force . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 curry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149

D

debugfile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 debugmode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 decr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 default_nblank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 default_nblank_quoted . . . . . . . . . . . . . . . . . . . . . . 143 default_quoted . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 define . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 defn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 DIRNAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 divert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 divert_once . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 divert_pop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 divert_push . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 divert_text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 divnum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 dnl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 do. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 dquote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 dquote_elt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 dumpdef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 dumpdefs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

E

echo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 ECHO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 ECHO_N . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 errprint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 errprintn. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 escape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 ESCAPE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 esyscmd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 esyscmd_s. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 EXIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 expand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150

F

fatal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 flatten . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 for . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 foreach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 foreach_w. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

H

HELP_STRING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274

I

if. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 IF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 ifblank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 ifdef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 ifnblank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 ifndef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 ifset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 ifval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 ifvaln . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 ignore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 incr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 indir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 INIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 INIT_GENERATED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171

J

join . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 joinall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155

L

len . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 LINENO_PREPARE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 list_cmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 LITERAL_IF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 LITERAL_WORD_IF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 location . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139

M

make_list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 maketemp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 map_args . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 map_args_pair . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 map_args_sep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 map_args_w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 map_sep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 mapall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 mapall_sep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 ME_PREPARE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 MESSAGE_FD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 MESSAGE_LOG_FD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 MKDIR_P . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 mkstemp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

N

n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 newline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 normalize. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155

O

ORIGINAL_STDIN_FD . . . . . . . . . . . . . . . . . . . . . . . . . . 173

P

pattern_allow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 pattern_forbid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 popdef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 pushdef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

Q

quote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151

R

re_escape. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 rename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 rename_force . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 reverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152

S

set_add . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 set_add_all . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 SET_CATFILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 set_contains . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 set_contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 set_delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 set_difference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 set_dump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 set_empty. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 set_foreach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 set_intersection. . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 set_list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 set_listc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 set_map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 set_map_sep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 set_remove . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 set_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 SET_STATUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 set_union. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 SHELL_SANITIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 shift . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 shift2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 shift3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 shiftn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 sign . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 sinclude . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 split . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 stack_foreach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 stack_foreach_lifo . . . . . . . . . . . . . . . . . . . . . . . . . 148 stack_foreach_sep . . . . . . . . . . . . . . . . . . . . . . . . . . 148 stack_foreach_sep_lifo. . . . . . . . . . . . . . . . . . . . . 148 strip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 substr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 syscmd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 sysval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

T

text_box . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 text_wrap. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 tolower . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 toupper . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 TR_CPP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 TR_SH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 traceoff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 traceon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 translit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

U

undefine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 undivert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 unquote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 UNSET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167

V

VAR_APPEND . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 VAR_ARITH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 VAR_COPY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 VAR_IF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 VAR_POPDEF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 VAR_PUSHDEF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 VAR_SET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 VAR_SET_IF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 VAR_TEST_SET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 version_compare. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 VERSION_COMPARE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 version_prereq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158

W

warn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 wrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 wrap_lifo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138

B.7 Autotest Macro Index

This is an alphabetical list of the Autotest macros. To make the list easier to use, the macros are listed without their preceding ‘AT_’.

A

ARG_OPTION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 ARG_OPTION_ARG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319

B

BANNER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319

C

CAPTURE_FILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320 CHECK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 CHECK_EUNIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 CHECK_UNQUOTED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 CLEANUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 COLOR_TESTS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 COPYRIGHT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318

D

DATA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321

F

FAIL_IF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320

I

INIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318

K

KEYWORDS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320

P

PACKAGE_BUGREPORT . . . . . . . . . . . . . . . . . . . . . . . . . . 325 PACKAGE_NAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 PACKAGE_STRING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 PACKAGE_TARNAME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 PACKAGE_URL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 PACKAGE_VERSION. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325

S

SETUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320 SKIP_IF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320

T

TESTED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319

X

XFAIL_IF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320

B.8 Program and Function Index

This is an alphabetical list of the programs and functions whose portability is discussed in this document.

!_{! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211}  ._{. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210}

/

/usr/bin/ksh on Solaris . . . . . . . . . . . . . . . . . . . . . . 188 /usr/dt/bin/dtksh on Solaris . . . . . . . . . . . . . . . . 189 /usr/xpg4/bin/sh on Solaris . . . . . . . . . . . . . . . . . 188

{

{...} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211

A

alloca . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 ‘alloca.h’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 ‘assert.h’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224

B

basename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 break . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212

C

case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 cat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 cc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 cd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 chgrp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 chmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 chown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 chown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 closedir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 cmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 cp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 ‘ctype.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

D

date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 diff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 ‘dirent.h’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 dirname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228

E

echo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 egrep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 error_at_line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 exec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 export . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 expr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 expr (‘|’) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229

F

false . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 fgrep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 find . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 ‘float.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 fnmatch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57, 62 ‘fnmatch.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 for . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 fork . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 fseeko . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 ftello . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

G

getgroups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 getloadavg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 getmntent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 getpgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 getpgrp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 grep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231

I

if. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 ‘inttypes.h’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 76 isinf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 isnan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

J

join . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232

K

‘ksh’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 ‘ksh88’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 ‘ksh93’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188

L

‘linux/irda.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 ‘linux/random.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 ln. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 ls. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 lstat. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 61

M

make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 malloc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53, 59 mbrtowc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 memcmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 mkdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 mkfifo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 mknod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 mktemp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 mktime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 mmap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 mv. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234

N

‘ndir.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 ‘net/if.h’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 ‘netinet/if_ether.h’ . . . . . . . . . . . . . . . . . . . . . . . . . 65 ‘nlist.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

O

od. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234

P

‘pdksh’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 printf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 putenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 pwd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218

R

read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 realloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54, 60 ‘resolv.h’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 rm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 rmdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234

S

sed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 sed (‘t’). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 select . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 setpgrp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 setvbuf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 shift . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 sigaction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 signal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 ‘signal.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 sleep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 snprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 source . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 sprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 sscanf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 stat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 ‘stdarg.h’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 ‘stdbool.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 ‘stdint.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 76 ‘stdlib.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65, 68, 76 strcoll . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 strerror_r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54, 61 strftime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 ‘string.h’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 ‘strings.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 strnlen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54, 62 strtod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 strtold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 ‘sys/dir.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 ‘sys/ioctl.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 ‘sys/mkdev.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 ‘sys/mount.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 ‘sys/ndir.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 ‘sys/ptem.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 ‘sys/socket.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 ‘sys/stat.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 ‘sys/sysmacros.h’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 ‘sys/time.h’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70, 74 ‘sys/types.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 ‘sys/ucred.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 ‘sys/wait.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 sysconf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 ‘system.h’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

T

tar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 ‘termios.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 ‘time.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70, 74 touch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 tr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 trap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 true . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223

U

‘unistd.h’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 unlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 unset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 unsetenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 utime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

V

va_copy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 va_list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 vfork . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 ‘vfork.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 vprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 vsnprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 vsprintf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54, 62

W

wait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 wait3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 ‘wchar.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

X

                                                ‘X11/extensions/scrnsaver.h’ . . . . . . . . . . . . . . . . 66

B.9 Concept Index

This is an alphabetical list of the files, tools, and concepts introduced in this document.

"

‘"$@"’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196

$

$((expression)) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 $(commands). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 $<, explicit rules, and VPATH . . . . . . . . . . . . . . . . . . 248 ${#var} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 ${var##word}. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 ${var#word} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 ${var%%word}. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 ${var%word} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 ${var:-value} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 ${var=expanded-value} . . . . . . . . . . . . . . . . . . . . . . 197 ${var=literal}. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197

@

‘@%:@’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 ‘@&t@’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 ‘@:>@’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 ‘@:}@’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 ‘@<:@’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 ‘@{:@’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 ‘@S|@’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126

^

^quoting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200

_m4_divert_diversion . . . . . . . . . . . . . . . . . . . . . . . 309

‘

‘commands‘ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198

6

64-bit libraries. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279

A

absolute file names, detect . . . . . . . . . . . . . . . . . . . . 193 ‘acconfig.h’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 ‘aclocal.m4’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Ash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 autoconf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Autoconf upgrading . . . . . . . . . . . . . . . . . . . . . 306, 308 Autoconf version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 autoheader. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Autoheader macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 autom4te debugging tips. . . . . . . . . . . . . . . . . . . . . . 162 Autom4te Library . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 ‘autom4te.cache’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 ‘autom4te.cfg’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 Automake . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Automatic remaking . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 automatic rule rewriting and VPATH . . . . . . . . . . . 248 autopoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 autoreconf. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 autoscan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Autotest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 AUTOTEST_PATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 autoupdate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292

B

Back trace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14, 131 balancing parentheses . . . . . . . . . . . . . . . . . . . . . . . . 127 Bash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 Bash 2.05 and later. . . . . . . . . . . . . . . . . . . . . . . . . . . 188 Bootstrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 BSD make and ‘obj/’ . . . . . . . . . . . . . . . . . . . . . . . . . 247 buffer overruns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 Build directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

C

C function portability . . . . . . . . . . . . . . . . . . . . . . . . . 52 C types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Cache . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 Cache variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 Cache, enabling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 Canonical system type. . . . . . . . . . . . . . . . . . . . . . . . 268 carriage return, deleting . . . . . . . . . . . . . . . . . . . . . . 239 changequote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 Coding style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 Command Substitution . . . . . . . . . . . . . . . . . . . . . . . 198 Commands for configuration . . . . . . . . . . . . . . . . . . . 39 Comments in ‘Makefile’ rules . . . . . . . . . . . . . . . . 246 Common autoconf behavior . . . . . . . . . . . . . . . . . . . . 43 Compilers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 composing variable names . . . . . . . . . . . . . . . . . . . . 170 ‘config.h’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 ‘config.h.bot’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 ‘config.h.in’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 ‘config.h.top’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 config.site . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 config.status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 config.sub . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 Configuration actions . . . . . . . . . . . . . . . . . . . . . . . . . . 23 Configuration commands . . . . . . . . . . . . . . . . . . . . . . 39 Configuration file creation . . . . . . . . . . . . . . . . . . . . . 25 Configuration Header . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Configuration Header Template . . . . . . . . . . . . . . . . 36 Configuration links . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 configure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 281 Configure subdirectories . . . . . . . . . . . . . . . . . . . . . . . 40 ‘configure.ac’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ‘configure.in’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Copyright Notice . . . . . . . . . . . . . . . . . . . . . . . . . 20, 318 Creating configuration files . . . . . . . . . . . . . . . . . . . . 25 Creating temporary files . . . . . . . . . . . . . . . . . . . . . . 233 Cross compilation . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310

D

Darwin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 Data structure, set . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 ‘datarootdir’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 debugging tips. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 Declaration, checking . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Default includes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 deleting carriage return . . . . . . . . . . . . . . . . . . . . . . . 239 Dependencies between macros . . . . . . . . . . . . . . . . 178 descriptors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 Descriptors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 Directories, build . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 Directories, installation . . . . . . . . . . . . . . . . . . . . . . . . 29 division, integer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 dnl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175, 183 double-colon rules and VPATH . . . . . . . . . . . . . . . . . 248

E

Endianness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 Erlang. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 Erlang, Library, checking . . . . . . . . . . . . . . . . . . . . . . 96 exiting portably . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 expanded before required . . . . . . . . . . . . . . . . . . . . . 334 explicit rules, $<, and VPATH . . . . . . . . . . . . . . . . . . 248 External software. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271

F

F77 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 FHS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 file descriptors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 File descriptors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 File system conventions. . . . . . . . . . . . . . . . . . . . . . . 193 File, checking. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Filesystem Hierarchy Standard . . . . . . . . . . . . . . . 279 floating point. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 Forbidden patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 Fortran. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 Function, checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

G

Gettext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 GNU build system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Gnulib . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

H

Header portability . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Header templates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 Header, checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Help strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 Here-documents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 History of autoconf . . . . . . . . . . . . . . . . . . . . . . . . . . . 339

I

ifnames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Imake . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 330 Includes, default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 indirection, variable name . . . . . . . . . . . . . . . . . . . . 168 input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 Install prefix. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Installation directories . . . . . . . . . . . . . . . . . . . . . . . . . 29 Instantiation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 integer overflow . . . . . . . . . . . . . . . . . . . . . . . . . . 258, 261 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

K

Korn shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 Ksh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188

L

Language. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 Large file support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 LFS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 lib64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 Library, checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Libtool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 329 Limitations of make . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 Limitations of shell builtins . . . . . . . . . . . . . . . . . . . 210 Limitations of usual tools . . . . . . . . . . . . . . . . . . . . . 224 Links. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 Links for configuration. . . . . . . . . . . . . . . . . . . . . . . . . 40 Listing directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 loop induction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 low-level output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172

M

M4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 M4 quotation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 M4sugar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 m4sugar debugging tips. . . . . . . . . . . . . . . . . . . . . . . 162 Macro invocation stack . . . . . . . . . . . . . . . . . . . 14, 131 Macros, called once . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 Macros, obsoleting. . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 Macros, ordering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 Macros, prerequisites . . . . . . . . . . . . . . . . . . . . . . . . . 178 make -k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 make and MAKEFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . 244 make and SHELL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 ‘Makefile’ rules and comments . . . . . . . . . . . . . . . 246 ‘Makefile’ rules and newlines . . . . . . . . . . . . . . . . . 246 Makefile substitutions . . . . . . . . . . . . . . . . . . . . . . . . . 25 MAKEFLAGS and make . . . . . . . . . . . . . . . . . . . . . . . . . . 244 Making directories. . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 Messages, from autoconf . . . . . . . . . . . . . . . . . . . . . 177 Messages, from configure . . . . . . . . . . . . . . . . . . . . 118 Messages, from M4sugar . . . . . . . . . . . . . . . . . . . . . . 138 Moving open files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234

N

newline, deleting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 Newlines in ‘Makefile’ rules . . . . . . . . . . . . . . . . . . 246 Notices in configure . . . . . . . . . . . . . . . . . . . . . . . . . . 20 null pointers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262

O

‘obj/’, subdirectory . . . . . . . . . . . . . . . . . . . . . . . . . . 247 Obsolete constructs. . . . . . . . . . . . . . . . . . . . . . . . . . . 291 Obsoleting macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 obstack. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 One-shot macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 Options, package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 Options, Package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 Ordering macros. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 Output variables . . . . . . . . . . . . . . . . . . . . . . . . . . 25, 112 Output variables, special characters in . . . . . . . . 114 output, low-level . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 Outputting files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 overflow, signed integer . . . . . . . . . . . . . . . . . . 258, 261

P

Package options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 ‘package.m4’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 Parallel make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 parentheses, balancing . . . . . . . . . . . . . . . . . . . . . . . . 127 Patterns, forbidden . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 polymorphic variable name . . . . . . . . . . . . . . . . . . . 168 portability . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 Portability of C functions . . . . . . . . . . . . . . . . . . . . . . 52 Portability of headers . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Portable C and C++ programming . . . . . . . . . . . . 257 Portable shell programming. . . . . . . . . . . . . . . . . . . 187 positional parameters . . . . . . . . . . . . . . . . . . . . . . . . . 197 Posix termios headers. . . . . . . . . . . . . . . . . . . . . . . . . . 95 Precious Variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 Prefix for install . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 preprocessor arithmetic . . . . . . . . . . . . . . . . . . . . . . . 262 Preprocessors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 prerequisite directories and VPATH. . . . . . . . . . . . . 251 Prerequisite macros. . . . . . . . . . . . . . . . . . . . . . . . . . . 178 Program names, transforming. . . . . . . . . . . . . . . . . 276 Programs, checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

Q

QNX 4.25 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 quadrigraphs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 quotation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9, 121

R

Remaking automatically . . . . . . . . . . . . . . . . . . . . . . . 34 Revision. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Rule, Single Suffix Inference . . . . . . . . . . . . . . . . . . 254

S

Separated Dependencies . . . . . . . . . . . . . . . . . . . . . . 254 set -b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 set -e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 set -m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 Set manipulation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 SHELL and make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 Shell assignments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 Shell builtins . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 Shell file descriptors . . . . . . . . . . . . . . . . . . . . . . . . . . 190 Shell Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 Shell here-documents . . . . . . . . . . . . . . . . . . . . . . . . . 189 Shell parentheses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 Shell pattern matching . . . . . . . . . . . . . . . . . . . . . . . 195 Shell slashes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 Shell substitutions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 Shell variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 Shellology . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 signed integer overflow . . . . . . . . . . . . . . . . . . . 258, 261 Single Suffix Inference Rule . . . . . . . . . . . . . . . . . . . 254 Site defaults . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 Site details . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 Special shell variables. . . . . . . . . . . . . . . . . . . . . . . . . 202 standard input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 Standard symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 Structure, checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Subdirectory configure . . . . . . . . . . . . . . . . . . . . . . . . . 40 Substitutions in makefiles . . . . . . . . . . . . . . . . . . . . . . 25 Symbolic links. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 System type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267, 268 Systemology. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108

T

Target triplet. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 termios Posix headers. . . . . . . . . . . . . . . . . . . . . . . . . . 95 test group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315 testsuite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 315, 323 timestamp resolution. . . . . . . . . . . . . . . . 227, 238, 254 Transforming program names . . . . . . . . . . . . . . . . . 276 Tru64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

U

unbalanced parentheses, managing . . . . . . . . . . . . 127 undefined macro. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 Unix version 7. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 Unordered set manipulation . . . . . . . . . . . . . . . . . . 158 Upgrading autoconf. . . . . . . . . . . . . . . . . . . . . . 306, 308

V

V7 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 variable name indirection . . . . . . . . . . . . . . . . . . . . . 168 variable names, composing. . . . . . . . . . . . . . . . . . . . 170 Variable, Precious . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 variables and VPATH . . . . . . . . . . . . . . . . . . . . . . . . . . 248 Version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 version, Autoconf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 volatile objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 VPATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 VPATH and automatic rule rewriting . . . . . . . . . . . 248 VPATH and double-colon rules . . . . . . . . . . . . . . . . . 248 VPATH and prerequisite directories. . . . . . . . . . . . . 251 VPATH and variables . . . . . . . . . . . . . . . . . . . . . . . . . . 248 VPATH, explicit rules, and $< . . . . . . . . . . . . . . . . . . 248 VPATH, resolving target pathnames . . . . . . . . . . . . 251

W

wraparound arithmetic. . . . . . . . . . . . . . . . . . . 258, 261

X

X Window System. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

Z

Zsh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189